A Complete Bibliography of Publications in American Mathematical Monthly: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.11

Title word cross-reference

(1 − 1)^{n} = 0 [2483]. (C^{∞}) [79]. (l^{p})′ ⊂ l^{q} [1315]. (Z/2Z) ∗ (Z/3Z) [1344]. $1 [1832]. 1 [2357]. 1 + (1/2)^{2} + (1/3)^{2} + ⋯ = π^{2}/6 [658]. 1 + 1/2^{2} + 1/3^{2} + ⋯ = π^{2}/6 [1040]. 1/π [2670]. 12 [927, 73]. 2 [772]. 2 + ε [2553]. 3 [729, 106, 74, 948]. 4 [2141]. ^{2} [42]. ^{3} [683]. ^{n} [77, 53]. ^{x} [283]. ^{x2} [283]. _{1}ψ_{1} [1399]. _{2} [581]. a(a + d)(a + 2d)(a + 3d) = x2 [63]. a^{4} + b^{4} + c^{4} + d^{4} = (a + b + c + d)^{4} [2313]. A^{n} − 1 [1232]. AB = ωBA [1368]. ax + by [2347]. β [1121]. C(K) [2182]. C^{∞} [1591]. C^{n} [103]. cos \sqrt{x} [1341]. Δ [1485, 454, 39]. e [1726, 1727, 1889]. ℓ_{p} [2758]. ϵ [454, 39]. f ∘ g = f [748]. GM ≤ AMF [2219]. ∫_{a}^{b} x^{p} dx [361]. k [668, 2471, 35]. L^{1} [1425]. λ [1759].

( \begin{matrix}n & d \\ m & d \end{matrix} ) ≡ ( \begin{matrix}n \\ m \end{matrix} r) (mod q) [2527]. ln(4/π) [1460]. ln 2 [388]. ln π/2

[1643]. m [1035, 829]. C [1266]. E^{4} [1462]. F_{1} [1312]. R^{n} [1516]. n! [177]. N [956, 76, 101, 1402, 2770, 637, 2373, 2357, 543, 1825, 1052, 820, 1417]. nx mod 1 [887]. p [1512, 534]. p ≡ 3 (mod 8) [147]. ϕ [424]. π [2649, 1838, 2475, 388, 2758, 2652, 2551, 2447, 1878, 2498, 1643, 2233]. π/2 [54]. π/4 [1104]. π_{p} [2758]. ψ(x) [712]. PSL(2, 7) ≅ GL(3, 2) [2704]. q [533, 2124]. R^{3} [1318]. sec θ [816]. σk(n) [1825]. sinπz = πz ∏_{k=1}^{∞}(1 − z^{2}/k^{2}) [542]. SO_{3}(R) [1344]. \sqrt{2} [2525]. ∑ 1/n^{2} [2448]. ∑_{k=−∞}^{∞}(4k + 1)^{−n} [1060]. ∑ \frac{1}{n} = ∞ [2205]. t [1847]. x [2523, 39, 62]. x^{2} + 2y^{2} [147]. x^{y} = y^{mx} [1173]. y = x^{1/x} [594]. y = x^{α} sin(1/x) [360]. y′ = f(x, y) [2523]. ζ(2) [388]. ζ(2m) [1288]. ζ(2p) [1179]. ζ(3) [388].

-1 [295]. -3 [295]. -Adic [534, 1512]. -Analogue [2124]. -Balls [76, 101, 543]. -Boundedness [1425]. -by- [772]. -Catalan [1847]. -Colored [106, 74]. -Cube [1402]. -Cycles [948]. -Dependence [2471]. -Dimensional [2373]. -Divergent [1417]. -Eulerian [533]. -Expansions [1121]. -Function [1485]. -Gons [1052, 956]. -Permutations [1759]. -Ring [1594]. -Rowed [2141]. -Simplex [820]. -Space [1035, 729]. -Spaces [2182]. -Tone [927]. -Tuples [668].

0.999 [2074].

1/M [1727]. 11212/11220 [2354]. 1794-1850 [1513]. 18 [2225]. 1809 [1839]. 1890-1950 [506]. 19th [2591].

2000-Year-Old [141]. 2001 [782]. 2003 [1337]. 20th [512]. 21st [321]. 2d [42]. 2i [2550].

3-D [2510]. 3d [42]. 3n [2420].

90s [232].

= [42, 228, 1012].

A-G [2195]. A. [1867]. A5 [2457]. ABC [724]. ABCD [1115]. Abel [2289, 2674]. Abelian [2755, 2234, 448, 933]. Abigail [351]. Abnormal [537]. Abreast [957]. Absolutely [2076, 537]. Abstract [1584]. Acceleration [72]. Accidental [2008]. Ace [351]. Actions [2647]. Actual [2346]. Acute [159]. Acyclic [2706, 331]. Ad [1397]. Adams [1535, 351, 2616]. Adding [892, 693]. Additive [858, 1385, 1715]. Additivity [1546]. Adic [534, 1512]. Adjacency [1301]. Adjoint [1604]. Adjugate [2235]. Admissions [1006]. Advanced [2696, 1200, 2524]. Affine [2365, 356, 597]. After [1095, 1256, 2414]. Again [2707, 536, 958]. Age [1937]. AGM [2746]. Ago [641]. Aid [758]. Alan [1908]. Alex [2591]. Alexander [1962, 1948]. Alexanderson [1510]. Alfred [417, 1828]. Algebra [1941, 417, 2152, 1064, 1487, 1535, 1808, 1200, 2752, 1640, 1149, 756, 2524, 268, 1920, 420, 1914, 2730, 540, 1463, 1535, 2362, 1584]. Algebraic [2445, 2142]. Algebraically [247]. Algebro [1940]. Algebro-Geometric [1940]. Algorithm [1775]. Algorithmic [1233]. Algorithms [1803, 2475, 67]. Alice [1256]. Alien [1993]. All-Ones [2008]. Allman [1688]. Almost [1617, 2671, 1912, 1121, 610, 362]. Along [2022, 374, 1969]. Alternating [2354, 774, 2625, 2098, 1912, 944, 1012, 654]. Alternations [2596]. Alternative [2175, 568, 447, 482, 7, 34]. Altitudes [1088, 280, 1634, 2718]. Altmann [2198]. Always [931, 419]. AM [2166]. AM-GM [2166]. Amazing [2728]. American [506]. Amicable [670]. among [289]. Amy [2464]. Anagrammatical [1165]. Analog [1460]. Analogue [2420, 2124]. Analysis [1700, 1145, 90, 2289, 2696, 1718, 2038, 2468, 2340, 925, 2524, 2120, 2091, 2491, 2696, 1718, 2567, 2038]. Analytic [2127, 2549, 980, 2421, 79, 266, 1646, 1742, 1635]. Analyticity [450]. Ancient [1810]. André [2346]. Angle [750, 2345, 820, 1354]. Angles [1469, 2463, 2705, 1077]. Anglesio [1427]. Angleworms [2261]. Animals [970]. Anita [1828]. Anne [1747]. Answers [1368]. Anthony [2696]. Anti [70, 44, 65]. Anti-Social [44, 65]. Anti-Squares [70]. Antichains [1245, 1246]. Antiderivative [510]. Antiquity [1117]. Antonis [1718]. Ants [1989]. Appear [1182]. Appearance [1404, 1079]. Appell [1134]. Appell-ing [1134]. Application [1849, 1655, 979, 1300, 2732, 1500, 2722, 130, 2346]. Applications [1535, 706, 2340, 2617]. Applied [2312, 2616, 2362]. Apply [2037]. Appolonius [698]. Apportionment [2339, 1257]. Approach [655, 1459, 2174, 230, 1145, 1545, 2393, 1535, 143, 1619, 2100, 2343, 1541, 1827]. Approaches [172, 2175]. Approximated [1833]. Approximating [1960, 1754, 994]. Approximation [419, 1143, 177]. Approximations [1050, 2551, 292]. Arbelos [1778, 916]. Arbitrariness [1836]. Arc [1361]. Archimedean [299]. Archimedes [1313, 2547]. Arctan [1162]. Arctangent [2649, 1754]. Arctangents [917, 1263]. Area [322, 1063, 1486, 844, 2412]. Areas [871, 2672, 481]. Arena [2228]. Argument [929]. Arithmetic [2317, 424, 2138, 884, 2397, 306, 1716, 1474, 957, 842, 2677, 2149]. Arithmetic-Geometric [2149]. Arrays [583]. Art [102, 1609]. Arthur [2747]. Arzelà [1147]. Ascending [1594]. Ascoli [1147]. Ash [2279]. Assessment [2497]. Assistance [2331]. Associated [2688, 1863]. Associativeness [1848]. Associativity [2759, 678]. Assumptions [396]. Aste [1746]. Asymmetric [1966]. Asymptotic [1835, 2259, 1150, 1423]. Asymptotics [131, 214, 1684, 1135]. Atom [2170]. Atoms [858]. Attacking [2612]. Attainable [1993]. Attending [1968]. Attribution [1165]. Aubel [2695]. Authority [1421]. Autobiography [1988]. Automorphism [1613]. Automorphisms [2234]. Average [407, 1657, 2411, 2496]. Averaging [200]. Avner [2279]. Avoidance [1389]. Avoiding [1566, 127]. Award [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Awards [225, 528]. Axiom [760, 2283]. Axis [2753]. Ayoub [1882].

B [2116, 1535, 2484]. BA [2484]. Back [28, 319, 59, 99, 138, 168, 196, 226, 257, 287, 353, 616, 383, 415, 445, 473, 502, 529, 560, 588, 646, 896, 674, 702, 727, 754, 780, 810, 839, 867, 923, 1170, 951, 977, 1003, 1029, 1058, 1085, 1112, 1139, 1198, 1453, 1224, 1254, 1282, 1309, 1335, 1365, 1393, 1419, 1481, 1720, 1508, 1536, 1560, 1586, 1611, 1636, 1664, 1690, 1748, 1990, 1772, 1799, 1830, 1857, 1883, 1910, 1938, 1963, 2019, 2253, 2039, 2065, 2092, 2118, 2143, 2172, 2199, 2226, 2281, 2517, 2307, 2336, 2363, 2390, 2416, 2441]. Back [2466, 2492, 2544, 2774, 2568, 2592, 2619, 2645, 2668, 2697, 2725, 2749, 2302, 786]. Backward [1429]. Baer [2425]. Baire [2626, 1499]. Bakerian [1376]. Baking [2743]. Bakshee [102]. Balanced [2770, 2596, 947]. Ball [103, 77, 1343]. Balloon [1114]. Ballot [2486, 2346]. Balls [76, 101, 2161, 2566, 2220, 1150, 543, 375]. Bamin [2415]. Banach [1920]. Banded [356]. Bang [485]. Bar [2731, 1032]. Barrett [2309]. Barycentric [1945]. Base [408, 2384]. Based [2384]. Bases [2311, 1506]. Bashful [1395]. Basic [654, 2696]. Basis [2425, 420]. Bass [1774]. BC [1115]. Be [365, 1833, 343, 1007, 106, 67, 74, 109]. Beads [2459]. Bears [1443]. Beaten [35, 6]. Beauregard [1535]. Beauty [102]. before [359]. Behaved [1258]. Behavior [1670]. Behaviour [1423]. behind [789, 1262]. Being [163]. Bell [744, 1116, 447]. Belshoff [2142]. Ben [475]. Benedict [1480]. Benford [2147]. Benson [2773]. Bent [817]. Bergman [1316]. Bernd [2644]. Bernoulli [277, 2314, 891, 669, 95, 2691, 1289, 819, 2496, 394]. Bernstein [2538, 2651, 1449, 1755, 1305, 1965]. Berriozábal [355]. Bertrand [2675]. Bessel [2328]. Best [1050, 179, 164]. Better [112, 2135, 2190, 1888]. Bettye [1747]. between [2594, 437, 757, 2357, 2283, 2344, 756, 746, 1316]. Beyond [2464, 705, 219, 677, 919]. Biased [2204]. Bidiagonal [1297, 532]. Bifurcations [2312]. Big [900]. Bighash [2415]. Bijection [2678]. Bijective [2700]. Bikes [1989]. Billiard [494]. Billiards [2392, 2231, 2575]. Binary [2317, 862, 913, 2687, 1863]. Binomial [1159, 2642, 1312, 995, 1903, 2732, 744, 1852, 291, 749, 861, 994, 816]. Biochemistry [1801]. Biology [2752, 1688]. Bipartite [437, 1108]. Biquadratic [2146]. Birkhoff [324]. Birthday [214]. Bisector [2719]. Blackjack [2462]. Blaschke [841]. Blet [814]. Block [237]. Blunder [2367]. Board [2485]. Bodies [1404]. Body [2440]. Bolzano [269]. Book [194, 501, 1137, 413, 837, 1223, 1798, 1222, 1363, 585, 1362, 2116, 643, 255, 315, 285, 500, 894, 752, 1167, 1027, 1082, 1392, 1480, 2306, 2142, 136, 2091, 2334, 2491, 1829, 2696, 1280, 1306, 920, 1988, 135, 97, 1718, 1535, 2038, 380, 195, 1252, 2170, 1989, 1937, 1609, 1827, 1168, 2017, 699, 1663, 1610, 808, 223, 2667, 350, 1962, 2252, 2773, 975, 1747, 382, 165, 442, 836, 2362, 2724, 2644, 1081, 166, 2198, 2464, 1110, 2747, 1418, 284, 1479, 1882, 2389, 1026, 612, 586, 2516, 950, 1391, 558, 777, 1251, 316, 726, 1688, 1584, 443, 526, 351, 559, 1000]. Book [974, 222, 2225, 470, 807, 1057, 1507, 1856, 2279, 919, 444, 256, 26, 1450, 1746, 613, 1195, 673, 1001, 865, 1559, 1083, 2388, 1136, 2440, 1056, 381, 2117, 471, 1334, 2280, 412, 2064, 1635, 2591, 67, 525, 778, 1585, 2543, 2617, 1111, 2415, 2567, 2616, 642, 1689, 1166, 1506, 949, 1828, 349, 1771, 1908, 1281]. Books [199]. Boole [2302, 2621]. Bootstrapping [160]. Borsuk [2400].

Borwein [2475]. Both [343]. Bound [2248, 2487, 112, 1710, 1248, 2386, 2718]. Boundary [296, 482, 1440]. Bounded [201, 2046, 1967]. Boundedness [1425]. Bounds [1577, 2288, 2084]. Boxes [2311, 1564, 1659]. Branched [2673]. Brannan [1507]. Brauer [1087]. Breaking [1877]. Bridge [756]. Bridges [536]. Brief [925]. Briggs [1989]. British [2591]. Brocard [750, 1505]. Broken [1865]. Brouwer [2400]. Brown [1962]. Brownian [1965]. Bruijn [2745]. Brunnian [1395, 2285]. Bruno [1484, 676]. Bubble [385]. Bubbles [2071]. Budget [2021]. Buffon [982]. Building [2139]. Built [2436]. Burdman [1828]. Burger [2116]. Burgiel [2667]. Burk [2543]. Bus [803].

C [1663, 1337, 782, 77]. CaB [228]. Cage [970]. Cages [1940]. Cake [2648, 233, 67]. Cake-Cutting [67]. Calculate [871]. Calculating [234, 848]. Calculation [820]. Calculational [453]. Calculus [1943, 758, 729, 1455, 843, 2038, 450, 1864, 1062, 351, 2524, 1888, 387, 2567, 2145, 351]. Calendar [26, 32]. Called [2547, 448]. Calls [926]. Cambridge [1663, 2591]. Camouflage [1446, 1215]. Can [2247, 1833, 957, 343, 1007, 2272, 106, 74, 1958, 109, 1425, 67]. Candy [899]. Cannot [365]. Canons [1966]. Cantor [1755, 453, 2572, 2372]. Car [2301]. Carathéodory [1610, 2340]. Cardinality [411]. Careless [857]. Carl [2464]. Carleman [1010]. Carlo [979]. Carries [2728]. Case [294, 2228, 1747]. Cassini [835]. Catalan [1666, 578, 1847, 671, 640, 2112, 1079]. Catalyst [2468]. Catch [1373]. Catch-up [1373]. Cauchy [269, 1609, 2246, 2423, 1205, 1399, 2179]. Causal [3]. Causing [1469]. Cavalieri [1402, 847]. Cayley [2627, 1144]. CD [1115]. Ceiling [762]. Ceilings [1107, 204]. Celtic [180]. Centered [1475]. Centers [330]. Central [2642, 1426, 2522, 2599, 1023]. Centroid [309, 1190]. Centroid-Circumcenter [1190]. Centroids [1795]. Century [512, 2626, 321, 457, 531, 2591]. Century-Old [2626]. Certain [2152, 946, 2677, 482, 2385]. Cevian [1836]. Cevians [129, 1795]. Chaim [2667]. Chain [1729, 979, 1594, 1054]. Challenge [593]. Challenges [2228]. Champions [263]. Chancy [2694]. Change [1668, 357, 926]. Chaos [449, 649]. Chaotic [1284, 1834]. Characteristic [1397]. Characterization [1918, 2182, 1119, 1426, 216, 2626, 2615, 2474, 1999, 159, 2577, 568, 638]. Characterizing [134, 1638]. Characters [1945]. Charles [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774, 1585]. Chebyshev [1615, 92]. Cheese [2623]. Chehrzad [2362]. Chessboard [2612, 1458]. Choice [760, 2283, 2361, 2084]. Chord [1969]. Christensen [1506]. Christine [1226]. Christos [1479]. Chromatic [555, 2263]. Chronicles [2747]. Chung [483]. Circle [698, 1563, 377, 1164, 2099, 705, 1266, 475, 1555, 192, 1475, 1065, 298, 496, 723]. Circles [1422, 1340, 218, 2243, 455, 709, 2444, 973, 2043, 1829]. Circuit [1145]. Circulant [565]. Circular [2062, 787]. Circum [2167]. Circum-to-In-Radius [2167]. Circumcenter [1444, 1190]. Circumcenters [804, 723]. Circumscribing [1422, 1861]. Cirumscribing [441]. Clarence [953]. Class [2000, 2626, 1609, 450, 2045, 1499, 482]. Classic [1386, 593]. Classical [2398, 760, 104, 78, 2175, 198, 2567]. Clocks [1989]. Close [54, 1543, 170]. Closed [2366, 2247, 2109, 2290, 193, 819, 504, 109]. Closure [1741, 2109]. Clothes [1395]. Clovers [1639]. Cochran [328]. Codes [2365]. Coefficient [2708, 995, 1476, 1903, 1320, 2661, 749]. Coefficients [2397, 1159, 1312, 2394, 2732, 1833, 1852, 291, 1996, 861, 816]. Cohen [1672]. Cohesive [465]. Cohomological [842]. Cohomology [417]. Coin [2756]. Coins [2566]. Colin [351, 2616]. Collaboration [2571]. College [327]. Collegiate [926]. Collide [2071]. Collinear [56, 1505]. Collinearity [635]. Colored [106, 74]. Colorful [1944]. Coloring [2203, 1824, 2303, 2224]. Colors [1009]. Combination [1105, 607]. Combinatorial [1517, 1630, 2072, 2086, 2413]. Combinatorics [680, 1945]. Come [452]. Commandant [2]. Comments [80]. Common [309, 720, 2088, 2732, 1531]. Community [979, 782]. Commutative [40]. Commutativity [1709, 125, 2664]. Commutator [2271]. Commuting [2563, 1217, 1742]. Compact [1695, 775, 68, 41]. Compact-Open [1695]. Compared [1657]. Comparing [439, 1330]. Comparison [1353]. Compass [651, 325]. Competition [1371, 231, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703, 1689]. Complement [830, 1741]. Completely [1934, 552]. Complex [1446, 980, 140, 1972, 1820, 2091, 1215, 2468, 2340, 2653, 1053, 925, 2524, 1635, 1388, 1711]. Complexes [1945]. Complexities [1747]. Complexity [2516]. Component [95]. Composite [2744, 1274]. Composition [2217]. Compositions [2088]. Computation [954, 1479]. Compute [2448]. Computer [1487, 2752, 2331]. Computers [1908]. Computing [1566, 1402]. Concave [1477]. Concept [1667]. Concerning [760, 1315, 681, 1837]. Concurrence [1065]. Concurrent [347, 129, 496]. Condition [1709, 1594, 125, 2664, 2271, 1818]. Conditionally [904, 1429]. Conditions [341, 1049, 253, 1728, 1543, 480]. Cones [2095]. Configuration [650, 2042]. Conflicts [4]. Congressional [2339]. Congruence [1107, 521, 2527, 1893, 2124, 435]. Congruences [297]. Congruent [1802, 376, 1132]. Conic [998]. Conics [2469, 2445, 2075, 1398, 2117]. Conjecture [955, 874, 30, 10, 385, 2548, 1746, 1862, 1730, 176, 2623]. Conjugate [1444, 2419, 1330]. Connected [68, 41]. Connection [757, 1151, 2283]. Consecutive [1850, 943, 343, 2411, 2318, 1905, 732]. Consequence [1879, 1713]. Consequences [860]. Considerations [2595]. Constant [148, 1458, 2245, 164, 815, 1860, 2367, 469, 1460, 848]. Constant-Calculating [848]. Constantin [1610]. Constrained [93]. Constraint [436]. Constructed [708]. Constructible [292]. Constructing [1285, 229]. Construction [1315, 2345, 35, 6, 2400, 2576]. Constructions [651, 1397]. Constructive [1545, 1867, 2628]. Constructivism [326]. Continued [2126, 666, 1589, 1726, 1162, 265, 1080, 1727, 2382, 1878, 2343, 2498].

Continuity [930, 1646]. Continuous [1935, 1385, 1715, 1728, 2678, 39, 62, 2076, 2462, 1595, 2025, 2097, 149, 708]. Continuously [454]. Continuum [760, 1121, 684]. Contour [510]. Contractions [1866]. Converge [2140]. Convergence [2395, 1564, 2191, 253, 2038, 1997, 2589, 419, 583, 556, 2686, 725, 2193]. Convergent [904, 1024, 1175, 567, 2396, 1429]. Convergents [1191]. Convex [1930, 1477, 290, 176, 2344, 1052, 1007, 2472]. Convexity [1645, 805, 623, 1341]. Convoluted [281]. Convolution [1159, 410]. Convolutions [551]. Conway [1937, 2667, 2044, 2393]. Cooperation [202]. Copositive [1569]. Core [1968]. Core-Plus [1968]. Corner [1290]. Corollary [2]. Coronas [176]. Correct [38]. Correction [1211, 593]. Cosine [2194, 1194, 254]. Cosines [1077, 1906]. Cosmological [2367]. Cotangent [788, 1361]. Could [396]. Count [1689]. Countable [132, 775]. Countably [1932]. Counterexample [959, 425]. Counting [2637, 2089, 1588, 1249, 373, 971, 1853, 1036, 2661, 2653, 1929, 694, 1804]. Couples [2577]. Course [2017, 2120]. Cousin [1983]. Covered [466]. Covergence [2305]. Covering [2612, 2553, 1103]. Covers [411, 2261]. Cox [36]. Coxeter [2389]. Craik [2591]. Cramer [902]. Created [2064]. Creates [536]. Creating [1175, 2043]. Criterion [1095, 1518]. Critical [2127]. Crofton [37]. Crooked [386]. Cross [2599]. Cross-Sections [2599]. Crust [2623]. Crux [2192]. Crystallographic [955]. Cube [1402, 1819, 2369, 1501, 282]. Cube-Free [1819]. Cubes [434, 1090, 493]. Cubic [310, 1569, 75, 105, 2443]. Cubical [345]. Cuboids [1287]. Curiosities [1837]. Curious [2602, 2499, 757, 676, 762, 2396]. Curriculum [1062]. Curvature [2366, 1424, 1150, 1062]. Curve [2366, 931, 2230, 1998, 109]. Curves [2095, 786, 1802, 2368, 314, 158, 1141, 2075, 2290, 1940, 636, 19, 1661, 2142]. Cute [159]. Cuts [2485]. Cutting [2648, 844, 67]. Cybernetics [1937]. Cycle [2625]. Cycles [2561, 406, 1338, 2402, 2745, 1021, 1824, 948]. Cyclic [347, 128, 331, 2166]. Cycling [1531]. Cycloid [2099]. Cycloidal [2672, 733]. Cyclotomic [1320, 1105, 2207, 1220]. Cylinder [1595]. Cylinders [2095, 2469].

D [1718, 2724, 2464, 2591, 42, 2510]. D. [424]. Dan [1856]. Daniel [1962]. Darboux [2046, 2626, 1375]. Dark [1937]. Data [2596, 1958]. David [1535, 2773, 2747, 1507]. Days [477]. Deahna [477]. Deal [1885, 900]. Decay [1054]. Decimal [391, 605]. Decomposition [1545]. Dedekind [682]. Defendens [198]. Defined [1104, 524, 2152, 696, 1656, 1503]. Definite [2259, 1163, 882, 221]. Definitely [806]. Definitions [1286]. Deflation [266]. Degrees [750]. Delannoy [1023]. Delta [269]. Demise [1116]. Denis [1746]. Dense [2676]. Densely [2273]. Density [657, 1797, 2548, 425]. Departments [321]. Dependence [2471]. Derangements [88, 1078, 538]. Derbyshire [1856]. Derivation [322, 1374, 818]. Derivative [1935, 2278, 763, 1617, 1724, 2587, 2288, 1461, 311]. Derivatives [1261, 618, 234, 2046, 293, 680, 2374]. Derived [2551]. Deriving [279]. Descartes [2596, 1947, 705, 1229, 1206, 1317]. Descending [2640]. Descent [1141, 2047]. Description [1023]. Descriptions [2469]. Deserves [521, 1578]. Design [1200]. Designing [453]. Designs [2365, 1206]. Detection [2394]. Determinant [800, 409, 1020, 1243, 639, 2137, 2072, 1075, 178, 1397, 2327, 681, 2613, 2611, 344]. Determinants [1588, 1095]. Determine [3]. Determined [1447, 1219]. Developing [2497]. Development [2724]. Devil [2501]. Diagonalizable [2121]. Diagonals [1738]. Diameter [290]. Diamond [2017]. Dice [1851]. Dickman [410]. Did [2520, 1775]. Difference [497, 1564, 1162, 1779, 128, 1823, 1892, 1818]. Differences [249, 2596, 1513]. Differentiability [2650]. Differentiable [495, 1790, 1582, 2273, 149, 708]. Differentiablity [1728]. Differential [618, 45, 61, 134, 1457, 2523, 37, 2289, 2038, 1668, 252, 1794, 477, 2202, 71, 2617]. Differential-like [71]. Differentiate [591]. Differentiating [480]. Differentiation [2125, 1120, 476]. Difficult [326]. Digital [1355]. Digits [2590, 1803, 408]. Dihedral [2647, 820]. Dimension [1673, 2123]. Dimensional [608, 2373, 1228, 358, 361, 425]. Dimensions [843, 925]. Dini [1724]. Diophantine [263, 2160, 1428, 1011, 969]. Diophantus [786]. Diploma [327]. Direct [2229, 787, 2400]. Directly [2490]. Directors [2469]. Dirichlet [596, 374, 2000, 1879, 296, 818]. Disc [392, 1486]. Disciplines [1888]. Discontinuities [1358]. Discontinuous [2273]. Discussions [1284]. Disentangling [2573]. Disguise [2486, 2115]. Disguised [2191, 2246, 1354]. Disjoint [2663]. Disk [325, 1415, 621]. Disks [2469]. Disorder [1333, 1356]. Disorienting [2753]. Displacement [2637]. Dissecting [2114, 1287]. Dissection [111]. Distance [1790, 1440, 746, 1190]. Distances [1035, 1131, 2344, 1353, 1902, 1740]. Distinct [1245, 1982]. Distinction [1172]. Distinguished [355, 96, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Distinguishing [437]. Distortion [2549]. Distributed [2665]. Distribution [188, 1579, 2049, 133, 994, 220]. Distributions [682, 538]. Divergence [2373, 1660]. Divergent [1417, 451, 1876]. Divides [1825, 520]. Dividing [2110]. Divine [2252]. Divisibility [2399, 2642, 1049, 1821, 942, 695, 607, 1247, 1359]. Divisible [1777, 2708, 1331]. Division [2201, 1848, 1546, 233]. Divisor [2767, 1246, 1248]. Divisors [637, 1329, 1850, 671, 1022, 1160, 829, 2100, 1439, 1531, 522]. DNA [1801]. Do [1565, 1009]. Does [2140, 931, 2523, 89, 1141, 1997]. Doesn’t [581]. Domain [1203]. Domains [729, 1613]. Dominate [89]. Domination [2471]. Dominoes [1277, 2141]. Donald [2389]. Don’t [1078, 1629]. Double [2438, 2330, 385, 2513, 1460, 2689]. Douglas [956]. Dove [2731]. Down [452, 263, 1689, 1095]. Downhill [1141]. Dr [355, 2041, 2309, 2570, 69, 592, 953, 1226, 1510, 475, 1774]. Draft [722]. Dragon [812]. Drazin [1368]. Drop [1425]. Dual [2575]. Dungeons [2640]. Dyadic [2678]. Dyck [1502, 2061]. Dynamical [1345, 143, 2345, 535, 1970]. Dynamics [569, 1751, 1512, 1228].

E. [32, 447, 531]. Early [1181, 36, 531, 477, 1311]. Earth [1969]. Easy [1204, 873, 1605, 2475, 1093, 8, 33, 1711, 1456]. Ebitees [1629]. Eclectic [2069]. Ecole [1513]. Ecology [979]. Economic [652]. Edge [2394, 595, 1697, 849]. Edge-Labeled [1697]. Edge-Length [595]. Edges [2561, 1767]. Editor [2335, 2171, 317, 137, 614, 2465, 645, 701, 838, 922, 1002, 1451, 1308, 1719, 1909, 2018, 2618, 2748]. Eduardo [2198]. Education [1286, 1172, 1337, 531, 2571]. Educational [1513]. Effectively [396]. Effros [1400]. Egyptian [30, 10]. Ehrenpreis [2628]. Ehrhart [2700]. Eidelman [1718]. Eigenvalue [882]. Eigenvalues [1566, 1297, 1617, 1205, 2671]. Eighth [2446]. EKG [1216]. Elaborate [725]. Elbert [36]. Element [89, 2625]. Elementary [2651, 93, 1618, 1757, 2755, 569, 1120, 1401, 2038, 2010, 842, 2341, 1177, 1180, 759, 1674, 820, 1265, 113, 1620, 1541, 2120, 1288, 2233, 2576, 654, 2449]. Elements [248, 1826, 692, 1203]. Elevated [1853]. Elias [2491]. Eliminate [2284]. Elizabeth [1688]. Ellipse [2089, 1791]. Ellipses [2232, 841, 1999, 2231, 2443]. Ellipsoids [1805]. Elliptic [786, 1802, 2368, 535, 19]. Embeddings [1234]. Emotional [396]. Empirical [1667]. Encryption [1200]. End [1987, 2515, 2772]. Endnotes [2335, 2171, 317, 137, 614, 2465, 645, 701, 838, 922, 1002, 1451, 1308, 1719, 1909, 2018, 2618, 2748]. Endpoints [746]. Energy [2035]. Engine [2464]. Engineering [2251]. Enhanced [1005]. Enough [170]. Ensuring [253]. Entire [2048, 1530, 2458]. Entries [653, 344, 509]. Entropy [1396]. Entry [1839]. Enumerating [1285]. Enumeration [1333, 1356]. Enumerative [2445]. Envelope [1414, 1262]. Envy [233]. Envy-Free [233]. Epsilon [269]. Epstein [1585]. Equal [2139, 1469, 25, 1063, 408, 1416, 2074, 844, 123, 2562, 2610, 2636]. Equal-Area [844]. Equality [1901, 2435]. Equation [1261, 2588, 2312, 1557, 2160, 885, 2523, 2723, 1854, 37, 2289, 42, 63, 1779, 1368, 748, 252, 2036, 1011, 1794, 1904, 784, 1540, 969]. Equations [618, 45, 61, 134, 1457, 1633, 2433, 565, 1428, 2202, 71]. Equicevian [217]. Equidistribution [789]. Equilateral [2392, 872, 1007, 724]. Equilibria [2208]. Equivalent [332, 759, 1684]. Erdos [2707, 1488, 364, 1698, 1515, 499, 1092]. Ernst [2142]. Errata [891, 892, 893, 888, 889, 890, 1445, 1446, 1447, 1448, 1449, 1987, 2249, 2515, 2772, 1715, 1716]. Erratum [1165]. Error [1304, 593]. Escape [142, 362]. Esplen [1507]. Essays [2225]. Establishing [423]. Estimate [293, 749, 133]. Euclid [919, 929, 1775, 2552, 1973]. Euclidean [1946, 1033, 787, 2108, 1775]. Eudoxus [1144]. Euler [424, 1776, 2621, 551, 2757, 2068, 419, 2245, 945, 2067, 2414, 2753, 2552, 681, 1460, 1475, 1288, 848, 496, 2388]. Eulerian [533, 2374]. Evaluate [1687, 2387]. Evaluating [2652, 108]. Evaluation [2073, 596, 918, 1258, 1581]. Even [891, 669, 408, 1036, 259, 948]. Even-Indexed [891, 669]. Eventually [452]. Every [1618, 2048, 1502, 454, 147, 1121, 35, 6]. Everything [1151]. Evolute [2150]. Evolution [11]. Ex [2304]. Ex-To-In-Radius [2304]. Exact [1567, 411, 1753]. Exactly [2046, 829]. Example [1835, 45, 61, 2045, 1769]. Examples [620]. Exams [1260]. Excenteric [1437]. Excenters [523, 57]. Exclusion [1580, 1929]. Excursion [1173]. Exercise [2755]. Exhaustive [2334]. Exists [454]. Expansion [1835, 1726, 1580, 2687, 1727]. Expansions [1887, 605, 1121]. Expectation [1608]. Expected [722]. Experience [2136]. Experimental [1859]. Explicit [1289]. Explorations [2091]. Explorers [2747]. Exploring [913]. Exponential [178, 2664, 1801, 2036, 681, 2032, 1913, 969, 986, 1694]. Exponentials [1369, 1373]. Exponentiation [2643]. Exponents [1259, 1024]. Exposing [2279]. Expression [891, 669]. Expressions [968]. Exradii [280]. Extension [902, 297]. Extensions [2146, 75, 105]. Extrema [93]. Extremal [2127, 553, 8, 33]. Extreme [1424, 2371]. Extremum [831]. Eyed [679].

F [36, 953, 1507]. Faà [1484, 484, 676]. Faces [1767, 1132]. Factor [1614]. Factorial [2110, 260, 1913]. Factorials [942, 2743]. Factoring [2757, 107]. Factorization [1545, 1578, 1775, 624]. Factorizations [144, 189, 1693, 532]. Factors [2088, 1614, 1232]. Factory [650]. Fagnano [2575]. Fair [2284, 1546, 67]. Familiar [2115]. Family [312, 1413, 862, 201, 1228, 2500]. Famous [1619]. Far [89]. Farkas [1204]. Fast [1078]. Faster [1643]. Fastest [1729]. Fat [2372]. Father [1937]. Fearless [2279]. Feferman [1828, 1828]. Feit [2343]. Feller [2522, 483]. Fermat [2638, 143, 2602, 1655, 1517, 1259, 2111, 297, 757, 730, 1644, 2175, 44, 65, 901, 1229, 1439]. Fermatian [2222]. Fertile [2470]. Few [2561, 1824, 620]. Fewest [948]. Fibonacci [1615, 1446, 172, 2768, 606, 788, 2509, 1215, 634, 1355, 831, 1331, 2485, 667]. Field [2769, 539, 1604, 1782]. Fields [625, 144, 2424, 2383, 2207, 534, 75, 105, 2653, 1562, 1034, 2639, 422, 1076]. Fifth [1641]. Figures [1422, 1458, 506]. Financial [618]. Find [1141]. Finding [650, 2248, 1614, 1179, 1823]. Fingertips [1994]. Finite [438, 625, 144, 2424, 2152, 841, 1514, 757, 2769, 2207, 411, 2234, 679, 358, 653, 1266, 933, 539, 422, 649, 1076]. Finite-Dimensional [358]. Finitely [859]. First [89, 1371, 2017, 592, 1264, 1331, 566]. Fitzroy [2403]. Fixed [132, 783, 143, 128, 610]. Fixture [1161]. Flat [1758]. Flaw [2068]. Flight [709]. Flips [2756]. Flo [1937]. Floor [2588, 762]. Floors [204]. Flows [2070]. Flying [2721]. Focal [2469]. Folding [2345, 1176]. Folium [1229]. Football [2201]. Force [478]. Forcing [536]. Ford [225, 528, 484]. Forest [1367, 2059]. Forests [245]. Form [2247, 2109, 25, 791, 147, 2347]. Formal [2397]. Formally [2230]. Formation [1667]. Formative [2497]. Forming [957]. Forms [2317, 1569, 1088, 1668, 493, 477, 1863, 2017]. Formula [1066, 1402, 551, 1484, 126, 410, 2278, 2068, 1372, 1619, 254, 1668, 676, 141, 1374, 2652, 711, 2473, 2447, 2049, 1289, 819, 1460, 1288, 2233, 847]. Formulas [1838, 2463]. Formulations [7, 34]. Forward [1429]. Found [2346]. Foundations [2773, 457]. Four [96, 365, 104, 78, 1009, 408, 2472, 264, 1032, 1708]. Four-Bar [1032]. Four-Point [365]. Fourier [2038, 2394, 1880, 507]. Fourth [1502, 898, 1370].

Fractal [45, 61]. Fractals [1034, 982, 2257]. Fraction [1726, 1162, 1727, 1878, 2498]. Fractions [2126, 666, 1589, 1231, 30, 10, 265, 104, 78, 1985, 391, 2382, 2343, 816]. Fraleigh [1535]. Frames [1506]. Framework [1942]. France [2198]. Frank [2170, 2543]. Franklin [475, 1285, 475]. Franzén [2334, 2334]. Franzosa [2616]. Fréchet [1582]. Fred [2017]. Fredholm [568]. Free [1819, 233, 1344]. French [90, 1151]. Fresh [1313]. Freshman [2393]. Friend [1291]. Friendship [656]. Frobenius [1490]. Front [29, 320, 60, 100, 139, 169, 197, 227, 258, 288, 354, 617, 384, 416, 446, 474, 503, 530, 561, 589, 647, 897, 675, 703, 728, 755, 781, 811, 840, 868, 924, 1171, 952, 978, 1004, 1030, 1059, 1086, 1113, 1140, 1199, 1454, 1225, 1255, 1283, 1310, 1336, 1366, 1394, 1420, 1482, 1721, 1509, 1537, 1561, 1587, 1612, 1637, 1665, 1691, 1749, 1991, 1773, 1800, 1831, 1858, 1884, 1911, 1939, 1964, 2020, 2254, 2040, 2066, 2093, 2119, 2144, 2173, 2200, 2227, 2282, 2518, 2308, 2337, 2364, 2391, 2417, 2442]. Front [2467, 2493, 2545, 2775, 2569, 2593, 2620, 2646, 2669, 2698, 2726, 2750]. FTA [1807]. Fubini [1261, 1459, 2315]. Fubinito [1807]. Full [2625]. Function [424, 495, 2366, 2741, 1666, 2650, 524, 260, 2125, 2286, 2276, 551, 930, 1887, 1485, 2091, 410, 39, 62, 2474, 2328, 2572, 1961, 1724, 2076, 1934, 640, 2694, 1577, 2015, 190, 677, 348, 522, 1012, 2069, 2771, 2168, 149, 425, 708]. Functional [1557, 885, 1633, 2723, 1854, 1718, 748, 2036, 1718]. Functionally [1191]. Functions [93, 2073, 2048, 1304, 2127, 2549, 980, 1477, 2046, 1385, 1715, 594, 1628, 696, 2626, 1617, 2038, 581, 1885, 79, 346, 883, 467, 1595, 535, 641, 2025, 2097, 762, 2014, 2382, 552, 621, 1530, 2729, 311, 620, 2038]. Functor [1061]. Fundamental [230, 1563, 1064, 843, 1864, 1808, 1640, 1149, 933, 2524, 268, 1914, 2730, 540, 1463]. Funding [869]. Furstenberg [2600]. Further [2747]. Future [2283]. Fuzzy [2494].

G [2142, 2091, 1882, 26, 32, 1746, 2195]. Gain [2496]. Gallery [1455]. Galois [734]. Gambler [721, 2528]. Gambling [1752]. Game [2203, 2566, 58, 1877, 1302, 1739, 619, 2176]. Games [2094, 175, 2496]. Gamma [2717, 1304, 2474, 2197, 2720, 1577, 2015, 677, 2069]. Garden [2543]. Gareth [2773]. Gärtner [2644]. Gauss [1568, 1839, 1562]. Gauss-Winckler [1568]. Gaussian [1930, 2649, 1312, 1614, 291, 1500, 2022]. GCD [2399]. General [2526, 423, 1662]. Generalization [342, 611]. Generalizations [260, 1644, 1967, 1065, 483]. Generalized [230, 2598, 2358, 578, 363, 671, 467, 903, 1904, 1542, 706, 2346]. Generalizing [956, 2483]. Generate [2257, 948]. Generated [2291, 859, 2096, 2625, 1007]. Generating [76, 101, 91, 640, 190, 2382, 1697, 522]. Generators [1076]. Generic [1339]. Generous [1248]. Geometriæ [2145]. Geometric [884, 267, 2070, 1535, 1827, 1464, 1940, 563, 1860, 1053, 423, 1889, 1032, 1998, 2149, 2091]. Geometrically [1290]. Geometry [495, 1376, 2445, 1312, 1639, 1490, 1141, 261, 2421, 787, 756, 5, 1284, 619, 2617, 1341, 2252, 2389, 1507, 919]. George [1962, 1746]. Georgiadou [1610].

Gerald [1510]. Gerechte [2365]. Gerhard [1635]. Gerrymandering [2284]. Gersgorin [1829]. Getting [407, 2490]. Ghys [2075]. Gibbs [751]. Gift [345]. Gilbert [1535]. Gin [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Ginzburg [2035]. Girth [2263]. Given [290, 1881, 1178, 1740]. Glass [1256]. Global [1376]. Glory [1689]. GM [2166]. God [2064]. Gödel [2334]. Godement [2038, 2038]. Goes [1630, 452, 2302, 1692]. Going [973]. Goldbach [955, 1776, 1862]. Golden [2436]. Gons [956, 1052]. Good [521, 1578]. Goodman [2667]. Goodman-Strauss [2667]. Google [2464]. Gossard [1412]. Goursat [2423]. Govern [1398]. Graceful [1328]. Grant [351]. Graph [1338, 437]. Graphica [102]. Graphical [1725]. Graphs [2561, 1338, 1755, 553, 2471, 2402, 1108, 465, 1824, 1034, 331, 2224]. Grauert [450]. Graves [1118]. Gravity [1031, 1540]. Gray [1507]. Greater [2767]. Greatest [2248, 2367]. Greats [389]. Green [1149]. Gregory [2145]. Grid [2231]. Grier [1908]. Gronwall [1983]. Gross [1480, 2279]. Grothendieck [1948]. Group [1447, 735, 230, 2162, 2152, 2699, 1563, 1514, 859, 1619, 2433, 1644, 248, 1219, 1826, 1103, 593, 692]. Grouping [668]. Groups [438, 2755, 2622, 2383, 2647, 1695, 757, 2625, 2234, 1022, 2402, 1613, 653, 110, 2206, 933, 2202, 1456]. Growing [507]. Growth [2691, 2032]. Grunbaum [176]. Guessing [1739]. Guide [2334, 351]. Gung [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Gyrogroups [2287].

H [2667, 1479, 531]. Habit [2731]. Hadamard [2342]. Hadjicostas [1460]. Hairy [1343]. Half [608, 1354]. Half-Angle [1354]. Half-Space [608]. Hall [736, 342, 611]. Halmos [69]. Halving [1340]. Ham [1181, 1972, 390]. Hamilton [2627]. Hamming [2365]. Hand [763]. Handbook [1885]. Hankel [178, 681]. Hannu [2440]. Hard [1605]. Hardy [863, 1915, 1037, 1316]. Harmonic [2540, 1881, 1008, 621, 2124, 1554, 2149]. Harmonious [1714]. Harnack [1457]. Harris [1480]. Hart [237]. Hartshorne [919]. Harvested [2312]. Hass [351]. Hausdorff [132, 2123]. Hawking [1663, 2064]. Head [2367]. Heads [2135]. Heesch [1314]. Heidi [2667]. Height [2139]. Helical [109]. Heritage [1172]. Hermite [2342]. Hermitian [1205, 960, 1607, 1556]. Hero [1937]. Heron [901, 1374]. Heronian [2058, 395]. Hersh [2225]. Herstein [2755]. Hex [2094]. Hexagonal [652]. Hidden [917, 2279]. High [2263, 926, 667]. High-Powered [667]. Higher [234, 293, 1051, 680, 1337, 925]. Highly [912, 465]. Hijab [2567]. Hilbert [1345, 2174, 2286, 2146, 332, 1837, 1891, 236, 898]. Histograms [1958]. Historical [2067]. History [1181, 1663, 1172, 676, 32]. History. [26]. Hitting [2449]. Hoc [1397]. Hölder [1971, 2037, 930, 332]. Holding [1049]. Holey [176]. Holomorphic [2038, 2444]. Home [2706]. Homeomorphism [1695, 1970]. Homogeneous [1516]. Homology [1338]. Homomorphisms [999]. Hook [1372]. Hopkin [2591]. Horn [417]. House [2403]. Hu [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Hull [1930]. Human [1908]. Hurwitz [873]. Hydra [2180]. Hydrogen [2170]. Hyman [1774].

Hyperbolic [2073, 229, 1628]. Hyperbolic-Looking [1628]. Hypercubes [1090, 1021]. Hypergeometric [2163, 1589, 1193, 2382]. Hypergeometrics [609]. Hypergraphs [1396]. Hyperplane [53]. Hypothesis [2578, 2368, 760, 759, 1856].

I-A [2201]. Ian [2306]. Ibadur [1635]. Icosahedron [1714]. Ideal [1594, 1822, 1203]. Ideals [1594, 9, 31, 1076]. Ideas [1464]. Idempotent [1618, 236, 1607]. Identify [2090]. Identities [1849, 172, 2649, 1487, 658, 448, 958, 2115, 1607]. Identity [2163, 745, 569, 88, 1820, 995, 94, 1903, 2732, 1852, 2328, 2641, 2059, 694, 1554, 2771, 667]. If [2484, 67, 1782]. Igor [102]. II [2038, 357]. III.36 [1033]. Ill [1258]. Ill-Behaved [1258]. Illuminating [959]. Images [102]. Imaginary [1089, 892, 102, 693]. Imaging [1585]. Immediately [1807]. Impact [1311]. Imperium [198]. Implications [926]. Implicit [1935]. Implies [1807, 846, 1039, 2024, 2400]. Imply [2523, 1997]. Important [1172]. Impossible [1818]. Improper [2073, 2589]. Incenter [309, 1902]. Incenters [998, 57]. Incidence [1301]. Incircle [313]. Inclusion [1580, 1797, 1929]. Inclusion-Exclusion [1580]. Incognito [609]. Incomplete [1304, 2334, 533]. Incorrectness [2693]. Increasing [346]. Indentity [2327]. Independence [1887]. Independent [553, 1932]. Indeterminates [1243]. Index [1235]. Indexed [891, 669]. India [199, 1260]. Indirectly [2490]. Indivisibility [2063]. Inductive [712]. Inequalities [283, 1204, 1194, 2195, 2746, 417, 1396, 332, 363, 1780, 1944, 92, 280, 1501, 423, 1092, 1192, 1959, 985, 1316, 1609]. Inequality [2439, 1568, 2332, 1933, 1457, 2342, 2286, 2276, 2085, 1645, 2330, 2244, 2164, 364, 1476, 2768, 1010, 1796, 2218, 2192, 1055, 2564, 2013, 1515, 2539, 2692, 164, 557, 2370, 1915, 2722, 863, 1109, 2015, 1386, 2413, 960, 379, 1792, 623, 2542, 8, 33, 1534, 1134, 2359, 2166, 2331, 2277, 2719, 1711, 1037, 2565, 1516, 2169, 1906, 1936, 1427, 1983]. Inequivalent [915]. Inexhaustibility [2334]. Infinite [2194, 2399, 1658, 1338, 2389, 1860, 583, 776, 2537, 2425, 1039, 2384, 2333, 708]. Infinitely [1777, 1428, 1203]. Infinitude [929, 2600]. Infinity [2022]. Inflating [2369]. Influence [1885]. Informally [2230]. Information [318, 615, 895, 1169, 1452, 1722, 1992, 2255, 2519, 2776, 1937, 2516]. ing [1134]. Ingelstam [1180]. Inheritance [90]. Initial [2009, 1728, 419]. Injective [1388]. Inradii [1981]. Inradius [290]. Inscribed [218, 2232, 498, 1791, 1555, 441, 2412]. Inscription [1478]. Insight [2672]. Insolvable [1428]. Instruments [421]. Integer [2397, 146, 2643, 1833, 54, 1683, 2690, 2671, 2356, 997, 1130, 1862]. Integers [668, 745, 2649, 249, 1614, 378, 2121, 695, 1263, 653, 943, 1584, 2425, 1905, 2347, 732, 2064]. Integral [2538, 806, 2438, 2085, 293, 341, 2196, 763, 1617, 2038, 2589, 2010, 1163, 310, 2423, 2360, 1441, 1500, 1687, 1685, 2179, 2751, 2652, 2587, 677, 1606, 2318, 2513, 1793, 1643, 2689, 480, 221, 816, 191, 1581]. Integrals [312, 2073, 2717, 2495, 2000, 2259, 2572, 357, 448, 818, 2551, 162, 1460, 108, 507, 451, 848, 2543]. Integrands [2551]. Integrate [591, 456].

Integrated [1880]. Integration [1616, 1459, 1519, 510, 1864, 620]. Interest [4]. Interesting [1338, 2150, 1228, 1319]. Interlace [1205]. Intermediate [883, 2512]. Interpolation [930]. Interpretation [291, 1860, 2149]. Interrupted [22]. Intersecting [314, 158]. Intersection [53, 1265]. Intersections [1805]. Interval [637, 193, 362]. Introduction [1718, 1535, 1609, 1962, 2644, 5, 2616, 1506, 2142, 1585, 2567]. Intuition [570]. Intuitive [1374]. Invariance [1608]. Inverse [594, 2229, 476, 1228, 2097, 1793, 1855]. Inverses [234, 680, 448]. Inversion [2137, 1591]. Inverted [2009]. Invertibility [1095]. Invertible [2162, 1218, 554]. Inverting [707, 2087]. Investigate [396]. Investigation [1667]. Investing [655]. Invitation [2415]. Involve [2247]. Involving [2073, 720, 1107, 189, 88, 4, 2732, 2463, 1852, 2572, 695, 2245, 762, 1011, 280, 944, 2059, 1554, 1959]. Irene [1885]. Irrational [391, 2345, 1889, 2742]. Irrationality [267, 2525, 388, 1889]. Irreducibility [2708, 1518]. Irreducible [2638, 2424, 1693, 731, 1203]. Irving [1584]. Isodynamic [886]. Isogonal [1444]. Isohedral [176]. Isolating [783]. Isometric [1234]. Isometries [110]. Isometry [110]. Isomorphic [2014, 1344]. Isoparametric [1201]. Isoperimetric [1201, 1596, 8, 33, 1037]. Isosceles [24]. Issues [1337]. Iterated [1369, 1373, 128, 1891]. Iterates [2651]. Iteration [1437, 1354, 2686]. Iterational [2305]. Iterative [1448, 1390].

J [424, 1609, 1962, 2252, 2773, 2362, 2724, 1507, 2415]. Jack [67]. Jacobi [1401, 535]. James [1962, 1337, 782, 2145]. January [1839]. Jeremy [1507]. Jerry [2017, 2747]. Jim [1937]. Jirí [2644]. Joe [1480]. Joel [351]. John [1718, 1535, 2667, 1688, 1856, 2617, 1771]. Jordan [1971, 2230]. Joris [1674]. Jorma [2516]. Jr. [2724]. Judith [2724]. Jug [1106]. Juggling [2310, 1483]. Jump [2706]. Just [1590]. Justification [1667].

Kakeya [539]. Kaleidoscopes [1256]. Kaleidoscopic [229]. Kaplansky [790]. Karrunen [2440]. Keeping [141]. Keith [2117]. Kelper [1746]. Kemeny [815]. Kempner [2499]. Kendig [2117]. Kepler [1725, 1264, 564]. Kevin [1663]. Kids [1689]. King [2577, 2389]. Kings [2471]. Kirkman [957]. KKM [2024]. Klee [1148]. Knapp [2696, 2696]. Kneser [874]. Knew [374]. Knockout [38]. Knoebel [2747]. Knot [107]. Know [2490]. Known [1485, 712]. Knox [1663]. Kobayashi [2340]. Koch [1998]. Komlós [1425]. Komornik [148]. Korovkin [1917, 985]. Krantz [2091]. Kronecker [1959, 2193]. Kronecker-Like [2193]. Krull [1673, 1265]. Kunz [2142].

L [2198, 1510, 1559, 1585, 1771, 971, 581]. L-Tiles [971]. L1 [2076]. Labeled [2161, 1697]. Lagrange [1066, 2520, 1591, 2526, 1914, 2730]. Lambert [2247]. Lamps [144]. Landau [2035]. Landen [2045]. Lane [1988, 1988]. Langville [2464]. Laplace [677]. Large [323, 2708, 536]. Largest [882]. Last [1259]. Late [2190, 359]. Latin [2221]. Lattice [2089, 1901, 124, 608, 2550, 1246, 73, 636, 483, 1782, 395]. Lattice-Ordered [1782]. Lattices [1266]. Laubenbacher [2747]. Law [2147, 90, 2398, 1177, 1264, 903, 2759]. Laws [564, 396]. Lawson [1827]. Lax [1718, 1668]. Lay [1535]. LCM [1332]. Leading [2708]. Learned [782]. Learning [2136]. Least [720, 2732, 2357, 829, 2677]. Leaves [580, 579]. Lebesgue [1120, 657, 1864, 2751]. Lecture [1376, 1337, 782]. Lectures [2491]. Lee [2041]. Leftovers [390]. Leggett [1747]. Lehmus [541]. Leitzel [782]. Lemma [1204, 1343, 449, 1562, 2024, 2576, 1583]. Lemniscate [1860]. Length [306, 406, 1372, 835, 2553, 595, 982]. Lengths [1661]. Lesser [1485]. Lesser-Known [1485]. Lesson [2205]. Lester [225, 528]. Letter [2403]. Letters [11, 80, 2306]. Lévy [812]. L’Hospital [1427]. Lida [2309]. Lie [931, 2152]. Life [1087, 1948, 447, 1828, 849]. Light [476, 2676, 648]. Lighthouse [2021]. Like [2574, 2193, 71]. Lily [619]. Limit [720, 980, 250, 468, 2196, 2197, 2720, 2522, 1768, 2245, 2011, 1279, 2458, 773]. Limiting [538]. Limits [1490, 1228]. Lindeberg [2522]. Line [2048, 1866, 1061, 2414, 2025, 1012]. Linear [1743, 417, 146, 1887, 1064, 1487, 1535, 1779, 2362, 2644, 505, 2419, 1863, 71, 509, 1535]. Linearity [478, 2170]. Linearly [1932]. Lines [347, 56, 1233, 466, 496, 1708]. Lingering [1885]. Link [2595, 984, 2099]. Linkages [1032]. Linked [1039]. Linking [1005]. Liouville [2418]. Lipschitz [2523]. Lists [912]. Literacy [2497]. Little [2602, 1655, 1659, 297, 1644, 805]. Littlewood [1037]. Local [2125, 1116]. Localized [1145]. Location [652, 113]. Locations [1886]. Lodder [2747]. Log [2717, 2438, 1985]. Logic [1142, 1798, 1828]. Logistic [2312, 362]. Logs [2011]. Long [465, 641, 941]. Longchamps [219]. Longer [2250]. Longest [746]. Look [1313, 2547, 1801, 2753, 2388]. Look-and-Say [1801]. Looking [1628, 1256]. Loops [171]. Lorch [2041]. Loreti [148]. Lost [2259, 1367, 261, 475, 2346]. Lottery [722]. Low [1692]. Low-Periodic [1692]. Lowell [231, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703]. Lower [2248, 2487, 2718]. Loy [2773]. LSB [2024]. Lucas [1517, 455, 831]. Lucasian [1663]. Lucy [619]. Lunes [206]. Luzin [11, 570, 80]. Lyapunov [1620].

M [2491, 1747, 1727]. M. [11, 1518, 80]. Mac [1988]. Maclaurin [1421]. Made [1204, 102, 1659, 1093, 1456, 386]. Magic [1090, 1851, 1480, 259, 475]. Magnecurve [180]. Majority [2284]. Majorization [1713]. Majorizations [932]. Make [1338, 2599]. Makers [2148]. Making [2116]. Malcolm [1535]. Malfatti [2021]. Malgrange [2628]. Man [2389]. Manifolds [2548]. Manipulability [704]. Manuel [355]. Many [1285, 1049, 252, 1428, 478, 1007, 127, 1958, 1203, 1928]. Map [2203, 1415, 362]. Map-Coloring [2203]. Mapping [1339, 2047, 32, 26]. Maps [1582, 1228, 2444, 1742]. Marden [2341]. Maria [1610]. Mark [2280]. Markov [1729, 979, 1867, 1054]. Maryland [191]. Master [1609]. Masterpieces [2747]. Masters [1882]. Matched [289]. Matchings [2663]. Math [1689]. Mathematic [1337]. Mathematical [1798, 1286, 1988, 1609, 1885, 2773, 2724, 231, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703, 562, 869, 2043, 814, 2747, 1688]. Mathematician [1771, 2306]. Mathematicians [1559]. Mathematics [512, 1859, 1286, 1256, 1421, 1610, 1172, 592, 1, 1117, 2228, 421, 1962, 321, 1810, 1867, 457, 1675, 202, 531, 926, 2571, 2591, 1585, 782, 506, 355, 2041, 2309, 2570, 1747, 953, 1226, 1510, 2225, 1774, 1663, 2198]. Mathesis [1810]. Matousek [2644]. Matrices [389, 324, 1618, 200, 1777, 2399, 2162, 554, 1227, 799, 532, 2615, 55, 2223, 2688, 1205, 565, 653, 563, 2671, 772, 435, 1388, 1607, 1178, 1556, 1076, 1301]. Matrix [1566, 707, 2439, 892, 1618, 639, 1075, 1297, 2728, 2662, 511, 2121, 1368, 801, 833, 640, 228, 2690, 2100, 882, 1579, 2613, 2611, 2087, 68, 41, 693, 2060, 2385, 1822, 509, 2235]. Matter [28, 319, 59, 99, 138, 168, 196, 226, 257, 287, 29, 320, 60, 100, 139, 169, 197, 227, 258, 288, 353, 616, 383, 415, 445, 473, 502, 529, 560, 588, 354, 617, 384, 416, 446, 474, 503, 530, 561, 589, 646, 896, 674, 702, 727, 754, 780, 810, 839, 867, 647, 897, 675, 703, 728, 755, 781, 811, 840, 868, 923, 1170, 951, 977, 1003, 1029, 1058, 1085, 1112, 1139, 924, 1171, 952, 978, 1004, 1030, 1059, 1086, 1113, 1140, 1198, 1453, 1224, 1254, 1282, 1309, 1335, 1365]. Matter [1393, 1419, 1199, 1454, 1225, 1255, 1283, 1310, 1336, 1366, 1394, 1420, 1481, 1720, 1508, 1536, 1560, 1586, 1611, 1636, 1664, 1690, 1482, 1721, 1537, 1587, 1612, 1665, 1691, 1748, 1990, 1772, 1799, 1830, 1857, 1883, 1910, 1938, 1963, 1749, 1991, 1773, 1800, 1831, 1858, 1884, 1911, 1939, 1964, 2019, 2253, 2039, 2065, 2092, 2118, 2143, 2172, 2199, 2226, 2020, 2254, 2040, 2066, 2093, 2119, 2144, 2173, 2200, 2227, 2281, 2517, 2307, 2336, 2363, 2390, 2416, 2441, 2466, 2492, 2282, 2518, 2308, 2337, 2364, 2391, 2417, 2442, 2467]. Matter [2493, 2544, 2774, 2568, 2592, 2619, 2645, 2668, 2697, 2725, 2749, 2545, 2775, 2569, 2593, 2620, 2646, 2669, 2698, 2726, 2750, 1031, 1509, 1561, 1637]. Matthew [1507]. Mauri [2440]. Max [216]. Maximal [438, 2125, 1108, 9, 31]. Maximization [173]. Maximizing [493]. Maximum [1245, 306, 1848, 1246, 377, 2086, 1657, 2663, 2727, 1592, 2301, 2016, 2496]. Maximum-Sized [2663]. Mazes [913]. Mazur [1067]. McCarthy [332]. Mean [1051, 251, 1997, 1150, 2601, 1756, 2460, 563, 379, 1053, 1592, 2149, 1558]. Meaner [2601]. Meanest [2601]. Means [884]. Measurable [1512]. Measure [2125, 1491, 205, 1889, 2149]. Measures [76, 101]. Measuring [2311, 2691]. Mechanical [322, 1692]. Medial [2272]. Median [698, 834]. Medians [1634]. Medical [1585]. Meets [1144, 698, 2707, 593, 1202]. Megginson [2570]. Mellin [1300, 1685]. Men [2591]. Menagerie [1809]. Mercator [479]. Meromorphic [1961]. Mersenne [2175]. Messer [2616]. Metaphor [1616]. Method [111, 1313, 2147, 1996, 2097, 678, 1752, 1970, 2346, 482, 423, 1311]. Methodologies [2497]. Methods [2574]. Metric [2702, 68, 41]. Metrical [265]. Metrics [2340, 563]. Meyer [2464]. Michael [102, 1609]. Michigan [1968]. Micro [2205]. Micro-Lesson [2205]. Middle [1159, 1320]. Military [198]. Millennial [771]. Millennium [1, 1675]. Miller [2415]. Milman [1718]. Min [216, 801, 833]. Min-Plus [801, 833]. Mina [869]. Minimal [870, 387]. Minimax [1567, 1544]. Minimizing [692]. Minimum [1115, 834, 2514]. Minkowski [1645, 2370, 1593, 2181]. Miracles [333]. Mirrors [1256]. Mis [1286]. Miss [2546]. Missed [1399]. Missing [475, 420]. Mistake [2067]. Misuse [2334]. Mixed [1261]. Mixing [1729]. Möbius [2137, 1398, 2771, 2287]. Model [1692, 2043]. Modeling [2516]. Models [1688]. Modern [1810, 2415]. Modifications [2650]. Modified [761]. Modular [1669, 1903, 2017, 436, 1532, 1656, 333]. Moduli [1699]. Modulo [554, 376, 2112, 1693, 1893]. Modulus [235]. Moment [918, 1863]. Moments [220]. Money [721]. Monic [392]. Monochrome [1233]. Monotonic [1934]. Monotonicity [1943, 1427]. Monster [1541, 1558, 2280]. Monte [979]. Monthly [1629]. Moore [1771, 531, 1311]. Mordell [786, 364, 1515, 2707, 499, 1092]. Mordell-Type [1092]. Morley [2021, 1632, 298, 1583]. Morrey [450]. Morse [2126, 1538]. Mortality [511]. Moser [2097]. Motion [2470, 1965, 787]. Motzkin [747, 1233]. Moving [1290, 1404]. Mr. [1285, 2591]. mth [2395, 607]. Much [1217]. Mullikin [2546]. Multi [361]. Multi-Dimensional [361]. Multidimensional [266]. Multimagic [2177]. Multiple [2732, 54, 357, 2528]. Multiples [720, 91, 2767, 943]. Multiplication [801, 833, 40]. Multiplicative [552]. Multipliers [2730]. Multiply [902]. Multistranded [1801]. Munshi [928]. Murty [1518]. Music [2773, 2773]. Musical [2647, 421, 927, 1892]. Musimathics [2773]. Musings [1882]. Must [452, 1928]. My [744]. Mylar [1114]. Mysteries [351]. Myth [2333].

N [11, 2252, 2464, 80, 554, 2435, 943, 2473, 1654, 2676, 1007]. N. [11, 80]. Napkin [2044, 2393]. Nash [2097]. Nation [1200]. Natural [2292, 2678, 1489]. Nature [2225, 396]. Nazis [1559]. NBA [722]. NCAA [2201]. Near [2745, 1670]. Near-de [2745]. Necessary [480]. Need [1775]. Needle [982]. Negative [1143, 692]. Neighborhood [947]. Neighborly [278]. Nested [2011, 1671, 983]. Neumann [956]. Never [2489, 2190]. Newman [424]. Newton [1421, 1663, 1089, 1849, 2147, 2574, 958, 2097, 2367, 564, 386]. Newton-Like [2574]. Nikodym [2401, 1038, 626]. Nilpotent [203]. Nine [1065, 298, 496]. Nine-Point [1065, 298, 496]. Ninth [2703]. No [2598, 683, 2250, 1767, 2671, 1318, 2425]. Noakes [1663]. Noetherian [981]. Non [2612, 2334, 1033, 787]. Non-Attacking [2612]. Non-Euclidean [1033, 787]. Non-Exhaustive [2334]. Noncommutative [981]. Nonconstant [311]. Noncrossing [1886]. Nondivisibility [861]. Nonholomorphic [1415]. Nonholonomy [2122]. Nonlinear [2012, 64, 43, 1692, 1994, 1423]. Nonmeasurable [2756]. Nonnegative [532, 1357, 2551]. Nonresidues [439]. Nonseparating [465]. Nonsingular [799]. Nonstandard [1700, 1640]. Nontransitive [1851]. Nontrivial [1278]. Noodle [1865]. Norbert [1937]. Norm [1780, 960, 236]. Normal [1835, 1565, 735, 55, 1403, 205, 133]. Normed [2151, 469]. Norms [110]. Notched [651]. Note [2757, 1234, 265, 1779, 2553, 2098, 2527, 1011, 819, 2262, 1759, 115, 1400]. Notes [1987, 2515, 2772, 101, 61, 63, 30, 64, 103, 104, 105, 65, 31, 33, 106, 34, 68, 35]. Noting [1892]. Notion [1967]. Novel [1479, 2257]. Nowhere [495, 79, 149, 708]. Nowhere-Differentiable [495]. Nth [1694]. Null [1822]. Nullstellen [1389]. Nullstellensatz [1757, 928]. Number [1340, 1090, 2770, 1173, 555, 2424, 590, 145, 637, 1485, 1639, 143, 2357, 1723, 2338, 2550, 104, 78, 1767, 44, 65, 670, 1739, 1562, 710, 2263, 829, 40, 73, 1331, 619, 1121, 420, 68, 41, 2415]. Number-Guessing [1739]. Number-Theoretic [1173]. Numbers [2314, 2578, 2540, 21, 891, 893, 1666, 189, 902, 2744, 1972, 1802, 669, 2292, 1881, 578, 297, 606, 757, 791, 534, 671, 697, 788, 2175, 247, 291, 2653, 1274, 1750, 1005, 343, 1355, 537, 2112, 731, 205, 203, 1489, 2577, 1439, 944, 681, 1289, 2374, 819, 393, 1554, 1079, 2087, 1023, 2485, 394, 1303, 2422, 1480, 2279]. Numerical [476, 620, 174].

O.K [813]. Obeys [2147]. Observation [1262]. Obsession [1856]. Obtained [640]. Obvious [419]. Odd [1933, 902, 791, 1036, 1824, 1330, 816]. Oddly [2614]. ODEs [1728]. Offering [2773]. Oh [744]. Old [2626, 1368, 141, 1291, 2388, 2222]. Oldie [1770]. Ole [1506]. Olinde [2198]. Olson [1689]. Olver [2362]. Olympiad [1360, 2358]. Olympics [263]. Omar [2567]. Once [958]. One [707, 521, 2135, 2626, 1605, 1228, 1578, 1632, 1499, 1875, 2272, 1262, 1038, 396, 1425]. One-Dimensional [1228]. Ones [2008]. Open [1695, 411]. Operation [2706]. Operators [2651, 332, 128, 960, 236, 568]. Oprea [2617]. Optimal [2339, 142, 1739]. Optimization [1567, 2597, 2644, 436, 1942]. Orbits [2392, 1202, 1769, 2575]. Orbs [1202]. Order [293, 248, 1826, 259, 1530, 1456]. Ordered [580, 579, 1804, 1782]. Ordering [145, 249, 2292]. Orders [2755, 692]. Ordinary [45, 61, 1631, 1457]. Orientations [331]. Origami [1511]. Origin [1311]. Orthic [813]. Orthocenter [1717, 1444, 1387]. Orthogonal [1615, 2300, 2221, 1855]. Ortiz [2198]. Oscillating [1712]. Oscillations [536]. Oscillators [1994]. Osculating [2075]. Ostrowski [2574]. Ostrowski-Like [2574]. Other [389, 541, 1193, 2623, 2094]. Our [1994]. Outcomes [2228]. Over-Iterates [2651]. Overhang [2727, 2521]. Overshooting [2574].

P [355, 1693, 376, 1331]. Pablo [2644]. Packing [1488, 1202, 1746]. PageRank [2464]. Pair [1039]. Pairings [1890]. Pairs [972, 123]. Palindromes [1539, 2371]. Panstochastic [324]. Papadimitriou [1479]. Paper [80, 2345, 1176]. Paper-Folding [2345]. Pappus [756]. Parabola [2754]. Parabolas [299]. Paradox [1756, 504]. Paradoxes [4, 2021]. Parallel [2624, 2520]. Parallelogram [804, 498, 903]. Parallelograms [2232]. Parameters [188, 2589]. Parity [2437]. Parker [1771]. Pars [2145].

Part [892, 2016, 693]. Partial [1261, 618, 2702, 104, 78, 816, 986, 1694]. Partite [437]. Partition [1332, 1619]. Partitioning [1716, 1474, 193, 2562, 2610]. Partitions [21, 20, 745, 2435, 1036, 2062, 1886, 22, 1330, 2636, 483]. Partner [1888]. Partnerships [313]. Parts [2435, 1332, 1330, 1804]. Pascal [389, 707, 1638, 1227, 435]. Pascal-Type [1638]. Path [1729, 1901, 124, 578, 1141, 494, 746]. Pathos [1302]. Paths [1502, 747, 1853, 1847, 2061, 1755, 1108, 465, 483]. Patrial [1985]. Pattern [102]. Patterns [2338, 2437, 1993, 2279, 1752]. Paul [69, 2724]. Peace [1616]. Peaks [2061]. Peculiar [2283]. Pedregal [2644]. Pell [126, 1904]. Pencil [1995]. Pencils [2444]. Pendulum [733]. Pengelley [2747]. Penrose [1202, 106, 74]. Pentagonal [21]. Pentagons [2042]. Perelman [2548]. Perennis [1810]. Perfect [2138, 902, 791, 2437, 343, 1957, 279, 1746, 1984]. Perimeter [2360, 2344, 870, 1593, 1486]. Period [846]. Periodic [111, 2366, 2392, 2312, 245, 1854, 2217, 1617, 1692, 582, 1769, 1794, 620]. Periodically [2312]. Periodicity [830, 1834]. Periods [405]. Permanent [2487, 1161]. Permutation [406, 188, 2615, 948]. Permutations [2637, 955, 2706, 1890, 1333, 1356, 2098, 404, 1328, 1892, 1759]. Permute [509]. Permuting [1804]. Perron [1490]. Perspective [1668, 2067]. Perspector [1412]. Peter [1718, 2362]. Petr [956]. Phase [2419]. Phenomenon [789]. Philip [2616]. Philosophies [1962]. Philosophy [1962, 1962]. Pi [1250, 1803, 2743]. Pick [900, 2181]. Pictures [534, 648, 1962]. Pie [2648]. Piece [2648]. Piles [2139]. Pillai [761]. Pinching [200]. Pioneer [36]. Pizza [2623]. Place [1231]. Planar [290, 2042, 2075, 331]. Plane [1805, 2366, 229, 140, 931, 2142, 2151, 2256, 1593, 2303, 1538]. Planets [564]. Platform [2501]. Player [2528, 35, 6]. Players [289, 35, 6]. Playground [1117]. Playing [1373]. Please [1414]. Plimpton [648]. Plucking [1303]. Plus [707, 1968, 801, 833]. Poesis [1941]. Poetry [1941]. Poincaré [325, 2548, 1730, 2576]. Point [132, 1502, 1426, 1469, 365, 535, 1670, 192, 1065, 1740, 219, 298, 496]. Pointed [1386]. Points [96, 523, 2127, 783, 245, 802, 143, 25, 1444, 1683, 128, 1164, 886, 582, 56, 1007, 217, 1505, 1065]. Pointwise [980, 2025]. Poisson [1519, 2049, 220]. Politics [1610]. Poly [819]. Poly-Bernoulli [819]. Polygon [2114, 1478, 2344, 1738]. Polygonal [393]. Polygons [229, 2070, 356, 1515, 2550, 1063, 2614, 171, 1543, 1806, 73, 1538, 264, 292, 2575, 174]. Polyhedral [2068]. Polyhedron [466]. Polymers [2673]. Polynomial [2538, 1050, 569, 771, 1673, 2165, 1320, 456, 1397, 2420, 2641, 377, 565, 1699, 2691, 1986, 2260, 1942, 2288, 1389, 1461, 2016, 638, 2562, 2610, 344]. Polynomials [2638, 1615, 1449, 1413, 1745, 2424, 2708, 1512, 263, 235, 2291, 392, 1693, 1982, 1424, 178, 1105, 597, 1833, 1357, 2207, 296, 2688, 1305, 1965, 1200, 1584, 1656, 1754, 2443, 2204, 731, 681, 533, 262, 1130, 1862, 1635, 394, 734, 624, 1516, 1635]. Polyominos [844]. Polytechnique [1513]. Poncelet [2231, 1543].

Popovidu [2722]. Portraits [2419]. Position [931, 1007]. Positional [145]. Positive [1173, 1986, 882, 982, 2347, 1782]. Positively [115]. Positivity [946]. Postulate [2520, 2675]. Potentials [2576]. Potter [1368]. Power [2397, 2708, 2383, 901, 1258, 946, 394, 816]. Powered [1442, 1660, 667]. Powerful [306, 215]. Powers [1566, 2138, 405, 2009, 2110, 606, 2661, 2690, 1929, 607, 692, 605, 1905, 1928]. Practice [1513]. Practicing [2228]. Praise [360]. Predictability [1834]. Predicting [2283]. Prediction [2170]. Prehistory [1484, 1915]. Presentation [1662]. Preserver [505]. Preservers [509]. Preserves [2484]. Preserving [434]. Prime [2110, 2708, 1329, 2643, 2357, 1723, 2338, 1232, 147, 2434, 1235, 901, 731, 1893, 2577, 1439, 1331, 2677, 1654, 7, 34, 2422, 1856]. Primes [929, 91, 263, 1638, 2096, 2488, 2022, 901, 634, 2600, 1039, 2675]. Primitive [23, 2310, 2488]. Prince [1174]. Princeton [2491]. Principal [1594, 1203]. Principle [2315, 789, 1592, 1646, 2176]. Probabilistic [1145, 2522, 818, 2447]. Probabilities [2705, 1483]. Probability [2205, 2121, 163, 2652, 982]. Problem [2056, 1976, 242, 2130, 15, 491, 2054, 793, 2736, 155, 2078, 2185, 2633, 2656, 2132, 2238, 1346, 2106, 1183, 1466, 1678, 598, 238, 271, 2379, 83, 514, 603, 1470, 52, 1345, 875, 398, 518, 910, 339, 574, 1325, 1523, 2240, 1951, 1702, 2635, 2426, 2429, 2454, 2605, 2507, 1126, 1326, 1601, 1813, 1897, 2188, 2682, 2709, 16, 118, 1472, 1844, 631, 1153, 1295, 1349, 1384, 2325, 632, 1785, 2030, 181, 243, 366, 427, 1576, 1653, 1679, 1787, 1816, 1927, 1952, 2008, 2209, 2486, 2295, 2738, 798, 1786, 2299, 2556, 2174, 1521, 1761, 2502, 1652, 1433, 1528, 187, 689, 2480, 547, 1596, 2712, 659, 2606, 2026]. Problem [2264, 1495, 1731, 991, 1766, 1871, 1188, 1494, 1571, 2505, 1873, 2757, 146, 1681, 1979, 1072, 2002, 1949, 1649, 663, 767, 1014, 1323, 1380, 2102, 2155, 2183, 2375, 1488, 151, 771, 2506, 2533, 2630, 1598, 2737, 2760, 48, 369, 401, 460, 549, 599, 909, 939, 1045, 1158, 1207, 1292, 1533, 1821, 1764, 82, 337, 794, 713, 880, 1360, 2456, 1626, 1812, 1682, 1123, 1187, 2270, 1213, 2044, 1410, 797, 1271, 1189, 370, 1269, 686, 1703, 2297, 2380, 464, 1406, 1648, 1898, 1926, 2377, 2358, 2716, 2050, 821, 156, 468, 335, 430, 1096, 1124, 968, 1647, 1677, 2321, 50, 116, 207, 399, 463, 573, 855, 1239, 1382]. Problem [1552, 1599, 1625, 1895, 2408, 2585, 688, 85, 372, 664, 990, 1602, 1704, 1843, 2216, 2269, 2452, 2657, 2693, 1431, 2055, 1872, 2597, 1436, 1529, 1525, 1784, 275, 2531, 1925, 2329, 629, 1106, 1381, 796, 180, 2634, 2393, 273, 49, 2582, 627, 2133, 2186, 2455, 2481, 1650, 1734, 2082, 2376, 84, 154, 852, 2409, 1347, 212, 1896, 2378, 1725, 575, 765, 828, 1016, 1868, 2156, 2187, 2268, 2406, 2530, 2680, 2294, 1989, 1377, 553, 1575, 987, 1405, 1042, 1099, 1550, 2005, 2604, 2714, 2405, 936, 1924, 1125, 1298, 1383, 1629, 1789, 2265, 1814, 1549, 661, 966, 214, 716, 876, 630, 730]. Problem [2431, 878, 761, 1922, 1100, 120, 2004, 2031, 2080, 2159, 2654, 1900, 1956, 2420, 685, 1348, 1763, 1841, 1870, 2081, 2211, 2298, 2427, 2583, 2679, 1155, 1572, 1623, 2581, 1955, 2131, 2215, 2355, 2558, 2609, 2764, 1299, 87, 1157, 2503, 419, 516, 1705, 2220, 47, 121, 2658, 1237, 1551, 1570, 1765, 717, 208, 244, 486, 737, 1522, 2101, 2381, 2681, 209, 431, 1874, 1953, 2128, 17, 371, 544, 304, 489, 769, 992, 2158, 2532, 2603, 600, 1148, 1975, 428, 18, 492, 1071, 2711, 1869, 1980, 660, 274, 400, 766, 1321, 1275, 572, 272, 397, 433, 461, 577, 604, 906, 1043, 1270, 1524, 1553, 1921, 2001, 2029, 2210, 2266, 2404, 2685, 2214]. Problem [2478, 2535, 2555, 1573, 210, 633, 2053, 213, 759, 665, 719, 764, 2348, 2476, 795, 963, 993, 1214, 1240, 1597, 2579, 787, 340, 2079, 239, 300, 690, 2482, 2632, 2735, 1185, 519, 826, 1296, 2213, 935, 86, 1322, 714, 1923, 2083, 2184, 2293, 2320, 2353, 2410, 2560, 2607, 2715, 911, 964, 1048, 2236, 334, 490, 1156, 1324, 2765, 1314, 817, 1651, 854, 1152, 1845, 2453, 182, 1017, 1379, 602, 241, 550, 739, 792, 1624, 1733, 823, 691, 1154, 1327, 2027, 2580, 240, 877, 962, 2584, 211, 2153, 2352, 432, 2430, 1603, 652, 1273, 517, 934, 1497, 2713, 1832, 2006, 1676, 119, 1465, 718, 1430, 2450, 1621, 2077, 2684]. Problem [2451, 1894, 488, 967, 1435, 302, 1735, 301, 152, 1600, 1817, 1977, 2534, 2766, 965, 1443, 184, 1074, 2349, 2479, 2683, 2734, 2761, 367, 403, 738, 1013, 1350, 2739, 206, 1018, 908, 940, 336, 122, 51, 402, 768, 1352, 1788, 1046, 1279, 1241, 1378, 1498, 1622, 2134, 429, 198, 1496, 2323, 2129, 2242, 1069, 513, 305, 2504, 571, 185, 1432, 539, 1102, 2154, 2528, 1811, 628, 851, 1760, 2007, 2103, 2350, 13, 1041, 662, 462, 186, 487, 770, 850, 1044, 1184, 1272, 937, 1208, 1526, 2324, 687, 1627, 853, 822, 1294, 1434, 2660, 2105, 2655, 740, 1019, 1209, 1680, 1954, 2051, 2631]. Problem [2762, 153, 1015, 879, 1127, 1351, 1242, 46, 1408, 1978, 1128, 1068, 1493, 1840, 2212, 387, 2296, 2052, 576, 827, 1574, 2319, 2710, 1706, 368, 601, 741, 684, 338, 1815, 1974, 2529, 458, 881, 988, 545, 7, 34, 961, 1238, 1520, 1899, 2241, 2432, 81, 1707, 1732, 1846, 1950, 2003, 2028, 2107, 2239, 2267, 2322, 2351, 2477, 2557, 2629, 2659, 2763, 898, 14, 905, 1122, 1407, 1471, 1548, 183, 1070, 1267, 1842, 938, 2428, 1212, 157, 2407, 2237, 1411, 459, 307, 856, 989, 715, 743, 515, 1097, 1293, 1467, 276, 546, 548, 825, 1210, 907, 1236, 2157, 2189, 2554, 1736, 303, 1098, 2608, 2740, 1762, 2222, 2508, 2559]. Problem [1186, 12, 1073, 1468, 1701, 1783, 1268, 1527, 1547, 1737, 2104, 1492, 824, 1409, 117, 270, 1047, 1101, 327, 2346, 2440]. Problems [1201, 1859, 1567, 476, 296, 505, 482, 2724]. Process [1768]. Product [1618, 2194, 569, 1329, 1244, 1985, 2421, 1860, 2652, 1826, 2067, 1179, 520, 2537, 1344, 2425, 1643, 2384, 1959, 1855, 2233]. Productive [2332]. Products [1946, 841, 2223, 1129, 1416, 2025, 1891, 2333, 246, 1556, 708, 732]. Professors [1663]. Program [531]. Programming [146, 2644]. Progression [2138, 306]. Progressions [2317, 424, 1716, 1474, 957, 2677]. Projectile [2470]. Projection [479, 1387]. Projections [1244, 2372, 1891]. Projective [1061, 1312, 1065]. Projectors [1855]. Prominent [2228]. Proof [2651, 2126, 2602, 267, 1618, 1757, 1568, 736, 625, 873, 2072, 2627, 2598, 1120, 2398, 1401, 1726, 1673, 453, 1342, 2023, 1142, 657, 2525, 2068, 1372, 2522, 874, 30, 10, 485, 2423, 1040, 658, 846, 1668, 147, 2341, 2370, 1177, 1544, 1947, 1180, 543, 1640, 1094, 2623, 2263, 928, 5, 2600, 2473, 2447, 385, 2548, 1674, 1318, 1672, 1375, 1727, 1265, 1592, 1233, 681, 568, 113, 177, 1620, 1289, 1973, 2700, 2524, 1038, 626, 268, 418, 1889, 1914, 1288, 1998, 1067, 712, 2628, 2422, 1317, 2233, 847, 237, 654, 542, 1463, 2449]. Proofreading [533]. Proofs [1517, 1339, 2475, 1962, 388, 2552, 1118, 175, 1032, 985]. Proper [735, 2622, 308, 1103]. Properties [299, 289, 2113, 2574, 479, 2123, 1388]. Property [438, 862, 132, 215, 2232, 2150, 942, 864, 55, 2536, 1403, 583, 1904, 1359, 1032]. Proportionality [597, 2229]. Proportions [2252]. Prove [2520, 1775]. Proved [2331]. Proving [1204, 1971]. Ptolemy [706, 2707, 469]. Public [2228]. Punctured [1984]. Pure [2616]. Purpose [1513]. Pursuers [142]. Pursuit [1746]. Putnam [231, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703, 2457, 232, 1371]. Putting [1231]. Puzzle [1146, 814]. Puzzles [1262]. Pythagoras [1669, 481]. Pythagorean [2759, 2373, 1781, 2113, 1957, 393, 278].

Qazi [1635]. Quadratic [2317, 666, 2398, 1514, 1088, 1177, 493, 439, 672, 638, 624, 295]. Quadrature [1402, 2033, 1229, 847]. Quadric [2599]. Quadrilateral [2666, 347, 2243, 498, 1473, 813, 2472, 441]. Quadrilaterals [2089, 218, 440, 24]. Qualitative [2312, 2289]. Quanta [1182]. Quantitative [2497]. Quantum [417, 954, 593, 649]. Quartic [1995, 1693, 2539]. Quasigroups [999]. Quaternions [1462]. Queens [2612]. Quest [2494, 981]. Question [790]. Questions [761, 1368]. Quick [2164, 543]. Quickie [2168]. Quintics [1753]. Quite [440, 1957]. Quotient [981, 2676].

R [225, 528, 1535, 69, 1337, 351, 782, 2386, 683, 2076, 2386, 2100, 1012]. R. [1771]. r/R [2386]. Raabe [2316]. Rabin [1233]. Races [1723]. Radical [75, 105, 2067]. Radicals [1011]. Radii [1634]. Radius [2191, 2304, 2167]. Rado [932]. Radon [2401, 1038, 626]. Rahman [1635]. Raindrops [2721]. Ram [1518]. Ramanujan [20, 2670, 199, 1260, 1399, 2675]. Rami [2491]. Random [409, 1666, 172, 2201, 406, 695, 815, 2094, 1008, 622]. Random-Turn [2094]. Randomly [973]. Range [1607]. Ranges [174]. Rank [722, 1959]. Ranking [2201, 1595]. Rankings [2464]. Rapidly [507]. Rare [927]. Rate [2305, 2691]. Ratio [2395, 2304, 1850, 2436, 2167]. Rational [1259, 1173, 94, 1163, 2252, 1077, 522, 2729]. Rationally [1686]. Rationals [114, 1250, 1080]. Ratios [774]. Raymond [1535, 1882]. Read [1629]. Real [2182, 1511, 1866, 102, 510, 2696, 450, 247, 582, 2288, 2025, 2016, 2120]. Reality [1116]. Rearing [2367]. Rearranged [1744]. Rearrangement [904, 363, 1429]. Rearranging [1888]. Recalling [2457]. Reciprocal [1960, 1416, 734]. Reciprocity [2314, 2398, 1514, 1177]. Recognize [2754]. Recognizing [1417]. Recounting [114, 1080]. Recover [2272]. Recovering [1724]. Recruitment [1006]. Rectangle [1290]. Rectangles [1174, 996, 2141]. Rectifying [931]. Recurrence [91, 2012, 2096, 2687, 190, 1359, 1135]. Recurrences [1743, 914]. Recurrent [2090, 131]. Recursion [281, 1319]. Recursive [968]. Recursively [1104, 696, 1503]. Reducibility [2207, 1263]. Reduction [228, 8, 33]. Rees [869]. Referential [2594]. Refinement [2316]. Reflections [1778, 1946, 494]. Reform [359, 531, 2591]. Regions [652]. Regular [1918, 1478, 259, 171, 1714, 820, 1738, 292, 2575]. Regularity [356, 2057]. Reguli [2365]. Reinhard [2747]. Related [2499, 761, 658, 479, 672, 1191, 1878, 1876, 298]. Relation [860]. Relations [2314, 1447, 1219]. Relationships [3]. Relative [522]. Relatives [499]. Relativity [72]. Relevance [321]. Remainder [1066]. Remark [1698, 1264]. Remarkable [426, 448]. Remarks [1675, 1540]. Reminiscent [751]. Renaissance [1810]. Rendezvous [68, 41]. Renewal [679]. Reorganizing [323]. Repeated [1823]. Repeating [1443]. Repetition [2008]. Representability [1696]. Representation [145, 2699, 1489]. Representations [1193]. Rescue [2111, 945, 19]. Rescued [2693]. Research [1286, 2724]. Residues [439, 672, 295]. Resistance [2470]. Resolutions [2365]. Respect [2523, 1728]. Responding [926]. Response [1337, 536]. Responses [1692]. Responsibility [321]. Restricting [248]. Restriction [955]. Result [2461, 635, 2499, 607, 1247, 681, 2343, 2425, 712]. Results [902, 2623]. Rethinking [2751, 2120]. Retraction [2598]. Return [2595]. Reuben [2225]. Reversal [2586]. Reverse [2251]. Review [194, 501, 1137, 413, 837, 1223, 1798, 1222, 1363, 585, 1362, 2116, 643, 255, 315, 285, 500, 894, 752, 1167, 1027, 1082, 1392, 1480, 2306, 2142, 136, 2091, 2334, 2491, 1829, 2696, 1280, 1306, 920, 1988, 135, 97, 1718, 1535, 2038, 380, 195, 1252, 2170, 1989, 1937, 1609, 1827, 1168, 2017, 699, 1663, 1610, 808, 223, 2667, 350, 1962, 2252, 2773, 975, 1747, 382, 165, 442, 836, 2362, 2724, 2644, 1081, 166, 2198, 2464, 1110, 2747, 1418, 284, 1479, 1882, 2389, 1026, 612, 586, 2516, 950, 1391, 558, 777, 1251, 316, 726, 1688, 1584, 443, 526, 351, 559, 1000]. Review [974, 222, 2225, 470, 807, 1057, 1507, 1856, 2279, 919, 444, 256, 26, 1450, 1746, 613, 1195, 673, 1001, 865, 1559, 1083, 2388, 1136, 2440, 1056, 381, 2117, 471, 1334, 2280, 412, 2064, 1635, 2591, 67, 525, 778, 1585, 2543, 2617, 1111, 2415, 2567, 2616, 642, 1689, 1166, 1506, 949, 1828, 349, 1771, 1908, 1281]. Reviews [809, 866, 921, 976, 1028, 1084, 1138, 27, 66, 98, 167, 224, 286, 352, 414, 472, 527, 587, 644, 700, 753, 1197, 1253, 1307, 1364]. Revisited [2621, 2733, 2023, 1291, 1114, 1642, 533, 1328, 1178, 2400, 2422, 540]. Revisiting [2626]. Revolution [1234, 2599]. Rhodes [1688]. Rhombic [106, 74]. Rhythms [1966]. Richard [2142, 1829, 1663, 1087]. Richards [32, 26]. Riemann [1856, 2578, 2741, 2368, 1464, 759, 2047]. Riesz [1506]. Right [763]. Right-Hand [763]. Rigid [161]. Ring [2397, 1673, 1594, 860]. Rings [2383, 981, 1614, 1584, 9, 31, 2014, 2639, 328]. Rino [709]. Rise [1078]. Risk [2595]. Rissanen [2516]. RNA [1801]. Robert [1798, 2116, 2570, 1962, 2279, 2616]. Roberts [2389]. Robertson [67]. Roberval [386]. Robin [919]. Rockmore [1856]. Rodrigues [2198]. Roger [2038]. Rogers [20]. Role [1142, 321]. Rolle [625]. Rolling [2122]. Romanian [2358]. Romanum [198]. Ronald [1584]. Ronan [2280]. Root [267, 683, 2013, 1109, 1318, 2385]. Rooted [1249, 2059]. Roots [1089, 497, 1745, 235, 392, 2488, 1355, 1671, 1182, 1461, 2318, 983, 2724].

Rotating [1969]. Rotation [2624, 2151, 2753]. Rotational [1805]. Rotations [2510, 1462]. Roughest [404]. Roulette [533]. Row [228, 1579]. Row-Stochastic [1579]. Rowed [2141]. Rows [927]. Ruin [2528]. Rule [1089, 1947, 2067, 422, 1206, 1753, 1542, 1317, 2596]. Rules [1943, 1427]. Running [142]. Runway [1395]. Rupert [2536, 1174]. Russian [533].

S [1798, 1829, 1688, 1584, 2388, 1012]. Saccheri [1473]. Saddle [2127]. Sally [2724]. Same [407, 188, 1767, 2060]. Sample [2741]. Sandwich [1181, 1972, 390]. Sanford [1559]. Sara [2731]. Satisfying [1594]. Saunders [1988, 1988]. Saved [2389]. Say [2460, 1801]. Scales [1892]. Schmeisser [1635]. School [842, 926]. Schoolgirls [957]. Schools [2571]. Schreier [1971]. Schröder [1853]. Schur [1545]. Schwarz [1609]. Science [2464]. Sciences [869]. SDRs [736]. Search [1937, 2464, 564, 1592]. Searches [533]. Seating [2577]. Secant [678]. Second [845]. Secondary [1981]. Secret [475]. Secrets [141]. Sections [2599]. Seductive [2422]. Seed [38]. Segal [1559]. Segments [916]. Selection [2094]. Self [2594, 108]. Self-Referential [2594]. Self-Similarity [108]. Semi [2148, 2075]. Semi-Osculating [2075]. Semi-Tile-Makers [2148]. Semidiscrete [2070]. Semigroups [511]. Semiregular [1806]. Semistable [1670]. Senates [2770]. Sequence [1448, 2246, 1390, 1881, 1712, 126, 696, 914, 131, 581, 1274, 1216, 1754, 1130, 1888, 2496, 773, 2636, 622]. Sequences [591, 2310, 2291, 2678, 2096, 2436, 788, 1656, 762, 1417, 1039]. Sequential [533]. Serendipitous [1025]. Series [2395, 2540, 1104, 1666, 596, 2670, 189, 524, 1776, 2397, 1589, 2000, 341, 2383, 1887, 2733, 774, 904, 2499, 378, 2038, 2511, 1024, 1744, 1093, 584, 1175, 567, 1442, 1258, 946, 2382, 2034, 2387, 348, 2541, 776, 2396, 1008, 2124, 1503, 1429, 556, 2500, 986, 1361, 1694, 654, 1876, 1907]. Servais [902]. Serve [1888]. Service [355, 2041, 2309, 2570, 69, 1200, 953, 1226, 1510, 1774]. Set [1249, 290, 365, 1797, 1683, 2756, 1768, 129, 2062, 2661, 8, 33, 1958]. Set-Valued [1768]. Sets [2594, 862, 201, 1245, 2612, 437, 2596, 553, 2676, 695, 193, 2372, 2025, 2562, 2610, 2636, 2060, 610, 115]. Setting [2259]. Seventh [2178]. Several [2424, 1944]. Shadows [2372]. Shakarchi [2491]. Shakiban [2362]. Shaking [1692]. Shape [2549]. Shapiro [1962]. Sharing [899]. Sharkovski [2292]. Sharkovskii [294]. Sharkovsky [1342, 2023]. Sharp [679, 932]. Sharp-Eyed [679]. Sharpening [2474]. Sheil [1635]. Sheil-Small [1635]. Shephard [176]. Shifrin [1535]. Shift [830, 2706]. Shifts [679, 1429]. Shooting [482]. Short [2651, 2126, 736, 2561, 1726, 1673, 406, 657, 874, 147, 1947, 681, 177, 1289]. Shortened [2398]. Shortest [1141]. Shots [2472]. Shrinking [1582, 636]. Shuffles [1218]. Shuffling [2728]. Shurman [2017]. Side [2553]. Sides [1131, 901]. Siegelman [1937]. Sign [480, 246]. Signed [1890, 2541]. Signs [1089, 1947, 1317, 2596, 1206]. Similarities [388, 2108]. Similarity [108]. Simon [2198]. Simple [1568, 1020, 625, 294, 1745, 1971, 134, 2598, 1512, 1726, 2678, 2597, 1342, 2023, 1372, 1040, 658, 846, 1544, 1180, 1780, 2045, 2263, 2260, 1358, 1672, 568, 2528, 2524, 422, 1032, 35, 6, 1317, 237]. Simpler [1459, 542]. Simplest [1835]. Simplex [1717, 820, 1606]. Simplices [1918, 2342, 25]. Simplicial [1945]. Simplifying [2125]. Simpson [1753, 1542]. Sinc [2495]. Sine [1710]. Sines [639, 1504, 1179, 2034, 162, 1906]. Singer [2170]. Single [914, 35, 6]. Singular [1618, 1545, 1519, 259]. Singularities [2180]. Sinh [2438, 2511]. Sink [1106]. Siobhan [2389]. Sion [1544]. Sites [323]. Six [998, 1401, 1371, 846, 192, 1689, 1708]. Sixteen [1456]. Sixth [2457, 1916]. Sixtieth [231]. Sixty [2457, 1371, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703]. Sixty-Eighth [2446]. Sixty-Fifth [1641]. Sixty-First [566]. Sixty-Fourth [1370]. Sixty-Ninth [2703]. Sixty-Second [845]. Sixty-Seventh [2178]. Sixty-Six [1371]. Sixty-Sixth [2457, 1916]. Sixty-Third [1091]. Sized [2663]. Sizes [1245]. Sketches [1087]. Skew [1782]. Skipping [1502]. Slide [422]. Sliding [1277, 1969, 2459]. Slowly [2396]. Small [1150, 1635]. Smallest [1604]. Smith [327]. Smooth [930, 79]. Smoothest [404]. Snakes [679]. Snapshots [2274]. So-Called [2547]. Soap [2071]. Social [2198, 44, 65]. Solid [2536]. Solids [1861]. Solomon [1584, 1828]. Solution [2755, 2275, 146, 1854, 2597, 176, 787, 817, 1794, 2032, 2528, 387, 2221, 1025]. Solutions [134, 2312, 1173, 1567, 1728, 373, 1725, 252, 2036, 282, 1670]. Solvability [1011]. Solvable [2433]. Solve [652]. Solving [1995, 1204, 1989, 296, 2687, 2448, 2202, 71]. Some [2073, 666, 1449, 1743, 1725, 2572, 55, 2475, 2228, 761, 141, 1305, 567, 448, 391, 1675, 762, 1967, 479, 1353, 1344, 451]. Sometimes [1825, 2356]. Source [1106]. Space [1035, 2365, 132, 729, 931, 608, 332, 1837, 1683, 2389, 1891, 2108, 469, 236, 772, 68, 41, 109]. Spaces [650, 2182, 2702, 1887, 2042, 534, 1538, 1316]. Spaghetti [1865]. Span [581]. Spanning [1021, 115]. Sparse [800]. Special [523, 294, 637, 1628, 747, 1683, 2488, 2434, 641, 72]. Specified [249]. Spectral [1339, 358, 420]. Spectrum [1920]. Sperner [1343, 130]. Sphere [1221, 1119, 555, 802, 456, 322, 2122, 2033, 1202, 1758, 8, 33]. Spheres [1861, 330, 2290, 2285]. Spherical [1631, 310, 2272]. Spigot [1803]. Spinning [2721]. Spiral [1230, 2613, 2611]. Split [365]. Splitting [246]. Sponges [2224]. Spreads [2365]. Squarable [206]. Square [2489, 497, 267, 409, 1488, 356, 864, 2769, 2013, 58, 2553, 2411, 1318, 1671, 279, 983, 2385]. Square-Banded [356]. Square-Packing [1488]. Squared [1329, 378, 1353]. Squares [1285, 1035, 668, 1090, 1851, 1401, 2177, 971, 70, 2300, 2063, 972, 2553, 408, 947, 996, 2411, 259, 1489, 475, 262, 123, 2221, 1984, 1782, 732]. Squaring [2256]. Squashing [892, 693]. Stable [1670, 1284]. Stacked [2206]. Stalking [1856]. Standard [1200]. Stanford [1276]. Start [5]. State [1968, 163]. States [2770]. Stationary [1579, 1670]. Statistical [1918, 2516]. Statistics [202, 3]. Steele [1609]. Steepest [1141, 2047]. Stegun [1885]. Stein [2491]. Steiner [730, 541]. Stephanie [2170]. Stephen [2064]. Stephens [953]. Stepping [2057, 2022]. Steve [1689]. Steven [2091, 2415].

Stevens [1226]. Stewart [1962, 2306]. Stick [1877]. Stick-Breaking [1877]. Sticky [1275]. Still [141, 1038]. Stirling [277, 2733, 297, 744, 1093, 584, 291, 711, 2473, 2049, 944, 177, 1303]. Stochastic [1257, 1579]. Stokes [1221]. Stone [508, 1642]. Stories [477]. Story [417, 2571, 941]. Straffin [2616]. Straight [1565, 386]. Straightedge [651, 325]. Strang [1535]. Strategies [984]. Strategy [142, 1739, 2701, 198]. Strauss [2667]. Streetwise [351]. Strengthening [508]. Stretching [2369]. Striling [2087]. String [145]. Strip [2329]. Strong [2418]. Structure [2701]. Structures [590]. Student [1286]. Students [1968]. Studied [199]. Studies [2228]. Study [2497, 1968]. Studying [2312]. Subcovers [411]. Subdivision [1945]. Subgroups [2755, 1565, 735, 2622, 1235, 1103, 1344]. Submultiplicativity [524]. Subsequence [1425]. Subset [1792]. Subset-Sum [1792]. Subsets [2676, 1266, 775, 127]. Subspaces [308, 478, 772]. Subtangents [758, 871]. Successful [160]. Successors [2767]. Sudoku [2365, 2300, 2701, 2221]. Suffice [1009]. Sufficient [1709, 125, 1543, 480]. Sum [434, 1630, 668, 862, 2354, 2590, 2330, 2642, 1658, 1796, 2767, 1244, 2223, 2063, 832, 95, 2509, 2361, 1160, 346, 2344, 567, 2690, 2112, 92, 520, 439, 672, 1501, 1792, 944, 776, 692, 1634, 2718, 2562, 2610, 2636, 394, 1220, 1694, 1936, 1907]. Summability [64, 43]. Summandification [1654]. Summation [2621, 1852, 557, 1399, 2500]. Summing [2009, 1982, 2396]. Sums [405, 1035, 2495, 682, 1060, 799, 376, 1819, 2661, 1248, 1416, 1912, 1489, 262, 2318, 1863, 2347, 986, 848]. Superabundant [2578]. Superposition [1345]. Surds [666]. Sure [362]. Surface [2043]. Surfaces [1234, 2599]. Surfer [815]. Surprise [173]. Surprises [1286, 760, 2123, 1708]. Surprising [2495, 250, 2099, 1886]. Surprisingly [1020, 2597]. Surround [710]. Survey [2670]. Survival [1117]. Survive [1928]. Suspension [536]. Swaps [2586]. Symmetric [93, 283, 1545, 2699, 1569, 2625, 2688, 882, 2359, 68, 41, 2169]. Symmetries [2667]. Symmetry [2205, 1402, 1426, 2170, 927, 2279, 2202, 2742, 150, 394, 2422, 2280]. Systems [1345, 143, 679, 2752, 1423, 2345, 2419, 535, 1970, 1834, 704, 2060]. Szekeres [1698]. Szpiro [1746].

T [1226, 447]. Table [2577]. Tacky [1659]. Tail [1539]. Tails [1565, 1589]. Takagi [930]. Take [1178]. Taken [1260]. Takes [592]. Takloo [2415]. Takloo-Bighash [2415]. Tamed [1558]. Taming [2180]. Tangent [1630, 2089, 330, 56, 1793]. Tangents [871, 386]. tanh [774]. Tarski [1828, 2262]. Tasty [2623]. Taylor [1066]. Teacher [1771]. Teaching [1142, 1513]. TEAM [655]. Teams [1752]. Technique [1519]. Telegraphic [809, 866, 921, 976, 1028, 1084, 1138, 27, 66, 98, 167, 224, 286, 352, 414, 472, 527, 587, 644, 700, 753, 1197, 1253, 1307, 1364]. Telescoping [2509]. Ten [1859]. Tend [2048]. Tension [202]. Terminal [1919]. Terms [1881, 1442, 2541]. Terry [1827, 1635]. Test [2395, 2316]. Tests [2395]. Tetrahedra [161]. Tetrahedron [1088, 595, 1132, 1115, 329]. Tetron [426]. th [2473]. Thébault [418]. Their [668, 1231, 1804]. Theme [251, 1176]. Themes [2388]. Theodore [1535]. Theodorus [1230]. Theorem [2638, 1261, 424, 1221, 2602, 324, 1655, 736, 625, 294, 1459, 873, 1259, 2627, 2598, 1120, 2494, 2259, 1401, 1051, 2292, 2334, 1064, 1755, 453, 297, 1490, 1342, 2023, 1698, 2373, 843, 1339, 657, 2075, 440, 342, 611, 1781, 1879, 956, 1591, 2021, 485, 2230, 2423, 2418, 791, 2674, 656, 1205, 1644, 1343, 2341, 1544, 1864, 2601, 358, 1808, 2401, 1180, 2179, 1640, 2047, 1094, 1917, 1149, 756, 2263, 1116, 2181, 1674, 1672, 933, 1375, 2753, 2526, 1265, 1118, 130, 481, 1970, 1620, 1973, 2700, 477, 2524, 1038, 626, 268, 2024, 1914, 264, 1147, 2695, 1037, 328, 1425]. Theorem [932, 237, 2730, 540, 1463, 1400, 2449, 1971, 1181, 508, 1972, 390, 2522, 450, 706, 541, 705, 900, 1233, 1642, 418, 1147, 1067, 2628]. Theorems [1517, 104, 78, 261, 2552, 1143, 1592, 985]. Theoretic [1173, 1619]. Theory [590, 1485, 1639, 2091, 2699, 143, 265, 104, 78, 1644, 1667, 205, 1143, 593, 8, 33, 619, 1538, 420, 1635, 2176, 734, 2415]. There [1285, 290, 454, 789, 1203]. Thin [2372]. Things [2667]. Think [1686]. Thinking [1962]. Third [2174, 1091]. Thomae [2650]. Thomas [2624]. Thompson [351]. Thomsen [784]. Thread [1491]. Three [434, 1426, 608, 914, 1982, 2589, 1767, 1441, 1103, 2440]. Three-Dimensional [608]. Three-Point [1426]. Thrillers [351]. Thue [2126]. Ticky [1659]. Ticky-Tacky [1659]. Tile [2148, 710]. Tile-Makers [2148]. Tiles [971, 1993]. Tiling [229, 70, 1966, 996, 1314, 2141]. Tilings [111, 971, 1403, 176, 1202, 106, 74]. Tilling [2485]. Time [26, 32, 2388, 2449]. Timely [1533]. Times [1049, 1810, 1752, 1610, 2198]. Timing [1151]. Tlirn [1675]. Tolstoy [1616]. Tomaso [1746]. Tone [927]. Too [1078, 359, 1217]. Tools [2312]. Topological [1146, 1809]. Topologies [915]. Topology [650, 140, 729, 1695, 1837, 2208, 2616, 1827, 2616]. Torkel [2334, 2334]. Torsional [536]. Total [406, 595]. Totally [1511, 532, 331, 1782]. Totient [1329, 832, 670]. Totients [943]. Touchard [791]. Toughest [1689]. Tour [1798, 925, 2069]. Touring [1455]. Tournament [38]. Tournaments [289, 35, 6]. Tower [1276]. Towering [506]. Trace [2223]. Traceless [1088]. Tracking [2661]. Trammel [2547]. Transcendence [2126, 2116]. Transcendental [148]. Transform [1300, 1685, 1880]. Transformation [2045]. Transformations [2594, 597, 1398, 2563]. Transforms [293]. Transitivity [846]. Translating [1316]. Translation [2346]. Translations [2510]. Transparent [2116]. Transport [2624]. Transposition [2260]. Transpositions [1696, 1129]. Trapezoid [2666, 1791]. Treatment [2334]. Tree [1669, 694]. Tree-Counting [694]. Trees [1666, 1249, 580, 1697, 579, 1804, 1303]. Trial [341]. Triangle [96, 523, 2392, 2244, 1638, 330, 2218, 2463, 1504, 1836, 455, 2692, 886, 2573, 1593, 24, 1386, 813, 217, 2272, 423, 1092, 1192, 1534, 2412, 2277, 1583]. Triangles [309, 1631, 2058, 635, 1055, 1063, 872, 870, 901, 159, 1632, 2443, 1007, 1957, 2448, 844, 1555, 2542, 395, 298]. Triangular [2311, 2461, 204, 435]. Triangulated [2614]. Triangulation [1902]. Tricky [2514]. Tridiagonal [1579]. Trig [2564, 2565]. Trig-Of-Trig [2565].

Trigonometric [1487, 467, 2204, 451]. Trigonometries [785]. Trigonometry [2252, 1148]. Trimming [200]. Trios [1714]. Trip [2301]. Triperfect [893, 697]. Triple [1532, 246]. Triples [2113, 204, 393, 278]. Triplets [40]. Trisections [1511]. Trivia [232]. Trivial [1751, 1563]. Trott [102]. Truncated [774]. Truncating [200]. Tsolomitis [1718]. Tubbs [2116]. Tumbling [329]. Tuples [668]. Turbulent [1610]. Turing [1479]. Turn [1, 2094]. Twentieth [457, 531]. Twentieth-Century [531]. Twenty [898]. Twenty-Fourth [898]. Twice [651, 466]. Twice-Notched [651]. Twin [1638, 634, 7, 34, 504, 916]. Twisted [2469, 109]. Two [2637, 407, 2717, 309, 267, 2139, 1260, 1075, 1539, 1633, 11, 2135, 843, 1339, 760, 706, 376, 2223, 80, 448, 343, 1118, 1262, 1534, 1945, 1065, 425, 654]. Two-Complexes [1945]. Two-Dimensional [425]. Two-Envelope [1262]. Two-Triangle [1534]. Twos [448]. Tychonoff [2494, 1147]. Type [124, 1638, 679, 1092].

Ugly [2367]. Ukrainian [1360]. Ulam [1067, 2400]. Ultimate [177]. Ultrafilters [2258]. Ultrametric [534]. Umbrella [2721]. Unbounded [1803]. Uncertainty [2176]. Unconventional [2225]. Uncountably [127]. Undergraduates [1989]. Understanding [2644]. Unexpected [1692, 997, 1079]. Unified [291]. Uniform [682, 1997, 592, 787, 483, 725]. Uniformly [2665, 2025]. Union [735, 1741, 308]. Unions [2622]. Unipotent [2662]. Unique [2275, 1775, 624]. Uniqueness [2523]. Unit [1119, 392, 103, 77, 2553, 377, 2344, 1007, 1486, 282, 621, 362]. Unity [235]. Universal [2311, 2252]. Universalis [2145]. Universe [504]. University [1663, 1968]. Unlabeled [2161]. Unsigned [944]. Unsolved [2326, 2250]. untitled [194, 501, 1137, 413, 779, 1196, 837, 1223, 1222, 1363, 585, 1362, 643, 255, 315, 285, 500, 894, 752, 1167, 1027, 1082, 1392, 136, 1280, 1306, 920, 135, 97, 380, 195, 1252, 1168, 699, 808, 223, 350, 975, 382, 165, 442, 836, 1081, 166, 1110, 1418, 284, 1026, 612, 586, 950, 1391, 558, 777, 1251, 316, 726, 443, 526, 559, 1000, 974, 222, 470, 807, 1057, 444, 256, 1450, 613, 1195, 673, 1001, 865, 1083, 1136, 1056, 381, 471, 1334, 412, 525, 778, 1111, 642, 1166, 949, 349, 1281]. Unusual [468]. Unwrapping [2095]. Upper [608]. Urn [968, 2161, 2220]. Urns [375, 721]. Use [2334, 1286]. Used [871]. Useful [250, 508]. Using [1221, 651, 293, 249, 2394, 2644, 1200, 2652, 3, 108]. Ut [1941]. Utopias [2198].

V [2388]. Valleys [2665]. Valtonen [2440]. Value [23, 1051, 251, 1545, 296, 419, 2758, 883, 2601, 2512, 482]. Valued [534, 1768]. Values [2637, 1592, 2562, 2610]. Vandermonde [2072]. Vanishing [2354, 1658]. Varadarajan [2388]. Varga [1829]. Variable [912]. Variables [1075, 2424, 1668, 357]. Variance [112, 1608]. Variant [2512]. Variation [1360, 251, 1847, 1967]. Variations [1917, 1176]. Veblen [592]. Vector [729, 931, 485, 772, 1178, 2576]. Velleman [1962]. Velocities [2274].

Velocity [1919, 2301]. Venn [2422]. Version [2418, 361]. Versions [541]. versus [2701]. Vertex [309, 1063, 1505, 264]. Vertical [2724, 536]. Vertices [2344, 1714]. Very [2744, 1180, 161, 681]. via [2469, 2373, 1644, 1343, 2181, 2049, 1531, 985, 1147]. Vie [1689]. View [535]. Viewpoint [842]. Virtual [1667]. Visible [590]. Visited [536]. Visual [758]. Vitaly [1718]. Volterra [477]. Volume [318, 615, 895, 1169, 1452, 1722, 1992, 2255, 2519, 2776, 25, 1491, 103, 77, 1150, 543, 387]. Voting [2284, 2699, 704]. vs [1913]. Vygodskii [11, 80].

W [2696]. W1 [2076]. Waerden [930]. Waiting [803, 1752]. Walk [957, 857]. Walker [2201]. Walks [1666, 608]. Wallis [2652, 2447, 2233]. Want [1049]. War [1616, 592]. Was [2546]. Water [1141, 2043]. Wave [1261]. Waves [2043]. Way [1971, 2257, 38]. Ways [407, 1441]. Weaire [1746]. Weak [1837]. Weakest [2595, 984]. Web [323]. Webb [67]. Wedderburn [1094]. Wedge [2421]. Weierstrass [508, 1642, 570]. Weighted [407, 2741, 364, 1515]. Well [289]. Well-Matched [289]. Were [1908]. Where [5, 1438]. Which [1061, 2046, 2433, 1178, 396]. Who [2546, 2389]. Whose [2767, 1063, 653, 943]. Width [290, 1458, 1115]. Wielandt [2474, 1368]. Wiener [1937]. Wigner [2624]. Wild [1750]. Wildberger [2252]. Wiley [1718]. Wilker [1427]. William [1989, 231, 566, 845, 1091, 1370, 1641, 1916, 2178, 2446, 2703, 67]. Wilson [1517]. Winckler [1568]. Wire [817]. Without [31, 2369, 2331, 1546, 2088, 9, 205, 396]. Witten [2286]. Wolf [1798]. Wolstenholme [2124]. Women [1747]. Won’t [1160]. Woods [351]. Wooley [1750]. Words [23, 648]. Work [1087, 1948]. Working [2571]. World [263, 592, 1962, 709, 1689]. Wrapped [1661]. Wunderkammern [562].

x [283, 392, 774, 42, 2100].

Ya [11, 80]. Year [1931, 141]. Years [1371]. Yielding [378]. Young [2306, 332]. Yueh [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Yueh-Gin [355, 2041, 2309, 2570, 69, 953, 1226, 1510, 1774]. Yuli [1718].

Z [392, 53]. Zero [2048, 1476, 1982, 1278, 2100, 982, 1530, 2016, 857, 311]. Zero-Divisors [2100]. Zeroes [1694]. Zerofinders [1942]. Zeroless [145]. Zeros [235, 2165, 1982, 1699, 1595, 2288, 113, 1012, 1438, 986]. Zeta [2741, 1960, 2286, 551, 348, 1012, 2168]. Zeta-Function [2286]. Zetafunctionology [1133]. Zoo [1940].

References

Griffiths:2000:MTM

[1] Phillip A. Griffiths. Mathematics at the turn of the millennium. American Mathematical Monthly, 107(1):1–14, January 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589372.

Groetsch:2000:CC

[2] C. W. Groetsch. The Commandant’s corollary. American Mathematical Monthly, 107(1):15–23, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589373.

Reiter:2000:USD

[3] Jerome Reiter. Using statistics to determine causal relationships. American Mathematical Monthly, 107(1):24–32, January 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589374.

Case:2000:PIC

[4] James Case. Paradoxes involving conflicts of interest. American Mathematical Monthly, 107(1):33–43, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589375.

McClure:2000:SWT

[5] J. E. McClure. Start where they are: Geometry as an introduction to proof. American Mathematical Monthly, 107(1):44–52, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589376.

Tyszkiewicz:2000:SCT

[6] Jerzy Tyszkiewicz. A simple construction for tournaments with every k players beaten by a single player. American Mathematical Monthly, 107 (1):53–54, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589377.

Suzuki:2000:AFT

[7] Maria Suzuki. Alternative formulations of the twin prime problem. American Mathematical Monthly, 107(1):55–56, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589378.

Russell:2000:ERI

[8] Alexander Russell. An easy reduction of an isoperimetric inequality on the sphere to extremal set theory. American Mathematical Monthly, 107 (1):57–59, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589379.

Malcolmson:2000:RMI

[9] Peter Malcolmson and Frank Okoh. Rings without maximal ideals. American Mathematical Monthly, 107(1):60–61, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589380.

Hagedorn:2000:PCE

[10] Thomas R. Hagedorn. A proof of a conjecture on Egyptian fractions. American Mathematical Monthly, 107(1):62–63, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589381.

Demidov:2000:ELT

[11] S. S. Demidov and A. Shenitzer. The evolution of … two letters by N. N. Luzin to M. Ya. Vygodskii. American Mathematical Monthly, 107 (1):64–82, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589382.

Zara:2000:P

[12] Catalin Zara. Problem 10774. American Mathematical Monthly, 107(1): 83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589383.

Salmasian:2000:P

[13] Hadi Salmasian. Problem 10775. American Mathematical Monthly, 107 (1):83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589384.

Tian:2000:P

[14] Yongge Tian. Problem 10776. American Mathematical Monthly, 107(1): 83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589385.

Ahmed:2000:P

[15] Zafar Ahmed. Problem 10777. American Mathematical Monthly, 107(1): 83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589386.

Bateman:2000:Pa

[16] Paul Bateman and Dennis Eichhorn. Problem 10778. American Mathematical Monthly, 107(1):83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589387.

Jorza:2000:P

[17] Andrei Jorza. Problem 10779. American Mathematical Monthly, 107(1): 83, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589388.

Kedlaya:2000:P

[18] Kiran Kedlaya. Problem 10780. American Mathematical Monthly, 107 (1):84, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589389.

Robertson:2000:ECR

[19] John P. Robertson and GCHQ Problems Group. Elliptic curves to the rescue: 10612. American Mathematical Monthly, 107(1):84–86, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589390.

Andrews:2000:RRP

[20] George E. Andrews, John Henry Steelman, G. L. Body, D. Callan, R. J. Chapman, K.-W. Lau, J. T. Lewis, J. H. Lindsey II, G. Lord, J. G. Merickel, and Anchorage Math Solutions Group. Rogers–Ramanujan partitions: 10627. American Mathematical Monthly, 107(1):86, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589391.

Andrews:2000:PPN

[21] George E. Andrews, Kee-Wai Lau, G. L. Body, D. Callan, R. J. Chapman, G. Lord, H.-J. Seiffert, and D. Zeilberger. Partitions and pentagonal numbers: 10628. American Mathematical Monthly, 107(1):86–87, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589392.

Schmidt:2000:IP

[22] Frank Schmidt, Peter Mork, D. Beckwith, D. Callan, D. J. Chapman, G. Lord, C. C. Rousseau, P. Simeonov, and J. H. Steelman. Interrupted partitions: 10629. American Mathematical Monthly, 107(1):87–88, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589393.

Callan:2000:VPW

[23] David Callan, Robin J. Chapman, J. H. Lindsey, II, and D. Huang. The value of primitive words: 10636. American Mathematical Monthly, 107 (1):88–89, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589394.

Mazur:2000:QIT

[24] Marcin Mazur, Roy Barbara, J. Anglesio, R. J. Chapman, N. Lakshmanan, J. Melville, G. Peng, C. R. Pranesachar, and GCHQ Problems Group. Quadrilaterals in an isosceles triangle: 10653. American Mathematical Monthly, 107(1):89–91, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589395.

Golomb:2000:PFS

[25] Michael Golomb, GCHQ Problems Group, G. D. Chakerian, R. J. Chapman, J. H. Lindsey II, O. P. Lossers, and A. Nijenhuis. Points that form simplices of equal volume: 10685. American Mathematical Monthly, 107 (1):91, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589396.

Pratt:2000:BRB

[26] John P. Pratt. Book review: Mapping Time: The Calendar and Its History., by E. G. Richards. American Mathematical Monthly, 107(1): 92–99, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589397.

Ostebee:2000:TRa

[27] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(1):100–102, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589398.

Anonymous:2000:BMa

[28] Anonymous. Back matter. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589399.

Anonymous:2000:FMa

[29] Anonymous. Front matter. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589371.

Hagedorn:2000:NPC

[30] Thomas R. Hagedorn. Notes: a proof of a conjecture on Egyptian fractions. American Mathematical Monthly, 107(1):??, January 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Malcomson:2000:NRM

[31] Peter Malcomson and Frank Okoh. Notes: Rings without maximal ideals. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pratt:2000:MTC

[32] John P. Pratt. Mapping Time: The Calendar and Its History, by E. G. Richards. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Russell:2000:NER

[33] Alexander Russell. Notes: An easy reduction of an isoperimetric inequality on the sphere to extremal set theory. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Suzuki:2000:NAF

[34] Maria Suzuki. Notes: Alternative formulations of the twin prime problem. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tsyzkiewicz:2000:NSC

[35] Jerzy Tsyzkiewicz. Notes: a simple construction for tournaments with every k players beaten by a single player. American Mathematical Monthly, 107(1):??, January 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Donaldson:2000:EFC

[36] James A. Donaldson and Richard J. Fleming. Elbert F. Cox: An early pioneer. American Mathematical Monthly, 107(2):105–128, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589433.

Eisenberg:2000:CDE

[37] Bennett Eisenberg and Rosemary Sullivan. Crofton’s differential equation. American Mathematical Monthly, 107(2):129–139, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589434.

Schwenk:2000:WCW

[38] Allen J. Schwenk. What is the correct way to seed a knockout tournament? American Mathematical Monthly, 107(2):140–150, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589435.

Enayat:2000:CF

[39] Ali Enayat. δ as a continuous function of x and ϵ. American Mathematical Monthly, 107(2):151–155, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589436.

Pfaff:2000:CMN

[40] Frank R. Pfaff. A commutative multiplication of number triplets. American Mathematical Monthly, 107(2):156–162, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589437.

Thomassen:2000:RNS

[41] Carsten Thomassen. The rendezvous number of a symmetric matrix and a compact connected metric space. American Mathematical Monthly, 107(2):163–166, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589438.

Erdelyi:2000:EDX

[42] Tamas Erdelyi. On the equation a(a + d)(a + 2d)(a + 3d) = x^{2}. American Mathematical Monthly, 107(2):166–169, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589439.

Herzog:2000:NS

[43] Gerd Herzog. On nonlinear summability. American Mathematical Monthly, 107(2):170–171, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589440.

Luca:2000:ASF

[44] Florian Luca. The anti-social Fermat number. American Mathematical Monthly, 107(2):171–173, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589441.

Beck:2000:FEO

[45] Anatole Beck. A fractal example in ordinary differential equations. American Mathematical Monthly, 107(2):174–175, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589442.

Smiley:2000:P

[46] Leonard Smiley. Problem 10781. American Mathematical Monthly, 107 (2):176, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589443.

Iannucci:2000:Pa

[47] Douglas Iannucci. Problem 10782. American Mathematical Monthly, 107 (2):176, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589444.

Chao:2000:P

[48] Wu Wei Chao. Problem 10783. American Mathematical Monthly, 107 (2):176, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589445.

Facchini:2000:P

[49] Alberto Facchini and Francesco Barioli. Problem 10784. American Mathematical Monthly, 107(2):176, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589446.

Deutsch:2000:Pa

[50] Emeric Deutsch. Problem 10785. American Mathematical Monthly, 107 (2):176, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589447.

Quet:2000:P

[51] Leroy Quet. Problem 10786. American Mathematical Monthly, 107(2): 177, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589448.

Arias-de-Reyna:2000:P

[52] Juan Arias de Reyna. Problem 10787. American Mathematical Monthly, 107(2):177, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589449.

Koepp:2000:IZH

[53] Warren Koepp, Robin Chapman, J. H. Lindsey II, and NSA Problems Group. On the intersection of Z^{n} with a hyperplane: 10639. American Mathematical Monthly, 107(2):177, February 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589450.

Filaseta:2000:WMC

[54] Michael A. Filaseta, Allen Stenger, D. Callan, and Z. Franco. When a multiple of π/2 is close to an integer: 10640. American Mathematical Monthly, 107(2):177–179, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589451.

Gross:2000:PSN

[55] J. Gross, G. Trenkler, S.-O. Troschke, William F. Trench, D. Callan, G. Wahba, R. J. Chapman, L. M. DeAlba, J. H. Lindsey II, and IUTS Problems Group. A property of some normal matrices: 10645. American Mathematical Monthly, 107(2):179–180, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589453.

Mazur:2000:TLC

[56] Marcin Mazur, Kit Hanes, J. Anglesio, M. Benedicty, S. B. Ekhad, N. Lakshmanan, A. Nijenhuis, and J. H. Smith. Tangent lines and collinear points: 10673. American Mathematical Monthly, 107(2):180– 181, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589454.

Chao:2000:IE

[57] Wu Wei Chao, Michael Reid, F. Bellot Rosado, R. J. Chapman, D. Donini, S. B. Ekhad, N. Lakshmanan, O. P. Lossers, A. Nijenhuis, P. E. Nuesch, and C. G. Petalas. Incenters and excenters: 10693. American Mathematical Monthly, 107(2):182–184, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589457.

Heuer:2000:SG

[58] Gerald Heuer, Kenneth F. Andersen, P. Fitzsimmons, K. Heuer, S. S. Kim, J. H. Lindsey II, O. P. Lossers, K. Poelladennou, K. Schiling, L. Scribani, P. Simeonov, GCHQ Problems Group, and NSA Problems Group. A square game: 10696. American Mathematical Monthly, 107 (2):184, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589458.

Anonymous:2000:BMc

[59] Anonymous. Back matter. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589460.

Anonymous:2000:FMc

[60] Anonymous. Front matter. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589432.

Beck:2000:NFE

[61] Anatole Beck. Notes: a fractal example in ordinary differential equations. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Enayat:2000:CFL

[62] Ali Enayat. As a continuous function of x and …. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Erdelyi:2000:NE

[63] Tamas Erdélyi. Notes: On the equation a(a + d)(a + 2d)(a + 3d) = x2. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Herzog:2000:NNS

[64] Gerd Herzog. Notes: On nonlinear summability. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Luca:2000:NAS

[65] Florian Luca. Notes: The anti-social Fermat number. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ostebee:2000:TRb

[66] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Su:2000:BRB

[67] Francis Edward Su. Book review: Cake-Cutting Algorithms: Be Fair If You Can, by Jack Robertson and William Webb. American Mathematical Monthly, 107(2):185–188, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589459.

Thomassen:2000:NRN

[68] Carsten Thomassen. Notes: The rendezvous number of a symmetric matrix and a compact connected metric space. American Mathematical Monthly, 107(2):??, February 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gleason:2000:YGG

[69] Andrew M. Gleason. Yueh-Gin Gung and Dr. Charles Y. Hu Award for Distinguished Service to Paul R. Halmos. American Mathematical Monthly, 107(3):193–194, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589312.

Freiling:2000:TSA

[70] Chris Freiling, Robert Hunter, Cynthia Turner, and Russell C. Wheeler. Tiling with squares and anti-squares. American Mathematical Monthly, 107(3):195–204, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589313.

Verde-Star:2000:SLD

[71] Luis Verde-Star. Solving linear differential-like equations. American Mathematical Monthly, 107(3):205–218, March 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2589314.

Helzer:2000:SRA

[72] Garry Helzer. Special Relativity with acceleration. American Mathematical Monthly, 107(3):219–237, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589315.

Poonen:2000:LPN

[73] Bjorn Poonen and Fernando Rodriguez-Villegas. Lattice polygons and the number 12. American Mathematical Monthly, 107(3):238–250, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589316.

Sibley:2000:RPT

[74] Tom Sibley and Stan Wagon. Rhombic Penrose tilings can be 3-colored. American Mathematical Monthly, 107(3):251–253, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589317.

Kang:2000:CFR

[75] Ming-Chang Kang. Cubic fields and radical extensions. American Mathematical Monthly, 107(3):254–256, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589318.

Badger:2000:GMB

[76] Lee Badger. Generating the measures of n-balls. American Mathematical Monthly, 107(3):256–258, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589319.

Hijab:2000:VUB

[77] Omar Hijab. The volume of the unit Ball in C^{n}. American Mathematical Monthly, 107(3):259, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589320.

Hirschhorn:2000:PFF

[78] Michael D. Hirschhorn. Partial fractions and four classical theorems of number theory. American Mathematical Monthly, 107(3):260–264, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589321.

Kim:2000:SNA

[79] Sung S. Kim and Kil H. Kwon. Smooth (C^{∞}) but nowhere analytic functions. American Mathematical Monthly, 107(3):264–266, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589322.

Laugwitz:2000:CPT

[80] Detlef Laugwitz. Comments on the paper “two letters by N. N. Luzin to M. Ya. Vygodskii”. American Mathematical Monthly, 107(3):267–276, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589323.

Taylor:2000:P

[81] Howard M. Taylor. Problem 10788. American Mathematical Monthly, 107(3):277, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589324.

Chapman:2000:P

[82] Robin Chapman. Problem 10789. American Mathematical Monthly, 107 (3):277, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589325.

Anglesio:2000:P

[83] Jean Anglesio. Problem 10790. American Mathematical Monthly, 107(3): 277, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589326.

Fekete:2000:Pa

[84] Antal Fekete. Problem 10791. American Mathematical Monthly, 107(3): 277, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589327.

Diamond:2000:P

[85] Harold G. Diamond and Ferrell S. Wheeler. Problem 10792. American Mathematical Monthly, 107(3):278, March 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2589328.

Luca:2000:P

[86] Florian Luca. Problem 10793. American Mathematical Monthly, 107(3): 278, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589329.

Hough:2000:P

[87] David S. Hough and Rodica E. Simion. Problem 10794. American Mathematical Monthly, 107(3):278, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589330.

Callan:2000:IID

[88] David Callan, Knut Dale, Ivar Skau, John Henry Steelman, GCHQ Problem Solving Group, J. C. Binz, R. J. Chapman, F. Herzig, A. Kundgen, R. Martin, J. H. Nieto, O. A. Saleh, S. Byrd, E. Schmeichel, I. Sofair, and M. Vowe. An identity involving derangements: 10643. American Mathematical Monthly, 107(3):278–279, March 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589331.

Deutsch:2000:HFD

[89] Emeric Deutsch, John H. Lindsey II, Michael H. Andreoli, R. A. Agnew, R. J. Chapman, M. P. Cohen, M. N. Deshpande, M. D. Ernst, E. Hertz, L. R. Ivory, W. Janous, G. Keselman, R. G. Mosier, J. H. Nieto, R. Pinkham, S. D. Sarti, P. Simeonov, W. R. Smythe, A. Stadler, J. H. Steelman, W. F. Trench, J. T. Ward, M. Woltermann, Anchorage Math Solutions Group, and GCHQ Problems Group. How far does the first element dominate?: 10649. American Mathematical Monthly, 107 (3):279–280, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589332.

Callan:2000:AFI

[90] David Callan, Robin Chapman, J. H. Lindsey II, J. Parkinson, V. Schindler, P. Simeonov, and GCHQ Problems Group. Analysis of a French inheritance law: 10652. American Mathematical Monthly, 107 (3):280–281, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589333.

Bencze:2000:RGM

[91] Mihaly Bencze, Dan Saracino, Allen Stenger, S. Amghibech, J. Anglesio, R. Bauer, A. Siegel, D. M. Bloom, G. L. Body, D. Callan, R. J. Chapman, K. Ford, S. M. Gagola, N. Gauthier, W. V. Grounds, T. Hagedorn, R. T. Koether, N. Komanda, R. N. Krishnan, K.-W. Lau, J. H. Lindsey II, L. E. Mattics, C. A. Minh, H. N. Ozsoylev, C. Y. Yildirim, P. G. Poonacha, N. R. Sanjeev, C. Popescu, J. Robertson, H. J. Sieffert, J. O. Shallit, N. C. Singer, A. Stadler, D. C. Terr, T. V. Triff, T. Trimble, P. Trojovsky, Anchorage Math Solutions Group, GCHQ Problems Group, and NSA Problems Group. A recurrence generating multiples of primes: 10655. American Mathematical Monthly, 107(3):281–282, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589334.

Owings:2000:SCI

[92] James C. Owings, Jr., Jeremy E. Dawson, Robin J. Chapman, Darryl K. Nester, R. A. Agnew, J. Beebee, D. Callan, R. Garcia-Pelayo, V. Hernandez, J. Martin, J. Hickman, D. Jones, S. S. Kim, O. Krafft, M. Schaefer, G. Letac, J. H. Lindsey II, O. P. Lossers, K. Poelladennou, K. Schilling, M. Shemesh, GCHQ Problems Group, and NSA Problems Group. A sum of Chebyshev inequalities: 10687. American Mathematical Monthly, 107 (3):282–283, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589335.

Ali:2000:ECE

[93] Hassan Ali Shah Ali, John H. Lindsey II, GCHQ Problems Group, and NSA Problems Group. Extrema for constrained elementary symmetric functions: 10694. American Mathematical Monthly, 107(3):283, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589336.

Diaz:2000:RI

[94] Jose L. Diaz, Tewodros Amdeberhan, and Erik I. Verriest. A rational identity: 10697. American Mathematical Monthly, 107(3):283–284, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589337.

Kagan:2000:BCS

[95] Abram Kagan, Colin Mallows, Larry Shepp, and John H. Lindsey II. Bernoulli component of a sum: 10699. American Mathematical Monthly, 107(3):284, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589338.

Anglesio:2000:FMD

[96] Jean Anglesio and Shalosh B. Ekhad. Four more distinguished points of a triangle: 10703. American Mathematical Monthly, 107(3):285, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589339.

Folland:2000:BRB

[97] Gerald B. Folland. Book review: [untitled]. American Mathematical Monthly, 107(3):286–291, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589340.

Ostebee:2000:TRc

[98] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(3):292–294, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589341.

Anonymous:2000:BMd

[99] Anonymous. Back matter. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589342.

Anonymous:2000:FMd

[100] Anonymous. Front matter. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589311.

Badger:2000:NGM

[101] Lee Badger. Notes: Generating the measures of n-balls. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boas:2000:GIM

[102] Harold P. Boas. Graphica 1. The Imaginary Made Real: The Images of Michael Trott; Graphica 2. The Pattern of Beauty: The Art of Igor Bakshee. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hijab:2000:NVU

[103] Omar Hijab. Notes: The volume of the unit ball in C^{n}. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hirschhorn:2000:NPF

[104] Michael D. Hirschhorn. Notes: Partial fractions and four classical theorems of number theory. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kang:2000:NCF

[105] Ming chang Kang. Notes: Cubic fields and radical extensions. American Mathematical Monthly, 107(3):??, March 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sibley:2000:NRP

[106] Tom Sibley and Stan Wagon. Notes: Rhombic Penrose tilings can be 3-colored. American Mathematical Monthly, 107(3):??, March 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sullivan:2000:KF

[107] Michael C. Sullivan. Knot factoring. American Mathematical Monthly, 107(4):297–315, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589175.

Strichartz:2000:EIU

[108] Robert S. Strichartz. Evaluating integrals using self-similarity. American Mathematical Monthly, 107(4):316–326, April 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589176.

Weiner:2000:HHC

[109] Joel L. Weiner. How helical can a closed, twisted space curve be? American Mathematical Monthly, 107(4):327–333, April 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589177.

Li:2000:NII

[110] Chi-Kwong Li. Norms, isometries, and isometry groups. American Mathematical Monthly, 107(4):334–340, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589178.

Aguilo:2000:PTD

[111] Francesc Aguilo, Miguel Angel Fiol, and Maria Lluisa Fiol. Periodic tilings as a dissection method. American Mathematical Monthly, 107 (4):341–352, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589179.

Bhatia:2000:BBV

[112] Rajendra Bhatia and Chandler Davis. A better bound on the variance. American Mathematical Monthly, 107(4):353–357, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589180.

Ridley:2000:EPL

[113] J. N. Ridley. An elementary proof on location of zeros. American Mathematical Monthly, 107(4):357–360, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589181.

Calkin:2000:RR

[114] Neil Calkin and Herbert S. Wilf. Recounting the rationals. American Mathematical Monthly, 107(4):360–363, April 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589182.

Wright:2000:NPS

[115] Stephen E. Wright. A note on positively spanning sets. American Mathematical Monthly, 107(4):364–366, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589183.

Deutsch:2000:Pb

[116] Emeric Deutsch. Problem 10795. American Mathematical Monthly, 107 (4):367, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589184.

vanLamoen:2000:Pa

[117] Floor van Lamoen. Problem 10796. American Mathematical Monthly, 107(4):367, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589185.

Bateman:2000:Pb

[118] Paul Bateman and Jeffrey Kalb. Problem 10797. American Mathematical Monthly, 107(4):367, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589186.

Neuman:2000:P

[119] Edward Neuman. Problem 10798. American Mathematical Monthly, 107 (4):367, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589187.

Herink:2000:P

[120] Curtis Herink and Gary Gruenhage. Problem 10799. American Mathematical Monthly, 107(4):367, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589188.

Iannucci:2000:Pb

[121] Douglas Iannucci. Problem 10800. American Mathematical Monthly, 107(4):368, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589189.

Pudaite:2000:P

[122] Paul R. Pudaite and Glen Ellyn. Problem 10801. American Mathematical Monthly, 107(4):368, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589190.

Stanley:2000:PES

[123] Richard P. Stanley, Richard Ehrenborg, Stephen Gagola, R. J. Chapman, and J. H. Lindsey II. Pairs with equal squares: 10654. American Mathematical Monthly, 107(4):368, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589191.

Deutsch:2000:ATL

[124] Emeric Deutsch, David Callan, M. Beck, D. Beckwith, W. Bohm, R. F. McCoart, and GCHQ Problems Group. Another type of lattice path: 10658. American Mathematical Monthly, 107(4):368–370, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589192.

Just:2000:SCC

[125] Erwin Just, A. Berele, J. Bergen, K. A. Kearnes, G. Marks, and G. P. Wene. A sufficient condition for commutativity: 10661. American Mathematical Monthly, 107(4):370, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589193.

Deutsch:2000:FPS

[126] Emeric Deutsch, Harris Kwong, Cecil C. Rousseau, and Paul K. Stockmeyer. A formula for the Pell sequence: 10663. American Mathematical Monthly, 107(4):370–371, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589194.

Sasvari:2000:AUM

[127] Zoltan Sasvari, Jerrold W. Grossman, and Sung Soo Kim. Avoiding uncountably many subsets: 10667. American Mathematical Monthly, 107 (4):371–372, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589195.

Isbell:2000:FPI

[128] John Isbell, Stephen Schanuel, Stephen M. Gagola, Jr., and Mowaffaq Hajja. Fixed points of iterated cyclic difference operators: 10676. American Mathematical Monthly, 107(4):372–373, April 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589196.

Kimberling:2000:ASC

[129] Clark Kimberling, Peter Yff, and Vasile Mihai. Another set of concurrent cevians: 10678. American Mathematical Monthly, 107(4):373, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589197.

Propp:2000:AST

[130] James G. Propp and Keith A. Kearnes. An application of Sperner’s theorem: 10679. American Mathematical Monthly, 107(4):374, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589198.

Ford:2000:ARS

[131] Kevin Ford and Kee-Wai Lau. Asymptotics of a recurrent sequence: 10682. American Mathematical Monthly, 107(4):374–375, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589199.

Brown:2000:CHS

[132] D. W. Brown and John Cobb. A countable Hausdorff space with the fixed point property: 10705. American Mathematical Monthly, 107(4): 375–376, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589200.

Sasvari:2000:END

[133] Zoltan Sasvari and John H. Lindsey II. An estimate for the normal distribution: 10709. American Mathematical Monthly, 107(4):376–377, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589201.

Bellamy:2000:CSS

[134] David P. Bellamy, Felix Lazebnik, and Ray Redheffer. Characterizing solutions to simple differential equations: 10729. American Mathematical Monthly, 107(4):377, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589202.

Fleron:2000:BRB

[135] Julian F. Fleron. Book review: [untitled]. American Mathematical Monthly, 107(4):378–384, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589203.

Cusick:2000:BRB

[136] Thomas W. Cusick. Book review: [untitled]. American Mathematical Monthly, 107(4):384–386, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589204.

Horn:2000:EEb

[137] Roger A. Horn. Editor’s endnotes. American Mathematical Monthly, 107(4):387, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589205.

Anonymous:2000:BMe

[138] Anonymous. Back matter. American Mathematical Monthly, 107(4): ??, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589206.

Anonymous:2000:FMe

[139] Anonymous. Front matter. American Mathematical Monthly, 107(4): ??, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589174.

Browder:2000:TCP

[140] Andrew Browder. Topology in the complex plane. American Mathematical Monthly, 107(5):393–401, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695294.

Kendig:2000:YOF

[141] Keith Kendig. Is a 2000-year-old formula still keeping some secrets? American Mathematical Monthly, 107(5):402–415, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695295.

Keller:2000:ORS

[142] Joseph B. Keller. Optimal running strategy to escape from pursuers. American Mathematical Monthly, 107(5):416–421, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695296.

Frame:2000:FPF

[143] Michael Frame, Brenda Johnson, and Jim Sauerberg. Fixed points and fermat: a dynamical systems approach to number theory. American Mathematical Monthly, 107(5):422–428, May 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695297.

Bartholdi:2000:LFF

[144] Laurent Bartholdi. Lamps, factorizations, and finite fields. American Mathematical Monthly, 107(5):429–436, May 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695298.

Boute:2000:ZPN

[145] Raymond T. Boute. Zeroless positional number representation and string ordering. American Mathematical Monthly, 107(5):437–444, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695299.

Broughan:2000:IPP

[146] Kevin Broughan and Nan Zhu. An integer programming problem with a linear programming solution. American Mathematical Monthly, 107 (5):444–446, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695300.

Jackson:2000:SPE

[147] Terence Jackson. A short proof that every prime p ≡ 3 (mod 8) is of the form x^{2} + 2y^{2}. American Mathematical Monthly, 107(5):447, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695301.

Allouche:2000:KLC

[148] Jean-Paul Allouche and Michel Cosnard. The Komornik–Loreti constant is transcendental. American Mathematical Monthly, 107(5):448– 449, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695302.

Wen:2000:NDC

[149] Liu Wen. A nowhere differentiable continuous function. American Mathematical Monthly, 107(5):450–453, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695303.

Tits:2000:S

[150] Jacques Tits and John Stillwell. Symmetry. American Mathematical Monthly, 107(5):454–461, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695304.

Caragea:2000:P

[151] Doru Caragea and Viviana Ene. Problem 10802. American Mathematical Monthly, 107(5):462, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695305.

Penrice:2000:P

[152] Stephen Penrice. Problem 10803. American Mathematical Monthly, 107 (5):462, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695306.

Sinefakopoulos:2000:P

[153] Achilleas Sinefakopoulos. Problem 10804. American Mathematical Monthly, 107(5):462, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695307.

Fekete:2000:Pb

[154] Antal Fekete. Problem 10805. American Mathematical Monthly, 107(5): 462, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695308.

Ali:2000:P

[155] Hassan Ali Shah Ali. Problem 10806. American Mathematical Monthly, 107(5):462, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695309.

Deleglise:2000:P

[156] Marc Deléglise. Problem 10807. American Mathematical Monthly, 107 (5):463, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695310.

Valdinoci:2000:P

[157] Enrico Valdinoci. Problem 10808. American Mathematical Monthly, 107 (5):463, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695311.

Deiermann:2000:ICb

[158] Paul Deiermann, Rick Mabry, and Alain Tissier. Intersecting curves: 10712. American Mathematical Monthly, 107(5):463–464, May 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695312.

Marquez:2000:CCA

[159] Juan-Bosco Romero Marquez and Heinz-Jurgen Seiffert. A cute characterization of acute triangles: 10713. American Mathematical Monthly, 107(5):464, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695313.

Cuculiere:2000:SB

[160] Roger Cuculière and Albert Stadler. Successful bootstrapping: 10715. American Mathematical Monthly, 107(5):465–466, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695314.

Mazur:2000:RVR

[161] Marcin Mazur, Robin Chapman, Petr Lisoněk, Robert B. Israel, and Murray S. Klamkin. Rigid and very rigid tetrahedra: 10717. American Mathematical Monthly, 107(5):466–467, May 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695315.

Sidney:2000:IS

[162] Daniel A. Sidney and National Security Agency Problems Group. Integrals and sines: 10721. American Mathematical Monthly, 107(5):468– 469, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695316.

Knuth:2000:PBS

[163] Donald Knuth and Darin Stephenson. The probability of being in a state: 10726. American Mathematical Monthly, 107(5):469–470, May 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695317.

Janous:2000:BCI

[164] Walther Janous and Enrico Valdinoci. The best constant in an inequality: 10730. American Mathematical Monthly, 107(5):470, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695318.

Hibbard:2000:BRB

[165] Allen C. Hibbard. Book review: [untitled]. American Mathematical Monthly, 107(5):471–474, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695319.

Jackson:2000:BRB

[166] Mic Jackson. Book review: [untitled]. American Mathematical Monthly, 107(5):474–478, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695320.

Ostebee:2000:TRd

[167] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(5):479–481, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695321.

Anonymous:2000:BMf

[168] Anonymous. Back matter. American Mathematical Monthly, 107(5): ??, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695322.

Anonymous:2000:FMf

[169] Anonymous. Front matter. American Mathematical Monthly, 107(5): ??, May 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695293.

Scheinerman:2000:WCE

[170] Edward R. Scheinerman. When close enough is close enough. American Mathematical Monthly, 107(6):489–499, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589344.

McLean:2000:LRP

[171] K. Robin McLean. Loops of regular polygons. American Mathematical Monthly, 107(6):500–510, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589345.

Benjamin:2000:RAF

[172] Arthur T. Benjamin, Gregory M. Levin, Karl Mahlburg, and Jennifer J. Quinn. Random approaches to Fibonacci identities. American Mathematical Monthly, 107(6):511–516, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589346.

Borwein:2000:SM

[173] D. Borwein, J. M. Borwein, and P. Marechal. Surprise maximization. American Mathematical Monthly, 107(6):517–527, June/July 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589347.

Wu:2000:PNR

[174] Pei Yuan Wu. Polygons and numerical ranges. American Mathematical Monthly, 107(6):528–540, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589348.

Pudlak:2000:PG

[175] Pavel Pudlak. Proofs as games. American Mathematical Monthly, 107(6): 541–550, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589349.

Kuperberg:2000:HCS

[176] Wlodzimierz Kuperberg. Holey coronas A solution of the Grunbaum– Shephard Conjecture on convex isohedral tilings. American Mathematical Monthly, 107(6):551–555, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589350.

Romik:2000:SAU

[177] Dan Romik. Stirling’s approximation for n!: The ultimate short proof? American Mathematical Monthly, 107(6):556–557, June/July 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589351.

Ehrenborg:2000:HDE

[178] Richard Ehrenborg. The Hankel determinant of exponential polynomials. American Mathematical Monthly, 107(6):557–560, June/July 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589352.

Benjamin:2000:WB

[179] Arthur T. Benjamin and Matthew T. Fluet. What’s best? American Mathematical Monthly, 107(6):560–562, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589353.

Engel:2000:CMP

[180] Doug Engel. A Celtic magnecurve problem. American Mathematical Monthly, 107(6):563–565, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589354.

Beckwith:2000:Pa

[181] David Beckwith. Problem 10809. American Mathematical Monthly, 107 (6):566, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589355.

Marquez:2000:P

[182] Juan-Bosco Romero Marquez. Problem 10810. American Mathematical Monthly, 107(6):566, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589356.

Tracy:2000:P

[183] Phil Tracy. Problem 10811. American Mathematical Monthly, 107(6):566, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589357.

Pinchover:2000:P

[184] Yehuda Pinchover and Simeon Reich. Problem 10812. American Mathematical Monthly, 107(6):566, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589358.

Richman:2000:P

[185] Fred Richman. Problem 10813. American Mathematical Monthly, 107 (6):567, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589359.

Satnoianu:2000:P

[186] Razvan Satnoianu. Problem 10814. American Mathematical Monthly, 107 (6):567, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589360.

Bertram:2000:P

[187] Barbara S. Bertram and Otto G. Ruehr. Problem 10815. American Mathematical Monthly, 107(6):567, June/July 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589361.

Deutsch:2000:PPS

[188] Emeric Deutsch, Ira M. Gessel, and David Callan. Permutation parameters with the same distribution: 10634. American Mathematical Monthly, 107(6):567–568, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589362.

Beckwith:2000:SIN

[189] David Beckwith, Tiberiu V. Trif, and Dan C. Popescu. A series involving numbers of factorizations: 10669. American Mathematical Monthly, 107(6):568–569, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589363.

Newman:2000:RIG

[190] David Newman. A recurrence inside a generating function: 10681. American Mathematical Monthly, 107(6):569–570, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2589364.

WMCPG:2000:WMI

[191] Western Maryland College Problems Group. A Western Maryland integral: 10708. American Mathematical Monthly, 107(6):570–572, June/ July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589365.

Suceava:2000:SPC

[192] Bogdan Suceava and Achilleas Sinefakopoulos. A six point circle: 10710. American Mathematical Monthly, 107(6):572–573, June/July 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589366.

Kim:2000:PIC

[193] Sung Soo Kim and K. P. Hart. Partitioning the interval into closed sets: 10733. American Mathematical Monthly, 107(6):573, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2589367.

Aldous:2000:BRB

[194] David Aldous. Book review: [untitled]. American Mathematical Monthly, 107(6):574–576, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589368.

Frame:2000:BRB

[195] Michael Frame. Book review: [untitled]. American Mathematical Monthly, 107(6):576–580, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589369.

Anonymous:2000:BMg

[196] Anonymous. Back matter. American Mathematical Monthly, 107(6):??, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589370.

Anonymous:2000:FMg

[197] Anonymous. Front matter. American Mathematical Monthly, 107(6):??, June/July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589343.

ReVelle:2000:DIR

[198] Charles S. ReVelle and Kenneth E. Rosing. Defendens imperium romanum: a classical problem in military strategy. American Mathematical Monthly, 107(7):585–594, August/September 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589113.

Berndt:2000:BSR

[199] Bruce C. Berndt and Robert A. Rankin. The books studied by Ramanujan in india. American Mathematical Monthly, 107(7):595–601, August/ September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589114.

Bhatia:2000:PTT

[200] Rajendra Bhatia. Pinching, trimming, truncating, and averaging of matrices. American Mathematical Monthly, 107(7):602–608, August/ September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589115.

Beer:2000:WFS

[201] Gerald Beer. When is a family of sets a family of bounded sets? American Mathematical Monthly, 107(7):609–614, August/September 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589116.

Moore:2000:SMT

[202] David S. Moore and George W. Cobb. Statistics and mathematics: Tension and cooperation. American Mathematical Monthly, 107(7):615–630, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589117.

Pakianathan:2000:NN

[203] Jonathan Pakianathan and Krishnan Shankar. Nilpotent numbers. American Mathematical Monthly, 107(7):631–634, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589118.

Jenkyns:2000:TTC

[204] Tom Jenkyns and Eric Muller. Triangular triples from ceilings to floors. American Mathematical Monthly, 107(7):634–639, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589119.

Nillsen:2000:NNM

[205] Rodney Nillsen. Normal numbers without measure theory. American Mathematical Monthly, 107(7):639–644, August/September 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589120.

Postnikov:2000:PSL

[206] M. M. Postnikov and Abe Shenitzer. The problem of squarable lunes. American Mathematical Monthly, 107(7):645–651, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589121.

Deutsch:2000:Pc

[207] Emeric Deutsch. Problem 10816. American Mathematical Monthly, 107 (7):652, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589122.

Janous:2000:Pa

[208] Walther Janous. Problem 10817. American Mathematical Monthly, 107 (7):652, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589123.

Joita:2000:P

[209] Cezar Joita. Problem 10818. American Mathematical Monthly, 107(7): 652, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589124.

Krafft:2000:P

[210] Olaf Krafft. Problem 10819. American Mathematical Monthly, 107(7): 652, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589125.

Mirzavaziri:2000:P

[211] M. Mirzavaziri. Problem 10820. American Mathematical Monthly, 107 (7):652, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589126.

Foschini:2000:P

[212] Gerard J. Foschini. Problem 10821. American Mathematical Monthly, 107(7):653, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589127.

Lagarias:2000:P

[213] Jeffrey Lagarias. Problem 10822. American Mathematical Monthly, 107 (7):653, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2589128.

Griggs:2000:ABP

[214] Jerrold R. Griggs, Darryl K. Nester, and Sung Soo Kim. The asymptotics of the birthday problem: 10665. American Mathematical Monthly, 107(7):653–654, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589129.

Cavachi:2000:PP

[215] Marius Cavachi. A powerful property: 10674. American Mathematical Monthly, 107(7):654, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589130.

Derksen:2000:CMM

[216] Harm Derksen and Jeremy Dawson. A characterization of min or max: 10677. American Mathematical Monthly, 107(7):654–655, August/ September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2589131.

Pranesachar:2000:EPT

[217] C. R. Pranesachar. Equicevian points of a triangle: 10686. American Mathematical Monthly, 107(7):656–657, August/September 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589132.

Chao:2000:WQI

[218] Wu Wei Chao and Plamen Simeonov. When quadrilaterals have inscribed circles: 10698. American Mathematical Monthly, 107(7):657–658, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589133.

vanLamoen:2000:BLP

[219] Floor van Lamoen and J. C. Binz. Beyond the de longchamps point: 10734. American Mathematical Monthly, 107(7):658–659, August/ September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2589134.

Theodorescu:2000:MPD

[220] Radu Theodorescu and Jonathan M. Borwein. Moments of the Poisson distribution: 10738. American Mathematical Monthly, 107(7):659, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589135.

Tersian:2000:DI

[221] Stepan Tersian and Richard Bagby. A definite integral: 10746. American Mathematical Monthly, 107(7):659–660, August/September 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589136.

Mulcahy:2000:BRB

[222] Colm Mulcahy. Book review: [untitled]. American Mathematical Monthly, 107(7):661–663, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589137.

Grossman:2000:BRB

[223] Jerrold W. Grossman. Book review: [untitled]. American Mathematical Monthly, 107(7):664–668, August/September 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589138.

Ostebee:2000:TRe

[224] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(7):669–671, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589139.

Anonymous:2000:LRF

[225] Anonymous. Lester R. Ford Awards for 1999. American Mathematical Monthly, 107(7):672, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589140.

Anonymous:2000:BMh

[226] Anonymous. Back matter. American Mathematical Monthly, 107(7):??, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589141.

Anonymous:2000:FMh

[227] Anonymous. Front matter. American Mathematical Monthly, 107(7):??, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2589112.

Lee:2000:RRM

[228] Steven L. Lee and Gilbert Strang. Row reduction of a matrix and A = CaB. American Mathematical Monthly, 107(8):681–688, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695466.

Broughton:2000:CKT

[229] S. Allen Broughton. Constructing kaleidoscopic tiling polygons in the hyperbolic plane. American Mathematical Monthly, 107(8):689–710, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695467.

Biss:2000:GAF

[230] Daniel K. Biss. A generalized approach to the fundamental group. American Mathematical Monthly, 107(8):711–720, October 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695468.

Klosinski:2000:SWL

[231] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixtieth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 107(8):721–732, October 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695469.

Gallian:2000:PT

[232] Joseph A. Gallian. Putnam trivia for the 90s. American Mathematical Monthly, 107(8):733–735, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695470.

Pikhurko:2000:EFC

[233] Oleg Pikhurko. On envy-free cake division. American Mathematical Monthly, 107(8):736–738, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695471.

Apostol:2000:CHD

[234] Tom M. Apostol. Calculating higher derivatives of inverses. American Mathematical Monthly, 107(8):738–741, October 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695472.

Carley:2000:MPZ

[235] Holly Carley and Xin Li. The modulus of polynomials with zeros at the roots of unity. American Mathematical Monthly, 107(8):742–748, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695473.

Rakocevic:2000:NIO

[236] Vladimir Rakočević. On the norm of idempotent operators in a Hilbert space. American Mathematical Monthly, 107(8):748–750, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695474.

Zhang:2000:SPT

[237] Guang Yuan Zhang. A simple proof of a theorem of block and hart. American Mathematical Monthly, 107(8):751, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695475.

Andrews:2000:Pa

[238] George E. Andrews. Problem 10823. American Mathematical Monthly, 107(8):752, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695476.

Lee:2000:Pa

[239] Ho joo Lee. Problem 10824. American Mathematical Monthly, 107(8): 752, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695477.

Miller:2000:P

[240] Carl Miller. Problem 10825. American Mathematical Monthly, 107(8): 752, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695478.

Martinez-Gimenez:2000:P

[241] Félix Martínez-Giménez. Problem 10826. American Mathematical Monthly, 107(8):752, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695479.

Abel:2000:P

[242] Ulrich Abel. Problem 10827. American Mathematical Monthly, 107(8): 752, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695480.

Beckwith:2000:Pb

[243] David Beckwith. Problem 10828. American Mathematical Monthly, 107 (8):753, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695481.

Janous:2000:Pb

[244] Walther Janous. Problem 10829. American Mathematical Monthly, 107 (8):753, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695482.

Callan:2000:PPF

[245] David Callan, David Beckwith, and Anchorage Math Solutions Group. Periodic points and forests: 10666. American Mathematical Monthly, 107 (8):753–754, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695483.

Tamvakis:2000:STP

[246] Harry Tamvakis and Joseph J. Rushanan. Splitting triple products by sign: 10683. American Mathematical Monthly, 107(8):754, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695484.

Knuth:2000:RNA

[247] Donald E. Knuth, Vaughan Pratt, and O. P. Lossers. The real numbers, algebraically: 10689. American Mathematical Monthly, 107(8):755, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695486.

Lenstra:2000:ROE

[248] H. W. Lenstra and National Security Agency Problems Group. Restricting the order of elements in a group: 10684. American Mathematical Monthly, 107(8):755, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695485.

Chappell:2000:OIU

[249] Glenn G. Chappell and Robin Chapman. Ordering the integers using specified differences: 10692. American Mathematical Monthly, 107(8): 756, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695487.

Bluskov:2000:USL

[250] Iliya Bluskov and Albert Nijenhuis. A useful and surprising limit: 10695. American Mathematical Monthly, 107(8):756–757, October 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695488.

Ciaurri:2000:VMV

[251] Oscar Ciaurri and G. E. Bilodeau. Variation on the mean value theme: 10739. American Mathematical Monthly, 107(8):757, October 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695489.

Kalakos:2000:DEM

[252] Athanasios Kalakos, Michel Bataille, and David M. Wells. A differential equation with many solutions: 10747. American Mathematical Monthly, 107(8):757–758, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695490.

Costovici:2000:CEC

[253] Gh. Costovici and Kenneth Schilling. Conditions ensuring convergence: 10752. American Mathematical Monthly, 107(8):758–759, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695491.

Iannucci:2000:CF

[254] Douglas Iannucci and Brian D. Beasley. A cosine formula: 10756. American Mathematical Monthly, 107(8):759, October 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695492.

Burgiel:2000:BRB

[255] H. Burgiel. Book review: [untitled]. American Mathematical Monthly, 107(8):760–763, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695493.

Pomeranz:2000:BRB

[256] Shirley B. Pomeranz. Book review: [untitled]. American Mathematical Monthly, 107(8):763–766, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695494.

Anonymous:2000:BMi

[257] Anonymous. Back matter. American Mathematical Monthly, 107(8):??, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695495.

Anonymous:2000:FMi

[258] Anonymous. Front matter. American Mathematical Monthly, 107(8):??, October 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695465.

Mattingly:2000:EOR

[259] R. Bruce Mattingly. Even order regular magic squares are singular. American Mathematical Monthly, 107(9):777–782, November 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695733.

Bhargava:2000:FFG

[260] Manjul Bhargava. The factorial function and generalizations. American Mathematical Monthly, 107(9):783–799, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695734.

Jeffers:2000:LTG

[261] Jason Jeffers. Lost theorems of geometry. American Mathematical Monthly, 107(9):800–812, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695735.

Rudin:2000:SSP

[262] Walter Rudin. Sums of squares of polynomials. American Mathematical Monthly, 107(9):813–821, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695736.

Burger:2000:DOW

[263] Edward B. Burger. Diophantine Olympics and world champions: Polynomials and primes down under. American Mathematical Monthly, 107(9):

822–829, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695737.

Tabachnikov:2000:FVT

[264] Serge Tabachnikov. A four vertex theorem for polygons. American Mathematical Monthly, 107(9):830–833, November 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695738.

Harman:2000:NMT

[265] Glyn Harman and Kam C. Wong. A note on the metrical theory of continued fractions. American Mathematical Monthly, 107(9):834–837, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695739.

Kravanja:2000:MAD

[266] Peter Kravanja and Ann Haegemans. Multidimensional analytic deflation. American Mathematical Monthly, 107(9):837–841, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695740.

Apostol:2000:ISR

[267] Tom M. Apostol. Irrationality of the square root of two – A geometric proof. American Mathematical Monthly, 107(9):841–842, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695741.

Sen:2000:FTA

[268] Anindya Sen. Fundamental theorem of algebra – yet another proof. American Mathematical Monthly, 107(9):842–843, November 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695742.

Felscher:2000:BCE

[269] Walter Felscher. Bolzano, cauchy, epsilon, delta. American Mathematical Monthly, 107(9):844–862, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695743.

vanLamoen:2000:Pb

[270] Floor van Lamoen. Problem 10830. American Mathematical Monthly, 107 (9):863, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695744.

Andrews:2000:Pb

[271] George E. Andrews, P. Paule, and A. Riese. Problem 10831. American Mathematical Monthly, 107(9):863, November 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695745.

Knuth:2000:P

[272] Donald E. Knuth. Problem 10832. American Mathematical Monthly, 107 (9):863, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695746.

Eynden:2000:P

[273] Charles Vanden Eynden. Problem 10833. American Mathematical Monthly, 107(9):863, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695747.

Kim:2000:P

[274] Sung Soo Kim. Problem 10834. American Mathematical Monthly, 107 (9):864, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695748.

Dyubina:2000:P

[275] Anna Dyubina and Pierre de la Harpe. Problem 10835. American Mathematical Monthly, 107(9):864, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695749.

Wellner:2000:P

[276] Jon A. Wellner. Problem 10836. American Mathematical Monthly, 107 (9):864, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695750.

Quet:2000:BSS

[277] Leroy Quet and Robin Chapman. Bernoulli, stirling, and stirling: 10700. American Mathematical Monthly, 107(9):864–865, November 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695751.

Spohn:2000:NPT

[278] William G. Spohn, Jr. and Jeremy E. Dawson. Neighborly Pythagorean triples: 10704. American Mathematical Monthly, 107(9):865–866, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695752.

Propp:2000:DPS

[279] James G. Propp and Tiberiu V. Trif. Deriving a perfect square: 10706. American Mathematical Monthly, 107(9):866–867, November 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695753.

Popescu:2000:IIA

[280] Călin Popescu and Achilleas Sinefakopoulos. Inequalities involving altitudes and exradii: 10759. American Mathematical Monthly, 107(9): 867–868, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695754.

Quet:2000:CR

[281] Leroy Quet and O. P. Lossers. A convoluted recursion: 10762. American Mathematical Monthly, 107(9):868, November 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695755.

Redheffer:2000:ASU

[282] Ray Redheffer and Robin Chapman. All solutions in the unit cube: 10764. American Mathematical Monthly, 107(9):868–869, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695756.

Andras:2000:SIX

[283] Szilárd András, Ronald A. Kopas, and Heinz-Jürgen Seiffert. Symmetric inequalities for x^{x} and x^{x2} : 10766. American Mathematical Monthly, 107(9):869–870, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695757.

Kennedy:2000:BRB

[284] Steve Kennedy. Book review: [untitled]. American Mathematical Monthly, 107(9):871–874, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695758.

Cohen:2000:BRBb

[285] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 107(9):874–876, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695759.

Ostebee:2000:TRf

[286] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 107(9):877–878, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695760.

Anonymous:2000:BMj

[287] Anonymous. Back matter. American Mathematical Monthly, 107(9):??, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695761.

Anonymous:2000:FMj

[288] Anonymous. Front matter. American Mathematical Monthly, 107(9):??, November 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695732.

Eschenbach:2000:PTA

[289] Carolyn Eschenbach, Frank Hall, Rohan Hemasinha, Stephen Kirkland, Zhongshan Li, Bryan Shader, Jeffrey Stuart, and James Weaver. Properties of tournaments among well-matched players. American Mathematical Monthly, 107(10):881–892, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695581.

Cifre:2000:TPC

[290] M. A. Hernández Cifre. Is there a planar convex set with given width, diameter, and inradius? American Mathematical Monthly, 107(10):893– 900, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695582.

Konvalina:2000:UIB

[291] John Konvalina. A unified interpretation of the binomial coefficients, the Stirling numbers, and the Gaussian coefficients. American Mathematical Monthly, 107(10):901–910, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695583.

Treatman:2000:CAR

[292] Stefan Treatman and Cameron Wickham. Constructible approximations of regular polygons. American Mathematical Monthly, 107(10):911–922, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695584.

Bradley:2000:UIT

[293] David M. Bradley. Using integral transforms to estimate higher order derivatives. American Mathematical Monthly, 107(10):923–931, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695585.

Barton:2000:SSC

[294] Reid Barton and Keith Burns. A simple special case of Sharkovskii’s theorem. American Mathematical Monthly, 107(10):932–933, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695586.

Watkins:2000:QR

[295] William Watkins. The quadratic residues -1 and -3. American Mathematical Monthly, 107(10):934, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695587.

Herzog:2000:PSD

[296] Gerd Herzog. Polynomials solving Dirichlet boundary value problems. American Mathematical Monthly, 107(10):934–936, December 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695588.

Dilcher:2000:EFL

[297] Karl Dilcher. An extension of Fermat’s little theorem, and congruences for Stirling numbers. American Mathematical Monthly, 107(10):936–940, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695589.

vanLamoen:2000:MRT

[298] Floor van Lamoen. Morley related triangles on the nine-point circle. American Mathematical Monthly, 107(10):941–945, December 2000. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695590.

Benyi:2000:APP

[299] Árpád Bényi, Pawel Szeptycki, and Fred Van Vleck. Archimedean properties and parabolas. American Mathematical Monthly, 107(10):945–949, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695591.

Lee:2000:Pb

[300] Ho joo Lee. Problem 10837. American Mathematical Monthly, 107(10): 950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695592.

Parvanescu:2000:P

[301] Florian S. Pârvânescu. Problem 10838. American Mathematical Monthly, 107(10):950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695593.

Parlett:2000:P

[302] Beresford N. Parlett. Problem 10839. American Mathematical Monthly, 107(10):950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695594.

Yang:2000:P

[303] Jiansheng Yang and Shulin Zhou. Problem 10840. American Mathematical Monthly, 107(10):950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695595.

Just:2000:P

[304] Erwin Just. Problem 10841. American Mathematical Monthly, 107(10): 950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695596.

Reznick:2000:P

[305] Bruce Reznick. Problem 10842. American Mathematical Monthly, 107 (10):950, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695597.

Boklan:2000:MLP

[306] Kent D. Boklan and John P. Robertson. The maximum length of a powerful arithmetic progression: 10702. American Mathematical Monthly, 107 (10):951, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695599.

Vince:2000:P

[307] Andrew Vince. Problem 10843. American Mathematical Monthly, 107 (10):951, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695598.

Isbell:2000:UPS

[308] John Isbell, Julio Kuplinsky, and Leon Mattics. A union of proper subspaces?: 10707. American Mathematical Monthly, 107(10):951–952, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695600.

Anglesio:2000:TTC

[309] Jean Anglesio and Volkhard Schindler. Two triangles with common vertex, incenter, and centroid: 10719. American Mathematical Monthly, 107(10):952–954, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695601.

Hajja:2000:CSI

[310] Mowaffaq Hajja, Peter Walker, Danny Arrigo, Xiaoshen Wang, and Zafar Ahmed. A cubic spherical integral: 10771. American Mathematical Monthly, 107(10):954–955, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695602.

Waterhouse:2000:NFD

[311] William C. Waterhouse and Steve Butcher. Nonconstant functions with derivative zero: 10772. American Mathematical Monthly, 107(10):956, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695603.

Ahmed:2000:FI

[312] Zafar Ahmed and Peter M. Jarvis. A family of integrals: 10777. American Mathematical Monthly, 107(10):956–957, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695604.

Kedlaya:2000:IP

[313] Kiran Kedlaya and Marcin Mazur. Incircle partnerships: 10780. American Mathematical Monthly, 107(10):957–958, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695605.

Deiermann:2000:ICa

[314] Paul Deiermann and Rick Mabry. Intersecting curves: 10712. American Mathematical Monthly, 107(10):958, December 2000. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695606.

Cohen:2000:BRBa

[315] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 107(10):959–963, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695607.

Logan:2000:BRB

[316] J. David Logan. Book review: [untitled]. American Mathematical Monthly, 107(10):964–967, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695608.

Horn:2000:EEa

[317] Roger A. Horn. Editor’s endnotes. American Mathematical Monthly, 107(10):968–969, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695609.

Anonymous:2000:VI

[318] Anonymous. Volume information. American Mathematical Monthly, 107 (10):970–980, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695579.

Anonymous:2000:BMb

[319] Anonymous. Back matter. American Mathematical Monthly, 107(10):??, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695610.

Anonymous:2000:FMb

[320] Anonymous. Front matter. American Mathematical Monthly, 107(10):??, December 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695580.

Kirwan:2001:MDC

[321] W. E. Kirwan. Mathematics departments in the 21st century: Role, relevance, and responsibility. American Mathematical Monthly, 108(1):

1–9, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695670.

Garber:2001:MDA

[322] David Garber and Boaz Tsaban. A mechanical derivation of the area of the sphere. American Mathematical Monthly, 108(1):10–15, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695671.

Aldous:2001:RLW

[323] David J. Aldous. Reorganizing large Web sites. American Mathematical Monthly, 108(1):16–27, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695672.

Alvis:2001:BTP

[324] Dean Alvis and Michael Kinyon. Birkhoff’s theorem for panstochastic matrices. American Mathematical Monthly, 108(1):28–37, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695673.

Goodman-Strauss:2001:CSP

[325] Chaim Goodman-Strauss. Compass and straightedge in the Poincaré disk. American Mathematical Monthly, 108(1):38–49, January 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695674.

Schechter:2001:CD

[326] Eric Schechter. Constructivism is difficult. American Mathematical Monthly, 108(1):50–54, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695675.

Gessel:2001:SCD

[327] Ira M. Gessel. The Smith College Diploma Problem. American Mathematical Monthly, 108(1):55–57, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695676.

Waterhouse:2001:CTR

[328] William C. Waterhouse. Cochran’s theorem for rings. American Mathematical Monthly, 108(1):58–59, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695677.

Simoson:2001:TT

[329] Andrew Simoson. The tumbling tetrahedron. American Mathematical Monthly, 108(1):59–63, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695678.

Eppstein:2001:TST

[330] David Eppstein. Tangent spheres and triangle centers. American Mathematical Monthly, 108(1):63–66, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695679.

Noy:2001:ATC

[331] Marc Noy. Acyclic and totally cyclic orientations in planar graphs. American Mathematical Monthly, 108(1):66–68, January 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695680.

Furuta:2001:HMY

[332] Takayuki Furuta. The Hölder–McCarthy and the Young inequalities are equivalent for Hilbert space operators. American Mathematical Monthly, 108(1):68–69, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695681.

Stillwell:2001:MM

[333] John Stillwell. Modular miracles. American Mathematical Monthly, 108 (1):70–76, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695682.

MacKinnon:2001:Pa

[334] Nick MacKinnon. Problem 10844. American Mathematical Monthly, 108(1):77, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695683.

Demeter:2001:P

[335] Ciprian Demeter. Problem 10845. American Mathematical Monthly, 108(1):77, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695684.

Przytycki:2001:P

[336] Józef Przytycki. Problem 10846. American Mathematical Monthly, 108 (1):77, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695685.

Chapman:2001:P

[337] Robin Chapman, Kimmo Eriksson, and Richard Stanley. Problem 10847. American Mathematical Monthly, 108(1):77, January 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695686.

Stolarsky:2001:P

[338] Kenneth B. Stolarsky. Problem 10848. American Mathematical Monthly, 108(1):77, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695687.

Bagby:2001:Pa

[339] Richard Bagby. Problem 10849. American Mathematical Monthly, 108 (1):78, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695688.

Lang:2001:P

[340] Wolfdieter Lang. Problem 10850. American Mathematical Monthly, 108 (1):78, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695689.

Bradley:2001:SIT

[341] David M. Bradley, Steven Levandosky, and A. Meir. Series and integral trial conditions: 10642. American Mathematical Monthly, 108(1):78–79, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695690.

Galvin:2001:GHTa

[342] Fred Galvin, Stephen C. Locke, Sung Soo Kim, and David Callan. A generalization of Hall’s theorem: 10701. American Mathematical Monthly, 108(1):79–80, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695691.

Luca:2001:CTC

[343] Florian Luca and Francis B. Coghlan. Can two consecutive numbers both be perfect?: 10711. American Mathematical Monthly, 108(1):80–81, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695692.

Wimp:2001:DPE

[344] Jet Wimp and Thomas Hermann. A determinant with polynomial entries: 10714. American Mathematical Monthly, 108(1):81, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695693.

Catalano-Johnson:2001:CG

[345] Michael L. Catalano-Johnson, Daniel Loeb, and John Beebee. A cubical gift: 10716. American Mathematical Monthly, 108(1):81–82, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695694.

Kim:2001:WSF

[346] Sung Soo Kim and Mark D. Meyerson. When a sum of functions is increasing: 10768. American Mathematical Monthly, 108(1):82, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695695.

Chao:2001:CLC

[347] Wu Wei Chao and Daniele Donini. Concurrent lines in a cyclic quadrilateral: 10783. American Mathematical Monthly, 108(1):83, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695696.

Quet:2001:ZFS

[348] Leroy Quet, Reiner Martin, and Otto G. Ruehr. A zeta function series: 10786. American Mathematical Monthly, 108(1):84, January 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695697.

Zeilberger:2001:BRB

[349] Doron Zeilberger. Book review: [untitled]. American Mathematical Monthly, 108(1):85–86, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695698.

Guckenheimer:2001:BRB

[350] John Guckenheimer. Book review: [untitled]. American Mathematical Monthly, 108(1):87–89, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695699.

McCallum:2001:BRB

[351] William G. McCallum. Book review: Calculus Mysteries and Thrillers, by R. Grant Woods; How to Ace Calculus: The Streetwise Guide by Colin Adams; Joel Hass; Abigail Thompson. American Mathematical Monthly, 108(1):90–93, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695700.

Ostebee:2001:TRa

[352] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 108(1):94–96, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695701.

Anonymous:2001:BMa

[353] Anonymous. Back matter. American Mathematical Monthly, 108(1):??, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695702.

Anonymous:2001:FMa

[354] Anonymous. Front matter. American Mathematical Monthly, 108(1):??, January 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695669.

Alder:2001:YGG

[355] Henry L. Alder. Yueh-Gin Gung and Dr. Charles Y. Hu Award to Manuel P. Berriozábal for Distinguished Service to Mathematics. American Mathematical Monthly, 108(2):97–99, February 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695522.

DeTemple:2001:SBP

[356] Duane DeTemple and Matthew Hudelson. Square-banded polygons and affine regularity. American Mathematical Monthly, 108(2):100–114, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695523.

Lax:2001:CVM

[357] Peter D. Lax. Change of variables in multiple integrals II. American Mathematical Monthly, 108(2):115–119, February 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695524.

Koranyi:2001:AFD

[358] Adam Korányi. Around the finite-dimensional spectral theorem. American Mathematical Monthly, 108(2):120–125, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695525.

Mueller:2001:RNB

[359] William Mueller. Reform now, before It’s too late! American Mathematical Monthly, 108(2):126–143, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695526.

Kaptanoglu:2001:P

[360] H. Turgay Kaptanoğlu. In praise of y = x^{α} sin(1/x). American Mathematical Monthly, 108(2):144–150, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695527.

Lasserre:2001:MDV

[361] Jean B. Lasserre and Konstantin E. Avrachenkov. The multi-dimensional version of ∫_{a}^{b} x^{p} dx. American Mathematical Monthly, 108(2):151–154, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695528.

Zeller:2001:ASE

[362] Sibylle Zeller and Maximilian Thaler. Almost sure escape from the unit interval under the logistic map. American Mathematical Monthly, 108(2): 155–158, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695529.

Geretschlager:2001:GRI

[363] Robert Geretschläger and Walther Janous. Generalized rearrangement inequalities. American Mathematical Monthly, 108(2):158–165, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695530.

Dar:2001:WEM

[364] Seannie Dar and Shay Gueron. A weighted Erdős–Mordell inequality. American Mathematical Monthly, 108(2):165–168, February 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695531.

Dijkstra:2001:FPS

[365] Jan J. Dijkstra. A four-point set that cannot be split. American Mathematical Monthly, 108(2):168–170, February 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695532.

Beckwith:2001:Pa

[366] David Beckwith. Problem 10851. American Mathematical Monthly, 108 (2):171, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695533.

Popescu:2001:Pa

[367] Călin Popescu. Problem 10852. American Mathematical Monthly, 108 (2):171, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695534.

Stevic:2001:Pa

[368] Stevo Stević. Problem 10853. American Mathematical Monthly, 108(2): 171, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695535.

Chao:2001:Pa

[369] Wu Wei Chao. Problem 10854. American Mathematical Monthly, 108 (2):171, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695536.

Cossi:2001:P

[370] Ernesto Bruno Cossi. Problem 10855. American Mathematical Monthly, 108(2):171, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695537.

Jorza:2001:Pa

[371] Andrei Jorza. Problem 10856. American Mathematical Monthly, 108 (2):172, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695538.

Diamond:2001:P

[372] Harold G. Diamond. Problem 10857. American Mathematical Monthly, 108(2):172, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695539.

Deutsch:2001:CS

[373] Emeric Deutsch, Murray S. Klamkin, and Denis Constales. Counting the solutions: 10540. American Mathematical Monthly, 108(2):172–173, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695540.

Bloom:2001:DKI

[374] David M. Bloom and Stephen M. Gagola, Jr. Dirichlet knew it all along: 10718. American Mathematical Monthly, 108(2):174, February 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695541.

McCoart:2001:MBU

[375] Richard F. McCoart and Lyle Ramshaw. More balls in urns: 10722. American Mathematical Monthly, 108(2):174–176, February 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695542.

Hillar:2001:TSC

[376] Christopher J. Hillar and Heinz-Jürgen Seiffert. Two sums that are congruent modulo p: 10723. American Mathematical Monthly, 108(2):176, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695543.

Jorza:2001:MPU

[377] Andrei Jorza and O. P. Lossers. Maximum of a polynomial on the unit circle: 10779. American Mathematical Monthly, 108(2):177, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695544.

Fekete:2001:SSY

[378] Antal Fekete and Gary Martin. Squared series yielding integers: 10791. American Mathematical Monthly, 108(2):177–178, February 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695545.

Newman:2001:MI

[379] Edward Newman and Li Zhou. A mean inequality: 10798. American Mathematical Monthly, 108(2):178, February 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695546.

Fontenot:2001:BRB

[380] Robert A. Fontenot. Book review: [untitled]. American Mathematical Monthly, 108(2):179–182, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695547.

Schattschneider:2001:BRB

[381] Doris Schattschneider. Book review: [untitled]. American Mathematical Monthly, 108(2):182–188, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695548.

Herbst:2001:BRB

[382] Ben Herbst. Book review: [untitled]. American Mathematical Monthly, 108(2):188–190, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695549.

Anonymous:2001:BMc

[383] Anonymous. Back matter. American Mathematical Monthly, 108(2):??, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695550.

Anonymous:2001:FMc

[384] Anonymous. Front matter. American Mathematical Monthly, 108(2):??, February 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695521.

Morgan:2001:PDB

[385] Frank Morgan. Proof of the double bubble conjecture. American Mathematical Monthly, 108(3):193–205, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695380.

Wolfson:2001:CMS

[386] Paul R. Wolfson. The crooked made straight: Roberval and Newton on tangents. American Mathematical Monthly, 108(3):206–216, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695381.

Spring:2001:SCP

[387] David Spring. Solution of a calculus problem on minimal volume. American Mathematical Monthly, 108(3):217–221, March 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695382.

Huylebrouck:2001:SIP

[388] Dirk Huylebrouck. Similarities in irrationality proofs for π, ln 2, ζ(2), and ζ(3). American Mathematical Monthly, 108(3):222–231, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695383.

Aceto:2001:MPO

[389] Lidia Aceto and Donato Trigiante. The matrices of Pascal and other greats. American Mathematical Monthly, 108(3):232–245, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695384.

Byrnes:2001:LHS

[390] Graham Byrnes, Grant Cairns, and Barry Jessup. Leftovers from the Ham Sandwich Theorem. American Mathematical Monthly, 108(3):246– 249, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695385.

Martinez:2001:SNI

[391] Patrick Martinez. Some new irrational decimal fractions. American Mathematical Monthly, 108(3):250–253, March 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2695386.

Damianou:2001:MPZ

[392] Pantelis A. Damianou. Monic polynomials in Z[x] with roots in the unit disc. American Mathematical Monthly, 108(3):253–257, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695387.

Scheffold:2001:PTP

[393] Egon Scheffold. Pythagorean triples of polygonal numbers. American Mathematical Monthly, 108(3):257–258, March 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695388.

Tuenter:2001:SPS

[394] Hans J. H. Tuenter. A symmetry of power sum polynomials and Bernoulli numbers. American Mathematical Monthly, 108(3):258–261, March 2001.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695389.

Yiu:2001:HTL

[395] Paul Yiu. Heronian triangles are lattice triangles. American Mathematical Monthly, 108(3):261–263, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695390.

Vizgin:2001:EAW

[396] Vl. P. Vizgin. On the emotional assumptions without which one could not effectively investigate the laws of nature. American Mathematical Monthly, 108(3):264–270, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695391.

Knuth:2001:Pa

[397] Donald E. Knuth. Problem 10858. American Mathematical Monthly, 108(3):271, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695392.

Bacher:2001:Pa

[398] Roland Bacher. Problem 10859. American Mathematical Monthly, 108 (3):271, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695393.

Deutsch:2001:Pa

[399] Emeric Deutsch. Problem 10860. American Mathematical Monthly, 108 (3):271, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695394.

Kim:2001:P

[400] Sung Soo Kim. Problem 10861. American Mathematical Monthly, 108 (3):271, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695395.

Chao:2001:Pb

[401] Wu Wei Chao. Problem 10862. American Mathematical Monthly, 108 (3):271, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695396.

Quet:2001:P

[402] Leroy Quet. Problem 10863. American Mathematical Monthly, 108(3): 272, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695397.

Popescu:2001:Pb

[403] Călin Popescu. Problem 10864. American Mathematical Monthly, 108 (3):272, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695398.

Mihai:2001:SRP

[404] Vasile Mihai and Michael Woltermann. The smoothest and roughest permutations: 10725. American Mathematical Monthly, 108(3):272–273, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695399.

Anglesio:2001:PSP

[405] Jean Anglesio and O. P. Lossers. Periods from sums of powers: 10727. American Mathematical Monthly, 108(3):273, March 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695400.

Deshpande:2001:TLS

[406] M. N. Deshpande and John H. Lindsey II. The total length of short cycles of a random permutation: 10732. American Mathematical Monthly, 108 (3):273–274, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695401.

Ali:2001:GSW

[407] Hassan Ali Shah Ali and National Security Agency Problems Group. Getting the same weighted average two ways: 10737. American Mathematical Monthly, 108(3):274, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695402.

Keller:2001:SFE

[408] Tim Keller and Javier Cilleruelo. Squares with four equal digits in an even base: 10741. American Mathematical Monthly, 108(3):275, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695403.

Arias-de-Reyna:2001:SRD

[409] Juan Arias de Reyna and Robin Chapman. The square of a random determinant: 10787. American Mathematical Monthly, 108(3):275–277, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695404.

Diamond:2001:CFD

[410] Harold G. Diamond and Ferrell S. Wheeler. A convolution formula for the Dickman function: 10792. American Mathematical Monthly, 108(3): 277–278, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695405.

Herink:2001:FSO

[411] Curtis Herink, Gary Gruenhage, and GCHQ Problems Group. Finite subcovers for open covers with exact cardinality: 10799. American Mathematical Monthly, 108(3):278, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695406.

Staples:2001:BRB

[412] Susan G. Staples. Book review: [untitled]. American Mathematical Monthly, 108(3):279–280, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695407.

Andrews:2001:BRB

[413] George E. Andrews. Book review: [untitled]. American Mathematical Monthly, 108(3):281–285, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695408.

Ostebee:2001:TRb

[414] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 108(3):286–288, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695409.

Anonymous:2001:BMd

[415] Anonymous. Back matter. American Mathematical Monthly, 108(3):??, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695410.

Anonymous:2001:FMd

[416] Anonymous. Front matter. American Mathematical Monthly, 108(3):??, March 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695379.

Bhatia:2001:LAQ

[417] Rajendra Bhatia. Linear algebra to quantum cohomology: The story of alfred Horn’s inequalities. American Mathematical Monthly, 108(4):289– 318, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695237.

Shail:2001:PTT

[418] R. Shail. A proof of Thébault’s Theorem. American Mathematical Monthly, 108(4):319–325, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695238.

Hubbard:2001:CEA

[419] John H. Hubbard, Samer S. Habre, and Beverly H. West. The convergence of an Euler approximation of an initial value problem is not always obvious. American Mathematical Monthly, 108(4):326–335, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695239.

Sobczyk:2001:MSB

[420] Garret Sobczyk. The missing spectral basis in algebra and number theory. American Mathematical Monthly, 108(4):336–346, April 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695240.

Hall:2001:MMI

[421] Rachel W. Hall and Krešimir Josić. The mathematics of musical instruments. American Mathematical Monthly, 108(4):347–357, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695241.

Schellwat:2001:SSR

[422] Holger Schellwat. A simple slide rule for finite fields. American Mathematical Monthly, 108(4):358–360, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695242.

Satnoianu:2001:GME

[423] Razvan Alin Satnoianu. A general method for establishing geometric inequalities in a triangle. American Mathematical Monthly, 108(4):360– 364, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695243.

Aldaz:2001:TDJ

[424] J. M. Aldaz, A. Bravo, S. Gutiérrez, and A. Ubis. A theorem of D. J. Newman on Euler’s ϕ function and arithmetic progressions. American Mathematical Monthly, 108(4):364–367, April 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695244.

Wen:2001:CTD

[425] Liu Wen. A counterexample for the two-dimensional density function. American Mathematical Monthly, 108(4):367–368, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695245.

Astapov:2001:RT

[426] N. S. Astapov and N. C. Noland. The remarkable tetron. American Mathematical Monthly, 108(4):368–370, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695246.

Beckwith:2001:Pb

[427] David Beckwith. Problem 10865. American Mathematical Monthly, 108 (4):371, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695247.

Kazdan:2001:P

[428] Jerry Kazdan and Herbert Wilf. Problem 10866. American Mathematical Monthly, 108(4):371, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695248.

Ramirez:2001:P

[429] Oscar Ciaurri Ramirez. Problem 10867. American Mathematical Monthly, 108(4):371, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695249.

Deshpande:2001:P

[430] M. N. Deshpande and A. V. Kharashikar. Problem 10868. American Mathematical Monthly, 108(4):371, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695250.

Jones:2001:P

[431] Lenny Jones. Problem 10869. American Mathematical Monthly, 108(4): 372, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695251.

Montgomery:2001:P

[432] Hugh L. Montgomery and Ulrike M. A. Vorhauer. Problem 10870. American Mathematical Monthly, 108(4):372, April 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695252.

Knuth:2001:Pb

[433] Donald E. Knuth. Problem 10871. American Mathematical Monthly, 108(4):372, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695253.

Andreescu:2001:PST

[434] Titu Andreescu and Gilbert N. Lewis. Preserving the sum of three cubes: 10728. American Mathematical Monthly, 108(4):372–373, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695254.

Simmons:2001:CPT

[435] Gustavus J. Simmons, John Henry Steelman, GCHQ Problem Solving Group, and O. P. Lossers. A congruence for Pascal triangular matrices: 10735. American Mathematical Monthly, 108(4):373–374, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695255.

Khan:2001:OMC

[436] Mizan R. Khan. An optimization with a modular constraint: 10736. American Mathematical Monthly, 108(4):374–375, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695256.

Eynden:2001:DBP

[437] Charles Vanden Eynden and Jerrold W. Grossman. Distinguishing between the partite sets of a bipartite graph: 10740. American Mathematical Monthly, 108(4):375, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695257.

Alkan:2001:MPF

[438] Emre Alkan and Stephen Gagola, Jr. The maximal property for finite groups: 10742. American Mathematical Monthly, 108(4):375–376, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695258.

Popescu:2001:CSQ

[439] Călin Popescu and O. P. Lossers. Comparing a sum over quadratic residues and nonresidues: 10743. American Mathematical Monthly, 108 (4):376–377, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695259.

Fukuta:2001:QTQ

[440] Jiro Fukuta. Not quite a theorem on quadrilaterals: 10755. American Mathematical Monthly, 108(4):377–378, April 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695260.

Sinefakopoulos:2001:CIQ

[441] Achilleas Sinefakopoulos and Daniele Donini. Cirumscribing an inscribed quadrilateral: 10804. American Mathematical Monthly, 108(4):378, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695261.

Hobbs:2001:BRB

[442] Arthur M. Hobbs. Book review: [untitled]. American Mathematical Monthly, 108(4):379–381, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695262.

Maor:2001:BRBa

[443] Eli Maor. Book review: [untitled]. American Mathematical Monthly, 108 (4):382–385, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695263.

Pittie:2001:BRB

[444] Harsh V. Pittie. Book review: [untitled]. American Mathematical Monthly, 108(4):386–387, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695264.

Anonymous:2001:BMe

[445] Anonymous. Back matter. American Mathematical Monthly, 108(4): ??, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695265.

Anonymous:2001:FMe

[446] Anonymous. Front matter. American Mathematical Monthly, 108(4): ??, April 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695236.

Reid:2001:ALB

[447] Constance Reid. The alternative life of E. T. Bell. American Mathematical Monthly, 108(5):393–402, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695793.

Lindqvist:2001:TRI

[448] Peter Lindqvist and Jaak Peetre. Two remarkable identities, called twos, for inverses to some Abelian integrals. American Mathematical Monthly, 108(5):403–410, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695794.

Kennedy:2001:CL

[449] Judy Kennedy, Sahin Koçak, and James A. Yorke. A chaos lemma. American Mathematical Monthly, 108(5):411–423, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695795.

Greene:2001:RAC

[450] Robert E. Greene. Real analyticity: From calculus class to the Grauert– Morrey Theorem. American Mathematical Monthly, 108(5):424–431, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695796.

Talvila:2001:SDT

[451] Erik Talvila. Some divergent trigonometric integrals. American Mathematical Monthly, 108(5):432–436, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695797.

Brauer:2001:WGM

[452] Fred Brauer. What goes up must come down, eventually. American Mathematical Monthly, 108(5):437–440, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695798.

Dijkstra:2001:DCP

[453] Edsger W. Dijkstra and Jayadev Misra. Designing a calculational proof of Cantor’s theorem. American Mathematical Monthly, 108(5):440–443, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695799.

DeMarco:2001:ETC

[454] Giuseppe De Marco. For every ϵ there continuously exists a δ. American Mathematical Monthly, 108(5):443–444, May 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695800.

Hatzipolakis:2001:LCT

[455] Antreas P. Hatzipolakis and Paul Yiu. The Lucas circles of a triangle. American Mathematical Monthly, 108(5):444–446, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695801.

Folland:2001:HIP

[456] Gerald B. Folland. How to integrate a polynomial over a sphere. American Mathematical Monthly, 108(5):446–448, May 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695802.

Marek:2001:FMT

[457] V. Wiktor Marek and Jan Mycielski. Foundations of mathematics in the twentieth century. American Mathematical Monthly, 108(5):449–468, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695803.

Suceava:2001:P

[458] Bogdan Suceavă. Problem 10872. American Mathematical Monthly, 108 (5):469, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695804.

Venkatachala:2001:P

[459] B. J. Venkatachala. Problem 10873. American Mathematical Monthly, 108(5):469, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695805.

Chao:2001:Pc

[460] Wu Wei Chao. Problem 10874. American Mathematical Monthly, 108(5): 469, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695806.

Knuth:2001:Pc

[461] Donald E. Knuth. Problem 10875. American Mathematical Monthly, 108(5):469, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695807.

Sasvari:2001:P

[462] Zoltán Sasvári. Problem 10876. American Mathematical Monthly, 108 (5):469, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695808.

Deutsch:2001:Pb

[463] Emeric Deutsch. Problem 10877. American Mathematical Monthly, 108 (5):470, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695809.

Dalyay:2001:P

[464] Pál Péter Dályay. Problem 10878. American Mathematical Monthly, 108(5):470, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695810.

Locke:2001:HCG

[465] S. C. Locke, Phil Tracy, and H.-J. Voss. Highly cohesive graphs have long nonseparating paths: 10647. American Mathematical Monthly, 108 (5):470–472, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695811.

Tabachnikov:2001:PCT

[466] Serge Tabachnikov. Polyhedron covered twice by lines: 10724. American Mathematical Monthly, 108(5):472–473, May 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695812.

Lindqvist:2001:GTF

[467] Peter Lindqvist, Jaak Peetre, and Jonathan M. Borwein. Generalized trigonometric functions: 10744. American Mathematical Monthly, 108 (5):473–474, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695813.

Deleglise:2001:ULP

[468] Marc Deléglise and Joe Flowers. An unusual limit problem: 10807. American Mathematical Monthly, 108(5):474–475, May 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695814.

Pinchover:2001:PCN

[469] Yehuda Pinchover, Simeon Reich, and Itai Shafrir. The Ptolemy constant of a normed space: 10812. American Mathematical Monthly, 108(5):475– 476, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695815.

Nievergelt:2001:BRB

[470] Yves Nievergelt. Book review: [untitled]. American Mathematical Monthly, 108(5):477–481, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695816.

Scott:2001:BRB

[471] David Scott. Book review: [untitled]. American Mathematical Monthly, 108(5):481–485, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695817.

Ostebee:2001:TRc

[472] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 108(5):486–487, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695818.

Anonymous:2001:BMf

[473] Anonymous. Back matter. American Mathematical Monthly, 108(5): ??, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695819.

Anonymous:2001:FMf

[474] Anonymous. Front matter. American Mathematical Monthly, 108(5): ??, May 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695792.

Pasles:2001:LSD

[475] Paul C. Pasles. The lost squares of dr. franklin: Ben Franklin’s missing squares and the secret of the magic circle. American Mathematical Monthly, 108(6):489–511, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695704.

Hanke:2001:IPL

[476] Martin Hanke and Otmar Scherzer. Inverse problems light: Numerical differentiation. American Mathematical Monthly, 108(6):512–521, June/ July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695705.

Samelson:2001:DFE

[477] Hans Samelson. Differential forms, the early days; or the stories of Deahna’s theorem and of Volterra’s theorem. American Mathematical Monthly, 108(6):522–530, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695706.

Maxson:2001:HMS

[478] C. J. Maxson and J. H. Meyer. How many subspaces force linearity? American Mathematical Monthly, 108(6):531–536, June/July 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695707.

Pijls:2001:SPR

[479] Wim Pijls. Some properties related to mercator projection. American Mathematical Monthly, 108(6):537–543, June/July 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695708.

Talvila:2001:NSC

[480] Erik Talvila. Necessary and sufficient conditions for differentiating under the integral sign. American Mathematical Monthly, 108(6):544–548, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695709.

Quadrat:2001:PTA

[481] Jean-P. Quadrat, Jean B. Lasserre, and Jean-B. Hiriart-Urruty. Pythagoras’ theorem for areas. American Mathematical Monthly, 108(6):549–551, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695710.

Sanchez:2001:ASM

[482] David A. Sánchez. An alternative to the shooting method for a certain class of boundary value problems. American Mathematical Monthly, 108(6):552–555, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695711.

Woan:2001:UPL

[483] Wen jin Woan. Uniform partitions of lattice paths and Chung–Feller generalizations. American Mathematical Monthly, 108(6):556–559, June/ July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695712.

Flanders:2001:FF

[484] Harley Flanders. From Ford to Faà. American Mathematical Monthly, 108(6):559–561, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695713.

Hajja:2001:VPT

[485] Mowaffaq Hajja. A vector proof of a theorem of bang. American Mathematical Monthly, 108(6):562–564, June/July 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695714.

Janous:2001:P

[486] Walther Janous. Problem 10879. American Mathematical Monthly, 108 (6):565, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695715.

Satnoianu:2001:P

[487] Razvan A. Satnoianu. Problem 10880. American Mathematical Monthly, 108(6):565, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695716.

Palacios:2001:P

[488] José Luis Palacios. Problem 10881. American Mathematical Monthly, 108 (6):565, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695717.

Just:2001:P

[489] Erwin Just and Norman Schaumberger. Problem 10882. American Mathematical Monthly, 108(6):565, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695718.

MacKinnon:2001:Pb

[490] Nick MacKinnon. Problem 10883. American Mathematical Monthly, 108 (6):565, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695719.

Ahmed:2001:P

[491] Zafar Ahmed. Problem 10884. American Mathematical Monthly, 108 (6):566, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695720.

Kedlaya:2001:P

[492] Kiran Kedlaya and Sasha Schwartz. Problem 10885. American Mathematical Monthly, 108(6):566, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695721.

Pelling:2001:MQF

[493] M. J. Pelling and GCHQ Problem Solving Group. Maximizing quadratic forms over cubes: 10731. American Mathematical Monthly, 108(6):566– 567, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695722.

Grigis:2001:RBP

[494] Alain Grigis and W. Weston Meyer. Reflections of a billiard path: 10749. American Mathematical Monthly, 108(6):567, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695723.

Anglesio:2001:NDF

[495] Jean Anglesio and Con Amore Problems Group. A nowhere-differentiable function from geometry: 10790. American Mathematical Monthly, 108 (6):568, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695724.

vanLamoen:2001:ELC

[496] Floor van Lamoen and W. Weston Meyer. Euler lines concurrent on the nine-point circle: 10796. American Mathematical Monthly, 108(6):569, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695725.

Ali:2001:DSR

[497] Hassan Ali Shah Ali, Keith A. Kearnes, Ágnes Szenderei, Zafar Ahmed, M. A. Prasad, and GCHQ Problems Group. A difference of square roots: 10806. American Mathematical Monthly, 108(6):570, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695726.

Marquez:2001:PIQ

[498] Juan-Bosco Romero Marquez and Mohammed Ali Salem. A parallelogram inscribed in a quadrilateral: 10810. American Mathematical Monthly, 108(6):571, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695727.

Satnoianu:2001:REM

[499] Razvan Satnoianu and Li Zhou. Relatives of Erdős–Mordell: 10814. American Mathematical Monthly, 108(6):572, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695728.

Cohen:2001:BRB

[500] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 108(6):573–577, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695729.

Alewine:2001:BRB

[501] J. Alan Alewine and Eric Schechter. Book review: [untitled]. American Mathematical Monthly, 108(6):577–582, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695730.

Anonymous:2001:BMg

[502] Anonymous. Back matter. American Mathematical Monthly, 108(6):??, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695731.

Anonymous:2001:FMg

[503] Anonymous. Front matter. American Mathematical Monthly, 108(6):??, June/July 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695703.

Weeks:2001:TPC

[504] Jeffrey R. Weeks. The twin paradox in a closed universe. American Mathematical Monthly, 108(7):585–590, August/September 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695267.

Li:2001:LPP

[505] Chi-Kwong Li and Stephen Pierce. Linear preserver problems. American Mathematical Monthly, 108(7):591–605, August/September 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695268.

Zitarelli:2001:TFA

[506] David E. Zitarelli. Towering figures in American mathematics, 1890-1950. American Mathematical Monthly, 108(7):606–635, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695269.

Talvila:2001:RGF

[507] Erik Talvila. Rapidly growing Fourier integrals. American Mathematical Monthly, 108(7):636–641, August/September 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695270.

Boel:2001:USS

[508] Soren Boel, Toke Meier Carlsen, and Niels Richard Hansen. A useful strengthening of the Stone–Weierstrass Theorem. American Mathematical Monthly, 108(7):642–643, August/September 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695271.

Zhan:2001:LPP

[509] Xingzhi Zhan. Linear preservers that permute the entries of a matrix. American Mathematical Monthly, 108(7):643–645, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695272.

Desbrow:2001:RAC

[510] Darrell Desbrow. A real antiderivative by contour integration. American Mathematical Monthly, 108(7):645–649, August/September 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695273.

Halava:2001:MMS

[511] Vesa Halava and Tero Harju. Mortality in matrix semigroups. American Mathematical Monthly, 108(7):649–653, August/September 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695274.

Atiyah:2001:MC

[512] Michael Atiyah. Mathematics in the 20th century. American Mathematical Monthly, 108(7):654–666, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695275.

Rey:2001:P

[513] Joaquín Gómez Rey. Problem 10886. American Mathematical Monthly, 108(7):667, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695276.

Anglesio:2001:Pa

[514] Jean Anglesio. Problem 10887. American Mathematical Monthly, 108 (7):667, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695277.

Wardlaw:2001:P

[515] William P. Wardlaw and William D. Withers. Problem 10888. American Mathematical Monthly, 108(7):667, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695278.

Huber:2001:P

[516] Klaus Huber. Problem 10889. American Mathematical Monthly, 108 (7):667, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695279.

Mortini:2001:P

[517] Raymond Mortini. Problem 10890. American Mathematical Monthly, 108 (7):668, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695280.

Bacher:2001:Pb

[518] Roland Bacher. Problem 10891. American Mathematical Monthly, 108 (7):668, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695281.

Locke:2001:P

[519] Stephen C. Locke. Problem 10892. American Mathematical Monthly, 108 (7):668, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695282.

Pelling:2001:SDP

[520] M. J. Pelling and F. W. Roush. The sum divides the product: 10745. American Mathematical Monthly, 108(7):668–669, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695283.

Borosh:2001:OGC

[521] Itshak Borosh, Douglas A. Hensley, Joel Zinn, and GCHQ Problem Solving Group. One good congruence deserves another: 10748. American Mathematical Monthly, 108(7):669, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695284.

Smiley:2001:RGF

[522] Leonard Smiley, David Callan, David M. Wells, and Said Amghibech. A rational generating function for relative divisors: 10750. American Mathematical Monthly, 108(7):670, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695285.

Anglesio:2001:SPT

[523] Jean Anglesio and Daniele Donini. Special points of the triangle of excenters: 10763. American Mathematical Monthly, 108(7):671, August/ September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695286.

Bertram:2001:SFD

[524] Barbara S. Bertram and Otto G. Ruehr. Submultiplicativity of a function defined by a series: 10815. American Mathematical Monthly, 108(7): 671–672, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695287.

Swetz:2001:BRB

[525] Frank Swetz. Book review: [untitled]. American Mathematical Monthly, 108(7):673–675, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695288.

Maor:2001:BRBb

[526] Eli Maor. Book review: [untitled]. American Mathematical Monthly, 108(7):675–677, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695289.

Ostebee:2001:TRd

[527] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 108(7):678–680, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695290.

Anonymous:2001:LRF

[528] Anonymous. Lester R. Ford Awards for 2000. American Mathematical Monthly, 108(7):681, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695291.

Anonymous:2001:BMh

[529] Anonymous. Back matter. American Mathematical Monthly, 108(7):??, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695292.

Anonymous:2001:FMh

[530] Anonymous. Front matter. American Mathematical Monthly, 108(7):??, August/September 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695266.

Roberts:2001:HME

[531] David Lindsay Roberts. E. H. Moore’s early Twentieth-Century program for reform in mathematics education. American Mathematical Monthly, 108(8):689–696, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695612.

Fallat:2001:BFT

[532] Shaun M. Fallat. Bidiagonal factorizations of totally nonnegative matrices. American Mathematical Monthly, 108(8):697–712, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695613.

Rawlings:2001:SSP

[533] Don Rawlings. Sequential searches: Proofreading, Russian roulette, and the incomplete q-Eulerian polynomials revisited. American Mathematical Monthly, 108(8):713–720, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695614.

Holly:2001:PUS

[534] Jan E. Holly. Pictures of ultrametric spaces, the p-adic numbers, and valued fields. American Mathematical Monthly, 108(8):721–728, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695615.

Meyer:2001:JEF

[535] Kenneth R. Meyer. Jacobi elliptic functions from a dynamical systems point of view. American Mathematical Monthly, 108(8):729–737, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695616.

McKenna:2001:LTO

[536] P. J. McKenna and Cillian Ó Tuama. Large torsional oscillations in suspension bridges visited again: Vertical forcing creates torsional response. American Mathematical Monthly, 108(8):738–745, October 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695617.

Martin:2001:AAN

[537] Greg Martin. Absolutely abnormal numbers. American Mathematical Monthly, 108(8):746–754, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695618.

Stuart:2001:LDD

[538] Christopher Stuart. Limiting distributions for derangements. American Mathematical Monthly, 108(8):755, October 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695619.

Rogers:2001:FFK

[539] Keith McKenzie Rogers. The finite field Kakeya problem. American Mathematical Monthly, 108(8):756–759, October 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695620.

deMedeiros:2001:FTA

[540] Airton von Sohsten de Medeiros. The fundamental theorem of algebra revisited. American Mathematical Monthly, 108(8):759–760, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695621.

Hajja:2001:OVS

[541] Mowaffaq Hajja. Other versions of the Steiner–Lehmus Theorem. American Mathematical Monthly, 108(8):760–767, October 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695622.

deAzevedoPribitkin:2001:SP

[542] Wladimir de Azevedo Pribitkin. A simpler proof of sinπz = πz ∏_{k=1}^{∞}(1− z^{2}/k^{2}). American Mathematical Monthly, 108(8):767–768, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695623.

Lasserre:2001:QPV

[543] Jean B. Lasserre. A quick proof for the volume of n-balls. American Mathematical Monthly, 108(8):768–769, October 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695624.

Jorza:2001:Pb

[544] Andrei Jorza. Problem 10893. American Mathematical Monthly, 108 (8):770, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695625.

Sulanke:2001:P

[545] Robert Sulanke. Problem 10894. American Mathematical Monthly, 108 (8):770, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695626.

Wendijk:2001:P

[546] Aūgūst Wendijk. Problem 10895. American Mathematical Monthly, 108 (8):770, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695627.

Bisgaard:2001:P

[547] Torben Maack Bisgaard, Zsolt Páles, and Zoltan Sasvari. Problem 10896. American Mathematical Monthly, 108(8):770, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695628.

Wene:2001:P

[548] Gregory P. Wene. Problem 10897. American Mathematical Monthly, 108 (8):771, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695629.

Chao:2001:Pd

[549] Wu Wei Chao. Problem 10898. American Mathematical Monthly, 108 (8):771, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695630.

Martini:2001:P

[550] R. Martini and G. F. Post. Problem 10899. American Mathematical Monthly, 108(8):771, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695631.

Bracken:2001:EFZ

[551] Paul Bracken and Bruce S. Burdick. Euler’s formula for zeta function convolutions: 10754. American Mathematical Monthly, 108(8):771–773, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695632.

Reznick:2001:CMF

[552] Bruce Reznick and Bruce Dearden. Completely multiplicative functions: 10760. American Mathematical Monthly, 108(8):773, October 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695633.

Galvin:2001:EPI

[553] Fred Galvin and Radhika Ramamurthi. An extremal problem for independent sets in graphs: 10761. American Mathematical Monthly, 108(8): 773–774, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695634.

Dearden:2001:IMM

[554] Bruce Dearden, Jerry Metzer, and Robert Gilmer. Invertible matrices modulo n: 10767. American Mathematical Monthly, 108(8):774, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695635.

Blatter:2001:CNS

[555] Christian Blatter and Josh Greene. The chromatic number of the sphere: 10769. American Mathematical Monthly, 108(8):774–775, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695636.

Valdinoci:2001:CS

[556] Enrico Valdinoci and Robin Chapman. Convergence of a series: 10808. American Mathematical Monthly, 108(8):775–777, October 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695637.

Janous:2001:SI

[557] Walther Janous and Anne-Maria Ernvall. A summation inequality: 10817. American Mathematical Monthly, 108(8):777, October 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695638.

Laudal:2001:BRB

[558] O. A. Laudal. Book review: [untitled]. American Mathematical Monthly, 108(8):778–781, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695639.

McGehee:2001:BRB

[559] O. Carruth McGehee. Book review: [untitled]. American Mathematical Monthly, 108(8):782–784, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695640.

Anonymous:2001:BMi

[560] Anonymous. Back matter. American Mathematical Monthly, 108(8):??, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695641.

Anonymous:2001:FMi

[561] Anonymous. Front matter. American Mathematical Monthly, 108(8):??, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695611.

Mueller:2001:MW

[562] William Mueller. Mathematical wunderkammern. American Mathematical Monthly, 108(9):785–796, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695552.

Lawson:2001:GMM

[563] Jimmie D. Lawson and Yongdo Lim. The geometric mean, matrices, metrics, and more. American Mathematical Monthly, 108(9):797–812, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695553.

Osserman:2001:KLN

[564] Robert Osserman. Kepler’s laws, Newton’s laws, and the search for new planets. American Mathematical Monthly, 108(9):813–820, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695554.

Kalman:2001:PEC

[565] Dan Kalman and James E. White. Polynomial equations and circulant matrices. American Mathematical Monthly, 108(9):821–840, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695555.

Klosinski:2001:SFW

[566] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-First William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 108(9):841–850, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695556.

Laszlo:2001:SSC

[567] Ákos László. The sum of some convergent series. American Mathematical Monthly, 108(9):851–855, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695557.

Ramm:2001:SPF

[568] A. G. Ramm. A simple proof of the Fredholm alternative and a characterization of the Fredholm operators. American Mathematical Monthly, 108(9):855–860, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695558.

Benedetto:2001:EPI

[569] Robert L. Benedetto. An elementary product identity in polynomial dynamics. American Mathematical Monthly, 108(9):860–864, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695559.

Gray:2001:WLI

[570] Jeremy Gray. Weierstrass, Luzin, and intuition. American Mathematical Monthly, 108(9):865–870, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695560.

Rice:2001:P

[571] Gordon Rice. Problem 10900. American Mathematical Monthly, 108 (9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695561.

Knight:2001:P

[572] M. J. Knight. Problem 10901. American Mathematical Monthly, 108 (9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695562.

Deutsch:2001:Pc

[573] Emeric Deutsch. Problem 10902. American Mathematical Monthly, 108 (9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695563.

Bagby:2001:Pb

[574] Richard Bagby. Problem 10903. American Mathematical Monthly, 108 (9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695564.

Furdui:2001:P

[575] Ovidiu Furdui. Problem 10904. American Mathematical Monthly, 108 (9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695565.

Stanley:2001:P

[576] Richard P. Stanley. Problem 10905. American Mathematical Monthly, 108(9):871, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695566.

Knuth:2001:Pd

[577] Donald E. Knuth. Problem 10906. American Mathematical Monthly, 108 (9):872, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695567.

Deutsch:2001:APG

[578] Emeric Deutsch, David Callan, Sabin Cautis, and Southwest Missouri Problems Group. Another path to generalized Catalan numbers: 10751. American Mathematical Monthly, 108(9):872–873, November 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695568.

Shapiro:2001:LOT

[579] Louis Shapiro, Richard Ehrenborg, John W. Moon, and Michael Reid. Leaves of ordered trees: 10753. American Mathematical Monthly, 108(9): 873–874, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695569.

Kidwell:2001:LOT

[580] Mark Kidwell, Donald Knuth, and O. P. Lossers. Leaves of ordered trees: 10757. American Mathematical Monthly, 108(9):874–875, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695570.

Foschini:2001:SFD

[581] Gerard J. Foschini and GCHQ Problems Group. A sequence of functions that doesn’t span L_{2} [0,1]: 10821. American Mathematical Monthly, 108(9):875–876, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695571.

Lagarias:2001:WPP

[582] Jeffrey Lagarias, Jade Vinson, and GCHQ Problems Group. When periodic points are real: 10822. American Mathematical Monthly, 108(9): 876–877, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695572.

Martinez-Gimenez:2001:CPI

[583] Félix Martínez-Giménez and John H. Lindsey II. A convergence property for infinite arrays: 10826. American Mathematical Monthly, 108(9):877, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695573.

Knuth:2001:SS

[584] Donald E. Knuth and Cecil C. Rousseau. A Stirling series: 10832. American Mathematical Monthly, 108(9):877–878, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695574.

Borel:2001:BRB

[585] Armand Borel. Book review: [untitled]. American Mathematical Monthly, 108(9):879–883, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695575.

Kon:2001:BRB

[586] Mark A. Kon. Book review: [untitled]. American Mathematical Monthly, 108(9):883–887, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695576.

Ostebee:2001:TRe

[587] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 108(9):888–890, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695577.

Anonymous:2001:BMj

[588] Anonymous. Back matter. American Mathematical Monthly, 108(9):??, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695578.

Anonymous:2001:FMj

[589] Anonymous. Front matter. American Mathematical Monthly, 108(9):??, November 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695551.

Borwein:2001:VSN

[590] Peter Borwein and Loki Jörgenson. Visible structures in number theory. American Mathematical Monthly, 108(10):897–910, December 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695413.

Chalice:2001:HDI

[591] Donald R. Chalice. How to differentiate and integrate sequences. American Mathematical Monthly, 108(10):911–921, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695414.

Grier:2001:DVT

[592] David Alan Grier. Dr. veblen takes a uniform mathematics in the first world war. American Mathematical Monthly, 108(10):922–931, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695415.

Pollatsek:2001:QEC

[593] Harriet Pollatsek. Quantum error correction: Classic group theory meets a quantum challenge. American Mathematical Monthly, 108(10):932–962, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695416.

Cho:2001:IF

[594] Yunhi Cho and Kyunghwan Park. Inverse functions of y = x^{1/x}. American Mathematical Monthly, 108(10):963–967, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695417.

Maehara:2001:TEL

[595] Hiroshi Maehara. On the total edge-length of a tetrahedron. American Mathematical Monthly, 108(10):967–969, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695418.

Balanzario:2001:EDS

[596] Eugenio P. Balanzario. Evaluation of Dirichlet series. American Mathematical Monthly, 108(10):969–971, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695419.

Feeman:2001:ATP

[597] Timothy G. Feeman and Osvaldo Marrero. Affine transformations, polynomials, and proportionality. American Mathematical Monthly, 108(10): 972–975, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695420.

Andras:2001:P

[598] Szilárd András. Problem 10907. American Mathematical Monthly, 108 (10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695421.

Chao:2001:Pe

[599] Wu Wei Chao. Problem 10908. American Mathematical Monthly, 108 (10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695422.

Kalman:2001:P

[600] Támás Kalmán. Problem 10909. American Mathematical Monthly, 108 (10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695423.

Stevic:2001:Pb

[601] Stevo Stević and Rade Todorović. Problem 10910. American Mathematical Monthly, 108(10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695424.

Martin:2001:P

[602] Greg Martin. Problem 10911. American Mathematical Monthly, 108 (10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695425.

Anglesio:2001:Pb

[603] Jean Anglesio. Problem 10912. American Mathematical Monthly, 108 (10):976, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695426.

Knuth:2001:Pe

[604] Donald E. Knuth. Problem 10913. American Mathematical Monthly, 108 (10):977, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695427.

Sapir:2001:DEP

[605] Mark Sapir, O. P. Lossers, and Hugh L. Montgomery. The decimal expansions of powers of 9: 10758. American Mathematical Monthly, 108(10): 977–978, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695428.

Ferraro:2001:FNP

[606] Peter J. Ferraro. Fibonacci numbers and powers: 10765. American Mathematical Monthly, 108(10):978–979, December 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695429.

Popescu:2001:DRC

[607] Călin Popescu, Jim Vandergriff, and Jim Delany. A divisibility result for a combination of mth powers: 10770. American Mathematical Monthly, 108(10):979–980, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695430.

Deutsch:2001:TDL

[608] Emeric Deutsch and Jim Brawner. Three-dimensional lattice walks in the upper half-space: 10795. American Mathematical Monthly, 108(10):980, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695431.

Wellner:2001:IH

[609] Jon A. Wellner and M. L. Glasser. Incognito hypergeometrics: 10836. American Mathematical Monthly, 108(10):980–981, December 2001. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695432.

Vince:2001:AFS

[610] Andrew Vince and Kenneth Schilling. Almost fixed sets: 10843. American Mathematical Monthly, 108(10):981–982, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695433.

Galvin:2001:GHTb

[611] Fred Galvin. A generalization of Hall’s theorem: 10701. American Mathematical Monthly, 108(10):982, December 2001. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2695434.

Koblitz:2001:BRB

[612] Neal Koblitz. Book review: [untitled]. American Mathematical Monthly, 108(10):983–988, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695435.

Ralston:2001:BRB

[613] Anthony Ralston. Book review: [untitled]. American Mathematical Monthly, 108(10):988–992, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695436.

Horn:2001:EE

[614] Roger A. Horn. Editor’s endnotes. American Mathematical Monthly, 108(10):993–994, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695437.

Anonymous:2001:VI

[615] Anonymous. Volume information. American Mathematical Monthly, 108 (10):995–1006, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695411.

Anonymous:2001:BMb

[616] Anonymous. Back matter. American Mathematical Monthly, 108(10):??, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695438.

Anonymous:2001:FMb

[617] Anonymous. Front matter. American Mathematical Monthly, 108(10):??, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695412.

Almgren:2002:FDP

[618] Robert Almgren. Financial derivatives and partial differential equations. American Mathematical Monthly, 109(1):1–12, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695763.

Schwartz:2002:LLG

[619] Richard Evan Schwartz. Lucy and Lily: a game of geometry and number theory. American Mathematical Monthly, 109(1):13–20, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695764.

Weideman:2002:NIP

[620] J. A. C. Weideman. Numerical integration of periodic functions: a few examples. American Mathematical Monthly, 109(1):21–36, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695765.

Shapiro:2002:HFU

[621] Victor L. Shapiro. Harmonic functions in the unit disk. American Mathematical Monthly, 109(1):37–45, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695766.

Volchan:2002:WRS

[622] Sérgio B. Volchan. What is a random sequence? American Mathematical Monthly, 109(1):46–63, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695767.

Roselli:2002:CI

[623] Paolo Roselli and Michel Willem. A convexity inequality. American Mathematical Monthly, 109(1):64–70, January 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695768.

Waterhouse:2002:QPU

[624] William C. Waterhouse. Quadratic polynomials and unique factorization. American Mathematical Monthly, 109(1):70–72, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695769.

Ballantine:2002:SPR

[625] C. Ballantine and J. Roberts. A simple proof of Rolle’s theorem for finite fields. American Mathematical Monthly, 109(1):72–74, January 2002.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695770.

Sellke:2002:YAP

[626] Thomas Sellke. Yet another proof of the Radon–Nikodým theorem. American Mathematical Monthly, 109(1):74–76, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695771.

Falcone:2002:P

[627] Giovanni Falcone. Problem 10914. American Mathematical Monthly, 109(1):77, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695772.

Rupert:2002:Pa

[628] C. P. Rupert. Problem 10915. American Mathematical Monthly, 109(1): 77, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695773.

Ehrlich:2002:P

[629] Gertrude Ehrlich. Problem 10916. American Mathematical Monthly, 109(1):77, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695774.

Gross:2002:P

[630] Jürgen Groß and Götz Trenkler. Problem 10917. American Mathematical Monthly, 109(1):77, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695775.

Beck:2002:P

[631] Matthias Beck. Problem 10918. American Mathematical Monthly, 109 (1):77, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695776.

Becker:2002:P

[632] Michael Becker. Problem 10919. American Mathematical Monthly, 109 (1):77–78, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695777.

Kundgen:2002:P

[633] André Kündgen and Radhika Ramamurthi. Problem 10920. American Mathematical Monthly, 109(1):78, January 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2695778.

MacKinnon:2002:FTP

[634] Nick MacKinnon and Stephen M. Gagola, Jr. Fibonacci twin primes: 10844. American Mathematical Monthly, 109(1):78, January 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695779.

Demeter:2002:ACR

[635] Ciprian Demeter and Li Zhou. Another collinearity result for triangles: 10845. American Mathematical Monthly, 109(1):78–79, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695780.

Przytycki:2002:SLC

[636] Józef Przytycki and Richard Stong. Shrinking lattice curves: 10846. American Mathematical Monthly, 109(1):79–80, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695781.

Chapman:2002:NDS

[637] Robin Chapman, Kimmo Ericksson, Richard P. Stanley, and Reiner Martin. On the number of divisors of n in a special interval: 10847. American Mathematical Monthly, 109(1):80, January 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/2695782.

Stolarsky:2002:CQP

[638] Kenneth Stolarsky, NSA Problems Group, and O. P. Lossers. A characterization of a quadratic polynomial: 10848. American Mathematical Monthly, 109(1):80–81, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695783.

Bagby:2002:DMS

[639] Richard Bagby and M. Battaile. The determinant of a matrix of sines: 10849. American Mathematical Monthly, 109(1):81–82, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695784.

Lang:2002:MOC

[640] Wolfdieter Lang and Said Amghibech. On a matrix obtained from the Catalan generating function: 10850. American Mathematical Monthly, 109(1):82–83, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695785.

Mitrinovic:2002:SFL

[641] D. S. Mitrinović and Slavko Simic. Special functions from long ago: 5626. American Mathematical Monthly, 109(1):83–84, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695786.

Wagon:2002:BRB

[642] Stan Wagon. Book review: [untitled]. American Mathematical Monthly, 109(1):85–92, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695787.

Buell:2002:BRB

[643] Duncan Buell. Book review: [untitled]. American Mathematical Monthly, 109(1):93–98, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695788.

Ostebee:2002:TRa

[644] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 109(1):99–101, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695789.

Palka:2002:EEa

[645] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 109(1):102, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695790.

Anonymous:2002:BMa

[646] Anonymous. Back matter. American Mathematical Monthly, 109(1):??, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695791.

Anonymous:2002:FMa

[647] Anonymous. Front matter. American Mathematical Monthly, 109(1):??, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695762.

Robson:2002:WPN

[648] Eleanor Robson. Words and pictures: New light on plimpton 322. American Mathematical Monthly, 109(2):105–120, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695324.

Terras:2002:FQC

[649] Audrey Terras. Finite quantum chaos. American Mathematical Monthly, 109(2):121–139, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695325.

Abrams:2002:FTF

[650] Aaron Abrams and Robert Ghrist. Finding topology in a factory: Configuration spaces. American Mathematical Monthly, 109(2):140–150, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695326.

Baragar:2002:CUC

[651] Arthur Baragar. Constructions using a compass and twice-notched straightedge. American Mathematical Monthly, 109(2):151–164, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695327.

Morgan:2002:HER

[652] Frank Morgan and Roger Bolton. Hexagonal economic regions solve the location problem. American Mathematical Monthly, 109(2):165–172, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695328.

Kuzmanovich:2002:FGM

[653] James Kuzmanovich and Andrey Pavlichenkov. Finite groups of matrices whose entries are integers. American Mathematical Monthly, 109(2):173– 186, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695329.

Zheng:2002:EPT

[654] Liu Zheng. An elementary proof for two basic alternating series. American Mathematical Monthly, 109(2):187–188, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695330.

Agnew:2002:TAI

[655] Robert A. Agnew. On the TEAM approach to investing. American Mathematical Monthly, 109(2):188–192, February 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695331.

Huneke:2002:FT

[656] Craig Huneke. The friendship theorem. American Mathematical Monthly, 109(2):192–194, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695332.

Faure:2002:SPL

[657] Claude-Alain Faure. A short proof of Lebesgue’s density theorem. American Mathematical Monthly, 109(2):194–196, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695333.

Hofbauer:2002:SPR

[658] Josef Hofbauer. A simple proof of 1 + (1/2)^{2} + (1/3)^{2} + ⋯ = π^{2}/6 and related identities. American Mathematical Monthly, 109(2):196–200, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695334.

Bloom:2002:P

[659] David M. Bloom. Problem 10921. American Mathematical Monthly, 109 (2):201, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695335.

Khan:2002:P

[660] Mizan R. Khan. Problem 10922. American Mathematical Monthly, 109 (2):201, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695336.

Gray:2002:P

[661] Stephen B. Gray. Problem 10923. American Mathematical Monthly, 109 (2):201, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695337.

Sasane:2002:P

[662] A. J. Sasane. Problem 10924. American Mathematical Monthly, 109 (2):201, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695338.

Callan:2002:Pa

[663] David Callan. Problem 10925. American Mathematical Monthly, 109(2): 201–202, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695339.

Diamond:2002:P

[664] Harold Diamond. Problem 10926. American Mathematical Monthly, 109 (2):202, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695340.

Lagarias:2002:Pa

[665] Jeffrey C. Lagarias, E. M. Rains, and N. J. A. Sloane. Problem 10927. American Mathematical Monthly, 109(2):202, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695341.

Anglesio:2002:CFS

[666] Jean Anglesio and Reiner Martin. Continued fractions for some quadratic surds: 10773. American Mathematical Monthly, 109(2):202–203, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695342.

Zara:2002:HPF

[667] Catalin Zara, Zachary Franco, T. S. Michael, and Todor E. Milev. A highpowered Fibonacci identity: 10774. American Mathematical Monthly, 109(2):203–204, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695343.

Bateman:2002:GTI

[668] Paul Bateman, Dennis Eichhorn, and Robin Chapman. Grouping ktuples of integers by the sum of their squares: 10778. American Mathematical Monthly, 109(2):204–205, February 2002. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695344.

Chapman:2002:EEI

[669] Robin Chapman and M. A. Prasad. An expression for even-indexed Bernoulli numbers: 10789. American Mathematical Monthly, 109(2): 205–206, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695345.

See [891].

Luca:2002:NTA

[670] Florian Luca and Nasha Komanda. A number and its totient are not amicable: 10793. American Mathematical Monthly, 109(2):206–207, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695346.

Hough:2002:DGC

[671] David S. Hough, Rodica E. Simion, and Ali Ekber Gürel. Divisors of generalized Catalan numbers: 10794. American Mathematical Monthly, 109(2):207–208, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695347.

Popescu:2002:SRQ

[672] Călin Popescu, Roy Barbara, and Omran Kouba. A sum related to quadratic residues: 10852. American Mathematical Monthly, 109(2):208, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695348.

Ratcliff:2002:BRB

[673] Gail Ratcliff. Book review: [untitled]. American Mathematical Monthly, 109(2):209–211, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695349.

Anonymous:2002:BMc

[674] Anonymous. Back matter. American Mathematical Monthly, 109(2):??, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695350.

Anonymous:2002:FMc

[675] Anonymous. Front matter. American Mathematical Monthly, 109(2):??, February 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695323.

Johnson:2002:CHF

[676] Warren P. Johnson. The curious history of Faà di Bruno’s formula. American Mathematical Monthly, 109(3):217–234, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695352.

Pribitkin:2002:LIG

[677] Wladimir de Azevedo Pribitkin. Laplace’s integral, the gamma function, and beyond. American Mathematical Monthly, 109(3):235–245, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695353.

Northshield:2002:ASM

[678] Sam Northshield. Associativity of the secant method. American Mathematical Monthly, 109(3):246–257, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695354.

Johnson:2002:RSS

[679] Aimee S. A. Johnson and Kathleen Madden. Renewal systems, sharpeyed snakes, and shifts of finite type. American Mathematical Monthly, 109(3):258–272, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695355.

Johnson:2002:CHD

[680] Warren P. Johnson. Combinatorics of higher derivatives of inverses. American Mathematical Monthly, 109(3):273–277, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695356.

Radoux:2002:HDE

[681] Christian Radoux. The Hankel determinant of exponential polynomials: a very short proof and a new result concerning Euler numbers. American Mathematical Monthly, 109(3):277–278, March 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695357.

Dilcher:2002:DSU

[682] Karl Dilcher and Kurt Girstmair. Dedekind sums and uniform distributions. American Mathematical Monthly, 109(3):279–284, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695358.

Fokkink:2002:RNR

[683] Robbert Fokkink. R^{3} has no root. American Mathematical Monthly, 109(3):285, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695359.

Stillwell:2002:CP

[684] John Stillwell. The continuum problem. American Mathematical Monthly, 109(3):286–297, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695360.

Hillar:2002:P

[685] Christopher J. Hillar. Problem 10928. American Mathematical Monthly, 109(3):298, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695361.

Csete:2002:P

[686] Lajos Csete. Problem 10929. American Mathematical Monthly, 109(3): 298, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695362.

Schweitzer:2002:P

[687] Paul Schweitzer. Problem 10930. American Mathematical Monthly, 109 (3):298, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695363.

DiSario:2002:P

[688] Robert DiSario. Problem 10931. American Mathematical Monthly, 109 (3):298, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695364.

Bey:2002:P

[689] Christian Bey and Jerrold R. Griggs. Problem 10932. American Mathematical Monthly, 109(3):298–299, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695365.

Lescott:2002:P

[690] Paul Lescott. Problem 10933. American Mathematical Monthly, 109(3): 299, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695366.

Mihai:2002:P

[691] Vasile Mihai. Problem 10934. American Mathematical Monthly, 109(3): 299, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695367.

Salmasian:2002:MSN

[692] Hadi Salmasian and NSA Problems Group. Minimizing the sum of negative powers of orders of group elements: 10775. American Mathematical Monthly, 109(3):299, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695368.

Tian:2002:SMA

[693] Yongge Tian and David W. Carter. Squashing a matrix by adding an imaginary part: 10776. American Mathematical Monthly, 109(3):300, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695369. See errata

[892].

Smiley:2002:TCI

[694] Leonard Smiley, GCHQ Problem Solving Group, Helmut Prodinger, and Robin Chapman. A tree-counting identity: 10781. American Mathematical Monthly, 109(3):300–301, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695370.

Iannucci:2002:DIR

[695] Douglas Iannucci, O. P. Lossers, and GCHQ Problem Solving Group. Divisibility involving random sets of integers: 10782. American Mathematical Monthly, 109(3):301–302, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695371.

Deutsch:2002:RDS

[696] Emeric Deutsch, John Henry Steelman, and David Callan. A recursively defined sequence of functions: 10785. American Mathematical Monthly, 109(3):302–304, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695372.

Iannucci:2002:TN

[697] Douglas Iannucci and Kee-Wai Lau. Triperfect numbers: 10800. American Mathematical Monthly, 109(3):304–305, March 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695373. See errata [893].

Chao:2002:MMC

[698] Wu Wei Chao and Geoffrey A. Kandall. The median meets the circle of appolonius: 10862. American Mathematical Monthly, 109(3):305, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695374.

Grabiner:2002:BRB

[699] Judith V. Grabiner. Book review: [untitled]. American Mathematical Monthly, 109(3):306–309, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695375.

Ostebee:2002:TRb

[700] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 109(3):310–312, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695376.

Palka:2002:EEb

[701] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 109 (3):313–314, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695377.

Anonymous:2002:BMd

[702] Anonymous. Back matter. American Mathematical Monthly, 109(3):??, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695378.

Anonymous:2002:FMd

[703] Anonymous. Front matter. American Mathematical Monthly, 109(3):??, March 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695351.

Taylor:2002:MVS

[704] Alan D. Taylor. The manipulability of voting systems. American Mathematical Monthly, 109(4):321–337, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695497.

Lagarias:2002:BDC

[705] Jeffrey C. Lagarias, Colin L. Mallows, and Allan R. Wilks. Beyond the Descartes Circle Theorem. American Mathematical Monthly, 109(4):338– 361, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695498.

Gueron:2002:TAG

[706] Shay Gueron. Two applications of the Generalized Ptolemy Theorem. American Mathematical Monthly, 109(4):362–370, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695499.

Aggarwala:2002:IPM

[707] Rita Aggarwala and Michael P. Lamoureux. Inverting the Pascal matrix plus one. American Mathematical Monthly, 109(4):371–377, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695500.

Wen:2002:NDC

[708] Liu Wen. A nowhere differentiable continuous function constructed by infinite products. American Mathematical Monthly, 109(4):378–380, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695501.

Huisman:2002:RFA

[709] Johannes Huisman. Rino’s flight around the world in circles. American Mathematical Monthly, 109(4):380–383, April 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695502.

Mann:2002:TSN

[710] Casey Mann. A tile with surround number 2. American Mathematical Monthly, 109(4):383–388, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695503.

Michel:2002:SF

[711] Reinhard Michel. On Stirling’s formula. American Mathematical Monthly, 109(4):388–390, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695504.

Vietri:2002:IPK

[712] Andrea Vietri. An inductive proof of a known result about ψ(x). American Mathematical Monthly, 109(4):390–391, April 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695505.

Chen:2002:Pa

[713] Zhibing Chen. Problem 10935. American Mathematical Monthly, 109(4): 392, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695506.

Lukacs:2002:P

[714] Andor Lukács and Szilard András. Problem 10936. American Mathematical Monthly, 109(4):392, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695507.

Wagner:2002:P

[715] David G. Wagner. Problem 10937. American Mathematical Monthly, 109(4):392, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695508.

Grivaux:2002:Pa

[716] Jean-Pierre Grivaux. Problem 10938. American Mathematical Monthly, 109(4):392, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695509.

Ivic:2002:P

[717] Aleksandar Ivic. Problem 10939. American Mathematical Monthly, 109 (4):392, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695510.

Nievergelt:2002:P

[718] Yves Nievergelt. Problem 10940. American Mathematical Monthly, 109 (4):393, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695511.

Lagarias:2002:Pb

[719] Jeffrey C. Lagarias. Problem 10941. American Mathematical Monthly, 109(4):393, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695512.

Bateman:2002:LIL

[720] Paul Bateman, Jeffrey Kalb, and Allen Stenger. A limit involving least common multiples: 10797. American Mathematical Monthly, 109(4):393– 394, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695513.

Pudaite:2002:GUH

[721] Paul R. Pudaite, Owen Byer, and Deirdre Smeltzer. A gambler urns his money: 10801. American Mathematical Monthly, 109(4):394–395, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695514.

Penrice:2002:ERN

[722] Stephen Penrice and GCHQ Problem Solving Group. Expected rank in an NBA draft lottery: 10803. American Mathematical Monthly, 109 (4):395–396, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695515.

vanLamoen:2002:CC

[723] Floor van Lamoen. Circumcenters on a circle: 10830. American Mathematical Monthly, 109(4):396–397, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695516.

Parvanescu:2002:WAE

[724] Florian S. Pârvânescu and Roy Barbara. When ABC is equilateral: 10838. American Mathematical Monthly, 109(4):397–398, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695517.

Yang:2002:EUC

[725] Jiansheng Yang, Shulin Zhou, and O. P. Lossers. An elaborate uniform convergence: 10840. American Mathematical Monthly, 109(4):398–399, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695518.

Logan:2002:BRB

[726] J. David Logan. Book review: [untitled]. American Mathematical Monthly, 109(4):400–402, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695519.

Anonymous:2002:BMe

[727] Anonymous. Back matter. American Mathematical Monthly, 109(4): ??, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695520.

Anonymous:2002:FMe

[728] Anonymous. Front matter. American Mathematical Monthly, 109(4): ??, April 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695496.

Cantarella:2002:VCT

[729] Jason Cantarella, Dennis DeTurck, and Herman Gluck. Vector calculus and the topology of domains in 3-space. American Mathematical Monthly, 109(5):409–442, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695643.

Gueron:2002:FSP

[730] Shay Gueron and Ran Tessler. The Fermat–Steiner problem. American Mathematical Monthly, 109(5):443–451, May 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695644.

Murty:2002:PNI

[731] M. Ram Murty. Prime numbers and irreducible polynomials. American Mathematical Monthly, 109(5):452–458, May 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695645.

vanderPoorten:2002:SPC

[732] Alf van der Poorten and Gerhard Woeginger. Squares from products of consecutive integers. American Mathematical Monthly, 109(5):459– 462, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695646.

Brooks:2002:CP

[733] Jeff Brooks and Satha Push. The cycloidal pendulum. American Mathematical Monthly, 109(5):463–465, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695647.

Viana:2002:GTR

[734] Paulo Viana and Paula Murgel Veloso. Galois theory of reciprocal polynomials. American Mathematical Monthly, 109(5):466–471, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695648.

Bhargava:2002:WGU

[735] Mira Bhargava. When is a group the union of proper normal subgroups? American Mathematical Monthly, 109(5):471–473, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695649.

Bachelis:2002:SPH

[736] Gregory F. Bachelis. A short proof of Hall’s theorem on SDRs. American Mathematical Monthly, 109(5):473–474, May 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2695650.

Janous:2002:P

[737] Walther Janous. Problem 10942. American Mathematical Monthly, 109 (5):475, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695651.

Popescu:2002:P

[738] Călin Popescu. Problem 10943. American Mathematical Monthly, 109(5): 475, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695652.

Mazur:2002:P

[739] Marcin Mazur. Problem 10944. American Mathematical Monthly, 109(5): 475, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695653.

Simic:2002:P

[740] Slavko Simic. Problem 10945. American Mathematical Monthly, 109 (5):475–476, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695654.

Stevic:2002:P

[741] Stevo Stević. Problem 10946. American Mathematical Monthly, 109(5): 476, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695655.

LPSG:2002:P

[742] LSU Problem Solving Group. Problem 10947. American Mathematical Monthly, 109(5):476, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695656.

Walker:2002:P

[743] Peter Walker. Problem 10948. American Mathematical Monthly, 109(5): 476, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695657. See errata

[888].

Fekete:2002:SBB

[744] Antal Fekete and Johann Cigler. Stirling and Bell and binomial, oh my!: 10805. American Mathematical Monthly, 109(5):476–477, May 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695658.

Beckwith:2002:IPI

[745] David Beckwith and Karl David. An identity for partitions of integers: 10809. American Mathematical Monthly, 109(5):477–478, May 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695659.

Tracy:2002:DBE

[746] Phil Tracy and NCCU Problems Group. The distance between the endpoints of a longest path: 10811. American Mathematical Monthly, 109(5): 478, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695660.

Deutsch:2002:SMP

[747] Emeric Deutsch and Wen jin Woan. Special Motzkin paths: 10816. American Mathematical Monthly, 109(5):478–479, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695661.

Joita:2002:FEF

[748] Cezar Joita and Robin Chapman. The functional equation f ∘ g = f: 10818. American Mathematical Monthly, 109(5):479–480, May 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2695662.

Krafft:2002:BCE

[749] Olaf Krafft, Gabriela Stănică, and Pantelimon Stănică. A binomial coefficient estimate: 10819. American Mathematical Monthly, 109(5):480–481, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695663.

Lee:2002:BAD

[750] Ho joo Lee and Marc A. P. Bernstein. Brocard angle 30 degrees: 10824. American Mathematical Monthly, 109(5):481–482, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695664.

Abel:2002:RG

[751] Ulrich Abel and Douglas B. Tyler. Reminiscent of gibbs: 10827. American Mathematical Monthly, 109(5):482, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695665.

Cohen:2002:BRBb

[752] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 109(5):483–485, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695666.

Ostebee:2002:TRc

[753] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 109(5):486–488, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695667.

Anonymous:2002:BMf

[754] Anonymous. Back matter. American Mathematical Monthly, 109(5): ??, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695668.

Anonymous:2002:FMf

[755] Anonymous. Front matter. American Mathematical Monthly, 109(5): ??, May 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695642.

Marchisotto:2002:TPB

[756] Elena Anne Marchisotto. The theorem of pappus: a bridge between algebra and geometry. American Mathematical Monthly, 109(6):497–516, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695440.

Finch:2002:CCB

[757] Carrie E. Finch and Lenny Jones. A curious connection between Fermat numbers and finite groups. American Mathematical Monthly, 109(6): 517–524, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695441.

Apostol:2002:SAV

[758] Tom M. Apostol and Mamikon A. Mnatsakanian. Subtangents – an aid to visual calculus. American Mathematical Monthly, 109(6):525–533, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695442.

Lagarias:2002:EPE

[759] Jeffrey C. Lagarias. An elementary problem equivalent to the Riemann hypothesis. American Mathematical Monthly, 109(6):534–543, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695443.

Gillman:2002:TCS

[760] Leonard Gillman. Two classical surprises concerning the axiom of choice and the continuum hypothesis. American Mathematical Monthly, 109(6): 544–553, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695444.

Hardy:2002:MPP

[761] G. E. Hardy and M. V. Subbarao. A modified problem of pillai and some related questions. American Mathematical Monthly, 109(6):554– 559, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695445.

Nyblom:2002:SCS

[762] M. A. Nyblom. Some curious sequences involving floor and ceiling functions. American Mathematical Monthly, 109(6):559–564, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/2695446.

Feeman:2002:RHD

[763] Timothy G. Feeman and Osvaldo Marrero. The right-hand derivative of an integral. American Mathematical Monthly, 109(6):565–568, June/ July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695447.

Lagarias:2002:Pc

[764] Jeffrey Lagarias. Problem 10949. American Mathematical Monthly, 109 (6):569, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695448.

Furdui:2002:Pa

[765] Ovidiu Furdui. Problem 10950. American Mathematical Monthly, 109 (6):569, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695449.

Kim:2002:P

[766] Sung Soo Kim. Problem 10951. American Mathematical Monthly, 109 (6):569, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695450.

Callan:2002:Pb

[767] David Callan. Problem 10952. American Mathematical Monthly, 109 (6):569, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695451.

Quet:2002:P

[768] Leroy Quet. Problem 10953. American Mathematical Monthly, 109(6): 570, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695452. See

errata [890].

Just:2002:P

[769] Erwin Just and Norman Schaumberger. Problem 10954. American Mathematical Monthly, 109(6):570, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695453.

Satnoianu:2002:Pa

[770] Razvan Satnoianu. Problem 10955. American Mathematical Monthly, 109 (6):570, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695454.

Caragea:2002:MPP

[771] Doru Caragea, Viviana Ene, Dean Alvis, and Nasha Komanda. A millennial polynomial problem: 10802. American Mathematical Monthly, 109(6):570–571, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

2695455.

Richman:2002:SVS

[772] Fred Richman and NCCU Problems Group. Subspaces of the vector space of 2-by-2 matrices: 10813. American Mathematical Monthly, 109(6): 571–572, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/2695456.

Stevic:2002:LS

[773] Stevo Stević and Alain Tissier. The limit of a sequence: 10853. American Mathematical Monthly, 109(6):572–573, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/2695457.

Diamond:2002:TSR

[774] Harold Diamond and Alain Tissier. Truncated series ratios alternating about tanh x: 10857. American Mathematical Monthly, 109(6):573–574, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695458.

Popescu:2002:CSC

[775] Călin Popescu, Xiaoshen Wang, and Thomas McMillan. Compact subsets countable: 10864. American Mathematical Monthly, 109(6):574, June/ July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695459.

Ramirez:2002:SIS

[776] Oscar Ciaurri Ramirez and Daniele Donini. Sum of an infinite series: 10867. American Mathematical Monthly, 109(6):575–576, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695460.

LeBrun:2002:BRB

[777] Claude LeBrun. Book review: [untitled]. American Mathematical Monthly, 109(6):577–578, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695461.

Swift:2002:BRB

[778] George H. Swift. Book review: [untitled]. American Mathematical Monthly, 109(6):579–580, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695462.

Anonymous:2002:U

[779] Anonymous. [untitled]. American Mathematical Monthly, 109(6):581– 583, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/2695463.

Anonymous:2002:BMg

[780] Anonymous. Back matter. American Mathematical Monthly, 109(6):??, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695464.

Anonymous:2002:FMg

[781] Anonymous. Front matter. American Mathematical Monthly, 109(6):??, June/July 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2695439.

Witte:2002:JRC

[782] Bob Witte. The 2001 James R. C. Leitzel Lecture: What I have learned from the mathematics community. American Mathematical Monthly, 109(7):585–594, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072424.

Brown:2002:IFP

[783] Robert F. Brown and Jack E. Girolo. Isolating fixed points. American Mathematical Monthly, 109(7):595–611, August/September 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072425.

Smith:2002:TE

[784] James T. Smith. Thomsen’s equation. American Mathematical Monthly, 109(7):612–622, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072426.

McCleary:2002:T

[785] John McCleary. Trigonometries. American Mathematical Monthly, 109 (7):623–638, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072427.

Brown:2002:ECM

[786] Ezra Brown and Bruce T. Myers. Elliptic curves from Mordell to Diophantus and back. American Mathematical Monthly, 109(7):639–649, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072428.

Lamphere:2002:SDP

[787] Robert L. Lamphere. Solution of the direct problem of uniform circular motion in non-Euclidean geometry. American Mathematical Monthly, 109(7):650–655, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072429.

Jamieson:2002:FNC

[788] M. J. Jamieson. Fibonacci numbers and cotangent sequences. American Mathematical Monthly, 109(7):655–657, August/September 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072430.

Pinelis:2002:EPT

[789] Iosif Pinelis. An equidistribution phenomenon: Is there a principle behind it? American Mathematical Monthly, 109(7):657–659, August/ September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072431.

Walsh:2002:QK

[790] P. G. Walsh. On a question of Kaplansky. American Mathematical Monthly, 109(7):660–661, August/September 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072432.

Holdener:2002:TTF

[791] Judy A. Holdener. A theorem of Touchard on the form of odd perfect numbers. American Mathematical Monthly, 109(7):661–663, August/ September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072433.

McCarthy:2002:P

[792] John McCarthy. Problem 10956. American Mathematical Monthly, 109 (7):664, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072434.

Alexandrov:2002:P

[793] Victor Alexandrov. Problem 10957. American Mathematical Monthly, 109(7):664, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072435.

Chapman:2002:P

[794] Robin Chapman. Problem 10958. American Mathematical Monthly, 109 (7):664, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072436.

Lamb:2002:P

[795] George L. Lamb. Problem 10959. American Mathematical Monthly, 109 (7):665, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072437.

Eng:2002:P

[796] Oliver Eng. Problem 10960. American Mathematical Monthly, 109(7): 665, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072438.

Conrad:2002:P

[797] Brian Conrad. Problem 10961. American Mathematical Monthly, 109 (7):665, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072439.

Bencze:2002:P

[798] Mihàly Bencze. Problem 10962. American Mathematical Monthly, 109 (7):665, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072440.

Facchini:2002:NSM

[799] Alberto Facchini, Francesco Barioli, Robin Chapman, and Ron Martin Carroll. Nonsingular sums of matrices: 10784. American Mathematical Monthly, 109(7):665–666, August/September 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072441.

Andrews:2002:SD

[800] George E. Andrews, Robin Chapman, and George Andrews. A sparse determinant: 10831. American Mathematical Monthly, 109(7):666–667, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072442.

Kim:2002:MPMa

[801] Sung Soo Kim, Richard Johnsonbaugh, Ronald E. Prather, and Donald Knuth. Min-plus matrix multiplication: 10834. American Mathematical Monthly, 109(7):667–668, August/September 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072443.

Cossi:2002:PS

[802] Ernesto Bruno Cossi and Robin Chapman. Points on a sphere: 10855. American Mathematical Monthly, 109(7):668–669, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072444.

Kazdan:2002:WB

[803] Jerry Kazdan, Herbert Wilf, and Darryl K. Nester. Waiting for the bus: 10866. American Mathematical Monthly, 109(7):669, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072445.

Chao:2002:PC

[804] Wu Wei Chao and Li Zhou. Parallelogram circumcenters: 10874. American Mathematical Monthly, 109(7):669–670, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072446.

Just:2002:LC

[805] Erwin Just, Norman Schaumberger, and Paul Budney. A little convexity: 10882. American Mathematical Monthly, 109(7):670, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072447.

Ahmed:2002:DI

[806] Zafar Ahmed, Knut Dale, and George L. Lamb Jr. Definitely an integral: 10884. American Mathematical Monthly, 109(7):670–671, August/ September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072448.

Nunemacher:2002:BRB

[807] Jeffrey Nunemacher. Book review: [untitled]. American Mathematical Monthly, 109(7):672–675, August/September 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072449.

Green:2002:BRB

[808] Mark Green. Book review: [untitled]. American Mathematical Monthly, 109(7):675–678, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072450.

Anonymous:2002:TRa

[809] Anonymous. Telegraphic reviews. American Mathematical Monthly, 109(7):679–681, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072451.

Anonymous:2002:BMh

[810] Anonymous. Back matter. American Mathematical Monthly, 109(7):??, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072452.

Anonymous:2002:FMh

[811] Anonymous. Front matter. American Mathematical Monthly, 109(7):??, August/September 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072423.

Bailey:2002:ILD

[812] Scott Bailey, Theodore Kim, and Robert S. Strichartz. Inside the Lévy Dragon. American Mathematical Monthly, 109(8):689–703, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072395.

Philippakis:2002:OTK

[813] Anthony Philippakis. The orthic triangle and the O.K. quadrilateral. American Mathematical Monthly, 109(8):704–728, October 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072396.

Villegas:2002:BMP

[814] F. Rodriguez Villegas, L. Sadun, and J. F. Voloch. Blet: a mathematical puzzle. American Mathematical Monthly, 109(8):729–740, October 2002.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072397.

Levene:2002:KCR

[815] Mark Levene and George Loizou. Kemeny’s constant and the random surfer. American Mathematical Monthly, 109(8):741–745, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072398.

Velleman:2002:PFB

[816] Daniel J. Velleman. Partial fractions, binomial coefficients, and the integral of an odd power of sec θ. American Mathematical Monthly, 109(8): 746–749, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072399.

Manne:2002:SBW

[817] Per E. Manne and Steven R. Finch. A solution to the bent wire problem. American Mathematical Monthly, 109(8):750–752, October 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072400.

Liseo:2002:PDD

[818] B. Liseo. A probabilistic derivation of Dirichlet integrals. American Mathematical Monthly, 109(8):753–754, October 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072401.

Sanchez-Peregrino:2002:NCF

[819] Roberto Sánchez-Peregrino. A note on a closed formula for poly-Bernoulli numbers. American Mathematical Monthly, 109(8):755–756, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072402.

Parks:2002:ECD

[820] Harold R. Parks and Dean C. Wills. An elementary calculation of the dihedral angle of the regular n-simplex. American Mathematical Monthly, 109(8):756–758, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072403.

Deiermann:2002:P

[821] Paul Deiermann. Problem 10963. American Mathematical Monthly, 109 (8):759, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072404.

ShahAli:2002:P

[822] H. A. ShahAli. Problem 10964. American Mathematical Monthly, 109 (8):759, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072405.

Melville:2002:P

[823] John Melville. Problem 10965. American Mathematical Monthly, 109 (8):759, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072406.

deKoninck:2002:P

[824] Jean-Marie de Koninck. Problem 10966. American Mathematical Monthly, 109(8):759, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072407.

West:2002:P

[825] Douglas B. West. Problem 10967. American Mathematical Monthly, 109 (8):759, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072408.

Locke:2002:P

[826] Stephen C. Locke. Problem 10968. American Mathematical Monthly, 109 (8):759–760, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072409.

Stanley:2002:P

[827] Richard Stanley. Problem 10969. American Mathematical Monthly, 109 (8):760, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072410.

Furdui:2002:Pb

[828] Ovidiu Furdui. Problem 10950. American Mathematical Monthly, 109 (8):760, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072411.

See errata [889].

Mirzavaziri:2002:LNE

[829] M. Mirzavaziri and GCHQ Problems Group. Least number with exactly m divisors: 10820. American Mathematical Monthly, 109(8):760–761, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072412.

Andrews:2002:PUS

[830] George E. Andrews, Karl David, S. J. Bernau, M. Getz, and D. Jones. Periodicity under shift and complement: 10823. American Mathematical Monthly, 109(8):761–762, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072413.

Miller:2002:FLE

[831] Carl Miller and Donald E. Knuth. A Fibonacci–Lucas extremum: 10825. American Mathematical Monthly, 109(8):762–763, October 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072414.

Janous:2002:TS

[832] Walther Janous, Rolf Richberg, and Raul A. Simon. A totient sum: 10829. American Mathematical Monthly, 109(8):763, October 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072415.

Kim:2002:MPMb

[833] Sung Soo Kim, Richard Johnsonbaugh, Ronald E. Prather, and Donald Knuth. Min-plus matrix multiplication: 10834. American Mathematical Monthly, 109(8):763–764, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072416.

Satnoianu:2002:MM

[834] Razvan A. Satnoianu and O. P. Lossers. The minimum median: 10880. American Mathematical Monthly, 109(8):764–765, October 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072417.

Huber:2002:CL

[835] Klaus Huber, O. P. Lossers, Norbert Ortner, and A. Serret. Cassini’s length: 10889. American Mathematical Monthly, 109(8):765–766, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072418.

Hobbs:2002:BRB

[836] Arthur M. Hobbs. Book review: [untitled]. American Mathematical Monthly, 109(8):767–770, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072419.

Bartle:2002:BRB

[837] Robert G. Bartle. Book review: [untitled]. American Mathematical Monthly, 109(8):770–773, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072420.

Palka:2002:EEc

[838] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 109 (8):774–776, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072421.

Anonymous:2002:BMi

[839] Anonymous. Back matter. American Mathematical Monthly, 109(8):??, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072422.

Anonymous:2002:FMi

[840] Anonymous. Front matter. American Mathematical Monthly, 109(8):??, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072394.

Daepp:2002:EFB

[841] Ulrich Daepp, Pamela Gorkin, and Raymond Mortini. Ellipses and finite Blaschke products. American Mathematical Monthly, 109(9):785–795, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072367.

Isaksen:2002:CVE

[842] Daniel C. Isaksen. A cohomological viewpoint on elementary school arithmetic. American Mathematical Monthly, 109(9):796–805, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072368.

Eisenberg:2002:FTC

[843] Bennett Eisenberg and Rosemary Sullivan. The fundamental theorem of calculus in two dimensions. American Mathematical Monthly, 109(9):

806–817, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072369.

Praton:2002:CPE

[844] Iwan Praton. Cutting polyominos into equal-area triangles. American Mathematical Monthly, 109(9):818–826, November 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072370.

Klosinski:2002:SSW

[845] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Second William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 109(9):827–839, November 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072371.

Hsu:2002:TIP

[846] Chun-Hung Hsu and Ming-Chia Li. Transitivity implies period six: a simple proof. American Mathematical Monthly, 109(9):840–843, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072372.

Wildberger:2002:NPC

[847] N. J. Wildberger. A new proof of Cavalieri’s quadrature formula. American Mathematical Monthly, 109(9):843–845, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072373.

Yingying:2002:ECC

[848] Li Yingying. On Euler’s constant-calculating sums by integrals. American Mathematical Monthly, 109(9):845–850, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072374.

vanderPoorten:2002:LE

[849] Alf van der Poorten. Life on the edge. American Mathematical Monthly, 109(9):850–853, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072375.

Satnoianu:2002:Pb

[850] Razvan Satnoianu. Problem 10970. American Mathematical Monthly, 109 (9):854, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072376.

Rupert:2002:Pb

[851] C. P. Rupert. Problem 10971. American Mathematical Monthly, 109 (9):854, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072377.

Ferguson:2002:P

[852] Tim Ferguson and Lenny Jones. Problem 10972. American Mathematical Monthly, 109(9):854, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072378.

Servi:2002:P

[853] L. D. Servi. Problem 10973. American Mathematical Monthly, 109(9): 854, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072379.

Marivani:2002:P

[854] Syrous Marivani. Problem 10974. American Mathematical Monthly, 109(9):854–855, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072380.

Deutsch:2002:P

[855] Emeric Deutsch. Problem 10975. American Mathematical Monthly, 109 (9):855, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072381.

Voloshin:2002:P

[856] Vitaly Voloshin. Problem 10976. American Mathematical Monthly, 109 (9):855, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072382.

Taylor:2002:CWT

[857] Howard M. Taylor, Richard A. Groeneveld, John Stufken, J. G. Kemeny, J. L. Snell, and A. W. Knapp. A careless walk toward zero: 10788. American Mathematical Monthly, 109(9):855–856, November 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072383.

Beckwith:2002:AA

[858] David Beckwith and NCCU Problems Group. Additive atoms: 10828. American Mathematical Monthly, 109(9):856–857, November 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072384.

Dyubina:2002:GFG

[859] Anna Dyubina and Stephen M. Gagola Jr. A group that is not finitely generated: 10835. American Mathematical Monthly, 109(9):857–858, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072385.

Just:2002:CRR

[860] Erwin Just and Charles Lanski. Consequences of a relation on a ring: 10841. American Mathematical Monthly, 109(9):858, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072386.

Reznick:2002:NBC

[861] Bruce Reznick, GCHQ Problems Group, and Li Zhou. Nondivisibility of binomial coefficients: 10842. American Mathematical Monthly, 109(9): 858–859, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072387.

Beckwith:2002:PBS

[862] David Beckwith, Li Zhou, and David Callan. A property of the binary sum of a family of sets: 10865. American Mathematical Monthly, 109(9): 859–860, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072388.

Martini:2002:IH

[863] R. Martini, G. F. Post, and Gerald Leibowitz. An inequality of hardy: 10899. American Mathematical Monthly, 109(9):860, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072389.

Furdui:2002:SP

[864] Ovidio Furdui and Albert Nijenhuis. A square property: 10904. American Mathematical Monthly, 109(9):860–861, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072390.

Richman:2002:BRB

[865] Fred Richman. Book review: [untitled]. American Mathematical Monthly, 109(9):862–865, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072391.

Anonymous:2002:TRb

[866] Anonymous. Telegraphic reviews. American Mathematical Monthly, 109(9):866–868, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072392.

Anonymous:2002:BMj

[867] Anonymous. Back matter. American Mathematical Monthly, 109(9):??, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072393.

Anonymous:2002:FMj

[868] Anonymous. Front matter. American Mathematical Monthly, 109(9):??, November 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072366.

Shell-Gellasch:2002:MRF

[869] Amy Shell-Gellasch. Mina Rees and the funding of the mathematical sciences. American Mathematical Monthly, 109(10):873–889, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072455.

Levi:2002:MPT

[870] Mark Levi. Minimal perimeter triangles. American Mathematical Monthly, 109(10):890–899, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072456.

Apostol:2002:TSU

[871] Tom M. Apostol and Mamikon A. Mnatsakanian. Tangents and subtangents used to calculate areas. American Mathematical Monthly, 109(10): 900–908, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072457.

Jerrard:2002:ETT

[872] Richard P. Jerrard and John E. Wetzel. Equilateral triangles and triangles. American Mathematical Monthly, 109(10):909–915, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072458.

Benito:2002:EPH

[873] Manuel Benito and J. Javier Escribano. An easy proof of Hurwitz’s theorem. American Mathematical Monthly, 109(10):916–918, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072459.

Greene:2002:NSP

[874] Joshua E. Greene. A new short proof of Kneser’s conjecture. American Mathematical Monthly, 109(10):918–920, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072460.

Axness:2002:P

[875] Carl Axness. Problem 10977. American Mathematical Monthly, 109(10): 921, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072461.

Grivaux:2002:Pb

[876] Jean-Pierre Grivaux. Problem 10978. American Mathematical Monthly, 109(10):921, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072462.

Minassian:2002:P

[877] Donald P. Minassian. Problem 10979. American Mathematical Monthly, 109(10):921, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072463.

Hao:2002:P

[878] Ye Zhong Hao. Problem 10980. American Mathematical Monthly, 109 (10):921, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072464.

Singer:2002:P

[879] Nicholas C. Singer. Problem 10981. American Mathematical Monthly, 109(10):922, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072465.

Chen:2002:Pb

[880] Yuming Chen. Problem 10982. American Mathematical Monthly, 109 (10):922, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072466.

Suceava:2002:P

[881] Bogdan Suceavă. Problem 10983. American Mathematical Monthly, 109 (10):922, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072467.

Parlett:2002:LEP

[882] Beresford N. Parlett, Olaf Krafft, and Martin Schaefer. The largest eigenvalue of a positive definite symmetric matrix: 10839. American Mathematical Monthly, 109(10):922–923, December 2002. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072468.

Kim:2002:IVF

[883] Sung Soo Kim, Charles Diminnie, and Roger Zarnowski. Intermediate value functions: 10861. American Mathematical Monthly, 109(10):923, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072469.

Andras:2002:AMG

[884] Szilárd András and Michael Reid. Arithmetic means and geometric means: 10907. American Mathematical Monthly, 109(10):924–925, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072470.

Chao:2002:FE

[885] Wu Wei Chao, Richard Stong, and Doyle Henderson. A functional equation: 10908. American Mathematical Monthly, 109(10):925–926, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072471.

Kalman:2002:IPT

[886] Támás Kalmán and O. P. Lossers. Isodynamic points of a triangle: 10909. American Mathematical Monthly, 109(10):926–927, December 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072472.

Stevic:2002:X

[887] Stevo Stević, Rade Todorović, and Peter M. Jarvis. nx mod 1: 10910. American Mathematical Monthly, 109(10):927–928, December 2002. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3072473.

Anonymous:2002:Ea

[888] Anonymous. Errata: 10948. American Mathematical Monthly, 109(10): 928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072474. See

[743].

Anonymous:2002:Eb

[889] Anonymous. Errata: 10950. American Mathematical Monthly, 109(10): 928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072475. See

[828].

Anonymous:2002:Ec

[890] Anonymous. Errata: 10953. American Mathematical Monthly, 109(10): 928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072476. See

[768].

Anonymous:2002:EEE

[891] Anonymous. Errata: An expression for even-indexed Bernoulli numbers: 10789. American Mathematical Monthly, 109(10):928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072478. See [669].

Anonymous:2002:ESM

[892] Anonymous. Errata: Squashing a matrix by adding an imaginary part: 10776. American Mathematical Monthly, 109(10):928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072477. See [693].

Anonymous:2002:ETN

[893] Anonymous. Errata: Triperfect numbers: 10800. American Mathematical Monthly, 109(10):928, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072479. See [697].

Cohen:2002:BRBa

[894] Marion Cohen. Book review: [untitled]. American Mathematical Monthly, 109(10):929–932, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072480.

Anonymous:2002:VI

[895] Anonymous. Volume information. American Mathematical Monthly, 109 (10):933–943, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072453.

Anonymous:2002:BMb

[896] Anonymous. Back matter. American Mathematical Monthly, 109(10): 944–945, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072481.

Anonymous:2002:FMb

[897] Anonymous. Front matter. American Mathematical Monthly, 109(10):??, December 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072454.

Thiele:2003:HTF

[898] Rüdger Thiele. Hilbert’s twenty-fourth problem. American Mathematical Monthly, 110(1):1–24, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072340.

Iba:2003:CS

[899] Glenn Iba and James Tanton. Candy sharing. American Mathematical Monthly, 110(1):25–35, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072341.

McCarthy:2003:PTW

[900] John E. McCarthy. Pick’s Theorem — what’s the big deal? American Mathematical Monthly, 110(1):36–45, January 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072342.

Luca:2003:FPH

[901] Florian Luca. Fermat primes and heron triangles with prime power sides. American Mathematical Monthly, 110(1):46–49, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072343.

Betcher:2003:ERS

[902] Jennifer T. Betcher and John H. Jaroma. An extension of the results of servais and Cramer on odd perfect and odd multiply perfect numbers. American Mathematical Monthly, 110(1):49–52, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072344.

Nash:2003:GPL

[903] Alan Nash. A generalized parallelogram law. American Mathematical Monthly, 110(1):52–57, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072345.

Elias:2003:RCC

[904] Uri Elias. Rearrangement of a conditionally convergent series. American Mathematical Monthly, 110(1):57, January 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/3072346.

Tian:2003:Pa

[905] Yongge Tian. Problem 10984. American Mathematical Monthly, 110(1): 58, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072347.

Knuth:2003:Pa

[906] Donald Knuth. Problem 10985. American Mathematical Monthly, 110 (1):58, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072348.

Wilf:2003:P

[907] Herbert Wilf. Problem 10986. American Mathematical Monthly, 110(1): 58, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072349.

Prasanna:2003:P

[908] Kumar N. Prasanna and B. G. Shenoy. Problem 10987. American Mathematical Monthly, 110(1):58, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072350.

Chao:2003:Pa

[909] Wu Wei Chao. Problem 10988. American Mathematical Monthly, 110(1): 59, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072351.

Bacher:2003:P

[910] Roland Bacher. Problem 10989. American Mathematical Monthly, 110 (1):59, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072352.

Mabry:2003:Pa

[911] Rick Mabry. Problem 10990. American Mathematical Monthly, 110(1): 59, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3072353.

Knuth:2003:HVL

[912] Donald E. Knuth and Allan Pedersen. Highly variable lists: 10691. American Mathematical Monthly, 110(1):59–60, January 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072354.

Knuth:2003:EAB

[913] Donald E. Knuth and O. P. Lossers. Exploring all binary mazes: 10720. American Mathematical Monthly, 110(1):60–61, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072355.

Deutsch:2003:TRS

[914] Emeric Deutsch, Doyle Henderson, and GCHQ Problem Group. Three recurrences for a single sequence: 10860. American Mathematical Monthly, 110(1):61–62, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3072356.

Mortini:2003:IT

[915] Raymond Mortini and Jeff Boersema. Inequivalent topologies: 10890. American Mathematical Monthly, 110(1):62–63, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3072357.

Wendijk:2003:TSA

[916] August Wendijk and Thomas Hermann. Twin segments in the arbelos: 10895. American Mathematical Monthly, 110(1):63–64, January 2003.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3072358.

Knight:2003:HA

[917] M. J. Knight and Li Zhou. Hidden arctangents: 10901. American Mathematical Monthly, 110(1):64, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072359.

Becker:2003:ME

[918] Michael Becker and Rolf Richberg. A moment evaluation: 10919. American Mathematical Monthly, 110(1):64–65, January 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3072360.

Pambuccian:2003:BRB

[919] Victor Pambuccian. Book review: Geometry: Euclid and Beyond, by Robin Hartshorne. American Mathematical Monthly, 110(1):66–70, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3072361.

Fenster:2003:BRB

[920] Della D. Fenster. Book review: [untitled]. American Mathematical Monthly, 110(1):70–74, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3072362.

Anonymous:2003:TRa

[921] Anonymous. Telegraphic reviews. American Mathematical Monthly, 110 (1):75–77, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072363.

Palka:2003:EEa

[922] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 110 (1):78–80, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3072364.

Anonymous:2003:BMa

[923] Anonymous. Back matter. American Mathematical Monthly, 110(1):??, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072365.

Anonymous:2003:FMa

[924] Anonymous. Front matter. American Mathematical Monthly, 110(1):??, January 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072339.

Range:2003:CAB

[925] R. Michael Range. Complex analysis: a brief tour into higher dimensions. American Mathematical Monthly, 110(2):89–108, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647769.

Schoen:2003:RCC

[926] Harold L. Schoen and Christian R. Hirsch. Responding to calls for change in high school mathematics: Implications for collegiate mathematics. American Mathematical Monthly, 110(2):109–123, February 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647770.

Hunter:2003:HRS

[927] David J. Hunter and Paul T. von Hippel. How rare is symmetry in musical 12-tone rows? American Mathematical Monthly, 110(2):124–132, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647771.

May:2003:MPN

[928] J. Peter May. Munshi’s proof of the Nullstellensatz. American Mathematical Monthly, 110(2):133–140, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647772.

Aldaz:2003:EAI

[929] J. M. Aldaz and A. Bravo. Euclid’s argument on the infinitude of primes. American Mathematical Monthly, 110(2):141–142, February 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647773.

Brown:2003:SIH

[930] Jack B. Brown and George Kozlowski. Smooth interpolation, Hölder continuity, and the Takagi–van der Waerden function. American Mathematical Monthly, 110(2):142–147, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647774.

Chen:2003:WDP

[931] Bang-Yen Chen. When does the position vector of a space curve always Lie in its rectifying plane? American Mathematical Monthly, 110(2): 147–152, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647775.

Zhan:2003:SRT

[932] Xingzhi Zhan. The sharp Rado theorem for majorizations. American Mathematical Monthly, 110(2):152–153, February 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647776.

Navarro:2003:FTF

[933] Gabriel Navarro. On the fundamental theorem of finite Abelian groups. American Mathematical Monthly, 110(2):153–154, February 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647777.

Mortini:2003:P

[934] Raymond Mortini. Problem 10991. American Mathematical Monthly, 110(2):155, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647778.

Lubell:2003:P

[935] David Lubell. Problem 10992. American Mathematical Monthly, 110 (2):155, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647779.

Goldoni:2003:P

[936] Giorgio Goldoni. Problem 10993. American Mathematical Monthly, 110 (2):155, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647780.

Savitt:2003:P

[937] David Savitt. Problem 10994. American Mathematical Monthly, 110(2): 155–156, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647781.

Trigiante:2003:P

[938] Donato Trigiante. Problem 10995. American Mathematical Monthly, 110 (2):156, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647782.

Chao:2003:Pb

[939] Wu Wei Chao. Problem 10996. American Mathematical Monthly, 110 (2):156, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647783.

Prothro:2003:P

[940] Ed T. Prothro. Problem 10997. American Mathematical Monthly, 110 (2):156, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647784.

Thebault:2003:LS

[941] V. Thebault and B. J. English. It’s a long story: 3887. American Mathematical Monthly, 110(2):156–158, February 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647785.

Eynden:2003:DPF

[942] Charles Vanden Eynden and Reiner Martin. A divisibility property of factorials: 10833. American Mathematical Monthly, 110(2):158, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647786.

Lee:2003:CIW

[943] Ho joo Lee and Gerald Myerson. Consecutive integers whose totients are multiples of n: 10837. American Mathematical Monthly, 110(2):158–159, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647787.

Quet:2003:ASI

[944] Leroy Quet and Richard Stong. An alternating sum involving unsigned Stirling numbers: 10863. American Mathematical Monthly, 110(2):159– 160, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647788.

Jones:2003:ER

[945] Lenny Jones and Florian Luca. Euler to the rescue: 10869. American Mathematical Monthly, 110(2):160, February 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647789.

Montgomery:2003:PCP

[946] Hugh L. Montgomery and John H. Lindsey II. Positivity of a certain power series: 10870. American Mathematical Monthly, 110(2):160–161, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647790.

Knuth:2003:BNS

[947] Donald E. Knuth and Gabriel D. Carroll. Balanced neighborhood squares: 10871. American Mathematical Monthly, 110(2):161–162, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647791.

Suceava:2003:FCG

[948] Bogdan Suceavǎ and Richard Stong. The fewest 3-cycles to generate an even permutation: 10872. American Mathematical Monthly, 110(2):162, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647792.

Wickerhauser:2003:BRB

[949] M. Victor Wickerhauser. Book review: [untitled]. American Mathematical Monthly, 110(2):163–167, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647793.

Krantz:2003:BRB

[950] Steven G. Krantz. Book review: [untitled]. American Mathematical Monthly, 110(2):167–171, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647794.

Anonymous:2003:BMc

[951] Anonymous. Back matter. American Mathematical Monthly, 110(2):??, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647795.

Anonymous:2003:FMc

[952] Anonymous. Front matter. American Mathematical Monthly, 110(2):??, February 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647768.

Megginson:2003:YGG

[953] Robert E. Megginson. Yueh-Gin Gung and Dr. Charles Y. Hu Award to Clarence F. Stephens for Distinguished Service to Mathematics. American Mathematical Monthly, 110(3):177–180, March 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647932.

Gudder:2003:QC

[954] Stan Gudder. Quantum computation. American Mathematical Monthly, 110(3):181–201, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647933.

Bamberg:2003:CRP

[955] John Bamberg, Grant Cairns, and Devin Kilminster. The crystallographic restriction, permutations, and Goldbach’s conjecture. American Mathematical Monthly, 110(3):202–209, March 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647934.

Gray:2003:GPD

[956] Stephen B. Gray. Generalizing the Petr–Douglas–Neumann theorem on N-gons. American Mathematical Monthly, 110(3):210–227, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647935.

Halbeisen:2003:CKS

[957] Lorenz Halbeisen. Can Kirkman’s schoolgirls walk abreast by forming arithmetic progressions? American Mathematical Monthly, 110(3):228– 231, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647936.

Minac:2003:NIO

[958] Ján Mináč. Newton’s identities once again! American Mathematical Monthly, 110(3):232–234, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647937.

Hardy:2003:IC

[959] Michael Hardy. An illuminating counterexample. American Mathematical Monthly, 110(3):234–238, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647938.

Nakamoto:2003:NIH

[960] Ritsuo Nakamoto. A norm inequality for Hermitian operators. American Mathematical Monthly, 110(3):238–239, March 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647939.

Tauraso:2003:P

[961] Roberto Tauraso. Problem 10998. American Mathematical Monthly, 110(3):240, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647940.

Minassian:2003:P

[962] D. P. Minassian. Problem 10999. American Mathematical Monthly, 110 (3):240, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647941.

Lamb:2003:Pa

[963] George Lamb. Problem 11000. American Mathematical Monthly, 110(3): 240, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647942.

Mabry:2003:Pb

[964] Rick Mabry. Problem 11001. American Mathematical Monthly, 110(3): 240, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647943.

Petermann:2003:P

[965] Y.-F. S. Pétermann. Problem 11002. American Mathematical Monthly, 110(3):240–241, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647944.

Gray:2003:P

[966] Stephen B. Gray. Problem 11003. American Mathematical Monthly, 110(3):241, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647945.

Pambuccian:2003:P

[967] Victor Pambuccian. Problem 11004. American Mathematical Monthly, 110(3):241, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647946.

Deshpande:2003:REU

[968] M. N. Deshpande, A. V. Kharashikar, and Walther Janous. Recursive expressions in an urn problem: 10868. American Mathematical Monthly, 110(3):241–243, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647947.

Venkatachala:2003:EDE

[969] B. J. Venkatachala and Doyle Henderson. An exponential Diophantine equation: 10873. American Mathematical Monthly, 110(3):243, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647948.

Knuth:2003:AC

[970] Donald E. Knuth and Richard Stong. Animals in a cage: 10875. American Mathematical Monthly, 110(3):243–245, March 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647949.

Deutsch:2003:CTT

[971] Emeric Deutsch and Daniele Donini. Counting tilings with L-tiles and squares: 10877. American Mathematical Monthly, 110(3):245–246, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647950.

Janous:2003:PS

[972] Walther Janous. Pairs of squares: 10879. American Mathematical Monthly, 110(3):246–247, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647951.

Palacios:2003:GAR

[973] José Luis Palacios and Richard Stong. Going around randomly in circles: 10881. American Mathematical Monthly, 110(3):247, March 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647952.

Mueller:2003:BRB

[974] William Mueller. Book review: [untitled]. American Mathematical Monthly, 110(3):248–254, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647953.

Henson:2003:BRB

[975] Shandelle M. Henson. Book review: [untitled]. American Mathematical Monthly, 110(3):254–258, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647954.

Anonymous:2003:TRb

[976] Anonymous. Telegraphic reviews. American Mathematical Monthly, 110 (3):259–261, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647955.

Anonymous:2003:BMd

[977] Anonymous. Back matter. American Mathematical Monthly, 110(3):??, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647956.

Anonymous:2003:FMd

[978] Anonymous. Front matter. American Mathematical Monthly, 110(3):??, March 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647931.

Cobb:2003:AMC

[979] George W. Cobb and Yung-Pin Chen. An application of Markov chain Monte Carlo to community ecology. American Mathematical Monthly, 110(4):265–288, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647877.

Beardon:2003:PLC

[980] Alan F. Beardon and David Minda. On the pointwise limit of complex analytic functions. American Mathematical Monthly, 110(4):289–297, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647878.

Coutinho:2003:QQR

[981] S. C. Coutinho and J. C. McConnell. The quest for quotient rings (of noncommutative Noetherian rings). American Mathematical Monthly, 110(4):298–313, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647879.

Peres:2003:FPL

[982] Yuval Peres, Károly Simon, and Boris Solomyak. Fractals with positive length and zero Buffon needle probability. American Mathematical Monthly, 110(4):314–325, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647880.

Servi:2003:NSR

[983] L. D. Servi. Nested square roots of 2. American Mathematical Monthly, 110(4):326–330, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647881.

Boston:2003:SWL

[984] Nigel Boston. Strategies for the weakest link. American Mathematical Monthly, 110(4):330–334, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647882.

Uchiyama:2003:PKT

[985] Mitsuru Uchiyama. Proofs of Korovkin’s theorems via inequalities. American Mathematical Monthly, 110(4):334–336, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647883.

Walker:2003:ZPS

[986] Peter Walker. The zeros of the partial sums of the exponential series. American Mathematical Monthly, 110(4):337–339, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647884.

Getz:2003:P

[987] Marty Getz and Dixon Jones. Problem 11005. American Mathematical Monthly, 110(4):340, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647885.

Suceava:2003:P

[988] Bogdan Suceavă. Problem 11006. American Mathematical Monthly, 110 (4):340, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647886.

WMCPG:2003:P

[989] Western Maryland College Problems Group. Problem 11007. American Mathematical Monthly, 110(4):340, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/3647887.

Diaz-Barrero:2003:P

[990] José Luis Díaz-Barrero and Juan José Egozcue. Problem 11008. American Mathematical Monthly, 110(4):340, April 2003. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/3647888.

Borwein:2003:P

[991] David Borwein, Jonathan M. Borwein, and Jamal Rooin. Problem 11009. American Mathematical Monthly, 110(4):341, April 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647889.

Just:2003:P

[992] Erwin Just and Norman Schaumberger. Problem 11010. American Mathematical Monthly, 110(4):341, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647890.

Lamb:2003:Pb

[993] George Lamb. Problem 11011. American Mathematical Monthly, 110(4): 341, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647891.

Sasvari:2003:ABD

[994] Zoltán Sasvári and Richard Stong. Approximating the binomial distribution: 10876. American Mathematical Monthly, 110(4):341–342, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647892.

Dalyay:2003:ABC

[995] Pál Péter Dályay, Said Amghibech, and David Callan. Another binomial coefficient identity: 10878. American Mathematical Monthly, 110(4):342– 343, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647893.

MacKinnon:2003:TSR

[996] Nick MacKinnon and John C. Cock. Tiling squares with rectangles: 10883. American Mathematical Monthly, 110(4):343–344, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647894.

Rey:2003:UI

[997] Joaquín Gómez Rey, Albert Nijenhuis, Doyle Henderson, Damuele Riccarelli, and Richard Stong. An unexpected integer: 10886. American Mathematical Monthly, 110(4):344–345, April 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647895.

Anglesio:2003:SIC

[998] Jean Anglesio, Said Amghibech, and GCHQ Problem Solving Group. Six incenters on a conic: 10887. American Mathematical Monthly, 110 (4):346–347, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647896.

Wardlaw:2003:HQ

[999] William P. Wardlaw, William D. Withers, Ellen Hertz, Horst Brunotte, and GCHQ Problem Solving Group. Homomorphisms of quasigroups: 10888. American Mathematical Monthly, 110(4):347, April 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647897.

Montgomery:2003:BRB

[1000] Richard Montgomery. Book review: [untitled]. American Mathematical Monthly, 110(4):348–353, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647898.

Rauch:2003:BRB

[1001] Jeffrey Rauch. Book review: [untitled]. American Mathematical Monthly, 110(4):353–357, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647899.

Palka:2003:EEb

[1002] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 110 (4):358–359, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647900.

Anonymous:2003:BMe

[1003] Anonymous. Back matter. American Mathematical Monthly, 110(4): ??, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647901.

Anonymous:2003:FMe

[1004] Anonymous. Front matter. American Mathematical Monthly, 110(4): ??, April 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647876.

Livingston:2003:ELN

[1005] Charles Livingston. Enhanced linking numbers. American Mathematical Monthly, 110(5):361–385, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647824.

Baiou:2003:AR

[1006] Mourad Baïou and Michel Balinski. Admissions and recruitment. American Mathematical Monthly, 110(5):386–399, May 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647825.

Pach:2003:HMU

[1007] János Pach and Rom Pinchasi. How many unit equilateral triangles can be generated by N points in convex position? American Mathematical Monthly, 110(5):400–406, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647826.

Schmuland:2003:RHS

[1008] Byron Schmuland. Random harmonic series. American Mathematical Monthly, 110(5):407–416, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647827.

Hudson:2003:FCD

[1009] Hud Hudson. Four colors do not suffice. American Mathematical Monthly, 110(5):417–423, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647828.

Duncan:2003:CI

[1010] John Duncan and Colin M. McGregor. Carleman’s inequality. American Mathematical Monthly, 110(5):424–431, May 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647829.

Nyblom:2003:NSD

[1011] M. A. Nyblom. A note on the solvability of a Diophantine equation involving radicals. American Mathematical Monthly, 110(5):431–434, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647830.

Sondow:2003:ZAZ

[1012] Jonathan Sondow. Zeros of the alternating zeta function on the line R(S) = 1. American Mathematical Monthly, 110(5):435–437, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647831.

Popescu:2003:P

[1013] Călin Popescu. Problem 11012. American Mathematical Monthly, 110(5): 438, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647832.

Callan:2003:P

[1014] David Callan. Problem 11013. American Mathematical Monthly, 110(5): 438, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647833.

Sinefakopoulos:2003:P

[1015] Achilleas Sinefakopoulos. Problem 11014. American Mathematical Monthly, 110(5):438, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647834.

Furdui:2003:P

[1016] Ovidiu Furdui. Problem 11015. American Mathematical Monthly, 110(5): 438, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647835.

Marquez:2003:P

[1017] Juan-Boscoa Romero Márquez. Problem 11016. American Mathematical Monthly, 110(5):439, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647836.

Pranesachar:2003:P

[1018] C. R. Pranesachar. Problem 11017. American Mathematical Monthly, 110(5):439, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647837.

Simic:2003:P

[1019] Slavko Simic. Problem 11018. American Mathematical Monthly, 110(5): 439, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647838.

Bacher:2003:SSD

[1020] Roland Bacher and Albert Nijenhuis. A surprisingly simple determinant: 10891. American Mathematical Monthly, 110(5):439–440, May 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647839.

Locke:2003:SCH

[1021] Stephen C. Locke and Richard Stong. Spanning cycles in hypercubes: 10892. American Mathematical Monthly, 110(5):440–441, May 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647840.

Jorza:2003:GD

[1022] Andrei Jorza, John H. Lindsey II, Anne-Maria Ernvall, and Reijo Ernvall. Groups of divisors: 10893. American Mathematical Monthly, 110(5):441– 443, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647841.

Sulanke:2003:ADC

[1023] Robert Sulanke, Jose Luis Arregui, Danijela Vujić, Nelson Castañeda, and David Callan. Another description of the central delannoy numbers: 10894. American Mathematical Monthly, 110(5):443–444, May 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647842.

Hillar:2003:CSE

[1024] Christopher J. Hillar and Patrick J. Fitzsimmons. Convergent series with exponents: 10928. American Mathematical Monthly, 110(5):444– 445, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647843.

Wimp:2003:SS

[1025] Jet Wimp and Kenneth Schilling. Serendipitous solution: 10428. American Mathematical Monthly, 110(5):445, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/3647844.

Knapp:2003:BRB

[1026] A. W. Knapp. Book review: [untitled]. American Mathematical Monthly, 110(5):446–455, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647845.

Cohen:2003:BRBb

[1027] Marion Cohen. Book review: [untitled]. American Mathematical Monthly, 110(5):455–458, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647846.

Anonymous:2003:TRc

[1028] Anonymous. Telegraphic reviews. American Mathematical Monthly, 110 (5):459–461, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647847.

Anonymous:2003:BMf

[1029] Anonymous. Back matter. American Mathematical Monthly, 110(5): ??, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647848.

Anonymous:2003:FMf

[1030] Anonymous. Front matter. American Mathematical Monthly, 110(5): ??, May 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647823.

Krantz:2003:MG

[1031] Steven G. Krantz. A matter of gravity. American Mathematical Monthly, 110(6):465–481, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647903.

Toussaint:2003:SPG

[1032] Godfried Toussaint. Simple proofs of a geometric property of four-bar linkages. American Mathematical Monthly, 110(6):482–494, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647904.

Hartshorne:2003:NEI

[1033] Robin Hartshorne. Non-Euclidean III.36. American Mathematical Monthly, 110(6):495–502, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647905.

Mendivil:2003:FGF

[1034] Franklin Mendivil. Fractals, graphs, and fields. American Mathematical Monthly, 110(6):503–515, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647906.

Apostol:2003:SSD

[1035] Tom M. Apostol and Mamikon A. Mnatsakanian. Sums of squares of distances in m-space. American Mathematical Monthly, 110(6):516–526, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647907.

Klazar:2003:CEO

[1036] Martin Klazar. Counting even and odd partitions. American Mathematical Monthly, 110(6):527–532, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647908.

Vukotic:2003:IIT

[1037] Dragan Vukotić. The isoperimetric inequality and a theorem of Hardy and Littlewood. American Mathematical Monthly, 110(6):532–536, June/ July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647909.

Schep:2003:SOM

[1038] Anton R. Schep. And still one more proof of the Radon–Nikodym theorem. American Mathematical Monthly, 110(6):536–538, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647910.

Somos:2003:LPS

[1039] Michael Somos and Robert Haas. A linked pair of sequences implies the primes are infinite. American Mathematical Monthly, 110(6):539–540, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647911.

Harper:2003:ASP

[1040] James D. Harper. Another simple proof of 1 + 1/2^{2} + 1/3^{2} + ⋯ = π^{2}/6. American Mathematical Monthly, 110(6):540–541, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647912.

Santos:2003:P

[1041] Bernardo Recamán Santos. Problem 11019. American Mathematical Monthly, 110(6):542, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647913.

Glasser:2003:Pa

[1042] M. L. Glasser. Problem 11020. American Mathematical Monthly, 110 (6):542, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647914.

Knuth:2003:Pb

[1043] Donald E. Knuth. Problem 11021. American Mathematical Monthly, 110 (6):542, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647915.

See correction [1165].

Satnoianu:2003:P

[1044] Razvan Satnoianu. Problem 11022. American Mathematical Monthly, 110 (6):542, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647916.

Chao:2003:Pc

[1045] Wu Wei Chao. Problem 11023. American Mathematical Monthly, 110 (6):542, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647917.

Radulescu:2003:P

[1046] Vicenţiu Rădulescu. Problem 11024. American Mathematical Monthly, 110(6):543, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647918.

vanLamoen:2003:Pa

[1047] Floor van Lamoen and Darij Grinberg. Problem 11025. American Mathematical Monthly, 110(6):543, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647919.

Mabry:2003:Pc

[1048] Rick Mabry. Problem 11001. American Mathematical Monthly, 110(6): 543, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647920.

Chao:2003:DCH

[1049] Wu Wei Chao and Gerry Myerson. Divisibility conditions holding as many times as you want: 10898. American Mathematical Monthly, 110 (6):543–544, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647921.

Bagby:2003:BPA

[1050] Richard Bagby and O. P. Lossers. Best polynomial approximations: 10903. American Mathematical Monthly, 110(6):544, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647922.

Chen:2003:HMV

[1051] Zhibing Chen and Shusen Ding. A higher mean value theorem: 10935. American Mathematical Monthly, 110(6):544–545, June/July 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647923.

Lucacs:2003:CG

[1052] Andor Lucács, Szilard András, and Nathan D. Cahill. Convex n-gons: 10936. American Mathematical Monthly, 110(6):545–546, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647924.

Nievergelt:2003:CGM

[1053] Yves Nievergelt and Achava Nakhash. The complex geometric mean: 10940. American Mathematical Monthly, 110(6):546–547, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647925.

Lagarias:2003:DMC

[1054] Jeffrey C. Lagarias, Thomas Hermann, and O. P. Lossers. Decay of a Markov chain: 10941. American Mathematical Monthly, 110(6):547–548, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647926.

Furdui:2003:IT

[1055] Ovidiu Furdui and O. P. Lossers. An inequality for triangles: 10950. American Mathematical Monthly, 110(6):548–549, June/July 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647927.

Schaeffer:2003:BRB

[1056] David G. Schaeffer. Book review: [untitled]. American Mathematical Monthly, 110(6):550–553, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647928.

Nunemacher:2003:BRB

[1057] Jeffrey L. Nunemacher. Book review: [untitled]. American Mathematical Monthly, 110(6):554–557, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647929.

Anonymous:2003:BMg

[1058] Anonymous. Back matter. American Mathematical Monthly, 110(6):??, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647930.

Anonymous:2003:FMg

[1059] Anonymous. Front matter. American Mathematical Monthly, 110(6):??, June/July 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647902.

Elkies:2003:S

[1060] Noam D. Elkies. On the sums ∑_{k=−∞}^{∞}(4k + 1)^{−n}. American Mathematical Monthly, 110(7):561–573, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647742.

Biss:2003:WFP

[1061] Daniel K. Biss. Which functor is the projective line? American Mathematical Monthly, 110(7):574–592, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647743.

Lodder:2003:CCC

[1062] Jerry Lodder. Curvature in the calculus curriculum. American Mathematical Monthly, 110(7):593–605, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647744.

Harel:2003:PWV

[1063] Guershon Harel and Jeffrey M. Rabin. Polygons whose vertex triangles have equal area. American Mathematical Monthly, 110(7):606–619, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647745.

Derksen:2003:FTA

[1064] Harm Derksen. The fundamental theorem of algebra and linear algebra. American Mathematical Monthly, 110(7):620–623, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647746.

Thas:2003:PGT

[1065] Charles Thas. Projective generalizations of two points of concurrence on the nine-point circle. American Mathematical Monthly, 110(7):624–627, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647747.

Abel:2003:LRT

[1066] Ulrich Abel. On the Lagrange remainder of the Taylor formula. American Mathematical Monthly, 110(7):627–633, August/September 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647748.

Vaisala:2003:PMU

[1067] Jussi Väisälä. A proof of the Mazur–Ulam Theorem. American Mathematical Monthly, 110(7):633–635, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647749.

Sondow:2003:P

[1068] Jonathan Sondow. Problem 11026. American Mathematical Monthly, 110 (7):636, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647750.

Reuvers:2003:P

[1069] H. Reuvers. Problem 11027. American Mathematical Monthly, 110(7): 636, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647751.

Trenkler:2003:P

[1070] Götz Trenkler. Problem 11028. American Mathematical Monthly, 110 (7):636, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647752.

Keller:2003:P

[1071] Tim Keller. Problem 11029. American Mathematical Monthly, 110(7): 636, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647753.

Butler:2003:P

[1072] Steven Butler. Problem 11030. American Mathematical Monthly, 110 (7):637, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647754.

Zhou:2003:P

[1073] Li Zhou. Problem 11031. American Mathematical Monthly, 110(7):637, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647755.

Pinsky:2003:P

[1074] Mark Pinsky. Problem 11032. American Mathematical Monthly, 110 (7):637, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647756.

Bisgaard:2003:DMT

[1075] Torben Maack Bisgaard, Zsolt Páles, Zoltán Sasvári, and Richard Bagby. Determinant of a matrix in two variables: 10896. American Mathematical Monthly, 110(7):637–638, August/September 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647757.

Wene:2003:GIM

[1076] Gregory P. Wene and Robin Chapman. Generators of ideals of matrices over finite fields: 10897. American Mathematical Monthly, 110(7):638– 639, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647758.

Rice:2003:ARC

[1077] Gordon Rice, Patrick A. Staley, Gerald Myerson, and Robin Chapman. Angles with rational cosines: 10900. American Mathematical Monthly, 110(7):639, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647759.

Deutsch:2003:DDR

[1078] Emeric Deutsch and Nasha Komanda. Derangements that don’t rise too fast: 10902. American Mathematical Monthly, 110(7):639–640, August/ September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647760.

Stanley:2003:UAC

[1079] Richard P. Stanley and Robin Chapman. An unexpected appearance of the Catalan numbers: 10905. American Mathematical Monthly, 110(7): 640–642, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647761.

Knuth:2003:RRC

[1080] Donald E. Knuth, C. P. Rupert, Alex Smith, and Richard Stong. Recounting the rationals, continued: 10906. American Mathematical Monthly, 110(7):642–643, August/September 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647762.

Hubbard:2003:BRB

[1081] John H. Hubbard. Book review: [untitled]. American Mathematical Monthly, 110(7):644–649, August/September 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647763.

Cohen:2003:BRBc

[1082] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 110(7):650–652, August/September 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647764.

Roy:2003:BRB

[1083] Ranjan Roy. Book review: [untitled]. American Mathematical Monthly, 110(7):652–657, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647765.

Anonymous:2003:TRd

[1084] Anonymous. Telegraphic reviews. American Mathematical Monthly, 110(7):658–660, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647766.

Anonymous:2003:BMh

[1085] Anonymous. Back matter. American Mathematical Monthly, 110(7):??, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647767.

Anonymous:2003:FMh

[1086] Anonymous. Front matter. American Mathematical Monthly, 110(7):??, August/September 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647741.

Curtis:2003:RBS

[1087] Charles W. Curtis. Richard Brauer: Sketches from his life and work. American Mathematical Monthly, 110(8):665–678, October 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647850.

Havlicek:2003:ATT

[1088] Hans Havlicek and Gunter Weiß. Altitudes of a tetrahedron and traceless quadratic forms. American Mathematical Monthly, 110(8):679–693, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647851.

Acosta:2003:NRS

[1089] Daniel J. Acosta. Newton’s rule of signs for imaginary roots. American Mathematical Monthly, 110(8):694–706, October 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647852.

Beck:2003:NMS

[1090] Matthias Beck, Moshe Cohen, Jessica Cuomo, and Paul Gribelyuk. The number of “magic” squares, cubes, and hypercubes. American Mathematical Monthly, 110(8):707–717, October 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647853.

Klosinski:2003:STW

[1091] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Third William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 110(8):718–726, October 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647854.

Satnoianu:2003:EMT

[1092] Razvan A. Satnoianu. Erdős–Mordell-type inequalities in a triangle. American Mathematical Monthly, 110(8):727–729, October 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647855.

Impens:2003:SSM

[1093] Chris Impens. Stirling’s series made easy. American Mathematical Monthly, 110(8):730–735, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647856.

Lichiardopol:2003:NPW

[1094] Nicolas Lichiardopol. A new proof of Wedderburn’s theorem. American Mathematical Monthly, 110(8):736–737, October 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647857.

Boulabiar:2003:ADC

[1095] Karim Boulabiar and Gerard Buskes. After the determinants are down: a criterion for invertibility. American Mathematical Monthly, 110(8): 737–741, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647858.

Deshpande:2003:Pa

[1096] M. N. Deshpande and R. M. Welukar. Problem 11033. American Mathematical Monthly, 110(8):742, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647859.

Wardlaw:2003:P

[1097] William P. Wardlaw. Problem 11034. American Mathematical Monthly, 110(8):742, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647860.

See correction [1211].

Yayenie:2003:P

[1098] Omer Yayenie. Problem 11035. American Mathematical Monthly, 110 (8):742, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647861.

Glasser:2003:Pb

[1099] M. L. Glasser. Problem 11036. American Mathematical Monthly, 110 (8):743, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647862.

Hautus:2003:P

[1100] M. L. J. Hautus. Problem 11037. American Mathematical Monthly, 110 (8):743, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647863.

vanLamoen:2003:Pb

[1101] Floor van Lamoen. Problem 11038. American Mathematical Monthly, 110(8):743, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647864.

Rosenberg:2003:P

[1102] Noah Rosenberg and Richard Stong. Problem 11039. American Mathematical Monthly, 110(8):743, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647865.

Martin:2003:TPS

[1103] Greg Martin, Jonathan Bloom, and Dan Saracino. Three proper subgroups covering a group: 10911. American Mathematical Monthly, 110 (8):743–744, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647866.

Anglesio:2003:RDS

[1104] Jean Anglesio and Kenneth F. Andersen. A recursively defined series for π/4: 10912. American Mathematical Monthly, 110(8):744–745, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647867.

Falcone:2003:CCP

[1105] Giovanni Falcone, Richard Stong, and Nasha Komanda. A combination of cyclotomic polynomials: 10914. American Mathematical Monthly, 110 (8):745–746, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647868.

Ehrlich:2003:JPS

[1106] Gertrude Ehrlich and Reiner Martin. The jug problem with a source and a sink: 10916. American Mathematical Monthly, 110(8):746, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647869.

Beck:2003:CIC

[1107] Mathias Beck and Nasha Komanda. A congruence involving ceilings: 10918. American Mathematical Monthly, 110(8):746–747, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647870.

Kundgen:2003:MPB

[1108] André Kündgen, Radhika Ramamurthi, and Reiner Martin. Maximal paths in bipartite graphs: 10920. American Mathematical Monthly, 110 (8):747, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647871.

Mazur:2003:RI

[1109] Marcin Mazur, Roberto Tauraso, and Michel Bataille. A root inequality: 10944. American Mathematical Monthly, 110(8):747–749, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647872.

Kasman:2003:BRB

[1110] Alex Kasman. Book review: [untitled]. American Mathematical Monthly, 110(8):750–754, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647873.

Tucker:2003:BRB

[1111] Warwick Tucker. Book review: [untitled]. American Mathematical Monthly, 110(8):754–759, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647874.

Anonymous:2003:BMi

[1112] Anonymous. Back matter. American Mathematical Monthly, 110(8):??, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647875.

Anonymous:2003:FMi

[1113] Anonymous. Front matter. American Mathematical Monthly, 110(8):??, October 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647849.

Mladenov:2003:MBR

[1114] Ivaïlo M. Mladenov and John Oprea. The mylar balloon revisited. American Mathematical Monthly, 110(9):761–784, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647797.

Sarkaria:2003:TAW

[1115] Karanbir S. Sarkaria. Tetrahedron ABCD of width 1 with minimum AB + BC + CD. American Mathematical Monthly, 110(9):785–799, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647798.

McAdam:2003:BTD

[1116] Stephen McAdam. Bell’s theorem and the demise of local reality. American Mathematical Monthly, 110(9):800–811, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647799.

Groer:2003:MSA

[1117] Chris Groër. The mathematics of survival: From antiquity to the playground. American Mathematical Monthly, 110(9):812–825, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647800.

Poorrezaei:2003:TPG

[1118] Kamal Poorrezaei. Two proofs of Graves’s theorem. American Mathematical Monthly, 110(9):826–830, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/3647801.

Baek:2003:CUS

[1119] Jeongseon Baek, Dong-Soo Kim, and Young Ho Kim. A characterization of the unit sphere. American Mathematical Monthly, 110(9):830–833, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647802.

Botsko:2003:EPL

[1120] Michael W. Botsko. An elementary proof of Lebesgue’s differentiation theorem. American Mathematical Monthly, 110(9):834–838, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/3647803.

Sidorov:2003:AEN

[1121] Nikita Sidorov. Almost every number has a continuum of β-expansions. American Mathematical Monthly, 110(9):838–842, November 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647804.

Tian:2003:Pb

[1122] Yongge Tian. Problem 11040. American Mathematical Monthly, 110 (9):843, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647805.

Ciaurri:2003:P

[1123] Oscar Ciaurri. Problem 11041. American Mathematical Monthly, 110 (9):843, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647806.

Deshpande:2003:Pb

[1124] M. N. Deshpande and Kavita Laghate. Problem 11042. American Mathematical Monthly, 110(9):843–844, November 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647807.

Golomb:2003:P

[1125] Michael Golomb. Problem 11043. American Mathematical Monthly, 110 (9):844, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647808.

Bataille:2003:P

[1126] Michel Bataille. Problem 11044. American Mathematical Monthly, 110 (9):844, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647809.

Singh:2003:P

[1127] Manoj Prakash Singh. Problem 11045. American Mathematical Monthly, 110(9):844, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647810.

Soland:2003:P

[1128] Christoph Soland. Problem 11046. American Mathematical Monthly, 110 (9):844, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647811.

Knuth:2003:PT

[1129] Donald E. Knuth and Reiner Martin. Products of transpositions: 10913. American Mathematical Monthly, 110(9):844–845, November 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647812.

Rupert:2003:SIP

[1130] C. P. Rupert and Richard Stong. A sequence of integer polynomials: 10915. American Mathematical Monthly, 110(9):845, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647813.

Csete:2003:DS

[1131] Lajos Csete and Li Zhou. Distances to the sides: 10929. American Mathematical Monthly, 110(9):845–846, November 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647814.

Mihai:2003:TAF

[1132] Vasile Mihai and Li Zhou. A tetrahedron with all faces congruent: 10934. American Mathematical Monthly, 110(9):846–847, November 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647815.

Ivic:2003:Z

[1133] Aleksandar Ivić. Zetafunctionology: 10939. American Mathematical Monthly, 110(9):847–848, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647816.

Simic:2003:AII

[1134] Slavko Simic and Said Amghibech. An appell-ing inequality: 10945. American Mathematical Monthly, 110(9):848–849, November 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647817.

Stevic:2003:AR

[1135] Stevo Stević and Allen Stenger. Asymptotics of a recurrence: 10946. American Mathematical Monthly, 110(9):849–850, November 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647818.

Rucker:2003:BRB

[1136] Rudy Rucker. Book review: [untitled]. American Mathematical Monthly, 110(9):851–861, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647819.

Alladi:2003:BRB

[1137] Krishnaswami Alladi. Book review: [untitled]. American Mathematical Monthly, 110(9):861–865, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647820.

Anonymous:2003:TRe

[1138] Anonymous. Telegraphic reviews. American Mathematical Monthly, 110(9):866–868, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647821.

Anonymous:2003:BMj

[1139] Anonymous. Back matter. American Mathematical Monthly, 110(9):??, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647822.

Anonymous:2003:FMj

[1140] Anonymous. Front matter. American Mathematical Monthly, 110(9):??, November 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647796.

Drucker:2003:WDW

[1141] Daniel Drucker and Stephen A. Williams. When does water find the shortest path downhill? the geometry of steepest descent curves. American Mathematical Monthly, 110(10):869–885, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647959.

Epp:2003:RLT

[1142] Susanna S. Epp. The role of logic in teaching proof. American Mathematical Monthly, 110(10):886–899, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/3647960.

Pinkus:2003:NTA

[1143] Allan Pinkus. Negative theorems in approximation theory. American Mathematical Monthly, 110(10):900–911, December 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647961.

Chandler:2003:EMC

[1144] Richard E. Chandler, Carl D. Meyer, and Nicholas J. Rose. Eudoxus meets Cayley. American Mathematical Monthly, 110(10):912–927, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647962.

Blume:2003:LPA

[1145] Frank Blume. A localized probabilistic approach to circuit analysis. American Mathematical Monthly, 110(10):928–936, December 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647963.

Bertuccioni:2003:TP

[1146] Inta Bertuccioni. A topological puzzle. American Mathematical Monthly, 110(10):937–939, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647964.

Ullrich:2003:AAT

[1147] David C. Ullrich. The Ascoli–Arzelà Theorem via Tychonoff’s theorem. American Mathematical Monthly, 110(10):939–940, December 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647965.

Kannappan:2003:KTP

[1148] PL. Kannappan. Klee’s trigonometry problem. American Mathematical Monthly, 110(10):940–944, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647966.

Loya:2003:GTF

[1149] Paul Loya. Green’s theorem and the fundamental theorem of algebra. American Mathematical Monthly, 110(10):944–946, December 2003. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/3647967.

Hulin:2003:MCA

[1150] Dominique Hulin and Marc Troyanov. Mean curvature and asymptotic volume of small balls. American Mathematical Monthly, 110(10):947–950, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647968.

Groetsch:2003:TEF

[1151] C. W. Groetsch. Timing is everything: The French connection. American Mathematical Monthly, 110(10):950–955, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647969.

Marivani:2003:P

[1152] Syrous Marivani. Problem 11047. American Mathematical Monthly, 110 (10):956, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647970.

Beck:2003:P

[1153] Matthias Beck, Marcin Mazur, and Shelemyahu Zacks. Problem 11048. American Mathematical Monthly, 110(10):956, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647971.

Mihai:2003:P

[1154] Vasile Mihai. Problem 11049. American Mathematical Monthly, 110 (10):956, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647972.

Holing:2003:P

[1155] Kent Holing. Problem 11050. American Mathematical Monthly, 110 (10):957, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647973.

MacKinnon:2003:P

[1156] Nick MacKinnon. Problem 11051. American Mathematical Monthly, 110 (10):957, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647974.

Huang:2003:P

[1157] Danrun Huang and Daniel Scully. Problem 11052. American Mathematical Monthly, 110(10):957, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647975.

Chao:2003:Pd

[1158] Wu Wei Chao. Problem 11053. American Mathematical Monthly, 110 (10):958, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/3647976.

Bloom:2003:CMB

[1159] David M. Bloom, Christopher N. Swanson, and Kenneth Schilling. A convolution of middle binomial coefficients: 10921. American Mathematical Monthly, 110(10):958–959, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647977.

Khan:2003:SDW

[1160] Mizan R. Khan and Michael R. Avidon. The sum of divisors won’t die: 10922. American Mathematical Monthly, 110(10):959, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/3647978.

Callan:2003:PF

[1161] David Callan and Harris Kwong. A permanent fixture: 10925. American Mathematical Monthly, 110(10):960, December 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/3647979.

Diamond:2003:CFA

[1162] Harold G. Diamond and Li Zhou. A continued fraction for an arctan difference: 10926. American Mathematical Monthly, 110(10):960–961, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647980.

Grivaux:2003:RDI

[1163] Jean-Pierre Grivaux, Allen Stenger, and J. Ernest Wilkins Jr. A rational definite integral: 10938. American Mathematical Monthly, 110(10):961– 962, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/3647981.

Just:2003:PC

[1164] Erwin Just, Norman Schaumberger, and Albert Stadler. Points on a circle: 10954. American Mathematical Monthly, 110(10):962–963, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647982.

Anonymous:2003:EAA

[1165] Anonymous. Erratum: Anagrammatical attribution: 11021. American Mathematical Monthly, 110(10):963, December 2003. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/3647983. See [1043], which was incorrectly

attributed to another.

Waymire:2003:BRB

[1166] Ed Waymire. Book review: [untitled]. American Mathematical Monthly, 110(10):964–967, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647984.

Cohen:2003:BRBa

[1167] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 110(10):968–969, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647985.

Gottesman:2003:BRB

[1168] Daniel Gottesman. Book review: [untitled]. American Mathematical Monthly, 110(10):969–972, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/3647986.

Anonymous:2003:VI

[1169] Anonymous. Volume information. American Mathematical Monthly, 110 (10):973–984, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

3647957.

Anonymous:2003:BMb

[1170] Anonymous. Back matter. American Mathematical Monthly, 110(10):??, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647987.

Anonymous:2003:FMb

[1171] Anonymous. Front matter. American Mathematical Monthly, 110(10):??, December 2003. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3647958.

Grattan-Guinness:2004:HHI

[1172] Ivor Grattan-Guinness. History or heritage? an important distinction in mathematics and for mathematics education. American Mathematical Monthly, 111(1):1–12, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145010.

Bennett:2004:PRS

[1173] Michael A. Bennett and Bruce Reznick. Positive rational solutions to x^{y} = y^{mx}: a number-theoretic excursion. American Mathematical Monthly, 111(1):13–21, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145011.

Jerrard:2004:PRR

[1174] Richard P. Jerrard and John E. Wetzel. Prince Rupert’s rectangles. American Mathematical Monthly, 111(1):22–31, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145012.

Krantz:2004:CMC

[1175] Steven G. Krantz and Jeffery D. McNeal. Creating more convergent series. American Mathematical Monthly, 111(1):32–38, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145013.

Polster:2004:VTP

[1176] Burkard Polster. Variations on a theme in paper folding. American Mathematical Monthly, 111(1):39–47, January 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145014.

Kim:2004:EPQ

[1177] Sey Y. Kim. An elementary proof of the quadratic reciprocity law. American Mathematical Monthly, 111(1):48–50, January 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145015.

Trenkler:2004:MWT

[1178] Götz Trenkler. Matrices which take a given vector into a given vector: Revisited. American Mathematical Monthly, 111(1):50–52, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145016.

Osler:2004:FPS

[1179] Thomas J. Osler. Finding ζ(2p) from a product of sines. American Mathematical Monthly, 111(1):52–54, January 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145017.

Kulkarni:2004:VSE

[1180] S. H. Kulkarni. A very simple and elementary proof of a theorem of ingelstam. American Mathematical Monthly, 111(1):54–58, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145018.

Beyer:2004:EHH

[1181] W. A. Beyer and Andrew Zardecki. The early history of the Ham Sandwich Theorem. American Mathematical Monthly, 111(1):58–61, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145019.

Perlis:2004:RAQ

[1182] Alexander R. Perlis. Roots appear in quanta. American Mathematical Monthly, 111(1):61–63, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145020.

Amrahov:2004:P

[1183] Shahin Amrahov. Problem 11054. American Mathematical Monthly, 111(1):64, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145021.

Satnoianu:2004:Pa

[1184] Razvan Satnoianu. Problem 11055. American Mathematical Monthly, 111(1):64, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145022.

Lindsey:2004:P

[1185] John H. Lindsey II. Problem 11056. American Mathematical Monthly, 111(1):64, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145023.

ZPG:2004:P

[1186] Zahlreich Problems Group. Problem 11057. American Mathematical Monthly, 111(1):64, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145024.

Ciaurri:2004:P

[1187] Óscar Ciaurri. Problem 11058. American Mathematical Monthly, 111(1): 64, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145025.

Bowron:2004:P

[1188] Mark Bowron. Problem 11059. American Mathematical Monthly, 111 (1):64–65, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145026.

Coppenbarger:2004:P

[1189] Matthew Coppenbarger. Problem 11060. American Mathematical Monthly, 111(1):65, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145027.

Walker:2004:CCD

[1190] Peter Walker and Tom Zerger. The centroid-circumcenter distance: 10948. American Mathematical Monthly, 111(1):65–66, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145028.

Quet:2004:FRC

[1191] Leroy Quet and GCHQ Problem Solving Group. Functionally related convergents: 10953. American Mathematical Monthly, 111(1):66–67, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145029.

Satnoianu:2004:MTI

[1192] Razvan Satnoianu and Isaac Sofair. More triangle inequalities: 10955. American Mathematical Monthly, 111(1):67–69, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145030.

Lamb:2004:HOR

[1193] George L. Lamb and Stephen J. Herschkorn. Hypergeometric and other representations: 10959. American Mathematical Monthly, 111(1):69–71, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145031.

Bencze:2004:CI

[1194] Mihály Bencze and Li Zhou. Cosine inequalities: 10962. American Mathematical Monthly, 111(1):71, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145032.

Ralston:2004:BRB

[1195] Anthony Ralston. Book review: [untitled]. American Mathematical Monthly, 111(1):72–75, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145033.

Anonymous:2004:U

[1196] Anonymous. [untitled]. American Mathematical Monthly, 111(1):75–81, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145034.

Ostebee:2004:TRa

[1197] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 111(1):82–84, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145035.

Anonymous:2004:BMa

[1198] Anonymous. Back matter. American Mathematical Monthly, 111(1):85– 88, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145036.

Anonymous:2004:FMa

[1199] Anonymous. Front matter. American Mathematical Monthly, 111(1):??, January 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145009.

Landau:2004:PNS

[1200] Susan Landau. Polynomials in the Nation’s service: Using algebra to design the advanced encryption standard. American Mathematical Monthly, 111(2):89–117, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145212.

Apostol:2004:IIP

[1201] Tom M. Apostol and Mamikon A. Mnatsakanian. Isoperimetric and isoparametric problems. American Mathematical Monthly, 111(2):118– 136, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145213.

Radin:2004:OOS

[1202] Charles Radin. Orbits of orbs: Sphere packing meets Penrose tilings. American Mathematical Monthly, 111(2):137–149, February 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145214.

Zanello:2004:WTI

[1203] Fabrizio Zanello. When are there infinitely many irreducible elements in a principal ideal domain? American Mathematical Monthly, 111(2): 150–152, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145215.

Avis:2004:SIP

[1204] David Avis and Bohdan Kaluzny. Solving inequalities and proving Farkas’s lemma made easy. American Mathematical Monthly, 111(2): 152–157, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145216.

Hwang:2004:CIT

[1205] Suk-Geun Hwang. Cauchy’s interlace theorem for eigenvalues of Hermitian matrices. American Mathematical Monthly, 111(2):157–159, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145217.

Schmitt:2004:NDD

[1206] Michael Schmitt. New designs for the Descartes Rule of Signs. American Mathematical Monthly, 111(2):159–164, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145218.

Chao:2004:Pa

[1207] Wu Wei Chao. Problem 11061. American Mathematical Monthly, 111 (2):165, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145219.

Schilling:2004:P

[1208] Kenneth Schilling. Problem 11062. American Mathematical Monthly, 111(2):165, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145220.

Simic:2004:P

[1209] Slavko Simic. Problem 11063. American Mathematical Monthly, 111 (2):165, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145221.

Wetzel:2004:P

[1210] John Wetzel. Problem 11064. American Mathematical Monthly, 111 (2):165, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145222.

Anonymous:2004:C

[1211] Anonymous. Correction: 11034. American Mathematical Monthly, 111 (2):166, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145226.

See [1097].

Tyler:2004:P

[1212] Douglas B. Tyler. Problem 11065. American Mathematical Monthly, 111 (2):166, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145223.

Cioaba:2004:P

[1213] Sebastian Cioabă. Problem 11066. American Mathematical Monthly, 111 (2):166, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145224.

Lamb:2004:Pa

[1214] George Lamb. Problem 11067. American Mathematical Monthly, 111 (2):166, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145225.

Knuth:2004:FCC

[1215] Donald E. Knuth and Robin Chapman. Fibonacci in complex camouflage: 10858. American Mathematical Monthly, 111(2):166–167, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145227. See errata

[1446].

Lagarias:2004:ES

[1216] Jeffrey C. Lagarias, E. M. Rains, N. J. A. Sloane, and Nasha Komanda. The EKG sequence: 10927. American Mathematical Monthly, 111(2):

167–168, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145228.

Schweitzer:2004:TMC

[1217] Paul Schweitzer, Xiaoshen Wang, and National Security Agency Problems Group. Too much commuting: 10930. American Mathematical Monthly, 111(2):168–169, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145229.

DeSario:2004:IS

[1218] Robert DeSario and LeRoy Wenstrom. Invertible shuffles: 10931. American Mathematical Monthly, 111(2):169–170, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145230.

Lescott:2004:GDR

[1219] Paul Lescott and Tom Jager. A group determined by relations?: 10933. American Mathematical Monthly, 111(2):170–171, February 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145231. See errata [1447].

Wagner:2004:CS

[1220] David G. Wagner, Patrick Corn, and Li Zhou. A cyclotomic sum: 10937. American Mathematical Monthly, 111(2):171–172, February 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145232.

Alexandrov:2004:UST

[1221] Victor Alexandrov and Knut Dale. Using Stokes’s theorem on the sphere: 10957. American Mathematical Monthly, 111(2):172, February 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145233.

Bell:2004:BRB

[1222] Steven R. Bell. Book review: [untitled]. American Mathematical Monthly, 111(2):173–176, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145234.

Bayer:2004:BRB

[1223] Margaret M. Bayer. Book review: [untitled]. American Mathematical Monthly, 111(2):177–180, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145235.

Anonymous:2004:BMc

[1224] Anonymous. Back matter. American Mathematical Monthly, 111(2): 181–184, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145236.

Anonymous:2004:FMc

[1225] Anonymous. Front matter. American Mathematical Monthly, 111(2):??, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145211.

Megginson:2004:YGG

[1226] Robert E. Megginson. Yueh-Gin Gung and Dr. Charles Y. Hu Award to T. Christine Stevens for Distinguished Service to Mathematics. American Mathematical Monthly, 111(3):185–188, March 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145126.

Edelman:2004:PM

[1227] Alan Edelman and Gilbert Strang. Pascal matrices. American Mathematical Monthly, 111(3):189–197, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145127.

Ingram:2004:IDI

[1228] William T. Ingram and William S. Mahavier. Interesting dynamics and inverse limits in a family of one-dimensional maps. American Mathematical Monthly, 111(3):198–215, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145128.

Paradis:2004:FQF

[1229] Jaume Paradís, Josep Pla, and Pelegrí Viader. Fermat and the quadrature of the folium of Descartes. American Mathematical Monthly, 111 (3):216–229, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145129.

Gronau:2004:ST

[1230] Detlef Gronau. The spiral of theodorus. American Mathematical Monthly, 111(3):230–237, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145130.

Galen:2004:PFT

[1231] Leslie Blackwell Galen. Putting fractions in their place. American Mathematical Monthly, 111(3):238–242, March 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/4145131.

Ishikawa:2004:PF

[1232] Tsuneo Ishikawa, Nobuhiko Ishida, and Yoshito Yukimoto. On prime factors of A^{n} − 1. American Mathematical Monthly, 111(3):243–245, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145132.

Pretorius:2004:APM

[1233] Lou M. Pretorius and Konrad J. Swanepoel. An algorithmic proof of the Motzkin–Rabin Theorem on monochrome lines. American Mathematical Monthly, 111(3):245–251, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145133.

Engman:2004:NIE

[1234] Martin Engman. A note on isometric embeddings of surfaces of revolution. American Mathematical Monthly, 111(3):251–255, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145134.

Lam:2004:SPI

[1235] T. Y. Lam. On subgroups of prime index. American Mathematical Monthly, 111(3):256–258, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145135.

Wilf:2004:P

[1236] Herbert Wilf. Problem 11068. American Mathematical Monthly, 111(3): 259, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145136.

Ivady:2004:P

[1237] Péter Ivády. Problem 11069. American Mathematical Monthly, 111(3): 259, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145137.

Tauraso:2004:P

[1238] Roberto Tauraso. Problem 11070. American Mathematical Monthly, 111(3):259, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145138.

Deutsch:2004:Pa

[1239] Emeric Deutsch. Problem 11071. American Mathematical Monthly, 111 (3):259, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145139.

Lamb:2004:Pb

[1240] George Lamb. Problem 11072. American Mathematical Monthly, 111(3): 259, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145140.

Radulescu:2004:Pa

[1241] Vicenţiu Rădulescu. Problem 11073. American Mathematical Monthly, 111(3):260, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145141.

Skordev:2004:P

[1242] Dimiter Skordev. Problem 11074. American Mathematical Monthly, 111 (3):260, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145142.

Bacher:2004:DI

[1243] Roland Bacher, Robin Chapman, and Reiner Martin. A determinant of indeterminates: 10859. American Mathematical Monthly, 111(3):260– 261, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145143.

Gross:2004:PSP

[1244] Jürgen Groß, Götz Trenkler, and Xiaoshen Wang. Product and sum of projections: 10917. American Mathematical Monthly, 111(3):261–262, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145144.

Bey:2004:MAS

[1245] Christian Bey and Garth Isaak. Maximum antichains of sets of distinct sizes: 10932. American Mathematical Monthly, 111(3):262–263, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145145.

Janous:2004:MAD

[1246] Walther Janous and J. P. V. Abad. Maximum antichains in the divisor lattice: 10942. American Mathematical Monthly, 111(3):263, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145146.

Popescu:2004:DR

[1247] Călin Popescu and GCHQ Problem Solving Group. A divisibility result: 10943. American Mathematical Monthly, 111(3):264, March 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145147.

Lagarias:2004:GBD

[1248] Jeffrey Lagarias and Walther Janous. A generous bound for divisor sums: 10949. American Mathematical Monthly, 111(3):264–265, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145148.

Callan:2004:CSR

[1249] David Callan and M. A. Prasad. Counting a set of rooted trees: 10952. American Mathematical Monthly, 111(3):265–266, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145149.

Conrad:2004:PR

[1250] Brian Conrad, Gheorge Simion, and Patrick Corn. Pi and rationals: 10961. American Mathematical Monthly, 111(3):266, March 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145150.

Little:2004:BRB

[1251] John B. Little. Book review: [untitled]. American Mathematical Monthly, 111(3):267–271, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145151.

Fulling:2004:BRB

[1252] Stephen A. Fulling. Book review: [untitled]. American Mathematical Monthly, 111(3):271–275, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145152.

Ostebee:2004:TRb

[1253] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 111(3):276–278, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145153.

Anonymous:2004:BMd

[1254] Anonymous. Back matter. American Mathematical Monthly, 111(3): 279–280, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145154.

Anonymous:2004:FMd

[1255] Anonymous. Front matter. American Mathematical Monthly, 111(3):??, March 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145125.

Goodman:2004:ATL

[1256] Roe Goodman. Alice through looking glass after looking glass: The mathematics of mirrors and kaleidoscopes. American Mathematical Monthly, 111(4):281–298, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145238.

Grimmett:2004:SA

[1257] Geoffrey Grimmett. Stochastic apportionment. American Mathematical Monthly, 111(4):299–307, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145239.

Mohazzabi:2004:EIB

[1258] Pirooz Mohazzabi and Thomas A. Fournelle. Evaluation of ill-behaved power series. American Mathematical Monthly, 111(4):308–321, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145240.

Bennett:2004:FLT

[1259] Curtis D. Bennett, A. M. W. Glass, and Gábor J. Székely. Fermat’s last theorem for rational exponents. American Mathematical Monthly, 111 (4):322–329, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145241.

Berndt:2004:TET

[1260] Bruce C. Berndt and C. A. Reddi. Two exams taken by Ramanujan in india. American Mathematical Monthly, 111(4):330–339, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145242.

Aksoy:2004:WEM

[1261] Asuman Aksoy and Mario Martelli. The wave equation, mixed partial derivatives, and Fubini’s theorem. American Mathematical Monthly, 111 (4):340–347, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145243.

Samet:2004:OOB

[1262] Dov Samet, Iddo Samet, and David Schmeidler. One observation behind two-envelope puzzles. American Mathematical Monthly, 111(4):347–351, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145244.

Kowalski:2004:RAI

[1263] E. Kowalski. On the “reducibility” of arctangents of integers. American Mathematical Monthly, 111(4):351–354, April 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145245.

Monsky:2004:KFL

[1264] Paul Monsky. Kepler’s first law: a remark. American Mathematical Monthly, 111(4):355–356, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145246.

Perdry:2004:EPK

[1265] Hervé Perdry. An elementary proof of Krull’s intersection theorem. American Mathematical Monthly, 111(4):356–357, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145247.

Mostovoy:2004:LFS

[1266] Jacob Mostovoy. Lattices in C and finite subsets of a circle. American Mathematical Monthly, 111(4):357–360, April 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145248.

Trenkler:2004:P

[1267] Götz Trenkler. Problem 11075. American Mathematical Monthly, 111(4): 361, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145249.

Zinn:2004:P

[1268] Joel Zinn. Problem 11076. American Mathematical Monthly, 111(4): 361, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145250.

Costovici:2004:P

[1269] Gheorge Costovici. Problem 11077. American Mathematical Monthly, 111(4):361, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145251.

Knuth:2004:P

[1270] Donald Knuth. Problem 11078. American Mathematical Monthly, 111(4): 361, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145252.

Cook:2004:P

[1271] Brian Cook and James Whitenton. Problem 11079. American Mathematical Monthly, 111(4):361–362, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145253.

Satnoianu:2004:Pb

[1272] Razvan Satnoianu. Problem 11080. American Mathematical Monthly, 111(4):362, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145254.

Morse:2004:P

[1273] Stephen Morse. Problem 11081. American Mathematical Monthly, 111 (4):362, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145255.

LPSG:2004:SCN

[1274] LSU Problem Solving Group and Marc A. Chamberland. A sequence of composite numbers: 10947. American Mathematical Monthly, 111 (4):362–363, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145256.

Kim:2004:SP

[1275] Sung Soo Kim and Li Zhou. A sticky problem: 10951. American Mathematical Monthly, 111(4):363–364, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145257.

McCarthy:2004:TS

[1276] John McCarthy, Lee Badger, and Marc Williams. The tower of stanford: 10956. American Mathematical Monthly, 111(4):364–365, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145258.

Eng:2004:SD

[1277] Oliver Eng and Li Zhou. Sliding dominoes: 10960. American Mathematical Monthly, 111(4):365–366, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145259.

Glosser:2004:NZ

[1278] M. L. Glosser and J. A. Grzesik. A nontrivial zero.: 11020. American Mathematical Monthly, 111(4):366–367, April 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145260.

Radulescu:2004:LP

[1279] Vicenţiu Rădulescu and Richard Stong. A limit problem: 11024. American Mathematical Monthly, 111(4):367–368, April 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145261.

Farkas:2004:BRB

[1280] Daniel R. Farkas. Book review: [untitled]. American Mathematical Monthly, 111(4):369–372, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145262.

Zorn:2004:BRB

[1281] Paul Zorn. Book review: [untitled]. American Mathematical Monthly, 111 (4):373–375, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145263.

Anonymous:2004:BMe

[1282] Anonymous. Back matter. American Mathematical Monthly, 111(4): 376, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145264.

Anonymous:2004:FMe

[1283] Anonymous. Front matter. American Mathematical Monthly, 111(4): ??, April 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145237.

Saari:2004:GCS

[1284] Donald G. Saari. Geometry of chaotic and stable discussions. American Mathematical Monthly, 111(5):377–393, May 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145266.

Ahmed:2004:HMS

[1285] Maya Mohsin Ahmed. How many squares are there, Mr. Franklin?: Constructing and enumerating Franklin squares. American Mathematical Monthly, 111(5):394–410, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145267.

Edwards:2004:SME

[1286] Barbara S. Edwards and Michael B. Ward. Surprises from mathematics education research: Student (mis)use of mathematical definitions. American Mathematical Monthly, 111(5):411–424, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145268.

Spandaw:2004:DCC

[1287] Jeroen Spandaw. Dissecting cuboids into cuboids. American Mathematical Monthly, 111(5):425–429, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145269.

Tsumura:2004:EPE

[1288] Hirofumi Tsumura. An elementary proof of Euler’s formula for ζ(2m). American Mathematical Monthly, 111(5):430–431, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145270.

Rzadkowski:2004:SPE

[1289] Grzegorz Rza̧dkowski. A short proof of the explicit formula for Bernoulli numbers. American Mathematical Monthly, 111(5):432–434, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145271.

Boute:2004:MRA

[1290] Raymond T. Boute. Moving a rectangle around a corner: Geometrically. American Mathematical Monthly, 111(5):435–437, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145272.

Leuenberger:2004:OFR

[1291] Christoph Leuenberger. An old friend revisited. American Mathematical Monthly, 111(5):437–439, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145273.

Chao:2004:Pb

[1292] Wu Wei Chao. Problem 11082. American Mathematical Monthly, 111(5): 440, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145274.

Wardlaw:2004:P

[1293] William P. Wardlaw. Problem 11083. American Mathematical Monthly, 111(5):440, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145275.

ShahAli:2004:Pa

[1294] H. A. ShahAli. Problem 11084. American Mathematical Monthly, 111(5): 440, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145276.

Beck:2004:Pa

[1295] Matthias Beck, Jesus DeLoera, Mike Develin, and Julian Pfeifle. Problem 11085. American Mathematical Monthly, 111(5):440, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145277.

Locke:2004:P

[1296] Stephen C. Locke. Problem 11086. American Mathematical Monthly, 111 (5):440–441, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145278.

Chapman:2004:EBM

[1297] Robin Chapman and Nicholas C. Singer. Eigenvalues of a bidiagonal matrix: 10958. American Mathematical Monthly, 111(5):441, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145281.

Golomb:2004:Pa

[1298] Michael Golomb. Problem 11087. American Mathematical Monthly, 111 (5):441, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145279.

Holzsager:2004:P

[1299] Richard Holzsager. Problem 11088. American Mathematical Monthly, 111(5):441, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145280.

Deiermann:2004:AMT

[1300] Paul Deiermann, Cape Girardeau, Joseph Chalmers, and George L. Lamb, Jr. An application of the Mellin transform: 10963. American Mathematical Monthly, 111(5):441–443, May 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145282.

West:2004:AMI

[1301] Douglas B. West and Csaba Megyeri. Adjacency matrices that are incidence matrices: 10967. American Mathematical Monthly, 111(5):443, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145283.

Locke:2004:GP

[1302] Stephen C. Locke and Philip D. Straffin. The game of pathos: 10968. American Mathematical Monthly, 111(5):443–444, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145284.

Voloshin:2004:PSN

[1303] Vitaly Voloshin, José Heber Nieto, Edward R. Scheinerman, and Robin Chapman. Plucking Stirling numbers from trees: 10976. American Mathematical Monthly, 111(5):444–445, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145285.

Axness:2004:IGE

[1304] Carl Axness, Bert Kerr, and Jean-François Chamayou. Incomplete gamma and error functions: 10977. American Mathematical Monthly, 111(5):445–447, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145286.

Knuth:2004:SBP

[1305] Donald Knuth and BSI Problems Group. Some Bernstein polynomials: 10985. American Mathematical Monthly, 111(5):447, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145287. See errata [1449].

Fasshauer:2004:BRB

[1306] G. E. Fasshauer. Book review: [untitled]. American Mathematical Monthly, 111(5):448–452, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145288.

Ostebee:2004:TRc

[1307] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 111(5):453–455, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145289.

Palka:2004:EEb

[1308] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 111 (5):456–460, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145290.

Anonymous:2004:BMf

[1309] Anonymous. Back matter. American Mathematical Monthly, 111(5):461– 464, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145291.

Anonymous:2004:FMf

[1310] Anonymous. Front matter. American Mathematical Monthly, 111(5): ??, May 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145265.

Zitarelli:2004:OEI

[1311] David E. Zitarelli. The origin and early impact of the Moore method. American Mathematical Monthly, 111(6):465–486, June/July 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145066.

Cohn:2004:PGG

[1312] Henry Cohn. Projective geometry over F_{1} and the Gaussian binomial coefficients. American Mathematical Monthly, 111(6):487–495, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145067.

Apostol:2004:FLM

[1313] Tom M. Apostol and Mamikon A. Mnatsakanian. A fresh look at the method of archimedes. American Mathematical Monthly, 111(6):496–508, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145068.

Mann:2004:HTP

[1314] Casey Mann. Heesch’s tiling problem. American Mathematical Monthly, 111(6):509–517, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145069.

Helmberg:2004:CC

[1315] Gilbert Helmberg. A construction concerning (l^{p})′ ⊂ l^{q}. American Mathematical Monthly, 111(6):518–520, June/July 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145070.

Zhu:2004:TIB

[1316] Kehe Zhu. Translating inequalities between Hardy and Bergman spaces. American Mathematical Monthly, 111(6):520–525, June/July 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145071.

Wang:2004:SPD

[1317] Xiaoshen Wang. A simple proof of Descartes’s rule of signs. American Mathematical Monthly, 111(6):525–526, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145072.

Nadler:2004:APR

[1318] Sam B. Nadler, Jr. Another proof that R^{3} has no square root. American Mathematical Monthly, 111(6):527–528, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145073.

Weintraub:2004:IR

[1319] Steven H. Weintraub. An interesting recursion. American Mathematical Monthly, 111(6):528–530, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145074.

Dresden:2004:MCC

[1320] Gregory P. Dresden. On the middle coefficient of a cyclotomic polynomial. American Mathematical Monthly, 111(6):531–533, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145075.

Kim:2004:P

[1321] Sun Soo Kim. Problem 11089. American Mathematical Monthly, 111 (6):534, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145076.

See errata [1445].

Luca:2004:P

[1322] Florian Luca. Problem 11090. American Mathematical Monthly, 111 (6):534, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145077.

Callan:2004:Pa

[1323] David Callan. Problem 11091. American Mathematical Monthly, 111 (6):534, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145078.

Magidin:2004:P

[1324] Arturo Magidin. Problem 11092. American Mathematical Monthly, 111 (6):534–535, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145079.

Bagby:2004:P

[1325] Richard Bagby. Problem 11093. American Mathematical Monthly, 111 (6):535, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145080.

Bataille:2004:P

[1326] Michel Bataille. Problem 11094. American Mathematical Monthly, 111 (6):535, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145081.

Mihai:2004:P

[1327] Vasile Mihai. Problem 11095. American Mathematical Monthly, 111 (6):535, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145082.

ShahAli:2004:GPR

[1328] H. A. ShahAli and Hamza Ahmad. Graceful permutations revisited: 10964. American Mathematical Monthly, 111(6):535–536, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145083.

DeKoninck:2004:WTP

[1329] Jean-Marie De Koninck and Stefan Gubo. When the totient is the product of the squared prime divisors: 10966. American Mathematical Monthly, 111(6):536, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145084.

Stanley:2004:COP

[1330] Richard Stanley and O. P. Lossers. Comparing odd parts in conjugate partitions: 10969. American Mathematical Monthly, 111(6):536–539, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145085.

Rupert:2004:FFN

[1331] C. P. Rupert and Richard Stong. The first Fibonacci number divisible by a prime p: 10971. American Mathematical Monthly, 111(6):539–540, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145086.

Ferguson:2004:LPP

[1332] Tim Ferguson, M. D. Frederick, Lenny Jones, and Stephen M. Gagola, Jr. The LCM of the parts of a partition: 10972. American Mathematical Monthly, 111(6):540–541, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145087.

Deutsch:2004:EPDa

[1333] Emeric Deutsch, Julio Kuplinsky, and Tina Garrett. Enumeration of permutations by disorder: 10975. American Mathematical Monthly, 111 (6):541, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145088.

Shier:2004:BRB

[1334] Douglas R. Shier. Book review: [untitled]. American Mathematical Monthly, 111(6):542–545, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145089.

Anonymous:2004:BMg

[1335] Anonymous. Back matter. American Mathematical Monthly, 111(6): 546–548, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145090.

Anonymous:2004:FMg

[1336] Anonymous. Front matter. American Mathematical Monthly, 111(6):??, June/July 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145065.

Leitzel:2004:JRC

[1337] Joan R. Leitzel. The 2003 James R. C. Leitzel Lecture: Mathematics’ response to issues in higher education. American Mathematical Monthly, 111(7):549–558, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145156.

Diestel:2004:ICG

[1338] Reinhard Diestel. On infinite cycles in graphs: Or how to make graph homology interesting. American Mathematical Monthly, 111(7):559–571, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145157.

Ekedahl:2004:TGP

[1339] Torsten Ekedahl and Dan Laksov. Two “generic” proofs of the spectral mapping theorem. American Mathematical Monthly, 111(7):572–585, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145158.

Ardila:2004:NHC

[1340] Federico Ardila. The number of halving circles. American Mathematical Monthly, 111(7):586–591, August/September 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145159.

Veljan:2004:GC

[1341] Darko Veljan. Geometry and convexity of cos \sqrt{x}. American Mathematical Monthly, 111(7):592–595, August/September 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145160.

Du:2004:SPS

[1342] Bau-Sen Du. A simple proof of Sharkovsky’s theorem. American Mathematical Monthly, 111(7):595–599, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145161.

Jarvis:2004:HBT

[1343] Tyler Jarvis and James Tanton. The hairy Ball theorem via Sperner’s lemma. American Mathematical Monthly, 111(7):599–603, August/ September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145162.

Sato:2004:SSI

[1344] Kenzi Satô. Some subgroups of SO_{3}(R) isomorphic to the free product (Z/2Z) ∗ (Z/3Z). American Mathematical Monthly, 111(7):603–607, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145163.

Arnold:2004:HSP

[1345] V. I. Arnol’d. From Hilbert’s superposition problem to dynamical systems. American Mathematical Monthly, 111(7):608–624, August/ September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145164.

Amghibech:2004:P

[1346] Said Amghibech. Problem 11096. American Mathematical Monthly, 111 (7):625, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145165.

Ford:2004:P

[1347] Kevin Ford. Problem 11097. American Mathematical Monthly, 111(7): 625, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145166.

Hillar:2004:P

[1348] Christopher Hillar and Darren Rhea. Problem 11098. American Mathematical Monthly, 111(7):625, August/September 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145167.

Beck:2004:Pb

[1349] Matthias Beck, Richard Ehrenborg, and Thomas Zaslavsky. Problem 11099. American Mathematical Monthly, 111(7):625–626, August/ September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145168.

Popescu:2004:P

[1350] Călin Popescu. Problem 11100. American Mathematical Monthly, 111 (7):626, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145169.

Skelton:2004:P

[1351] Earl F. Skelton. Problem 11101. American Mathematical Monthly, 111 (7):626, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145170.

Quet:2004:P

[1352] Leroy Quet. Problem 11102. American Mathematical Monthly, 111(7): 626, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145171.

Satnoianu:2004:CSS

[1353] Razvan Satnoianu and John G. Heuver. Comparison of some squared distances: 10970. American Mathematical Monthly, 111(7):626–627, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145172.

Servi:2004:DHA

[1354] L. D. Servi and Richard Stong. A disguised half-angle iteration: 10973. American Mathematical Monthly, 111(7):627–628, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145173.

Marivani:2004:DRF

[1355] Syrous Marivani and Bill Mixon. Digital roots of Fibonacci numbers: 10974. American Mathematical Monthly, 111(7):628, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145174.

Deutsch:2004:EPDb

[1356] Emeric Deutsch, Julio Kuplinsky, and Tina Garrett. Enumeration of permutations by disorder: 10975. American Mathematical Monthly, 111(7): 628–629, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145175.

Grivaux:2004:NP

[1357] Jean-Pierre Grivaux and Tsehaye Andebrhan. Nonnegative polynomials: 10978. American Mathematical Monthly, 111(7):629–630, August/ September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145176.

Minassian:2004:SD

[1358] Donald P. Minassian and K. P. Hart. Simple discontinuities: 10979. American Mathematical Monthly, 111(7):630, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145177.

Singer:2004:DPR

[1359] Nicholas C. Singer. A divisibility property of a recurrence: 10981. American Mathematical Monthly, 111(7):631, August/September 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145178.

Chen:2004:VUO

[1360] Yuming Chen, Olaf Krafft, and Martin Schaefer. Variation of a ukrainian olympiad problem: 10982. American Mathematical Monthly, 111(7):631– 632, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145179.

Wilf:2004:ACS

[1361] Herbert Wilf and J. A. Grzesik. An arc cotangent series: 10986. American Mathematical Monthly, 111(7):632, August/September 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145180.

Boyd:2004:BRB

[1362] David W. Boyd. Book review: [untitled]. American Mathematical Monthly, 111(7):633–636, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/4145181.

Benjamin:2004:BRB

[1363] Arthur T. Benjamin. Book review: [untitled]. American Mathematical Monthly, 111(7):636–638, August/September 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145182.

Ostebee:2004:TRd

[1364] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 111(7):639–641, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145183.

Anonymous:2004:BMh

[1365] Anonymous. Back matter. American Mathematical Monthly, 111(7): 642–644, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145184.

Anonymous:2004:FMh

[1366] Anonymous. Front matter. American Mathematical Monthly, 111(7):??, August/September 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145155.

Finch:2004:LF

[1367] Steven R. Finch and John E. Wetzel. Lost in a forest. American Mathematical Monthly, 111(8):645–654, October 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145038.

Holtz:2004:PWD

[1368] Olga Holtz, Volker Mehrmann, and Hans Schneider. Potter, Wielandt, and Drazin on the matrix equation AB = ωBA: New answers to old questions. American Mathematical Monthly, 111(8):655–667, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145039.

Anderson:2004:IE

[1369] Joel Anderson. Iterated exponentials. American Mathematical Monthly, 111(8):668–679, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145040.

Klosinski:2004:SFW

[1370] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Fourth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 111(8):680–690, October 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145041.

Gallian:2004:FSS

[1371] Joseph A. Gallian. The first sixty-six years of the Putnam Competition. American Mathematical Monthly, 111(8):691–699, October 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145042.

Glass:2004:SPH

[1372] Kenneth Glass and Chi-Keung Ng. A simple proof of the hook length formula. American Mathematical Monthly, 111(8):700–704, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145043.

Devaney:2004:PCI

[1373] R. L. Devaney, K. Josić, M. Moreno Rocha, P. Seal, Y. Shapiro, and A. T. Frumosu. Playing catch-up with iterated exponentials. American Mathematical Monthly, 111(8):704–709, October 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145044.

Klain:2004:IDH

[1374] Daniel A. Klain. An intuitive derivation of Heron’s formula. American Mathematical Monthly, 111(8):709–712, October 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145045.

Olsen:2004:NPD

[1375] Lars Olsen. A new proof of Darboux’s theorem. American Mathematical Monthly, 111(8):713–715, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145046.

Atiyah:2004:BLG

[1376] M. F. Atiyah. Bakerian lecture, 1975: Global geometry. American Mathematical Monthly, 111(8):716–723, October 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145047.

Galperin:2004:P

[1377] Gregory Galperin and Hillel Gauchman. Problem 11103. American Mathematical Monthly, 111(8):724, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145048.

Radulescu:2004:Pb

[1378] Vicenţiu Rădulescu. Problem 11104. American Mathematical Monthly, 111(8):724, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145049.

Marquez:2004:P

[1379] Juan-Bosco Romero Marquez, Rick Luttmann, Miguel de Guzman, and G. Donald Chakerian. Problem 11105. American Mathematical Monthly, 111(8):724–725, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145050.

Callan:2004:Pb

[1380] David Callan. Problem 11106. American Mathematical Monthly, 111 (8):725, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145051.

Emery:2004:P

[1381] Michel Émery. Problem 11107. American Mathematical Monthly, 111 (8):725, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145052.

Deutsch:2004:Pb

[1382] Emeric Deutsch. Problem 11108. American Mathematical Monthly, 111 (8):725, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145053.

Golomb:2004:Pb

[1383] Solomon W. Golomb. Problem 11109. American Mathematical Monthly, 111(8):725, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145054.

Beck:2004:Pc

[1384] Matthias Beck, Jesus DeLoera, Mike Develin, and Julian Pfeifle. Problem 11085. American Mathematical Monthly, 111(8):725, October 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145055.

Chao:2004:CAF

[1385] Wu Wei Chao and Doyle Henderson. Continuous additive functions: 10854. American Mathematical Monthly, 111(8):725–726, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145056. See errata [1715].

Melville:2004:CPT

[1386] John Melville, Jason Heller, and Stony Brook. A classic pointed triangle inequality: 10965. American Mathematical Monthly, 111(8):726, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145057.

Suceava:2004:OP

[1387] Bogdan Suceavă, C. G. Petalas, and T. P. Vidalis. Orthocenter projection: 10983. American Mathematical Monthly, 111(8):726–728, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145058.

Tian:2004:IPC

[1388] Yongge Tian and Carl Meyer. Injective properties of complex matrices: 10984. American Mathematical Monthly, 111(8):728, October 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145059.

Prasanna:2004:PNA

[1389] Kumar N. Prasanna, B. G. Shenoy, and J. A. Grzesik. Polynomial Nullstellen avoidance: 10987. American Mathematical Monthly, 111(8):729, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145060.

Chao:2004:IS

[1390] Wu Wei Chao, Doyle Henderson, and John H. Lindsey, II. An iterative sequence: 10988. American Mathematical Monthly, 111(8):729–731, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145061. See errata

[1448].

Krantz:2004:BRB

[1391] Steven G. Krantz. Book review: [untitled]. American Mathematical Monthly, 111(8):732–735, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145062.

Cohen:2004:BRB

[1392] Marion D. Cohen. Book review: [untitled]. American Mathematical Monthly, 111(8):735–738, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145063.

Anonymous:2004:BMi

[1393] Anonymous. Back matter. American Mathematical Monthly, 111(8): 739–740, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145064.

Anonymous:2004:FMi

[1394] Anonymous. Front matter. American Mathematical Monthly, 111(8):??, October 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145037.

Adams:2004:BCR

[1395] Colin Adams, Thomas Fleming, and Christopher Koegel. Brunnian clothes on the runway: Not for the bashful. American Mathematical Monthly, 111(9):741–748, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145186.

Friedgut:2004:HEI

[1396] Ehud Friedgut. Hypergraphs, entropy, and inequalities. American Mathematical Monthly, 111(9):749–760, November 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145187.

Garibaldi:2004:CPD

[1397] Skip Garibaldi. The characteristic polynomial and determinant are not ad hoc constructions. American Mathematical Monthly, 111(9):761–778, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145188.

Frantz:2004:HCG

[1398] Marc Frantz. How conics govern Möbius transformations. American Mathematical Monthly, 111(9):779–790, November 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145189.

Johnson:2004:HCM

[1399] Warren P. Johnson. How Cauchy missed Ramanujan’s _{1}ψ_{1} summation. American Mathematical Monthly, 111(9):791–800, November 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145190.

vanMill:2004:NET

[1400] Jan van Mill. A note on the effros theorem. American Mathematical Monthly, 111(9):801–806, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145191.

Chan:2004:EPJ

[1401] Song Heng Chan. An elementary proof of Jacobi’s six squares theorem. American Mathematical Monthly, 111(9):806–811, November 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145192.

Barth:2004:CCQ

[1402] Nils R. Barth. Computing Cavalieri’s quadrature formula by a symmetry of the n-cube. American Mathematical Monthly, 111(9):811–813, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145193.

Kazanci:2004:PNT

[1403] Deniz Kazanci and Andrew Vince. A property of normal tilings. American Mathematical Monthly, 111(9):813–816, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145194.

Nowojewski:2004:AMB

[1404] Andrzej Nowojewski, Jakub Kallas, Andrzej Dragan, Abe Shenitzer, Stefan Mykytiuk, and John Stillwell. On the appearance of moving bodies. American Mathematical Monthly, 111(9):817–821, November 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145195.

Ghergu:2004:P

[1405] Marius Ghergu. Problem 11110. American Mathematical Monthly, 111 (9):822, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145196.

Dalyay:2004:P

[1406] Pál Péter Dalyay. Problem 11111. American Mathematical Monthly, 111 (9):822, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145197.

Tian:2004:P

[1407] Yongge Tian. Problem 11112. American Mathematical Monthly, 111 (9):822, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145198.

Sofair:2004:P

[1408] Isaac Sofair. Problem 11113. American Mathematical Monthly, 111(9): 822, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145199.

deWeger:2004:P

[1409] Benne de Weger. Problem 11114. American Mathematical Monthly, 111 (9):823, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145200.

Clark:2004:P

[1410] Jack Clark. Problem 11115. American Mathematical Monthly, 111(9): 823, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145201.

Veljan:2004:P

[1411] Darko Veljan. Problem 11116. American Mathematical Monthly, 111 (9):823, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145202.

Hao:2004:GP

[1412] Ye Zhong Hao, Wu Wei Chao, and J. Schaer. Gossard perspector: 10980. American Mathematical Monthly, 111(9):823–824, November 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145203.

Bacher:2004:FP

[1413] Roland Bacher, Julio Kuplinsky, and Richard Stong. A family of polynomials: 10989. American Mathematical Monthly, 111(9):824–825, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145204.

Mabry:2004:EP

[1414] Rick Mabry and Bernardo M. Ábrego. The envelope, please: 10990. American Mathematical Monthly, 111(9):825–826, November 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145205.

Mortini:2004:NMD

[1415] Raymond Mortini and S. Amghibech. A nonholomorphic map of the disk: 10991. American Mathematical Monthly, 111(9):826–827, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145206.

Lubell:2004:ESR

[1416] David Lubell, Chris Hill, and Albert Nijenhuis. Equal sums of reciprocal products: 10992. American Mathematical Monthly, 111(9):827–829, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145207.

Savitt:2004:RDS

[1417] David Savitt and Chris Hill. Recognizing n-divergent sequences: 10994. American Mathematical Monthly, 111(9):829, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145208.

Kauffman:2004:BRB

[1418] Louis H. Kauffman. Book review: [untitled]. American Mathematical Monthly, 111(9):830–838, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145209.

Anonymous:2004:BMj

[1419] Anonymous. Back matter. American Mathematical Monthly, 111(9): 839–840, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145210.

Anonymous:2004:FMj

[1420] Anonymous. Front matter. American Mathematical Monthly, 111(9):??, November 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145185.

Grabiner:2004:NMA

[1421] Judith V. Grabiner. Newton, maclaurin, and the authority of mathematics. American Mathematical Monthly, 111(10):841–852, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/4145093.

Apostol:2004:FCC

[1422] Tom M. Apostol and Mamikon A. Mnatsakanian. Figures circumscribing circles. American Mathematical Monthly, 111(10):853–863, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145094.

Logemann:2004:ABN

[1423] Hartmut Logemann and Eugene P. Ryan. Asymptotic behaviour of nonlinear systems. American Mathematical Monthly, 111(10):864–889, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145095.

Edwards:2004:ECP

[1424] Stephanie Edwards and Russell A. Gordon. Extreme curvature of polynomials. American Mathematical Monthly, 111(10):890–899, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145096.

Weizsacker:2004:COD

[1425] Heinrich v. Weizsäcker. Can one drop L^{1}-boundedness in Komlós’s subsequence theorem? American Mathematical Monthly, 111(10):900–903, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145097.

Chakerian:2004:TPC

[1426] G. D. Chakerian and M. S. Klamkin. A three-point characterization of central symmetry. American Mathematical Monthly, 111(10):903–905, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145098.

Pinelis:2004:LRM

[1427] Iosif Pinelis. L’Hospital rules for monotonicity and the Wilker–Anglesio Inequality. American Mathematical Monthly, 111(10):905–909, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/4145099.

Kimura:2004:IMI

[1428] Noriaki Kimura and Kenneth S. Williams. Infinitely many insolvable Diophantine equations. American Mathematical Monthly, 111(10):909– 913, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/4145100.

Stefansson:2004:FSB

[1429] Jón R. Stefánsson. Forward shifts and backward shifts in a rearrangement of a conditionally convergent series. American Mathematical Monthly, 111(10):913–914, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145101.

Nyblom:2004:P

[1430] Michael Nyblom. Problem 11117. American Mathematical Monthly, 111 (10):915, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145102.

Dimitrov:2004:P

[1431] Vesselin Dimitrov. Problem 11118. American Mathematical Monthly, 111 (10):915, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145103.

Robertson:2004:P

[1432] John Robertson. Problem 11119. American Mathematical Monthly, 111 (10):915, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145104.

Berk:2004:P

[1433] Bob Berk and Larry Shepp. Problem 11120. American Mathematical Monthly, 111(10):915, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145105.

ShahAli:2004:Pb

[1434] H. A. ShahAli. Problem 11121. American Mathematical Monthly, 111 (10):916, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145106.

Pambuccian:2004:P

[1435] Victor Pambuccian. Problem 11122. American Mathematical Monthly, 111(10):916, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145107.

Dubickas:2004:P

[1436] Artūras Dubickas. Problem 11123. American Mathematical Monthly, 111 (10):916, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145108.

Goldoni:2004:EI

[1437] Giorgi Goldoni and Alan Berele. An excenteric iteration: 10993. American Mathematical Monthly, 111(10):916–917, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145109.

Trigiante:2004:WZ

[1438] Donato Trigiante and Albert Stadler. Where are the zeros?: 10995. American Mathematical Monthly, 111(10):917, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145110.

Prothro:2004:PDF

[1439] Ed T. Prothro and Chris Hill. Prime divisors of Fermat numbers: 10997. American Mathematical Monthly, 111(10):917–918, December 2004. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/4145111.

Tauraso:2004:DB

[1440] Roberto Tauraso and Jiri Lebl. Distance to the boundary: 10998. American Mathematical Monthly, 111(10):918, December 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145112.

Lamb:2004:ITW

[1441] George Lamb, Cornelius Stallmann, Gerald Thompson, Zafar Ahmed, and O. P. Lossers. An integral three ways: 11000. American Mathematical Monthly, 111(10):918–920, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/4145113.

Mabry:2004:PTS

[1442] Rick Mabry and Jody Lockhart. Powered terms in a series: 11001. American Mathematical Monthly, 111(10):920, December 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145114.

Petermann:2004:PBR

[1443] Y.-F. S. Pétermann, C. T. Stretch, and Jamie Simpson. A problem that bears repeating: 11002. American Mathematical Monthly, 111(10): 920–921, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145115.

Gray:2004:OCI

[1444] Stephen B. Gray and Clayton W. Dodge. The orthocenter and circumcenter are isogonal conjugate points: 11003. American Mathematical Monthly, 111(10):921–922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/4145116.

Anonymous:2004:E

[1445] Anonymous. Errata for 2004: 11089. American Mathematical Monthly, 111(10):922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/ 4145118. See [1321].

Anonymous:2004:EFC

[1446] Anonymous. Errata for 2004: Fibonacci in complex camouflage: 10858. American Mathematical Monthly, 111(10):922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145117. See errata [1215].

Anonymous:2004:EGD

[1447] Anonymous. Errata for 2004: a group determined by relations?: 10933. American Mathematical Monthly, 111(10):922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145120. See [1219].

Anonymous:2004:EIS

[1448] Anonymous. Errata for 2004: An iterative sequence: 10988. American Mathematical Monthly, 111(10):922, December 2004. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/4145121. See [1390].

Anonymous:2004:ESB

[1449] Anonymous. Errata for 2004: Some Bernstein polynomials: 10985. American Mathematical Monthly, 111(10):922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/4145119. See [1305].

Protter:2004:BRB

[1450] Philip Protter. Book review: [untitled]. American Mathematical Monthly, 111(10):923–926, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145122.

Palka:2004:EEa

[1451] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 111(10):927–929, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145123.

Anonymous:2004:VI

[1452] Anonymous. Volume information. American Mathematical Monthly, 111 (10):930–937, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

4145091.

Anonymous:2004:BMb

[1453] Anonymous. Back matter. American Mathematical Monthly, 111(10): 938–944, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/4145124.

Anonymous:2004:FMb

[1454] Anonymous. Front matter. American Mathematical Monthly, 111(10):??, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/4145092.

Dunham:2005:TCG

[1455] William Dunham. Touring the calculus gallery. American Mathematical Monthly, 112(1):1–19, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037380.

Wild:2005:GOS

[1456] Marcel Wild. The groups of order sixteen made easy. American Mathematical Monthly, 112(1):20–31, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037381.

Berhanu:2005:HIO

[1457] Shif Berhanu and Ahmed Mohammed. A Harnack inequality for ordinary differential equations. American Mathematical Monthly, 112(1):32–41, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037382.

Hernandez:2005:FCW

[1458] Janko Hernández and Leonel Robert. Figures of constant width on a chessboard. American Mathematical Monthly, 112(1):42–50, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037383.

Bear:2005:SAI

[1459] H. S. Bear and Dale Myers. A simpler approach to integration and the Fubini theorem. American Mathematical Monthly, 112(1):51–60, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037384.

Sondow:2005:DIE

[1460] Jonathan Sondow. Double integrals for Euler’s constant and in ln(4/π) and an analog of Hadjicostas’s formula. American Mathematical Monthly, 112(1):61–65, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037385.

Romik:2005:RDP

[1461] Dan Romik. Roots of the derivative of a polynomial. American Mathematical Monthly, 112(1):66–69, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037386.

Weiner:2005:QR

[1462] Joel L. Weiner and George R. Wilkens. Quaternions and rotations in E^{4}. American Mathematical Monthly, 112(1):69–76, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037387.

deSousa:2005:APF

[1463] José Carlos de Sousa and Oliveira Santos. Another proof of the fundamental theorem of algebra. American Mathematical Monthly, 112(1): 76–78, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037388.

Kagan:2005:RGI

[1464] V. F. Kagan. Riemann’s geometric ideas. American Mathematical Monthly, 112(1):79–86, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037389.

Nica:2005:P

[1465] Bogdan Nica. Problem 11124. American Mathematical Monthly, 112(1): 87, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037390.

Amrahov:2005:Pa

[1466] Shahin Amrahov. Problem 11125. American Mathematical Monthly, 112 (1):87, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037391.

Weber:2005:P

[1467] James S. Weber. Problem 11126. American Mathematical Monthly, 112 (1):87, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037392.

Zhou:2005:Pa

[1468] Li Zhou. Problem 11127. American Mathematical Monthly, 112(1):87, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037393.

Chao:2005:PCE

[1469] Wu Wei Chao and O. P. Lossers. Point causing equal angles: 10996. American Mathematical Monthly, 112(1):88, January 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037397.

Apostol:2005:P

[1470] Tom Apostol. Problem 11128. American Mathematical Monthly, 112(1): 88, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037394.

Tian:2005:Pa

[1471] Yongge Tian. Problem 11129. American Mathematical Monthly, 112(1): 88, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037395.

Bateman:2005:P

[1472] Paul Bateman. Problem 11130. American Mathematical Monthly, 112 (1):88, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037396.

Pambuccian:2005:SQ

[1473] Victor Pambuccian and Michael Wolterman. Saccheri quadrilateral: 11004. American Mathematical Monthly, 112(1):88–89, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037398.

Getz:2005:PAP

[1474] Marty Getz, Dixon Jones, and Ken Dutch. Partitioning by arithmetic progressions: 11005. American Mathematical Monthly, 112(1):89–90, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037399. See er-

rata [1716].

Suceava:2005:CCE

[1475] Bogdan Suceavă and Daniele Donini. A circle centered on the Euler circle: 11006. American Mathematical Monthly, 112(1):90–91, January 2005.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037400.

Diaz-Barrero:2005:ZCI

[1476] José Luis Díaz-Barrero, Juan José Egozcue, and Artūras Dubickas. Zero and coefficient inequality: 11008. American Mathematical Monthly, 112(1):91–92, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037401.

Borwein:2005:CCF

[1477] David Borwein and Julien Grivaux. Convex and concave functions: 11009. American Mathematical Monthly, 112(1):92–93, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037402.

Just:2005:RPI

[1478] Erwin Just, Norman Schaumberger, and GCHQ Problem Solving Group. Regular polygon inscription: 11010. American Mathematical Monthly, 112(1):93–94, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037403.

Kennedy:2005:BRB

[1479] Steve Kennedy. Book review: Turing: A Novel About Computation, by Christos H. Papadimitriou. American Mathematical Monthly, 112(1): 95–98, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037404.

Cohen:2005:BRB

[1480] Marion D. Cohen. Book review: The Magic of Numbers, by Benedict Gross; Joe Harris. American Mathematical Monthly, 112(1):98–101, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037405.

Anonymous:2005:BMa

[1481] Anonymous. Back matter. American Mathematical Monthly, 112(1):??, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037406.

Anonymous:2005:FMa

[1482] Anonymous. Front matter. American Mathematical Monthly, 112(1):??, January 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037379.

Warrington:2005:JP

[1483] Gregory S. Warrington. Juggling probabilities. American Mathematical Monthly, 112(2):105–118, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037409.

Craik:2005:PFB

[1484] Alex D. D. Craik. Prehistory of Faà di Bruno’s formula. American Mathematical Monthly, 112(2):119–130, February 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037410.

Codeca:2005:LKF

[1485] Paolo Codecà and Mohan Nair. The lesser-known Δ-function in number theory. American Mathematical Monthly, 112(2):131–140, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037411.

Paiva:2005:PAU

[1486] Juan Carlos Álvarez Paiva and Anthony Thompson. On the perimeter and area of the unit disc. American Mathematical Monthly, 112(2):141– 154, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037412.

Doust:2005:TIL

[1487] Ian Doust, Michael D. Hirschhorn, and Jocelyn Ho. Trigonometric identities, linear algebra, and computer algebra. American Mathematical Monthly, 112(2):155–164, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037413.

Campbell:2005:SPP

[1488] Connie Campbell and William Staton. A square-packing problem of Erdős. American Mathematical Monthly, 112(2):165–167, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037414.

Panaitopol:2005:RNN

[1489] Laureiţiu Panaitopol. On the representation of natural numbers as sums of squares. American Mathematical Monthly, 112(2):168–171, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037415.

Ding:2005:PFT

[1490] Jiu Ding and Temple H. Fay. The Perron–Frobenius theorem and limits in geometry. American Mathematical Monthly, 112(2):171–175, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037416.

Hajlasz:2005:HMV

[1491] Piotr Hajłasz and Paweł Strzelecki. How to measure volume with a thread. American Mathematical Monthly, 112(2):176–179, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037417.

deBruijn:2005:P

[1492] N. G. de Bruijn. Problem 11131. American Mathematical Monthly, 112(2):180, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037418.

Sondow:2005:P

[1493] Jonathan Sondow. Problem 11132. American Mathematical Monthly, 112(2):180, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037419.

Bracken:2005:Pa

[1494] Paul Bracken. Problem 11133. American Mathematical Monthly, 112(2): 180, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037420.

Boklan:2005:P

[1495] Kent D. Boklan. Problem 11134. American Mathematical Monthly, 112(2):180, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037421.

Redheffer:2005:P

[1496] Ray Redheffer. Problem 11135. American Mathematical Monthly, 112(2): 181, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037422.

Mortini:2005:P

[1497] Raymond Mortini. Problem 11136. American Mathematical Monthly, 112(2):181, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037423.

Radulescu:2005:Pa

[1498] Vicenţiu Rădulescu. Problem 11137. American Mathematical Monthly, 112(2):181, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037424.

Minassian:2005:BCO

[1499] D. P. Minassian. Baire class one: 10999. American Mathematical Monthly, 112(2):181–182, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037425.

Lamb:2005:AGI

[1500] George Lamb and Zafar Ahmed. An application of the Gaussian integral: 11011. American Mathematical Monthly, 112(2):182–183, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037426.

Popescu:2005:CSI

[1501] Călin Popescu and Chu Wenchang. Cube sum inequalities: 11012. American Mathematical Monthly, 112(2):183–184, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037427.

Callan:2005:DPS

[1502] David Callan and Marc Renault. Dyck paths skipping every fourth point: 11013. American Mathematical Monthly, 112(2):184, February 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037428.

Sinefakopoulos:2005:RDS

[1503] Achilleas Sinefakopoulos and Knut Dale. A recursively defined series: 11014. American Mathematical Monthly, 112(2):184–185, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037429.

Furdui:2005:ST

[1504] Ovidiu Furdui and Li Zhou. Sines in a triangle: 11015. American Mathematical Monthly, 112(2):185–186, February 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037430.

Pranesachar:2005:WBP

[1505] C. R. Pranesachar and S. Amghibech. When the brocard points are collinear with a vertex: 11017. American Mathematical Monthly, 112 (2):186–187, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037431.

Weber:2005:BRB

[1506] Eric S. Weber. Book review: An Introduction to Frames and Riesz Bases, by Ole Christensen. American Mathematical Monthly, 112(2):188–191, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037432.

Nunemacher:2005:BRB

[1507] Jeffrey L. Nunemacher. Book review: Geometry, by David A. Brannan; Matthew F. Esplen; Jeremy J. Gray. American Mathematical Monthly, 112(2):191–194, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037433.

Anonymous:2005:BMc

[1508] Anonymous. Back matter. American Mathematical Monthly, 112(2):??, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037434.

Anonymous:2005:FMc

[1509] Anonymous. Front matter. American Mathematical Monthly, 112(2):??, February 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037408.

Megginson:2005:YGG

[1510] Robert E. Megginson. Yueh-Gin Gung and Dr. Charles Y. Hu Award to Gerald L. Alexanderson for Distinguished Service to Mathematics. American Mathematical Monthly, 112(3):197–199, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037437.

Alperin:2005:TTR

[1511] Roger C. Alperin. Trisections and totally real origami. American Mathematical Monthly, 112(3):200–211, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037438.

Bryk:2005:MDS

[1512] John Bryk and Cesar E. Silva. Measurable dynamics of simple p-adic polynomials. American Mathematical Monthly, 112(3):212–232, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037439.

Grattan-Guinness:2005:EPD

[1513] Ivor Grattan-Guinness. The “ecole polytechnique”, 1794-1850: Differences over educational purpose and teaching practice. American Mathematical Monthly, 112(3):233–250, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037440.

Duke:2005:QRF

[1514] William Duke and Kimberly Hopkins. Quadratic reciprocity in a finite group. American Mathematical Monthly, 112(3):251–256, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037441.

Gueron:2005:WEM

[1515] Shay Gueron and Itai Shafrir. A weighted Erdős–Mordell inequality for polygons. American Mathematical Monthly, 112(3):257–263, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037442.

Xuebo:2005:IHP

[1516] Luo Xuebo and Zhu-Jun Zheng. An inequality for homogeneous polynomials on R^{n}. American Mathematical Monthly, 112(3):264–266, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037443.

Anderson:2005:CPF

[1517] Peter G. Anderson, Arthur T. Benjamin, and Jeremy A. Rouse. Combinatorial proofs of Fermat’s, Lucas’s, and Wilson’s theorems. American Mathematical Monthly, 112(3):266–268, March 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037444.

Girstmair:2005:ICM

[1518] Kurt Girstmair. On an irreducibility criterion of M. Ram Murty. American Mathematical Monthly, 112(3):269–270, March 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037445.

Dawson:2005:SIT

[1519] Robert J. MacG. Dawson. On a “singular” integration technique of Poisson. American Mathematical Monthly, 112(3):270–272, March 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037446.

Tauraso:2005:P

[1520] Roberto Tauraso. Problem 11138. American Mathematical Monthly, 112 (3):273, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037447.

Bennett:2005:P

[1521] Grahame Bennett. Problem 11139. American Mathematical Monthly, 112 (3):273, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037448.

Janous:2005:P

[1522] Walther Janous. Problem 11140. American Mathematical Monthly, 112 (3):273, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037449.

Bagby:2005:P

[1523] Richard Bagby. Problem 11141. American Mathematical Monthly, 112 (3):273, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037450.

Knuth:2005:Pa

[1524] Donald Knuth. Problem 11142. American Mathematical Monthly, 112(3):273–274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037451.

Dubinov:2005:P

[1525] Alexander Dubinov and Irina Dubinova. Problem 11143. American Mathematical Monthly, 112(3):274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037452.

Schulte-Geers:2005:P

[1526] Ernst Schulte-Geers. Problem 11144. American Mathematical Monthly, 112(3):274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037453.

Zinn:2005:Pa

[1527] Joel Zinn. Problem 11076. American Mathematical Monthly, 112(3):274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037454.

Berk:2005:P

[1528] Bob Berk and Larry Shepp. Problem 11120. American Mathematical Monthly, 112(3):274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037455.

Dubickas:2005:P

[1529] Artūras Dubickas. Problem 11123. American Mathematical Monthly, 112 (3):274, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037456.

Simic:2005:EFO

[1530] Slavko Simic and Richard Stong. Entire functions of order zero: 11018. American Mathematical Monthly, 112(3):274–277, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037457.

Santos:2005:CCD

[1531] Bernardo Recamán Santos, Michail Reid, and Kevin Ford. Cycling via common divisors: 11019. American Mathematical Monthly, 112(3):277– 278, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037458.

Knuth:2005:MT

[1532] Donald Knuth, Silvia Fernández, and Gerry Myerson. A modular triple: 11021. American Mathematical Monthly, 112(3):279–280, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037459.

Chao:2005:TP

[1533] Wu Wei Chao, Nadjib Bouzar, and Allan K. Swett. A timely problem: 11023. American Mathematical Monthly, 112(3):280, March 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037461.

Satnoianu:2005:TTI

[1534] Razvan Satnoianu and Walther Janous. A two-triangle inequality: 11022. American Mathematical Monthly, 112(3):280, March 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037460.

Folland:2005:BRBb

[1535] Gerald B. Folland and Jeffrey L. Stuart. Book review: Linear Algebra, by John B. Fraleigh; Raymond A. Beauregard; Linear Algebra and its Applications by David Lay; Linear Algebra: A Geometric Approach by Theodore Shifrin; Malcolm R. Adams; Introduction to Linear Algebra by Gilbert Strang. American Mathematical Monthly, 112(3):281–288, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037462.

Anonymous:2005:BMd

[1536] Anonymous. Back matter. American Mathematical Monthly, 112(3):??, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037463.

Anonymous:2005:FMd

[1537] Anonymous. Front matter. American Mathematical Monthly, 112(3):??, March 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037436.

Shimamoto:2005:SPP

[1538] Don Shimamoto and Catherine Vanderwaart. Spaces of polygons in the plane and Morse theory. American Mathematical Monthly, 112(4):289– 310, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037466.

Burger:2005:TTP

[1539] Edward B. Burger. A tail of two palindromes. American Mathematical Monthly, 112(4):311–321, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037467.

Stein:2005:RGE

[1540] Sherman Stein. Remarks on the gravity equation. American Mathematical Monthly, 112(4):322–333, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037468.

Simons:2005:EAM

[1541] Christopher S. Simons. An elementary approach to the monster. American Mathematical Monthly, 112(4):334–341, April 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037469.

Velleman:2005:GSR

[1542] Daniel J. Velleman. The generalized Simpson’s rule. American Mathematical Monthly, 112(4):342–350, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037470.

Mirman:2005:SCP

[1543] Boris Mirman. Sufficient conditions for Poncelet polygons not to close. American Mathematical Monthly, 112(4):351–356, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037471.

Kindler:2005:SPS

[1544] Jürgen Kindler. A simple proof of Sion’s minimax theorem. American Mathematical Monthly, 112(4):356–358, April 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037472.

Clifford:2005:CAS

[1545] John Clifford, David James, Michael Lachance, and Joan Remski. A constructive approach to singular value decomposition and symmetric Schur factorization. American Mathematical Monthly, 112(4):358–363, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037473.

DallAglio:2005:FDA

[1546] Marco Dall’Aglio and Fabio Maccheroni. Fair division without additivity. American Mathematical Monthly, 112(4):363–365, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037474.

Zinn:2005:Pb

[1547] Joel Zinn. Problem 11145. American Mathematical Monthly, 112(4): 366, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037475.

Tian:2005:Pb

[1548] Yongge Tian. Problem 11146. American Mathematical Monthly, 112(4): 366, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037476.

Gorkin:2005:P

[1549] Pamela Gorkin. Problem 11147. American Mathematical Monthly, 112 (4):366, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037477.

Glasser:2005:P

[1550] M. L. Glasser. Problem 11148. American Mathematical Monthly, 112(4): 366, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037478.

Ivan:2005:Pa

[1551] Mircea Ivan and Ioan Rasa. Problem 11149. American Mathematical Monthly, 112(4):367, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037479.

Deutsch:2005:Pa

[1552] Emeric Deutsch. Problem 11150. American Mathematical Monthly, 112 (4):367, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037480.

Knuth:2005:Pb

[1553] Donald Knuth. Problem 11151. American Mathematical Monthly, 112(4): 367, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037481.

Sondow:2005:IIH

[1554] Jonathan Sondow, Richard Stong, and Julien Grivaux. An identity involving harmonic numbers: 11026. American Mathematical Monthly, 112(4):367–369, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037482.

Reuvers:2005:TIC

[1555] H. Reuvers, Li Zhou, and A. Engle. Triangles inscribed in a circle: 11027. American Mathematical Monthly, 112(4):369–370, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037483.

Trenkler:2005:HPH

[1556] Götz Trenkler and G. P. Shannon. Hermitian products of Hermitian matrices: 11028. American Mathematical Monthly, 112(4):370–371, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037484.

Butler:2005:FE

[1557] Steven Butler and Bruce Ebanks. A functional equation: 11030. American Mathematical Monthly, 112(4):371–372, April 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037485.

Zhou:2005:MMT

[1558] Li Zhou and Robin Chapman. Monster mean tamed: 11031. American Mathematical Monthly, 112(4):372–373, April 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037486.

Rowe:2005:BRB

[1559] David E. Rowe. Book review: Mathematicians under the Nazis, by Sanford L. Segal. American Mathematical Monthly, 112(4):374–381, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037487.

Anonymous:2005:BMe

[1560] Anonymous. Back matter. American Mathematical Monthly, 112(4): ??, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037488.

Anonymous:2005:FMe

[1561] Anonymous. Front matter. American Mathematical Monthly, 112(4): ??, April 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037465.

Magidin:2005:GLN

[1562] Arturo Magidin and David McKinnon. Gauss’s lemma for number fields. American Mathematical Monthly, 112(5):385–416, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037491.

Deloup:2005:FGC

[1563] Florian Deloup. The fundamental group of the circle is trivial. American Mathematical Monthly, 112(5):417–425, May 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037492.

Behn:2005:CDB

[1564] Antonio Behn, Christopher Kribs-Zaleta, and Vadim Ponomarenko. The convergence of difference boxes. American Mathematical Monthly, 112(5):426–439, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037493.

Becker:2005:DNS

[1565] Paul E. Becker. Do normal subgroups have straight tails? American Mathematical Monthly, 112(5):440–449, May 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037494.

Abu-Saris:2005:AEC

[1566] Raghib Abu-Saris and Wajdi Ahmad. Avoiding eigenvalues in computing matrix powers. American Mathematical Monthly, 112(5):450–454, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037495.

Comellas:2005:ESM

[1567] Francesc Comellas and J. Luis A. Yebra. Exact solutions for minimax optimization problems. American Mathematical Monthly, 112(5):454– 458, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037496.

Avkhadiev:2005:SPG

[1568] F. G. Avkhadiev. A simple proof of the Gauss-Winckler inequality. American Mathematical Monthly, 112(5):459–462, May 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037497.

Hajja:2005:CSC

[1569] Mowaffaq Hajja. Copositive symmetric cubic forms. American Mathematical Monthly, 112(5):462–466, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037498.

Ivan:2005:Pb

[1570] Mircea Ivan. Problem 11152. American Mathematical Monthly, 112(5): 467, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037499.

Bracken:2005:Pb

[1571] Paul Bracken. Problem 11153. American Mathematical Monthly, 112(5): 467, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037500.

Holing:2005:Pa

[1572] Kent Holing. Problem 11154. American Mathematical Monthly, 112(5): 467, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037501.

Kowalski:2005:P

[1573] Emmanuel Kowalski. Problem 11155. American Mathematical Monthly, 112(5):467, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037502.

Stanley:2005:P

[1574] Richard Stanley. Problem 11156. American Mathematical Monthly, 112(5):467–468, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037503.

Gauchman:2005:P

[1575] Hillel Gauchman. Problem 11157. American Mathematical Monthly, 112 (5):468, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037504.

Beckwith:2005:Pa

[1576] David Beckwith. Problem 11158. American Mathematical Monthly, 112 (5):468, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037505.

Marquez:2005:GFB

[1577] Juan-Bosco Romero Márquez and Martin Muldoon. Gamma function bounds: 11016. American Mathematical Monthly, 112(5):468, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037506.

Keller:2005:OGF

[1578] Tim Keller and Doyle Henderson. One good factorization deserves another: 11029. American Mathematical Monthly, 112(5):469, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037507.

Pinsky:2005:SDT

[1579] Mark Pinsky and Nicholas Singer. The stationary distribution of a tridiagonal row-stochastic matrix: 11032. American Mathematical Monthly, 112(5):469–470, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037508.

Deshpande:2005:EIE

[1580] M. N. Deshpande, R. M. Welukar, Richard Stong, and BSI Problems Group. Expansion by inclusion-exclusion: 11033. American Mathematical Monthly, 112(5):470–472, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037509.

Yayenie:2005:EI

[1581] Omer Yayenie and Joe Geiman. Evaluation of an integral: 11035. American Mathematical Monthly, 112(5):472–473, May 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037510.

Hautus:2005:SFD

[1582] M. L. J. Hautus and Julien Grivaux. Shrinking Fréchet differentiable maps: 11037. American Mathematical Monthly, 112(5):473–474, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037511.

vanLamoen:2005:MTL

[1583] Floor van Lamoen and Michel Bataille. Morley’s triangle and lemma: 11038. American Mathematical Monthly, 112(5):474, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037512.

Madden:2005:BRB

[1584] Daniel J. Madden. Book review: Abstract Algebra, by Ronald Solomon; Integers, Polynomials and Rings by Ronald S. Irving. American Mathematical Monthly, 112(5):475–478, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037513.

Sylvester:2005:BRB

[1585] John Sylvester. Book review: Introduction to the Mathematics of Medical Imaging, by Charles L. Epstein. American Mathematical Monthly, 112(5):479–480, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037514.

Anonymous:2005:BMf

[1586] Anonymous. Back matter. American Mathematical Monthly, 112(5): ??, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037515.

Anonymous:2005:FMf

[1587] Anonymous. Front matter. American Mathematical Monthly, 112(5): ??, May 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037490.

Benjamin:2005:CD

[1588] Arthur T. Benjamin and Naiomi T. Cameron. Counting on determinants. American Mathematical Monthly, 112(6):481–492, June/July 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037518.

Borwein:2005:CFT

[1589] Jonathan Michael Borwein, Kwok-Kwong Stephen Choi, and Wilfried Pigulla. Continued fractions of tails of hypergeometric series. American Mathematical Monthly, 112(6):493–501, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037519.

Balinski:2005:WJ

[1590] Michel Balinski. What is just? American Mathematical Monthly, 112 (6):502–511, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037520.

Grossman:2005:LIT

[1591] Nathaniel Grossman. A C^{∞} Lagrange inversion theorem. American Mathematical Monthly, 112(6):512–514, June/July 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037521.

Pinsky:2005:MVM

[1592] Mark A. Pinsky. Mean values and the maximum principle: a proof in search of more theorems. American Mathematical Monthly, 112(6):515– 521, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037522.

Maehara:2005:PTM

[1593] H. Maehara and T. Zamfirescu. On the perimeter of a triangle in a Minkowski plane. American Mathematical Monthly, 112(6):521–522, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037523.

Frohn:2005:KRS

[1594] Daniel Frohn. A (K)-ring satisfying the ascending chain condition on principal ideals that is not a principal ideal ring. American Mathematical Monthly, 112(6):523–524, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037524.

Lorch:2005:CRZ

[1595] Lee Lorch. Continuous ranking of zeros of cylinder functions. American Mathematical Monthly, 112(6):524–525, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037525.

Blaasjo:2005:IP

[1596] Viktor Blåsjö. The isoperimetric problem. American Mathematical Monthly, 112(6):526–566, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037526.

Lamb:2005:P

[1597] George Lamb. Problem 11159. American Mathematical Monthly, 112(6): 567, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037527.

Chamberland:2005:P

[1598] Marc Chamberland. Problem 11160. American Mathematical Monthly, 112(6):567, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037528.

Deutsch:2005:Pb

[1599] Emeric Deutsch. Problem 11161. American Mathematical Monthly, 112(6):567, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037529.

Perfetti:2005:P

[1600] Paolo Perfetti. Problem 11162. American Mathematical Monthly, 112(6): 567, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037530.

Bataille:2005:P

[1601] Michel Bataille. Problem 11163. American Mathematical Monthly, 112(6):567, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037531.

Diaz-Barrero:2005:Pa

[1602] Luis Díaz-Barrero. Problem 11164. American Mathematical Monthly, 112(6):568, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037532.

More:2005:P

[1603] Yogesh More. Problem 11165. American Mathematical Monthly, 112(6): 568, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037533.

Wardlaw:2005:ASF

[1604] William P. Wardlaw and Richard Stong. Adjoint over the smallest field: 11034. American Mathematical Monthly, 112(6):568–569, June/ July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037534.

Glasser:2005:OEO

[1605] M. L. Glasser and George Lamb. One easy, one hard: 11036. American Mathematical Monthly, 112(6):569–572, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037535.

Rosenberg:2005:IS

[1606] Noah Rosenberg and Richard Bagby. Integral over a simplex: 11039. American Mathematical Monthly, 112(6):572–573, June/July 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037536.

Tian:2005:RIH

[1607] Yongge Tian and BSI Problems Group. Range identities for Hermitian idempotent matrices: 11040. American Mathematical Monthly, 112(6): 573, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037537.

Deshpande:2005:IVE

[1608] M. N. Deshpande and John Henry Steelman. An invariance of variance over expectation: 11042. American Mathematical Monthly, 112(6):573– 574, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037538.

Garling:2005:BRB

[1609] D. J. H. Garling. Book review: The Cauchy–Schwarz Master Class: An Introduction to the Art of Mathematical Inequalities, by J. Michael Steele. American Mathematical Monthly, 112(6):575–579, June/July 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037539.

Grafakos:2005:BRB

[1610] Loukas Grafakos. Book review: Constantin Carathéodory: Mathematics and Politics in Turbulent Times, by Maria Georgiadou. American Mathematical Monthly, 112(6):579–581, June/July 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037540.

Anonymous:2005:BMg

[1611] Anonymous. Back matter. American Mathematical Monthly, 112(6):??, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037541.

Anonymous:2005:FMg

[1612] Anonymous. Front matter. American Mathematical Monthly, 112(6):??, June/July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037517.

Kim:2005:AGD

[1613] Kang-Tae Kim and Steven G. Krantz. The automorphism groups of domains. American Mathematical Monthly, 112(7):585–601, August/ September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037544.

Dresden:2005:FFF

[1614] Greg Dresden and Wayne M. Dymàček. Finding factors of factor rings over the Gaussian integers. American Mathematical Monthly, 112(7): 602–611, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037545.

Aharonov:2005:FCO

[1615] Dov Aharonov, Alan Beardon, and Kathy Driver. Fibonacci, Chebyshev, and orthogonal polynomials. American Mathematical Monthly, 112(7): 612–630, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037546.

Ahearn:2005:TIM

[1616] Stephen T. Ahearn. Tolstoy’s integration metaphor from war and peace. American Mathematical Monthly, 112(7):631–638, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037547.

Finkelstein:2005:EAP

[1617] Mark Finkelstein and Robert Whitley. Eigenvalues, almost periodic functions, and the derivative of an integral. American Mathematical Monthly, 112(7):639–641, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037548.

Araujo:2005:EPE

[1618] J. Araújo and J. D. Mitchell. An elementary proof that every singular matrix is a product of idempotent matrices. American Mathematical Monthly, 112(7):641–645, August/September 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037549.

Grady:2005:GTA

[1619] Michael J. Grady. A group theoretic approach to a famous partition formula. American Mathematical Monthly, 112(7):645–651, August/ September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037550.

Ross:2005:EPL

[1620] David A. Ross. An elementary proof of Lyapunov’s theorem. American Mathematical Monthly, 112(7):651–653, August/September 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037551.

Oman:2005:P

[1621] Greg Oman. Problem 11166. American Mathematical Monthly, 112(7): 654, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037552.

Radulescu:2005:Pb

[1622] Vicenţiu Rădulescu. Problem 11167. American Mathematical Monthly, 112(7):654, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037553.

Holing:2005:Pb

[1623] Kent Holing. Problem 11168. American Mathematical Monthly, 112 (7):654, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037554.

Mehrabi:2005:P

[1624] Mohammad Hossein Mehrabi. Problem 11169. American Mathematical Monthly, 112(7):654, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037555.

Deutsch:2005:Pc

[1625] Emeric Deutsch. Problem 11170. American Mathematical Monthly, 112 (7):655, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037556.

Cheng:2005:P

[1626] Chu Cheng. Problem 11171. American Mathematical Monthly, 112(7): 655, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037557.

Seiffert:2005:P

[1627] Heinz-Jürgen Seiffert. Problem 11172. American Mathematical Monthly, 112(7):655, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037558.

Ciaurri:2005:HLS

[1628] Oscar Ciaurri and Florin Boca. Hyperbolic-looking special functions: 11041. American Mathematical Monthly, 112(7):655–657, August/ September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037559.

Golomb:2005:DPE

[1629] Michael Golomb. Don’t the problem ebitees read the monthly? 11043. American Mathematical Monthly, 112(7):657, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037560.

Bataille:2005:CSG

[1630] Michel Bataille and Li Zhou. A combinatorial sum goes on tangent: 11044. American Mathematical Monthly, 112(7):657–659, August/ September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037561.

Beck:2005:OST

[1631] Matthias Beck, Marcin Mazur, and Shelemyahu Zacks. “ordinary” spherical triangles: 11048. American Mathematical Monthly, 112(7):659– 660, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037562.

Mihai:2005:OMT

[1632] Vasile Mihai and Zoltán Vőrős. One of the Morley triangles: 11049. American Mathematical Monthly, 112(7):660–661, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037563.

Chao:2005:TFE

[1633] Wu Wei Chao and The University of Louisiana at Lafayette Math Club. Two functional equations: 11053. American Mathematical Monthly, 112(7):661–662, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037564.

Satnoianu:2005:MAS

[1634] Razvan Satnoianu and Daniele Donini. Medians and altitudes and the sum of the radii: 11055. American Mathematical Monthly, 112(7):662– 663, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037565.

Stolarsky:2005:BRB

[1635] Kenneth B. Stolarsky. Book review: Analytic Theory of Polynomials, by Qazi Ibadur Rahman; Gerhard Schmeisser; Complex Polynomials by Terry Sheil-Small. American Mathematical Monthly, 112(7): 664–671, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037566.

Anonymous:2005:BMh

[1636] Anonymous. Back matter. American Mathematical Monthly, 112(7): ??, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037567.

Anonymous:2005:FMh

[1637] Anonymous. Front matter. American Mathematical Monthly, 112(7): ??, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037543.

Dilcher:2005:PTT

[1638] Karl Dilcher and Kenneth B. Stolarsky. A Pascal-type triangle characterizing twin primes. American Mathematical Monthly, 112(8):673–681, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037570.

Cox:2005:GNT

[1639] David A. Cox and Jerry Shurman. Geometry and number theory on clovers. American Mathematical Monthly, 112(8):682–704, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037571.

Leibman:2005:NPF

[1640] George Leibman. A nonstandard proof of the fundamental theorem of algebra. American Mathematical Monthly, 112(8):705–712, October 2005.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037572.

Klosinski:2005:SFW

[1641] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Fifth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 112(8):713–725, October 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037573.

Rao:2005:SWT

[1642] N. V. Rao. The Stone–Weierstrass Theorem revisited. American Mathematical Monthly, 112(8):726–729, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037574.

Sondow:2005:FPN

[1643] Jonathan Sondow. A faster product for π and a new integral for ln π/2. American Mathematical Monthly, 112(8):729–734, October 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037575.

Isaacs:2005:GFL

[1644] I. M. Isaacs and M. R. Pournaki. Generalizations of Fermat’s little theorem via group theory. American Mathematical Monthly, 112(8):734–740, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037576.

Brown:2005:CMI

[1645] Geoffrey Brown. Convexity and Minkowski’s inequality. American Mathematical Monthly, 112(8):740–742, October 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037577.

Rosenfeld:2005:APC

[1646] B. A. Rosenfeld. The analytic principle of continuity. American Mathematical Monthly, 112(8):743–748, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037578.

Deshpande:2005:Pa

[1647] M. N. Deshpande. Problem 11173. American Mathematical Monthly, 112(8):749, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037579.

Dalyay:2005:P

[1648] Pál Péter Dályay. Problem 11174. American Mathematical Monthly, 112(8):749, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037580.

Calbeck:2005:P

[1649] William Calbeck. Problem 11175. American Mathematical Monthly, 112(8):749, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037581.

Faynshteyn:2005:P

[1650] Oleg Faynshteyn. Problem 11176. American Mathematical Monthly, 112(8):749, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037582.

Marghidanu:2005:P

[1651] Dorin Mărghidanu. Problem 11177. American Mathematical Monthly, 112(8):750, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037583.

Bentley:2005:P

[1652] Jon Bentley and Colin Mallows. Problem 11178. American Mathematical Monthly, 112(8):750, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037584.

Beckwith:2005:Pb

[1653] David Beckwith. Problem 11179. American Mathematical Monthly, 112(8):750, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037585.

Singh:2005:PS

[1654] Manoj Prakash Singh and Richard Stong. A prime summandification of n: 11045. American Mathematical Monthly, 112(8):750–751, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037586.

Amrabov:2005:AFL

[1655] Shahin Amrabov and Bernard M. Abrego. An application of Fermat’s little theorem: 11054. American Mathematical Monthly, 112(8):751, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037588.

Marivani:2005:MSD

[1656] Syrous Marivani and O. P. Lossers. Modular sequences defined by polynomials: 11047. American Mathematical Monthly, 112(8):751, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037587.

Lindsey:2005:MCA

[1657] John H. Lindsey II and Omran Kouba. Maximum compared to average: 11056. American Mathematical Monthly, 112(8):752, October 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037589.

Ciaurri:2005:VIS

[1658] Óscar Ciaurri and O. P. Lossers. A vanishing infinite sum: 11058. American Mathematical Monthly, 112(8):752–753, October 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037590.

Coppenbarger:2005:LBM

[1659] Matthew Coppenbarger and GCHQ Problem Solving Group. “little boxes made of ticky-tacky”: 11060. American Mathematical Monthly, 112(8):753–754, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037591.

Schilling:2005:PD

[1660] Kenneth Schilling and BSI Problems Group. Powered divergence: 11062. American Mathematical Monthly, 112(8):754–755, October 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037592.

Wetzel:2005:LWC

[1661] John Wetzel and Li Zhou. Lengths of wrapped curves: 11064. American Mathematical Monthly, 112(8):755, October 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037593.

Tyler:2005:MGP

[1662] Douglas B. Tyler and Robin Chapman. A more general presentation: 11065. American Mathematical Monthly, 112(8):755–756, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037594.

Grabiner:2005:BRB

[1663] Judith Grabiner. Book review: From Newton to Hawking: A History of Cambridge University’s Lucasian Professors of Mathematics, by Kevin C. Knox; Richard Noakes. American Mathematical Monthly, 112(8):757– 762, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037595.

Anonymous:2005:BMi

[1664] Anonymous. Back matter. American Mathematical Monthly, 112(8):??, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037596.

Anonymous:2005:FMi

[1665] Anonymous. Front matter. American Mathematical Monthly, 112(8):??, October 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037569.

Bajunaid:2005:FSC

[1666] Ibtesam Bajunaid, Joel M. Cohen, Flavia Colonna, and David Singman. Function series, Catalan numbers, and random walks on trees. American Mathematical Monthly, 112(9):765–785, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037599.

Kalman:2005:VEI

[1667] Dan Kalman. Virtual empirical investigation: Concept formation and theory justification. American Mathematical Monthly, 112(9):786–798, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037600.

Ivanov:2005:DFP

[1668] Nikolai V. Ivanov. A differential forms perspective on the Lax proof of the change of variables formula. American Mathematical Monthly, 112(9):799–806, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037601.

Alperin:2005:MTP

[1669] Roger C. Alperin. The modular tree of Pythagoras. American Mathematical Monthly, 112(9):807–816, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037602.

Redheffer:2005:BSN

[1670] Ray Redheffer. The behavior of solutions near a stable or semistable stationary point. American Mathematical Monthly, 112(9):817–822, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037603.

Nyblom:2005:MNS

[1671] M. A. Nyblom. More nested square roots of 2. American Mathematical Monthly, 112(9):822–825, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037604.

Naghipour:2005:SPC

[1672] A. R. Naghipour. A simple proof of Cohen’s theorem. American Mathematical Monthly, 112(9):825–826, November 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037605.

Coquand:2005:SPK

[1673] Thierry Coquand and Henri Lombardi. A short proof for the Krull dimension of a polynomial ring. American Mathematical Monthly, 112(9):826– 829, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037606.

Myers:2005:EPJ

[1674] Robert Myers. An elementary proof of Joris’s theorem. American Mathematical Monthly, 112(9):829–831, November 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037607.

Monastyrsky:2005:SRM

[1675] Michael Monastyrsky. Some remarks on “mathematics at the tlirn of the millennium”. American Mathematical Monthly, 112(9):832–838, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037608.

Namli:2005:P

[1676] Suat Namli. Problem 11180. American Mathematical Monthly, 112(9): 839, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037609.

Deshpande:2005:Pb

[1677] M. N. Deshpande. Problem 11181. American Mathematical Monthly, 112(9):839, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037610.

Amrahov:2005:Pb

[1678] Shahin Amrahov. Problem 11182. American Mathematical Monthly, 112(9):839, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037611.

Beckwith:2005:Pc

[1679] David Beckwith. Problem 11183. American Mathematical Monthly, 112(9):839, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037612.

Simic:2005:P

[1680] Slavko Simic. Problem 11184. American Mathematical Monthly, 112(9): 840, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037613.

Bruck:2005:P

[1681] Rainer Brück. Problem 11185. American Mathematical Monthly, 112(9): 840, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037614.

Christino:2005:P

[1682] Angelo J. Christino, Jr. and William C. Waterhouse. Problem 11186. American Mathematical Monthly, 112(9):840, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037615.

Holing:2005:SSI

[1683] Kent Holing and Richard Stong. A special set of integer points in space: 11050. American Mathematical Monthly, 112(9):840–841, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/30037616.

Simic:2005:EA

[1684] Slavko Simic and John H. Lindsey II. Equivalent asymptotics: 11063. American Mathematical Monthly, 112(9):841–843, November 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037617.

Lamb:2005:IMT

[1685] George Lamb and O. P. Lossers. Integral by Mellin transform: 11067. American Mathematical Monthly, 112(9):843–844, November 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037618.

Wilf:2005:TR

[1686] Herbert S. Wilf and Kenneth E. Schilling. Think rationally: 11068. American Mathematical Monthly, 112(9):844–845, November 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037619.

Lamb:2005:EI

[1687] George Lamb and T. K. Callahan. Evaluate an integral: 11072. American Mathematical Monthly, 112(9):845–846, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037620.

Logan:2005:BRB

[1688] J. David Logan. Book review: Mathematical Models in Biology, by Elizabeth S. Allman; John A. Rhodes. American Mathematical Monthly, 112(9):847–850, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037621.

Warfield:2005:BRB

[1689] Virginia M. Warfield. Book review: Count Down: Six Kids Vie for Glory at the World’s Toughest Math Competition, by Steve Olson. American Mathematical Monthly, 112(9):850–851, November 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037622.

Anonymous:2005:BMj

[1690] Anonymous. Back matter. American Mathematical Monthly, 112(9):??, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037623.

Anonymous:2005:FMj

[1691] Anonymous. Front matter. American Mathematical Monthly, 112(9):??, November 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037598.

Humphreys:2005:WMM

[1692] Lisa D. Humphreys and P. Joseph McKenna. When a mechanical model goes nonlinear: Unexpected responses to low-periodic shaking. American Mathematical Monthly, 112(10):861–875, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037627.

Driver:2005:IQP

[1693] Eric Driver, Philip A. Leonard, and Kenneth S. Williams. Irreducible quartic polynomials with factorizations modulo p. American Mathematical Monthly, 112(10):876–890, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037628.

Zemyan:2005:ZNP

[1694] Stephen M. Zemyan. On the zeroes of the nth partial sum of the exponential series. American Mathematical Monthly, 112(10):891–909, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037629.

Dijkstra:2005:HGC

[1695] Jan J. Dijkstra. On homeomorphism groups and the compact-open topology. American Mathematical Monthly, 112(10):910–912, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037630.

Chen:2005:TR

[1696] Zhibo Chen and Gary L. Mullen. Transpositions and representability. American Mathematical Monthly, 112(10):913–919, December 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037631.

Pikhurko:2005:GEL

[1697] Oleg Pikhurko. Generating edge-labeled trees. American Mathematical Monthly, 112(10):919–921, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037632.

Dumitrescu:2005:RES

[1698] Adrian Dumitrescu. A remark on the Erdős–Szekeres theorem. American Mathematical Monthly, 112(10):921–924, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037633.

Kim:2005:MZP

[1699] Seon-Hong Kim. On the moduli of the zeros of a polynomial. American Mathematical Monthly, 112(10):924–925, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037634.

Arkeryd:2005:NA

[1700] Leif Arkeryd. Nonstandard analysis. American Mathematical Monthly, 112(10):926–928, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037635.

Zhou:2005:Pb

[1701] Li Zhou. Problem 11187. American Mathematical Monthly, 112(10):929, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037636.

Bailey:2005:P

[1702] David Bailey. Problem 11188. American Mathematical Monthly, 112(10): 929, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037637.

Csete:2005:P

[1703] Lajos Csete. Problem 11189. American Mathematical Monthly, 112(10): 929, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037638.

Diaz-Barrero:2005:Pb

[1704] José Luis Díaz-Barrero. Problem 11190. American Mathematical Monthly, 112(10):929, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/30037639.

Huff:2005:P

[1705] Barthel Wayne Huff. Problem 11191. American Mathematical Monthly, 112(10):929–930, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037640.

Stevens:2005:P

[1706] Brett Stevens. Problem 11192. American Mathematical Monthly, 112(10):930, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037641.

Tetiva:2005:P

[1707] Marian Tetiva. Problem 11193. American Mathematical Monthly, 112(10):930, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037642.

vanLamoen:2005:SSF

[1708] Floor van Lamoen. Six surprises from four lines: 11025. American Mathematical Monthly, 112(10):930–932, December 2005. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/30037643.

Cioaba:2005:SCC

[1709] Sebastian Cioabă, Tim LeSaulnier, and Dan Saracino. A sufficient condition for commutativity: 11066. American Mathematical Monthly, 112 (10):932–933, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037644.

Ivady:2005:SB

[1710] Péter Ivády and Richard Stong. Sine bound: 11069. American Mathematical Monthly, 112(10):933, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037645.

Trenkler:2005:ECI

[1711] Götz Trenkler and Lawrence J. Baker. An easy complex inequality: 11075. American Mathematical Monthly, 112(10):933, December 2005.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037646.

Costovici:2005:OS

[1712] Gheorge Costovici and Uroš Milutinović. An oscillating sequence: 11077. American Mathematical Monthly, 112(10):934, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037647.

Satnoianu:2005:MC

[1713] Razvan Satnoianu and Li Zhou. A majorization consequence: 11080. American Mathematical Monthly, 112(10):934–935, December 2005. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037648.

Morse:2005:HTV

[1714] Stephen Morse and Robin Chapman. Harmonious trios are the vertices of a regular icosahedron: 11081. American Mathematical Monthly, 112(10): 935, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/30037649.

Chao:2005:ECA

[1715] Wu Wei Chao. Errata: Continuous additive functions. American Mathematical Monthly, 112(10):936, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/30037651. See [1385].

Getz:2005:EPA

[1716] Marty Getz and Dixon Jones. Errata: Partitioning by arithmetic progressions. American Mathematical Monthly, 112(10):936, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/30037652. See [1474].

Golomb:2005:OS

[1717] Michael Golomb and Kit Hanes. The orthocenter of a simplex: 11087. American Mathematical Monthly, 112(10):936, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/30037650.

Folland:2005:BRBa

[1718] Gerald B. Folland. Book review: Functional Analysis, by Peter D. Lax; John Wiley; Functional Analysis: An Introduction by Yuli Eidelman; Vitaly Milman; Antonis Tsolomitis. American Mathematical Monthly, 112(10):937–940, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037653.

Palka:2005:EE

[1719] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 112(10):941–946, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037654.

Anonymous:2005:BMb

[1720] Anonymous. Back matter. American Mathematical Monthly, 112(10):??, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037655.

Anonymous:2005:FMb

[1721] Anonymous. Front matter. American Mathematical Monthly, 112(10):??, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/30037626.

Anonymous:2005:VI

[1722] Anonymous. Volume information. American Mathematical Monthly, 112(10):??, December 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

30037625.

Granville:2006:PNR

[1723] Andrew Granville and Greg Martin. Prime number races. American Mathematical Monthly, 113(1):1–33, January 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641834.

Hagood:2006:RFD

[1724] John W. Hagood and Brian S. Thomson. Recovering a function from a Dini derivative. American Mathematical Monthly, 113(1):34–46, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641835.

Frantz:2006:SGS

[1725] Marc Frantz. Some graphical solutions of the Kepler problem. American Mathematical Monthly, 113(1):47–56, January 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641836.

Cohn:2006:SPS

[1726] Henry Cohn. A short proof of the simple continued fraction expansion of e. American Mathematical Monthly, 113(1):57–62, January 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641837.

Osler:2006:PCF

[1727] Thomas J. Osler. A proof of the continued fraction expansion of e1/M. American Mathematical Monthly, 113(1):62–66, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641838.

Dai:2006:CDS

[1728] Xiongping Dai. Continuous differentiablity of solutions of ODEs with respect to initial conditions. American Mathematical Monthly, 113(1): 66–70, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641839.

Boyd:2006:FMM

[1729] Stephen Boyd, Persi Diaconis, Jun Sun, and Lin Xiao. Fastest mixing Markov chain on a path. American Mathematical Monthly, 113(1):70–74, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641840.

Strzelecki:2006:PC

[1730] Paweł Strzelecki and Abe Shenitzer. The Poincaré conjecture? American Mathematical Monthly, 113(1):75–78, January 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641841.

Bornemann:2006:P

[1731] Folkmar Bornemann. Problem 11194. American Mathematical Monthly, 113(1):79, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641842.

Tetiva:2006:Pa

[1732] Marian Tetiva. Problem 11195. American Mathematical Monthly, 113(1): 79, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641843.

Mehrabi:2006:P

[1733] Mohammad Hossein Mehrabi. Problem 11196. American Mathematical Monthly, 113(1):79, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641844.

Faynshteyn:2006:P

[1734] Oleg Faynshteyn. Problem 11197. American Mathematical Monthly, 113 (1):79, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641845.

Parthasarathy:2006:P

[1735] P. R. Parthasarathy. Problem 11198. American Mathematical Monthly, 113(1):79, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641846.

Yakub:2006:P

[1736] Aliyev Yakub. Problem 11199. American Mathematical Monthly, 113(1): 80, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641847.

Zinn:2006:P

[1737] Joel Zinn. Problem 11200. American Mathematical Monthly, 113(1):80, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641848.

Sasane:2006:DRP

[1738] A. J. Sasane and Robin Chapman. Diagonals of a regular polygon: 10924. American Mathematical Monthly, 113(1):80–81, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641849.

MacKinnon:2006:OSN

[1739] Nick MacKinnon and Kenneth Schilling. Optimal strategy for a numberguessing game: 11051. American Mathematical Monthly, 113(1):81–82, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641850.

ZPG:2006:PGD

[1740] Zahlreich Problems Group and C. T. Stretch. A point with given distances: 11057. American Mathematical Monthly, 113(1):82–83, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641851.

Bowron:2006:CCU

[1741] Mark Bowron and Bruce Burdick. Closure, complement, and union: 11059. American Mathematical Monthly, 113(1):83, January 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641852.

Tauraso:2006:CAM

[1742] Roberto Tauraso and Richard Stong. Commuting analytic maps: 11070. American Mathematical Monthly, 113(1):83–84, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641853.

Beck:2006:SLR

[1743] Matthias Beck, Jesus DeLoera, Mike Develin, Julian Pfeifle, and Northwestern University Math Problem Solving Group. Some linear recurrences: 11085. American Mathematical Monthly, 113(1):84–85, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641854.

Holzsager:2006:RS

[1744] Richard Holzsager and Anton R. Schep. Rearranged series: 11088. American Mathematical Monthly, 113(1):85, January 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/27641855.

Bataille:2006:SRP

[1745] Michel Bataille and O. P. Lossers. Simple roots for polynomials: 11094. American Mathematical Monthly, 113(1):85–86, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641856.

Radin:2006:BRB

[1746] Charles Radin. Book review: The Pursuit of Perfect Packing, by Tomaso Aste; Denis Weaire; Kelper’s Conjecture by George G. Szpiro. American Mathematical Monthly, 113(1):87–90, January 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641857.

Henson:2006:BRB

[1747] Shandelle M. Henson. Book review: Complexities: Women in Mathematics, by Bettye Anne Case; Anne M. Leggett. American Mathematical Monthly, 113(1):91–93, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641858.

Anonymous:2006:BMa

[1748] Anonymous. Back matter. American Mathematical Monthly, 113(1):??, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641859.

Anonymous:2006:FMa

[1749] Anonymous. Front matter. American Mathematical Monthly, 113(1):??, January 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641833.

Lagarias:2006:WWN

[1750] Jeffrey C. Lagarias. Wild and wooley numbers. American Mathematical Monthly, 113(2):97–108, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641862.

Blaya:2006:TDT

[1751] Alejo Blaya and Víctor Jiménez López. Is trivial dynamics that trivial? American Mathematical Monthly, 113(2):109–133, February 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641863.

Pozdnyakov:2006:WTP

[1752] Vladimir Pozdnyakov and Martin Kulldorff. Waiting times for patterns and a method of gambling teams. American Mathematical Monthly, 113(2):134–143, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641864.

Talman:2006:SRE

[1753] Louis A. Talman. Simpson’s rule is exact for quintics. American Mathematical Monthly, 113(2):144–155, February 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641865.

Medina:2006:SPA

[1754] Herbert A. Medina. A sequence of polynomials for approximating arctangent. American Mathematical Monthly, 113(2):156–161, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641866.

Diestel:2006:CBT

[1755] Reinhard Diestel and Carsten Thomassen. A Cantor–Bernstein theorem for paths in graphs. American Mathematical Monthly, 113(2):161–165, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641867.

Konvalina:2006:MP

[1756] John Konvalina, Jack Heidel, and Jim Rogers. A mean paradox. American Mathematical Monthly, 113(2):166–169, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641868.

Arrondo:2006:AEP

[1757] Enrique Arrondo. Another elementary proof of the Nullstellensatz. American Mathematical Monthly, 113(2):169–171, February 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641869.

Robinson:2006:SF

[1758] P. L. Robinson. The sphere is not flat. American Mathematical Monthly, 113(2):171–173, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641870.

Velleman:2006:NP

[1759] Daniel J. Velleman. A note on λ-permutations. American Mathematical Monthly, 113(2):173–178, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641871.

Russell:2006:P

[1760] Robert Russell. Problem 11201. American Mathematical Monthly, 113(2):179, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641872.

Bennett:2006:P

[1761] Grahame Bennett. Problem 11202. American Mathematical Monthly, 113(2):179, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641873.

Yu:2006:P

[1762] Fuxiang Yu. Problem 11203. American Mathematical Monthly, 113(2): 179, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641874.

Hillar:2006:Pa

[1763] Christopher J. Hillar. Problem 11204. American Mathematical Monthly, 113(2):179, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641875.

Chao:2006:P

[1764] Wu Wei Chao and Guang Zhou. Problem 11205. American Mathematical Monthly, 113(2):179, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641876.

Ivan:2006:P

[1765] Mircea Ivan and Alexandru Lupaş. Problem 11206. American Mathematical Monthly, 113(2):180, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641877.

Botsko:2006:Pa

[1766] Michael W. Botsko. Problem 11207. American Mathematical Monthly, 113(2):180, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641878.

Jorza:2006:NTF

[1767] Andrei Jorza, Edward Ordman, and Richard Stong. No three faces with the same number of edges: 10856. American Mathematical Monthly, 113(2):180–183, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641879.

Huang:2006:LSV

[1768] Danrun Huang, Daniel Scully, Achava Nakhash, and Karl David. The limit of a set-valued process: 11052. American Mathematical Monthly, 113(2):183, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641880.

Radulescu:2006:EPO

[1769] Vicenţiu Rǎdulescu and Richard Stong. An example with periodic orbits: 11073. American Mathematical Monthly, 113(2):184–186, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641881.

Hillar:2006:O

[1770] Christopher Hillar, Darren Rhea, and Godfrey L. Isaacs. An oldie: 11098. American Mathematical Monthly, 113(2):186, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641882.

Zitarelli:2006:BRBa

[1771] David E. Zitarelli. Book review: R. L. Moore: Mathematician and Teacher, by John Parker. American Mathematical Monthly, 113(2):187– 192, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641883.

Anonymous:2006:BMc

[1772] Anonymous. Back matter. American Mathematical Monthly, 113(2):??, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641884.

Anonymous:2006:FMc

[1773] Anonymous. Front matter. American Mathematical Monthly, 113(2):??, February 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641861.

Roberts:2006:YGG

[1774] Wayne Roberts. Yueh-Gin Gung and Dr. Charles Y. Hu Award to Hyman Bass for Distinguished Service to Mathematics. American Mathematical Monthly, 113(3):193–195, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641887.

Pengelley:2006:DEN

[1775] David Pengelley and Fred Richman. Did Euclid need the Euclidean algorithm to prove unique factorization? American Mathematical Monthly, 113(3):196–205, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641888.

Bibiloni:2006:SGE

[1776] Lluís Bibiloni, Pelegrí Viader, and Jaume Paradís. On a series of Goldbach and Euler. American Mathematical Monthly, 113(3):206–220, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641889.

Bhatia:2006:IDM

[1777] Rajendra Bhatia. Infinitely divisible matrices. American Mathematical Monthly, 113(3):221–235, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641890.

Boas:2006:RA

[1778] Harold P. Boas. Reflections on the arbelos. American Mathematical Monthly, 113(3):236–249, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641891.

Hartwig:2006:NLD

[1779] Robert E. Hartwig. Note on a linear difference equation. American Mathematical Monthly, 113(3):250–256, March 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/27641892.

Maligranda:2006:SNI

[1780] Lech Maligranda. Simple norm inequalities. American Mathematical Monthly, 113(3):256–260, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641893.

Givental:2006:PTW

[1781] Alexander Givental. The Pythagorean theorem: What is it about? American Mathematical Monthly, 113(3):261–265, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641894.

Yang:2006:LOS

[1782] YiChuan Yang. A lattice-ordered skew field is totally ordered if squares are positive. American Mathematical Monthly, 113(3):265–266, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641895.

Zhou:2006:P

[1783] Li Zhou. Problem 11208. American Mathematical Monthly, 113(3):267, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641896.

Dubinov:2006:P

[1784] Alexander Dubinov and Irina Dubinova. Problem 11209. American Mathematical Monthly, 113(3):267, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27641897.

Becker:2006:P

[1785] Michael S. Becker. Problem 11210. American Mathematical Monthly, 113 (3):267, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641898.

Bencze:2006:P

[1786] Mihály Bencze. Problem 11211. American Mathematical Monthly, 113 (3):268, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641899.

Beckwith:2006:Pa

[1787] David Beckwith. Problem 11212. American Mathematical Monthly, 113 (3):268, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641900.

Rabinowitz:2006:P

[1788] Stanley Rabinowitz. Problem 11213. American Mathematical Monthly, 113(3):268, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641901.

Golomb:2006:P

[1789] Solomon W. Golomb. Problem 11214. American Mathematical Monthly, 113(3):268, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641902.

Bagby:2006:DD

[1790] Richard Bagby and GCHQ Problem Solving Group. Differentiable distance: 11093. American Mathematical Monthly, 113(3):268–269, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641903.

Mihai:2006:TIE

[1791] Vasile Mihai and O. P. Lossers. Trapezoid inscribed in ellipse: 11095. American Mathematical Monthly, 113(3):269, March 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641904.

Popescu:2006:SSI

[1792] Călin Popescu and Richard Stong. A subset-sum inequality: 11100. American Mathematical Monthly, 113(3):269–270, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641905.

Skelton:2006:ITI

[1793] Earl F. Skelton and Zoltán Vőrős. An inverse tangent integral: 11101. American Mathematical Monthly, 113(3):270–271, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641906.

Radulescu:2006:PSD

[1794] Vincentţiu Rǎdulescu and GCHQ Problem Solving Group. Periodic solution of a differential equation: 11104. American Mathematical Monthly, 113(3):271–273, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641907.

Marquez:2006:CC

[1795] Juan-Bosco Romero Marquez, Rick Luttmann, Miguel de Guzman, G. Donald Chakerian, and Li Zhou. Cevians and centroids:11105. American Mathematical Monthly, 113(3):273–274, March 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641908.

Emery:2006:SI

[1796] Michel Émery and Szeged Problem Solving Group Fejéntaláltuka. A sum inequality:11107. American Mathematical Monthly, 113(3):274, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641909.

Ghergu:2006:SID

[1797] Marius Ghergu and Microsoft Research Problem Group. Set inclusion and density:11110. American Mathematical Monthly, 113(3):274, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641910.

Beeson:2006:BRB

[1798] Michael Beeson. Book review: A Tour through Mathematical Logic, by Robert S. Wolf. American Mathematical Monthly, 113(3):275–280, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641911.

Anonymous:2006:BMd

[1799] Anonymous. Back matter. American Mathematical Monthly, 113(3):??, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641912.

Anonymous:2006:FMd

[1800] Anonymous. Front matter. American Mathematical Monthly, 113(3):??, March 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641886.

Martin:2006:LSB

[1801] Óscar Martín. Look-and-say biochemistry: Exponential RNA and multistranded DNA. American Mathematical Monthly, 113(4):289–307, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641915.

Chahal:2006:CNE

[1802] Jasbir S. Chahal. Congruent numbers and elliptic curves. American Mathematical Monthly, 113(4):308–317, April 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641916.

Gibbons:2006:USA

[1803] Jeremy Gibbons. Unbounded spigot algorithms for the digits of pi. American Mathematical Monthly, 113(4):318–328, April 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641917.

Vance:2006:COT

[1804] Bennet Vance. Counting ordered trees by permuting their parts. American Mathematical Monthly, 113(4):329–335, April 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641918.

Abramson:2006:PIR

[1805] Nils Abramson, Jan Boman, and Björn Bonnevier. Plane intersections of rotational ellipsoids. American Mathematical Monthly, 113(4):336– 339, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641919.

Mushkarov:2006:SP

[1806] Oleg Mushkarov and Nikolai Nikolov. Semiregular polygons. American Mathematical Monthly, 113(4):339–344, April 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641920.

Burckel:2006:FII

[1807] R. B. Burckel. Fubinito (immediately) implies FTA. American Mathematical Monthly, 113(4):344–347, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641921.

Korner:2006:FTA

[1808] T. W. Körner. On the fundamental theorem of algebra. American Mathematical Monthly, 113(4):347–348, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641922.

Melvin:2006:TM

[1809] Paul Melvin. A topological menagerie. American Mathematical Monthly, 113(4):348–351, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641923.

Knobloch:2006:MPM

[1810] Eberhard Knobloch. Mathesis perennis: Mathematics in ancient, renaissance, and modern times. American Mathematical Monthly, 113(4):352– 365, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641924.

Rosset:2006:P

[1811] Shmuel Rosset. Problem 11215. American Mathematical Monthly, 113 (4):366, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641925.

Chinburg:2006:P

[1812] Ted Chinburg and Shahriar Shahriari. Problem 11216. American Mathematical Monthly, 113(4):366, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641926.

Bataille:2006:Pa

[1813] Michel Bataille. Problem 11217. American Mathematical Monthly, 113 (4):366, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641927.

Gordon:2006:P

[1814] Gary Gordon. Problem 11218. American Mathematical Monthly, 113(4): 366–367, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641928.

Strubel:2006:P

[1815] R. A. Strubel. Problem 11219. American Mathematical Monthly, 113(4): 367, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641929.

Beckwith:2006:Pb

[1816] David Beckwith. Problem 11220. American Mathematical Monthly, 113 (4):367, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641930.

Perfetti:2006:Pa

[1817] Paolo Perfetti. Problem 11221. American Mathematical Monthly, 113(4): 367, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641931.

Skordev:2006:IDC

[1818] Dimiter Skordev and Gerry Myerson. An impossible difference condition: 11074. American Mathematical Monthly, 113(4):367–368, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641932.

Knuth:2006:CFS

[1819] Donald Knuth and Robin Chapman. Cube-free sums: 11078. American Mathematical Monthly, 113(4):368–369, April 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641933.

Cook:2006:CI

[1820] Brian Cook, James Whitenton, and Richard Stong. A complex identity: 11079. American Mathematical Monthly, 113(4):369–370, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641934.

Chao:2006:DP

[1821] Wu Wei Chao and National Security Agency Problems Group. A divisibility problem: 11082. American Mathematical Monthly, 113(4):370– 371, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641935.

Wardlaw:2006:NIM

[1822] William P. Wardlaw and Jean-Pierre Grivaux. The null ideal of a matrix: 11083. American Mathematical Monthly, 113(4):371, April 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641936.

ShahAli:2006:FRD

[1823] H. A. ShahAli and Timothy LeSaulnier. Finding a repeated difference: 11084. American Mathematical Monthly, 113(4):371–372, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641937.

Locke:2006:CGF

[1824] Stephen C. Locke and Richard Stong. Coloring graphs with few odd cycles: 11086. American Mathematical Monthly, 113(4):372, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641938.

Luca:2006:SD

[1825] Florian Luca and John Ferdinands. Sometimes n divides σk(n): 11090. American Mathematical Monthly, 113(4):372–373, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641939.

Magidin:2006:OPG

[1826] Arturo Magidin and John Cade. The order of a product of group elements: 11092. American Mathematical Monthly, 113(4):373, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27641940.

Goerss:2006:BRB

[1827] Paul G. Goerss. Book review: Topology: A Geometric Approach, by Terry Lawson. American Mathematical Monthly, 113(4):374–377, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641941.

Wood:2006:BRB

[1828] Carol Wood. Book review: Alfred Tarski: Life and Logic, by Anita Burdman Feferman; Solomon Feferman. American Mathematical Monthly, 113(4):377–379, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641942.

Elsner:2006:BRB

[1829] L. Elsner. Book review: Geršgorin and His Circles, by Richard S. Varga. American Mathematical Monthly, 113(4):379–381, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641943.

Anonymous:2006:BMe

[1830] Anonymous. Back matter. American Mathematical Monthly, 113(4): ??, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641944.

Anonymous:2006:FMe

[1831] Anonymous. Front matter. American Mathematical Monthly, 113(4): ??, April 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641914.

Mossinghoff:2006:P

[1832] Michael J. Mossinghoff. A $1 problem. American Mathematical Monthly, 113(5):385–402, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641947.

Ferguson:2006:WCA

[1833] Le Baron O. Ferguson. What can be approximated by polynomials with integer coefficients. American Mathematical Monthly, 113(5):403–414, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641948.

Sobottka:2006:PPC

[1834] Marcelo Sobottka and Luiz P. L. de Oliveira. Periodicity and predictability in chaotic systems. American Mathematical Monthly, 113(5):415–424, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641949.

Adell:2006:SEN

[1835] José Antonio Adell and Alberto Lekuona. The simplest example of a normal asymptotic expansion. American Mathematical Monthly, 113(5): 425–442, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641950.

Hajja:2006:ACT

[1836] Mowaffaq Hajja. The arbitrariness of the cevian triangle. American Mathematical Monthly, 113(5):443–447, May 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641951.

Helmberg:2006:CCW

[1837] Gilbert Helmberg. Curiosities concerning weak topology in Hilbert space. American Mathematical Monthly, 113(5):447–452, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641952.

Chan:2006:MFP

[1838] Hei-Chi Chan. More formulas for π. American Mathematical Monthly, 113(5):452–455, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641953.

Groger:2006:GEJ

[1839] Detlef Gröger. On Gauss’s entry from January 6, 1809. American Mathematical Monthly, 113(5):455–458, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641954.

Sondow:2006:P

[1840] Jonathan Sondow. Problem 11222. American Mathematical Monthly, 113 (5):459, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641955.

Hillar:2006:Pb

[1841] Christopher Hillar and Lionel Levine. Problem 11123. American Mathematical Monthly, 113(5):459, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641956.

Trenkler:2006:P

[1842] Dietrich Trenkler and Götz Trenkler. Problem 11224. American Mathematical Monthly, 113(5):459, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641957.

Diaz-Barrero:2006:P

[1843] José Luis Díaz-Barrero. Problem 11225. American Mathematical Monthly, 113(5):459, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641958.

Beaucoup:2006:P

[1844] Franck Beaucoup and Tamás Erdélyi. Problem 11226. American Mathematical Monthly, 113(5):460, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641959.

Marivani:2006:P

[1845] Syrous Marivani. Problem 11227. American Mathematical Monthly, 113 (5):460, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641960.

Tetiva:2006:Pb

[1846] Marian Tetiva. Problem 11228. American Mathematical Monthly, 113(5): 460, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641961.

Deutsch:2006:VCP

[1847] Emeric Deutsch and Marc Renault. A variation on t-Catalan paths: 11071. American Mathematical Monthly, 113(5):460–461, May 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641962.

Callan:2006:MAD

[1848] David Callan and GCHQ Problem Solving Group. The maximum associativeness of division: 11091. American Mathematical Monthly, 113(5): 462–463, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641963.

Amghibech:2006:ANI

[1849] Said Amghibech and José Heber Nieto. An application of Newton’s identities: 11096. American Mathematical Monthly, 113(5):463–464, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641964.

Ford:2006:RCD

[1850] Kevin Ford and Robin Chapman. Ratio of consecutive divisors: 11097. American Mathematical Monthly, 113(5):464, May 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641965.

Beck:2006:MSN

[1851] Matthias Beck, Richard Ehrenborg, Thomas Zaslavsky, Michael R. Avidon, and John H. Lindsey, II. Magic squares and nontransitive dice: 11099. American Mathematical Monthly, 113(5):464–465, May 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641966.

Galperin:2006:ASI

[1852] Gregory Galperin, Hillel Gauchman, and Jany C. Binz. Another summation identity involving binomial coefficients: 11103. American Mathematical Monthly, 113(5):465–466, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641967.

Deutsch:2006:CES

[1853] Emeric Deutsch and José Heber Nieto. Counting elevated Schröder paths: 11108. American Mathematical Monthly, 113(5):466–467, May 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641968.

Dalyay:2006:PSF

[1854] Pál Péter Dalyay and GCHQ Problem Solving Group. Periodic solution to a functional equation: 11111. American Mathematical Monthly, 113 (5):467, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641969.

Tian:2006:IPO

[1855] Yongge Tian. On the inverse of a product of orthogonal projectors: 11112. American Mathematical Monthly, 113(5):467–468, May 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641970.

Nunemacher:2006:BRB

[1856] Jeffrey Nunemacher. Book review: Prime Obsession, by John Derbyshire; Stalking the Riemann Hypothesis by Dan Rockmore. American Mathematical Monthly, 113(5):469–473, May 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27641971.

Anonymous:2006:BMf

[1857] Anonymous. Back matter. American Mathematical Monthly, 113(5): ??, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641972.

Anonymous:2006:FMf

[1858] Anonymous. Front matter. American Mathematical Monthly, 113(5): ??, May 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641946.

Bailey:2006:TPE

[1859] David H. Bailey, Jonathan M. Borwein, Vishaal Kapoor, and Eric W. Weisstein. Ten problems in experimental mathematics. American Mathematical Monthly, 113(6):481–509, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27641975.

Levin:2006:GII

[1860] Aaron Levin. A geometric interpretation of an infinite product for the lemniscate constant. American Mathematical Monthly, 113(6):510–520, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641976.

Apostol:2006:SCS

[1861] Tom M. Apostol and Mamikon A. Mnatsakanian. Solids circumscribing spheres. American Mathematical Monthly, 113(6):521–540, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641977.

Saidak:2006:GCI

[1862] Filip Saidak. On Goldbach’s conjecture for integer polynomials. American Mathematical Monthly, 113(6):541–545, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641978.

Shiu:2006:MSA

[1863] P. Shiu. Moment sums associated with binary linear forms. American Mathematical Monthly, 113(6):545–550, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641979.

Koliha:2006:FTC

[1864] J. J. Koliha. A fundamental theorem of calculus for Lebesgue integration. American Mathematical Monthly, 113(6):551–555, June/July 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641980.

DAndrea:2006:BSN

[1865] Carlos D’Andrea and Emiliano Gómez. The broken spaghetti noodle. American Mathematical Monthly, 113(6):555–557, June/July 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641981.

Beardon:2006:CRL

[1866] A. F. Beardon. Contractions of the real line. American Mathematical Monthly, 113(6):557–558, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641982.

Kushner:2006:CMM

[1867] Boris A. Kushner. The constructive mathematics of A. A. Markov. American Mathematical Monthly, 113(6):559–566, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641983.

Furdui:2006:P

[1868] Ovidiu Furdui. Problem 11229. American Mathematical Monthly, 113(6): 567, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27641984.

Keselman:2006:Pa

[1869] Gregory Keselman. Problem 11230. American Mathematical Monthly, 113(6):567, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641985.

Hillar:2006:Pc

[1870] Christopher Hillar. Problem 11231. American Mathematical Monthly, 113(6):567, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641986.

Botsko:2006:Pb

[1871] Michael W. Botsko. Problem 11232. American Mathematical Monthly, 113(6):567, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641987.

Downes:2006:P

[1872] Robert Downes. Problem 11233. American Mathematical Monthly, 113(6):568, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27641988.

Brennan:2006:P

[1873] Jim Brennan and Richard Ehrenborg. Problem 11234. American Mathematical Monthly, 113(6):568, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641989.

Jones:2006:Pa

[1874] Lenny Jones and Rachel Keller. Problem 11235. American Mathematical Monthly, 113(6):568, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27641990.

Rabinowitz:2006:O

[1875] Stanley Rabinowitz and Roberto Tauraso. One from 1995: 10483. American Mathematical Monthly, 113(6):568–569, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27641991.

Zinn:2006:RDS

[1876] Joel Zinn and Gagik Amirkhanyan. Related divergent series: 11076. American Mathematical Monthly, 113(6):569–571, June/July 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641992.

Kim:2006:ASB

[1877] Sung Soo Kim and C. Ray Rosentrater. Another stick-breaking game: 11089. American Mathematical Monthly, 113(6):571–572, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641993.

Quet:2006:CFR

[1878] Leroy Quet and Peter Ørno. A continued fraction related to π: 11102. American Mathematical Monthly, 113(6):572–573, June/July 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641994.

Golomb:2006:CDT

[1879] Solomon W. Golomb and Pat Costello. A consequence of Dirichlet’s theorem: 11109. American Mathematical Monthly, 113(6):573, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641995.

Sofair:2006:IFT

[1880] Isaac Sofair and Otto Ruehr. Integrated by Fourier transform: 11113. American Mathematical Monthly, 113(6):573–574, June/July 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27641996.

Clark:2006:SGT

[1881] Jack Clark and Jaime Vinuesa. A sequence given in terms of harmonic numbers: 11115. American Mathematical Monthly, 113(6):574, June/ July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641997.

Kennedy:2006:BRB

[1882] Steve Kennedy. Book review: Musings of the Masters, by Raymond G. Ayoub. American Mathematical Monthly, 113(6):575–580, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27641998.

Anonymous:2006:BMg

[1883] Anonymous. Back matter. American Mathematical Monthly, 113(6):??, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641999.

Anonymous:2006:FMg

[1884] Anonymous. Front matter. American Mathematical Monthly, 113(6):??, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27641974.

Grier:2006:ISH

[1885] David Alan Grier. Irene stegun, the “handbook of mathematical functions”, and the lingering influence of the new deal. American Mathematical Monthly, 113(7):585–597, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642002.

McCammond:2006:NPS

[1886] Jon McCammond. Noncrossing partitions in surprising locations. American Mathematical Monthly, 113(7):598–610, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642003.

Christensen:2006:LIS

[1887] Ole Christensen and Khadija L. Christensen. Linear independence and series expansions in function spaces. American Mathematical Monthly, 113(7):611–627, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642004.

Schroder:2006:RCS

[1888] Bernd S. W. Schröder. Rearranging the calculus sequence to better serve its partner disciplines. American Mathematical Monthly, 113(7):628– 636, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642005.

Sondow:2006:GPI

[1889] Jonathan Sondow. A geometric proof that e is irrational and a new measure of its irrationality. American Mathematical Monthly, 113(7):

637–641, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642006.

DeAngelis:2006:PSP

[1890] Valerio De Angelis. Pairings and signed permutations. American Mathematical Monthly, 113(7):642–644, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642007.

Netyanun:2006:IPP

[1891] Anupan Netyanun and Donald C. Solmon. Iterated products of projections in Hilbert space. American Mathematical Monthly, 113(7):644– 648, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642008.

Silverman:2006:NDM

[1892] Danielle Silverman and Jim Wiseman. Noting the difference: Musical scales and permutations. American Mathematical Monthly, 113(7):648– 651, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642009.

Pan:2006:CMP

[1893] Hao Pan. On a congruence modulo a prime. American Mathematical Monthly, 113(7):652–654, August/September 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642010.

Owings:2006:P

[1894] Jim Owings. Problem 11236. American Mathematical Monthly, 113(7): 655, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642011.

Deutsch:2006:P

[1895] Emeric Deutsch. Problem 11237. American Mathematical Monthly, 113 (7):655, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642012.

Fraenkel:2006:P

[1896] Aviezri S. Fraenkel. Problem 11238. American Mathematical Monthly, 113(7):655, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642013.

Bataille:2006:Pb

[1897] Michel Bataille. Problem 11239. American Mathematical Monthly, 113 (7):655, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642014.

Dalyay:2006:Pa

[1898] Pál Péter Dályay. Problem 11240. American Mathematical Monthly, 113 (7):655, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642015.

Tauraso:2006:P

[1899] Roberto Tauraso. Problem 11241. American Mathematical Monthly, 113 (7):656, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642016.

Herzog:2006:Pa

[1900] Gerd Herzog and Roland Lemmert. Problem 11242. American Mathematical Monthly, 113(7):656, August/September 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642017.

Callan:2006:LPE

[1901] David Callan and Kenneth Bernstein. A lattice path equality: 11106. American Mathematical Monthly, 113(7):656–657, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642018.

Veljan:2006:IDT

[1902] Darko Veljan and Philipp Lampe. Incenter distances in a triangulation: 11116. American Mathematical Monthly, 113(7):657, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642019.

Dimitrov:2006:MBC

[1903] Vesselin Dimitrov and Robin Chapman. A modular binomial coefficient identity: 11118. American Mathematical Monthly, 113(7):657– 658, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642020.

Robertson:2006:PGP

[1904] John Robertson and Microsoft Research Problems Group. A property of the generalized Pell equation: 11119. American Mathematical Monthly, 113(7):658–659, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642021.

ShahAli:2006:PCI

[1905] H. A. ShahAli and Michael Reid. Powers of consecutive integers: 11121. American Mathematical Monthly, 113(7):659–660, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642022.

Zhou:2006:ICS

[1906] Li Zhou and Michael Avidon. An inequality of cosines and sines: 11127. American Mathematical Monthly, 113(7):660–661, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642023.

deBruijn:2006:SS

[1907] N. G. de Bruijn and Rolf Richberg. Sum of a series: 11131. American Mathematical Monthly, 113(7):661–662, August/September 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642024.

Zitarelli:2006:BRBb

[1908] David E. Zitarelli. Book review: When Computers Were Human, by David Alan Grier. American Mathematical Monthly, 113(7):663– 668, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642025.

Palka:2006:EE

[1909] Bruce P. Palka. Editor’s endnotes. American Mathematical Monthly, 113(7):669–672, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642026.

Anonymous:2006:BMh

[1910] Anonymous. Back matter. American Mathematical Monthly, 113(7): ??, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642027.

Anonymous:2006:FMh

[1911] Anonymous. Front matter. American Mathematical Monthly, 113(7): ??, August/September 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642001.

OBryant:2006:AAS

[1912] Kevin O’Bryant, Bruce Reznick, and Monika Serbinowska. Almost alternating sums. American Mathematical Monthly, 113(8):673–688, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642030.

Velleman:2006:EVF

[1913] Daniel J. Velleman. Exponential vs. factorial. American Mathematical Monthly, 113(8):689–704, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642031.

Suzuki:2006:LPF

[1914] Jeff Suzuki. Lagrange’s proof of the fundamental theorem of algebra. American Mathematical Monthly, 113(8):705–714, October 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642032.

Kufner:2006:PHI

[1915] Alois Kufner, Lech Maligranda, and Lars-Erik Persson. The prehistory of the Hardy inequality. American Mathematical Monthly, 113(8):715–732, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642033.

Klosinski:2006:SSW

[1916] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Sixth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 113(8):733–743, October 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642034.

Lomeli:2006:VTK

[1917] Héctor E. Lomelí and César L. García. Variations on a theorem of korovkin. American Mathematical Monthly, 113(8):744–750, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642035.

Abramson:2006:SCR

[1918] Ian Abramson and Larry Goldstein. A statistical characterization of regular simplices. American Mathematical Monthly, 113(8):750–752, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642036.

Long:2006:HTT

[1919] Lyle N. Long and Howard Weiss. How terminal is terminal velocity? American Mathematical Monthly, 113(8):752–755, October 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642037.

Singh:2006:SBA

[1920] Dinesh Singh. The spectrum in a Banach algebra. American Mathematical Monthly, 113(8):756–758, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642038.

Knuth:2006:P

[1921] Donald Knuth. Problem 11243. American Mathematical Monthly, 113(8): 759, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642039.

Harris:2006:P

[1922] Jolene Harris and Bogdan Suceavǎ. Problem 11244. American Mathematical Monthly, 113(8):759–760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642040.

Lupu:2006:P

[1923] Cezar Lupu and Tudorel Lupu. Problem 11245. American Mathematical Monthly, 113(8):760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642041.

Goldstein:2006:P

[1924] Lee Goldstein. Problem 11246. American Mathematical Monthly, 113(8): 760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642042.

Eckhoff:2006:P

[1925] Jürgen Eckhoff. Problem 11247. American Mathematical Monthly, 113(8):760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642043.

Dalyay:2006:Pb

[1926] Pál Péter Dályay. Problem 11248. American Mathematical Monthly, 113(8):760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642044.

Beckwith:2006:Pc

[1927] David Beckwith. Problem 11249. American Mathematical Monthly, 113(8):760, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642045.

deWeger:2006:HMP

[1928] Benne de Weger and Microsoft Research Problems Group. How many powers of 2 must survive? 11114. American Mathematical Monthly, 113(8):760–761, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642046.

Nyblom:2006:CPI

[1929] Michael Nyblom, Giulio Francot, and Roberto Tauraso. Counting powers by inclusion and exclusion: 11117. American Mathematical Monthly, 113(8):762, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642047.

Berk:2006:GCH

[1930] Bob Berk and Larry Shepp. Gaussian convex hull: 11120. American Mathematical Monthly, 113(8):762–763, October 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642048.

Amrahov:2006:NWY

[1931] Shahin Amrahov and Andrew D. Loveless. Now, what year is it? 11125. American Mathematical Monthly, 113(8):763, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642049.

Mortini:2006:CLI

[1932] Raymond Mortini and Patrick J. Fitzsimmons. Countably linearly independent: 11136. American Mathematical Monthly, 113(8):763–764, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642050.

Bennett:2006:OI

[1933] Grahame Bennett and O. P. Lossers. An odd inequality: 11139. American Mathematical Monthly, 113(8):764, October 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642051.

Janous:2006:CMF

[1934] Walther Janous and Rolf Richberg. A completely monotonic function: 11140. American Mathematical Monthly, 113(8):764–765, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642052.

Bagby:2006:IDC

[1935] Richard Bagby and O. P. Lossers. Implicit derivative continuous: 11141. American Mathematical Monthly, 113(8):765–766, October 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642053.

Zinn:2006:SI

[1936] Joel Zinn and Grahame Bennett. A sum inequality: 11145. American Mathematical Monthly, 113(8):766, October 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642054.

Gangolli:2006:BRB

[1937] Ramesh Gangolli. Book review: Dark Hero of the Information Age: In Search of Norbert Wiener, the Father of Cybernetics, by Flo Conway; Jim Siegelman. American Mathematical Monthly, 113(8):767–774, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642055.

Anonymous:2006:BMi

[1938] Anonymous. Back matter. American Mathematical Monthly, 113(8):??, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642056.

Anonymous:2006:FMi

[1939] Anonymous. Front matter. American Mathematical Monthly, 113(8):??, October 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642029.

Katz:2006:CCA

[1940] Gabriel Katz. Curves in cages: An algebro-geometric zoo. American Mathematical Monthly, 113(9):777–791, November 2006. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642059.

Barbu:2006:UAP

[1941] Ion Barbu, Sarah Glaz, and JoAnne Growney. Ut algebra poesis (as algebra, so poetry). American Mathematical Monthly, 113(9):792–793, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642060.

Melman:2006:OFP

[1942] Aaron Melman and Bill Gragg. An optimization framework for polynomial zerofinders. American Mathematical Monthly, 113(9):794–804, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642061.

Anderson:2006:MRC

[1943] Glen Anderson, Mavina Vamanamurthy, and Matti Vuorinen. Monotonicity rules in calculus. American Mathematical Monthly, 113(9):805– 816, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642062.

Olkin:2006:SCI

[1944] Ingram Olkin and Larry Shepp. Several colorful inequalities. American Mathematical Monthly, 113(9):817–822, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642063.

Snyder:2006:CBS

[1945] David F. Snyder. Combinatorics of barycentric subdivision and characters of simplicial two-complexes. American Mathematical Monthly, 113(9):822–826, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642064.

Brady:2006:PER

[1946] Thomas Brady and Colum Watt. On products of Euclidean reflections. American Mathematical Monthly, 113(9):826–829, November 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642065.

Komornik:2006:ASP

[1947] Vilmos Komornik. Another short proof of Descartes’s rule of signs. American Mathematical Monthly, 113(9):829–830, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642066.

Pragacz:2006:LWA

[1948] Piotr Pragacz, Abe Shenitzer, John Stillwell, and Roman Duda. The life and work of Alexander Grothendieck. American Mathematical Monthly, 113(9):831–846, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642067.

Cai:2006:P

[1949] Sun Wen Cai. Problem 11250. American Mathematical Monthly, 113(9): 847, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642068.

Tetiva:2006:Pc

[1950] Marian Tetiva. Problem 11251. American Mathematical Monthly, 113(9): 847, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642069.

Bagdasar:2006:P

[1951] Ovidiu Bagdasar. Problem 11252. American Mathematical Monthly, 113(9):847, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642070.

Beckwith:2006:Pd

[1952] David Beckwith. Problem 11253. American Mathematical Monthly, 113(9):847, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642071.

Jones:2006:Pb

[1953] Lenny Jones. Problem 11254. American Mathematical Monthly, 113 (9):847–848, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642072.

Simic:2006:P

[1954] Slavko Simic. Problem 11255. American Mathematical Monthly, 113(9): 848, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642073.

Holland:2006:P

[1955] Finbarr Holland. Problem 11256. American Mathematical Monthly, 113(9):848, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642074.

Herzog:2006:Pb

[1956] Gerd Herzog and Roland Lemmert. Problem 11242. American Mathematical Monthly, 113(9):848, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642075.

Pambuccian:2006:PTQ

[1957] Victor Pambuccian and Robin Chapman. Pythagorean triangles are not quite perfect: 11122. American Mathematical Monthly, 113(9):848–850, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642076.

Weber:2006:HMH

[1958] James S. Weber. How many histograms can a data set have? 11126. American Mathematical Monthly, 113(9):850–851, November 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642077.

Tian:2006:IIR

[1959] Yongge Tian and Robin Chapman. Inequalities involving rank and Kronecker product: 11129. American Mathematical Monthly, 113(9):851, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642078.

Bateman:2006:ARZ

[1960] Paul Bateman and Michael R. Avidon. Approximating the reciprocal of zeta: 11130. American Mathematical Monthly, 113(9):852–854, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642079.

Gorkin:2006:MF

[1961] Pamela Gorkin, Raymond Mortini, and the Microsoft Research Problems Group. A meromorphic function: 11147. American Mathematical Monthly, 113(9):854, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642080.

Hampton:2006:BRB

[1962] Charles R. Hampton. Book review: Philosophy of Mathematics: An Introduction to the World of Proofs Pictures, by James Robert Brown; Philosophies of Mathematics by Alexander George; Daniel J. Velleman; Thinking about Mathematics: The Philosophy of Mathematics by Stewart Shapiro. American Mathematical Monthly, 113(9):855–860, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642081.

Anonymous:2006:BMj

[1963] Anonymous. Back matter. American Mathematical Monthly, 113(9):??, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642082.

Anonymous:2006:FMj

[1964] Anonymous. Front matter. American Mathematical Monthly, 113(9):??, November 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642058.

Kowalski:2006:BPB

[1965] Emmanuel Kowalski. Bernstein polynomials and Brownian motion. American Mathematical Monthly, 113(10):865–886, December 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642086.

Hall:2006:ART

[1966] Rachel W. Hall and Paul Klingsberg. Asymmetric rhythms and tiling canons. American Mathematical Monthly, 113(10):887–896, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642087.

Pierce:2006:SGN

[1967] Pamela B. Pierce and Daniel J. Velleman. Some generalizations of the notion of bounded variation. American Mathematical Monthly, 113 (10):897–904, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642088.

Hill:2006:SCP

[1968] Richard O. Hill and Thomas H. Parker. A study of core-plus students attending Michigan State University. American Mathematical Monthly, 113(10):905–921, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642089.

Simoson:2006:SAC

[1969] Andrew J. Simoson. Sliding along a chord through a rotating Earth. American Mathematical Monthly, 113(10):922–928, December 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642090.

Ramm:2006:DSM

[1970] A. G. Ramm. Dynamical systems method and a homeomorphism theorem. American Mathematical Monthly, 113(10):928–933, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642091.

Baumslag:2006:SWP

[1971] Benjamin Baumslag. A simple way of proving the Jordan–Hölder– Schreier Theorem. American Mathematical Monthly, 113(10):933–935, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642092.

Browder:2006:CNH

[1972] Andrew Browder. Complex numbers and the Ham Sandwich Theorem. American Mathematical Monthly, 113(10):935–937, December 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642093.

Saidak:2006:NPE

[1973] Filip Saidak. A new proof of Euclid’s theorem. American Mathematical Monthly, 113(10):937–938, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642094.

Struble:2006:P

[1974] Raimond Struble. Problem 11257. American Mathematical Monthly, 113(10):939, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642095.

Kauers:2006:P

[1975] Manuel Kauers. Problem 11258. American Mathematical Monthly, 113(10):939, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642096.

Abe:2006:P

[1976] Nobuhisa Abe. Problem 11259. American Mathematical Monthly, 113(10):939, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642097.

Perfetti:2006:Pb

[1977] Paolo Perfetti. Problem 11260. American Mathematical Monthly, 113(10):939, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642098.

Sofair:2006:P

[1978] Isaac Sofair. Problem 11261. American Mathematical Monthly, 113(10): 940, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642099.

Burungale:2006:P

[1979] Ashay Burungale. Problem 11262. American Mathematical Monthly, 113(10):940, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642100.

Keselman:2006:Pb

[1980] Gregory Keselman. Problem 11263. American Mathematical Monthly, 113(10):940, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642101.

Soland:2006:SI

[1981] Christoph Soland. Secondary inradii: 11046. American Mathematical Monthly, 113(10):940–941, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642102.

Dubickas:2006:PTD

[1982] Arturas Dubickas, Chris Smyth, and Richard Stong. Polynomials with three distinct zeros summing to zero: 11123. American Mathematical Monthly, 113(10):941–942, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642103.

Redheffer:2006:CGI

[1983] Ray Redheffer and Richard Stong. A cousin of Gronwall’s Inequality: 11135. American Mathematical Monthly, 113(10):942–943, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642104.

Tauraso:2006:PSP

[1984] Roberto Tauraso and O. P. Lossers. Perfect squares from punctured squares: 11138. American Mathematical Monthly, 113(10):943–944, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642105.

Ivan:2006:PFP

[1985] Mircea Ivan and Microsoft Research Problems Group. Patrial fractions, product log: 11152. American Mathematical Monthly, 113(10):945, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642106.

Kowalski:2006:PP

[1986] Emmanuel Kowalski and Tanguy Rivoall. A positive polynomial: 11155. American Mathematical Monthly, 113(10):945–946, December 2006. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642107.

Anonymous:2006:EEN

[1987] Anonymous. Errata and end notes for 2006. American Mathematical Monthly, 113(10):946, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642108.

Fenster:2006:BRB

[1988] Della D. Fenster. Book review: Saunders Mac Lane: A Mathematical Autobiography, by Saunders Mac Lane. American Mathematical Monthly, 113(10):947–951, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642109.

Galovich:2006:BRB

[1989] Steven Galovich. Book review: Ants, Bikes, and Clocks: Problem Solving for Undergraduates, by William Briggs. American Mathematical Monthly, 113(10):951–954, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642110.

Anonymous:2006:BMb

[1990] Anonymous. Back matter. American Mathematical Monthly, 113(10):??, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642111.

Anonymous:2006:FMb

[1991] Anonymous. Front matter. American Mathematical Monthly, 113(10):??, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642085.

Anonymous:2006:VI

[1992] Anonymous. Volume information. American Mathematical Monthly, 113(10):??, December 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642084.

Maier:2007:APA

[1993] Peter Maier and Werner Nickel. Attainable patterns in alien tiles. American Mathematical Monthly, 114(1):1–13, January 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642114.

Leise:2007:NOO

[1994] Tanya Leise and Andrew Cohen. Nonlinear oscillators at our fingertips. American Mathematical Monthly, 114(1):14–28, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642115.

Auckly:2007:SQP

[1995] Dave Auckly. Solving the quartic with a pencil. American Mathematical Monthly, 114(1):29–39, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642116.

Merlini:2007:MC

[1996] Donatella Merlini, Renzo Sprugnoli, and Maria Cecilia Verri. The method of coefficients. American Mathematical Monthly, 114(1):40–57, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642117.

Ford:2007:WDC

[1997] William F. Ford and James A. Pennline. When does convergence in the mean imply uniform convergence? American Mathematical Monthly, 114(1):58–60, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642118.

Ungar:2007:KCG

[1998] Šime Ungar. The Koch curve: a geometric proof. American Mathematical Monthly, 114(1):61–66, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642119.

Kim:2007:CE

[1999] Dong-Soo Kim and Young Ho Kim. A characterization of ellipses. American Mathematical Monthly, 114(1):66–70, January 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642120.

Borwein:2007:CDS

[2000] Jonathan M. Borwein. A class of Dirichlet series integrals. American Mathematical Monthly, 114(1):70–76, January 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642121.

Knuth:2007:Pa

[2001] Donald Knuth. Problem 11264. American Mathematical Monthly, 114 (1):77, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642122.

Butler:2007:P

[2002] Steve Butler and Jia Mao. Problem 11265. American Mathematical Monthly, 114(1):77, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642123.

Tetiva:2007:Pa

[2003] Marian Tetiva. Problem 11266. American Mathematical Monthly, 114(1): 77, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642124.

Herman:2007:Pa

[2004] Eugene A. Herman. Problem 11267. American Mathematical Monthly, 114(1):77, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642125.

Glasser:2007:P

[2005] M. L. Glasser. Problem 11268. American Mathematical Monthly, 114(1): 78, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642126.

Moulin:2007:P

[2006] Hervé Moulin. Problem 11269. American Mathematical Monthly, 114(1): 78, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642127.

Sadov:2007:Pa

[2007] Sergey Sadov. Problem 11270. American Mathematical Monthly, 114(1): 78, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642128.

Beckwith:2007:ARA

[2008] David Beckwith. An accidental repetition of the all-ones problem: 10851. American Mathematical Monthly, 114(1):78–79, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642129.

Apostol:2007:SII

[2009] Tom Apostol and Microsoft Research Problems Group. Summing the initial inverted powers: 11128. American Mathematical Monthly, 114(1): 79–80, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642130.

Glasser:2007:EI

[2010] M. L. Glasser and Richard Stong. An elementary integral: 11148. American Mathematical Monthly, 114(1):80–81, January 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642131.

Ivan:2007:LNL

[2011] Mircea Ivan, loan Rasa, and M. A. Prasad. Limit with nested logs: 11149. American Mathematical Monthly, 114(1):81, January 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642132.

Bracken:2007:NR

[2012] Paul Bracken, Michael Goldenberg, and Mark Kaplan. A nonlinear recurrence: 11153. American Mathematical Monthly, 114(1):81–82, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642133.

Gauchman:2007:SRI

[2013] Hillel Gauchman and Northwestern University Math Problem Solving Group. Square root inequality: 11157. American Mathematical Monthly, 114(1):82–83, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642134.

Oman:2007:IRF

[2014] Greg Oman and Michael W. Botsko. Isomorphic rings of functions: 11166. American Mathematical Monthly, 114(1):83, January 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642135.

Mehrabi:2007:GFI

[2015] Mohammad Hossein Mehrabi and O. P. Lossers. A gamma function inequality: 11169. American Mathematical Monthly, 114(1):83–84, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642136.

Simic:2007:MRP

[2016] Slavko Simic and GCHQ Problem Solving Group. Maximum real part of a zero of a polynomial: 11184. American Mathematical Monthly, 114(1): 84, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642137.

Gouvea:2007:BRB

[2017] Fernando Q. Gouvêa. Book review: A First Course in Modular Forms, by Fred Diamond; Jerry Shurman. American Mathematical Monthly, 114(1): 85–90, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642138.

Velleman:2007:EE

[2018] Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 114(1):91, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642139.

Anonymous:2007:BMa

[2019] Anonymous. Back matter. American Mathematical Monthly, 114(1):??, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642140.

Anonymous:2007:FMa

[2020] Anonymous. Front matter. American Mathematical Monthly, 114(1):??, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642113.

Guy:2007:LTM

[2021] Richard K. Guy. The lighthouse theorem, Morley & Malfatti: a budget of paradoxes. American Mathematical Monthly, 114(2):97–141, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642143.

Loh:2007:SIA

[2022] Po-Ru Loh. Stepping to infinity along Gaussian primes. American Mathematical Monthly, 114(2):142–151, February 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642144.

Du:2007:SPS

[2023] Bau-Sen Du. A simple proof of Sharkovsky’s theorem revisited. American Mathematical Monthly, 114(2):152–155, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642145.

Spencer:2007:LTI

[2024] Gwen Spencer and Francis Edward Su. The LSB theorem implies the KKM lemma. American Mathematical Monthly, 114(2):156–159, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642146.

Nadler:2007:PPU

[2025] Sam B. Nadler, Jr. and Donna M. Zitney. Pointwise products of uniformly continuous functions on sets in the real line. American Mathematical Monthly, 114(2):160–163, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642147.

Bluskov:2007:P

[2026] Iliya Bluskov. Problem 11271. American Mathematical Monthly, 114(2): 164, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642148.

Mihai:2007:P

[2027] Vasile Mihai. Problem 11272. American Mathematical Monthly, 114(2): 164, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642149.

Tetiva:2007:Pb

[2028] Marian Tetiva. Problem 11273. American Mathematical Monthly, 114 (2):164–165, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642150.

Knuth:2007:Pb

[2029] Donald Knuth. Problem 11274. American Mathematical Monthly, 114(2): 165, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642151.

Becker:2007:P

[2030] Michael S. Becker. Problem 11275. American Mathematical Monthly, 114(2):165, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642152.

Herman:2007:Pb

[2031] Eugene Herman. Problem 11276. American Mathematical Monthly, 114(2):165, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642153.

Radulescu:2007:EGS

[2032] Vicenţiu Rădulescu and Richard Stong. Exponential growth of a solution: 11137. American Mathematical Monthly, 114(2):165–166, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642154.

Lamb:2007:QS

[2033] George Lamb and J. A. Grzesik. A quadrature on a sphere: 11159. American Mathematical Monthly, 114(2):167, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642155.

Perfetti:2007:SS

[2034] Paolo Perfetti and Microsoft Research Problems Group. Series with sines: 11162. American Mathematical Monthly, 114(2):167–168, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642156.

Radulescu:2007:GLE

[2035] Vicenţiu Rădulescu and Microsoft Research Problems Group. Ginzburg– Landau energy: 11167. American Mathematical Monthly, 114(2):168– 170, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642157.

Namli:2007:FEE

[2036] Suat Namli and O. P. Lossers. A functional equation with exponential solutions: 11180. American Mathematical Monthly, 114(2):170–171, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642158.

Bennett:2007:AH

[2037] Grahame Bennett and Eugene A. Herman. Apply Hölder: 11202. American Mathematical Monthly, 114(2):171, February 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642159.

Folland:2007:BRB

[2038] Gerald B. Folland. Book review: Analysis I: Convergence, Elementary Functions, by Roger Godement; Analysis II: Differential and Integral Calculus, Fourier Series, Holomorphic Functions by Roger Godement. American Mathematical Monthly, 114(2):172–176, February 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642160.

Anonymous:2007:BMc

[2039] Anonymous. Back matter. American Mathematical Monthly, 114(2):??, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642161.

Anonymous:2007:FMc

[2040] Anonymous. Front matter. American Mathematical Monthly, 114(2):??, February 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642142.

Cohen-Corwin:2007:YGG

[2041] Amy Cohen-Corwin. Yueh-Gin Gung and Dr. Charles Y. Hu Award to Lee Lorch for Distinguished Service to Mathematics. American Mathematical Monthly, 114(3):181–182, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642164.

Curtis:2007:CSP

[2042] Robyn Curtis and Marcel Steiner. Configuration spaces of planar pentagons. American Mathematical Monthly, 114(3):183–201, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642165.

Socha:2007:CCC

[2043] Katherine Socha. Circles in circles: Creating a mathematical model of surface water waves. American Mathematical Monthly, 114(3):202–216, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642166.

Claesson:2007:CNP

[2044] Anders Claesson and T. Kyle Petersen. Conway’s napkin problem. American Mathematical Monthly, 114(3):217–231, March 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642167.

Manna:2007:SEN

[2045] Dante Manna and Victor H. Moll. A simple example of a new class of Landen transformation. American Mathematical Monthly, 114(3):232– 241, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642168.

Botsko:2007:EWB

[2046] Michael W. Botsko. Exactly which bounded Darboux functions are derivatives? American Mathematical Monthly, 114(3):242–246, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642169.

Levi:2007:RMT

[2047] Mark Levi. Riemann mapping theorem by steepest descent. American Mathematical Monthly, 114(3):246–251, March 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642170.

Armitage:2007:EFT

[2048] David H. Armitage. Entire functions that tend to zero on every line. American Mathematical Monthly, 114(3):251–256, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642171.

Pinsky:2007:SFP

[2049] Mark A. Pinsky. Stirling’s formula via the Poisson distribution. American Mathematical Monthly, 114(3):256–258, March 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642172.

De:2007:P

[2050] Prithwijit De. Problem 11277. American Mathematical Monthly, 114(3): 259, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642173.

Simic:2007:P

[2051] Slavko Simic. Problem 11278. American Mathematical Monthly, 114(3): 259, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642174.

Stakhovsky:2007:P

[2052] Vitaly Stakhovsky. Problem 11279. American Mathematical Monthly, 114(3):259, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642175.

Kwong:2007:P

[2053] Harris Kwong. Problem 11280. American Mathematical Monthly, 114(3): 259, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642176.

Alekseyev:2007:P

[2054] Max Alekseyev and Emeric Deutsch. Problem 11281. American Mathematical Monthly, 114(3):259, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642177.

Dkovic:2007:P

[2055] Dragomir Ž. – Dković and Kaiming Zhao. Problem 11282. American Mathematical Monthly, 114(3):260, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642178.

Abbot:2007:P

[2056] John Abbot and Umberto Zannier. Problem 11283. American Mathematical Monthly, 114(3):260, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642179.

Gray:2007:SR

[2057] Stephen B. Gray, Thomas McMillan, and Xiaoshen Wang. Stepping to regularity: 10923. American Mathematical Monthly, 114(3):260–261, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642180.

Boklan:2007:HT

[2058] Kent D. Boklan and Michael R. Avidon. Heronian triangles: 11134. American Mathematical Monthly, 114(3):261–262, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642181.

Schulte-Geers:2007:IIR

[2059] Ernst Schulte-Geers and Karl David. An identity involving rooted forest: 11144. American Mathematical Monthly, 114(3):262–264, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642182.

Tian:2007:MSS

[2060] Yongge Tian and Eugene A. Herman. Matrix systems with the same sets: 11146. American Mathematical Monthly, 114(3):264, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642183.

Deutsch:2007:PDP

[2061] Emeric Deutsch and O. P. Lossers. Peaks in Dyck paths: 11150. American Mathematical Monthly, 114(3):264–265, March 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642184.

Knuth:2007:PCS

[2062] Donald E. Knuth and O. P. Lossers. Partitions of a circular set: 11151. American Mathematical Monthly, 114(3):265–266, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642185.

Holing:2007:ISS

[2063] Kent Holing and Michael R. Avidon. Indivisibility of the sum of squares: 11154. American Mathematical Monthly, 114(3):266, March 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642186.

Stillwell:2007:BRB

[2064] John Stillwell. Book review: God Created the Integers, by Stephen Hawking. American Mathematical Monthly, 114(3):267–271, March 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642187.

Anonymous:2007:BMd

[2065] Anonymous. Back matter. American Mathematical Monthly, 114(3):??, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642188.

Anonymous:2007:FMd

[2066] Anonymous. Front matter. American Mathematical Monthly, 114(3):??, March 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642163.

Martinez:2007:EMR

[2067] Alberto A. Martínez. Euler’s “mistake”? the radical product rule in historical perspective. American Mathematical Monthly, 114(4):273–285, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642191.

Francese:2007:FEP

[2068] Christopher Francese and David Richeson. The flaw in Euler’s proof of his polyhedral formula. American Mathematical Monthly, 114(4):286– 296, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642192.

Srinivasan:2007:GFE

[2069] Gopala Krishna Srinivasan. The gamma function: An eclectic tour. American Mathematical Monthly, 114(4):297–315, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642193.

Chow:2007:SGF

[2070] Bennett Chow and David Glickenstein. Semidiscrete geometric flows of polygons. American Mathematical Monthly, 114(4):316–328, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642194.

Adams:2007:WSB

[2071] Colin Adams, Frank Morgan, and John M. Sullivan. When soap bubbles collide. American Mathematical Monthly, 114(4):329–337, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642195.

Benjamin:2007:CPV

[2072] Arthur T. Benjamin and Gregory P. Dresden. A combinatorial proof of Vandermonde’s determinant. American Mathematical Monthly, 114(4): 338–341, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642196.

Allen:2007:ESI

[2073] Michael A. Allen. Evaluation of some improper integrals involving hyperbolic functions. American Mathematical Monthly, 114(4):341–343, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642197.

Pemantle:2007:WE

[2074] Robin Pemantle and Carsten Schneider. When is 0.999… equal to 1? American Mathematical Monthly, 114(4):344–350, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642198.

Foreman:2007:GTS

[2075] Brendan Foreman. Ghys’s theorem and semi-osculating conics of planar curves. American Mathematical Monthly, 114(4):351–356, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642199.

Helmberg:2007:ACF

[2076] Gilbert Helmberg. An absolutely continuous function in L1 (R) W1,1 (R). American Mathematical Monthly, 114(4):356–357, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642200.

Oman:2007:P

[2077] Greg Oman. Problem 11284. American Mathematical Monthly, 114(4): 358, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642201.

Aliyev:2007:Pa

[2078] Yakub Aliyev. Problem 11285. American Mathematical Monthly, 114(4): 358, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642202.

LeBrun:2007:P

[2079] Marc LeBrun, David Applegate, and N. J. A. Sloane. Problem 11286. American Mathematical Monthly, 114(4):358, April 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642203.

Herschkorn:2007:Pa

[2080] Stephen J. Herschkorn. Problem 11287. American Mathematical Monthly, 114(4):359, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642204.

Hillar:2007:Pa

[2081] Christopher Hillar and Troels Windfeldt. Problem 11288. American Mathematical Monthly, 114(4):359, April 2007. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642205.

Faynshteyn:2007:P

[2082] Oleh Faynshteyn. Problem 11289. American Mathematical Monthly, 114 (4):359, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642206.

Lupu:2007:Pa

[2083] Cezar Lupu and Tudorel Lupu. Problem 11290. American Mathematical Monthly, 114(4):359, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642207.

Sondow:2007:CB

[2084] Jonathan Sondow and Richard Stong. Choice bounds: 11132. American Mathematical Monthly, 114(4):359–360, April 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642208.

Bracken:2007:II

[2085] Paul Bracken and Heinz-Jürgen Seiffert. An integral inequality: 11133. American Mathematical Monthly, 114(4):360–361, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642209.

Knuth:2007:CM

[2086] Donald Knuth and Roberto Tauraso. A combinatorial maximum: 11142. American Mathematical Monthly, 114(4):361–362, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642210.

Stanley:2007:IMS

[2087] Richard Stanley and GCHQ Problem Solving Group. Inverting a matrix of striling numbers: 11156. American Mathematical Monthly, 114(4): 362, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642211.

Deutsch:2007:CCF

[2088] Emeric Deutsch and Lafayette College Problem Group. Compositions without common factors: 11161. American Mathematical Monthly, 114 (4):363, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642212.

Bataille:2007:CLQ

[2089] Michel Bataille and Richard Stong. Counting lattice quadrilaterals tangent to an ellipse: 11163. American Mathematical Monthly, 114(4):363– 364, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642213.

Diaz-Barrero:2007:RI

[2090] José Luis Díaz-Barrero, Ulrich Abel, and Mircea Ivan. A recurrent identify: 11164. American Mathematical Monthly, 114(4):364–365, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642214.

DAngelo:2007:BRB

[2091] John P. D’Angelo. Book review: Geometric Function Theory: Explorations in Complex Analysis, by Steven G. Krantz. American Mathematical Monthly, 114(4):366–370, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642215.

Anonymous:2007:BMe

[2092] Anonymous. Back matter. American Mathematical Monthly, 114(4): ??, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642216.

Anonymous:2007:FMe

[2093] Anonymous. Front matter. American Mathematical Monthly, 114(4): ??, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642190.

Peres:2007:RTH

[2094] Yuval Peres, Oded Schramm, Scott Sheffield, and David B. Wilson. Random-turn hex and other selection games. American Mathematical Monthly, 114(5):373–387, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642219.

Apostol:2007:UCC

[2095] Tom M. Apostol and Mamikon A. Mnatsakanian. Unwrapping curves from cylinders and cones. American Mathematical Monthly, 114(5):388– 416, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642220.

Everest:2007:PGR

[2096] Graham Everest, Shaun Stevens, Duncan Tamsett, and Tom Ward. Primes generated by recurrence sequences. American Mathematical Monthly, 114(5):417–431, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642221.

Neuberger:2007:CNM

[2097] J. W. Neuberger. The continuous Newton’s method, inverse functions, and Nash–Moser. American Mathematical Monthly, 114(5):432–437, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642222.

Mendes:2007:NAP

[2098] Anthony Mendes. A note on alternating permutations. American Mathematical Monthly, 114(5):437–440, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642223.

Kerckhove:2007:CCS

[2099] Michael Kerckhove and Gary Lawlor. The circle and the cycloid: a surprising link. American Mathematical Monthly, 114(5):441–443, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642224.

Ohm:2007:MAZ

[2100] Jack Ohm. A matrix approach to zero-divisors in R[x]. American Mathematical Monthly, 114(5):444–450, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642225.

Jerrard:2007:P

[2101] Richard Jerrard and John Wetzel. Problem 11291. American Mathematical Monthly, 114(5):451, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642226.

Callan:2007:Pa

[2102] David Callan. Problem 11292. American Mathematical Monthly, 114(5): 451, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642227.

Sadov:2007:Pb

[2103] Sergey Sadov. Problem 11293. American Mathematical Monthly, 114(5): 451, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642228.

Zucker:2007:P

[2104] John Zucker and Ross McPhedran. Problem 11294. American Mathematical Monthly, 114(5):452, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642229.

Siboni:2007:P

[2105] Stefano Siboni. Problem 11295. American Mathematical Monthly, 114 (5):452, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642230.

Amghibech:2007:P

[2106] Said Amghibech. Problem 11296. American Mathematical Monthly, 114 (5):452, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642231.

Tetiva:2007:Pc

[2107] Marian Tetiva. Problem 11297. American Mathematical Monthly, 114(5): 452, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642232.

Nica:2007:SES

[2108] Bogdan Nica and S. Amghibech. Similarities of Euclidean space: 11124. American Mathematical Monthly, 114(5):452–453, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642233.

Dubinov:2007:CFC

[2109] Alexander Dubinov and Irina Dubinova. Closed form closure: 11143. American Mathematical Monthly, 114(5):453, May 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642234.

Beckwith:2007:PPD

[2110] David Beckwith and Harris Kwong. Prime powers dividing a factorial: 11158. American Mathematical Monthly, 114(5):453–454, May 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642235.

Chamberland:2007:FR

[2111] Marc Chamberland and C. R. Pranesachar. Fermat to the rescue: 11160. American Mathematical Monthly, 114(5):454, May 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642236.

More:2007:SCN

[2112] Yogesh More and Robin Chapman. The sum of Catalan numbers, modulo 3: 11165. American Mathematical Monthly, 114(5):454–455, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642237.

Holing:2007:PPT

[2113] Kent Holing and Microsoft Research Problems Group. Properties of Pythagorean triples: 11168. American Mathematical Monthly, 114(5): 455–456, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642238.

Deutsch:2007:DP

[2114] Emeric Deutsch and S. C. Locke. Dissecting a polygon: 11170. American Mathematical Monthly, 114(5):456–457, May 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642239.

Seiffert:2007:FID

[2115] Heinz-Jürgen Seiffert. Familiar identities in disguise: 11172. American Mathematical Monthly, 114(5):457–458, May 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642240.

Bressoud:2007:BRB

[2116] David M. Bressoud. Book review: Making Transcendence Transparent, by B. Burger; Robert Tubbs. American Mathematical Monthly, 114(5): 459–461, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642241.

Schwartz:2007:BRB

[2117] Mark Schwartz. Book review: Conics, by Keith Kendig. American Mathematical Monthly, 114(5):461–464, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642242.

Anonymous:2007:BMf

[2118] Anonymous. Back matter. American Mathematical Monthly, 114(5): ??, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642243.

Anonymous:2007:FMf

[2119] Anonymous. Front matter. American Mathematical Monthly, 114(5): ??, May 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642218.

Thomson:2007:RER

[2120] Brian S. Thomson. Rethinking the elementary real analysis course. American Mathematical Monthly, 114(6):469–490, June/July 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642246.

Hetzel:2007:PMI

[2121] Andrew J. Hetzel, Jay S. Liew, and Kent E. Morrison. The probability that a matrix of integers is diagonalizable. American Mathematical Monthly, 114(6):491–499, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642247.

Johnson:2007:NRS

[2122] Brody Dylan Johnson. The nonholonomy of the rolling sphere. American Mathematical Monthly, 114(6):500–508, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642248.

Schleicher:2007:HDP

[2123] Dierk Schleicher. Hausdorff dimension, its properties, and its surprises. American Mathematical Monthly, 114(6):509–528, June/July 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642249.

Shi:2007:AWH

[2124] Ling-Ling Shi and Hao Pan. A q-analogue of Wolstenholme’s harmonic series congruence. American Mathematical Monthly, 114(6):529–531, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642250.

Bliedtner:2007:LMF

[2125] Jürgen Bliedtner and Peter A. Loeb. A local maximal function simplifying measure differentiation. American Mathematical Monthly, 114 (6):532–536, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642251.

Adamczewski:2007:SPT

[2126] Boris Adamczewski and Yann Bugeaud. A short proof of the transcendence of Thue–Morse continued fractions. American Mathematical Monthly, 114(6):536–540, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642252.

Bak:2007:EPC

[2127] Joseph Bak, Pisheng Ding, and Donald Newman. Extremal points, critical points, and saddle points of analytic functions. American Mathematical Monthly, 114(6):540–546, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642253.

Jonsson:2007:P

[2128] Jakob Jonsson. Problem 11298. American Mathematical Monthly, 114(6): 547, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642254.

Refolio:2007:Pa

[2129] Pablo Fernàndez Refolio. Problem 11299. American Mathematical Monthly, 114(6):547, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642255.

Abel:2007:P

[2130] Ulrich Abel. Problem 11300. American Mathematical Monthly, 114(6): 547, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642256.

Holland:2007:Pa

[2131] Finbarr Holland. Problem 11301. American Mathematical Monthly, 114(6):547, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642257.

Alzer:2007:P

[2132] Horst Alzer. Problem 11302. American Mathematical Monthly, 114(6): 547, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642258.

Farrokhi:2007:Pa

[2133] M. Farrokhi. Problem 11303. American Mathematical Monthly, 114(6): 548, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642259.

Radulescu:2007:P

[2134] Teodora-Liliana Rădulescu. Problem 11304. American Mathematical Monthly, 114(6):548, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642260.

Deshpande:2007:THB

[2135] M. N. Deshpande, J. P. Shiwalkar, and GCHQ Problem Solving Group. Two heads are better than one: 11173. American Mathematical Monthly, 114(6):548–549, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642261.

Bentley:2007:LE

[2136] Jon Bentley, Colin Mallows, and Richard Stong. Learning from experience: 11178. American Mathematical Monthly, 114(6):549–550, June/ July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642262.

Beckwith:2007:DMI

[2137] David Beckwith and Li Zhou. A determinant by Möbius inversion: 11179. American Mathematical Monthly, 114(6):550, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642263.

Amrahov:2007:PPA

[2138] Shahin Amrahov and Nicholas C. Singer. Perfect powers in an arithmetic progression: 11182. American Mathematical Monthly, 114(6):550–551, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642264.

Beckwith:2007:BTP

[2139] David Beckwith, Vadim Ponomarenko, and Rob Pratt. Building two piles of equal height: 11183. American Mathematical Monthly, 114(6):551–552, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642265.

Bruck:2007:WDI

[2140] Rainer Brück and Robin Chapman. When does it converge: 11185. American Mathematical Monthly, 114(6):552–554, June/July 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642266.

Zhou:2007:TRR

[2141] Li Zhou, Northwestern University Math Problem Solving Group, Christopher Carl Heckman, and Jerry Minkus. Tiling 4-rowed rectangles with dominoes: 11187. American Mathematical Monthly, 114(6):554–556, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642267.

Colley:2007:BRB

[2142] Susan Jane Colley. Book review: Introduction to Plane Algebraic Curves, by Ernst Kunz; Richard G. Belshoff. American Mathematical Monthly, 114(6):557–561, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642268.

Anonymous:2007:BMg

[2143] Anonymous. Back matter. American Mathematical Monthly, 114(6):??, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642269.

Anonymous:2007:FMg

[2144] Anonymous. Front matter. American Mathematical Monthly, 114(6):??, June/July 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642245.

Velasco:2007:JGC

[2145] Enrique A. González-Velasco. James Gregory’s calculus in the “Geometriæ Pars Universalis”. American Mathematical Monthly, 114(7): 565–576, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642272.

Dwilewicz:2007:HBE

[2146] Roman Dwilewicz, Ján Mináč, Andrew Schultz, and John Swallow. Hilbert 90 for biquadratic extensions. American Mathematical Monthly, 114(7):577–587, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642273.

Berger:2007:NMO

[2147] Arno Berger and Theodore P. Hill. Newton’s method obeys Benford’s law. American Mathematical Monthly, 114(7):588–601, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642274.

Akiyama:2007:TMS

[2148] Jin Akiyama. Tile-makers and semi-tile-makers. American Mathematical Monthly, 114(7):602–609, August/September 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642275.

Walden:2007:HMI

[2149] Byron L. Walden and Lesley A. Ward. A harmonic measure interpretation of the arithmetic-geometric mean. American Mathematical Monthly, 114(7):610–622, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642276.

Escudero:2007:IPE

[2150] Carlos A. Escudero and Agustí Reventós. An interesting property of the evolute. American Mathematical Monthly, 114(7):623–628, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642277.

Cook:2007:RNP

[2151] Jack Cook, Jonathan Lovett, and Frank Morgan. Rotation in a normed plane. American Mathematical Monthly, 114(7):628–632, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642278.

Cohen:2007:CLA

[2152] Arjeh M. Cohen and D. E. Taylor. On a certain Lie algebra defined by a finite group. American Mathematical Monthly, 114(7):633–639, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642279.

Mkrtchyan:2007:P

[2153] Vahan Mkrtchyan. Problem 11305. American Mathematical Monthly, 114 (7):640, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642280.

Rosoiu:2007:P

[2154] Alexandru Rosoiu. Problem 11306. American Mathematical Monthly, 114 (7):640, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642281.

Callan:2007:Pb

[2155] David Callan. Problem 11307. American Mathematical Monthly, 114 (7):640, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/ 27642282. See errata [2249].

Furdui:2007:Pa

[2156] Ovidiu Furdui. Problem 11308. American Mathematical Monthly, 114 (7):640, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642283.

Witula:2007:Pa

[2157] Roman Wituła and Damian Słota. Problem 11309. American Mathematical Monthly, 114(7):640, August/September 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642284.

Just:2007:P

[2158] Erwin Just. Problem 11310. American Mathematical Monthly, 114(7): 641, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642285.

Herschkorn:2007:Pb

[2159] Stephen J. Herschkorn. Problem 11311. American Mathematical Monthly, 114(7):641, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642286.

Calbeck:2007:ADE

[2160] William Calbeck and Richard Stong. Another Diophantine equation: 11175. American Mathematical Monthly, 114(7):641–642, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642287.

Deshpande:2007:UPL

[2161] M. N. Deshpande and GCHQ Problem Solving Group. An urn problem with labeled and unlabeled balls: 11181. American Mathematical Monthly, 114(7):642, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642288.

Christino:2007:GIM

[2162] Angelo J. Christino, Jr., William C. Waterhouse, and the NSA Problems Group. A group of invertible matrices: 11186. American Mathematical Monthly, 114(7):642–643, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642289.

Bailey:2007:HI

[2163] David Bailey, M. L. Glasser, and Richard Stong. A hypergeometric identity: 11188. American Mathematical Monthly, 114(7):643–645, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642290.

Csete:2007:QI

[2164] Lajos Csete and O. P. Lossers. A quick inequality: 11189. American Mathematical Monthly, 114(7):645, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642291.

Diaz-Barrero:2007:PZ

[2165] José Luis Díaz-Barrero and Julien Grivaux. A polynomial and its zeros: 11190. American Mathematical Monthly, 114(7):645–647, August/ September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642292.

Tetiva:2007:CGI

[2166] Marian Tetiva. A cyclic AM-GM inequality: 11193. American Mathematical Monthly, 114(7):647, August/September 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642293.

Tetiva:2007:RCR

[2167] Marian Tetiva and Pál Péter Dályay. Ratio circum-to-in-radius: 11195. American Mathematical Monthly, 114(7):648–649, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642294.

Strubel:2007:ZFQ

[2168] R. A. Strubel and Donald R. Bridges. A zeta function quickie: 11219. American Mathematical Monthly, 114(7):649, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642296.

Yakub:2007:SI

[2169] Aliyev Yakub and David B. Leep. Symmetric inequality: 11199. American Mathematical Monthly, 114(7):649, August/September 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642295.

Fulling:2007:BRB

[2170] Stephen A. Fulling. Book review: Linearity, Symmetry, and Prediction in the Hydrogen Atom, by Stephanie Frank Singer. American Mathematical Monthly, 114(7):650–653, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642297.

Fey:2007:EE

[2171] James Fey, Eric Hart, Christian Hirsch, Harold Schoen, Ann Watkins, and Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 114(7):654–661, August/September 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642298.

Anonymous:2007:BMh

[2172] Anonymous. Back matter. American Mathematical Monthly, 114(7): ??, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642299.

Anonymous:2007:FMh

[2173] Anonymous. Front matter. American Mathematical Monthly, 114(7): ??, August/September 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642271.

Benko:2007:NAH

[2174] David Benko. A new approach to Hilbert’s third problem. American Mathematical Monthly, 114(8):665–676, October 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642302.

Jaroma:2007:CAA

[2175] John H. Jaroma and Kamaliya N. Reddy. Classical and alternative approaches to the Mersenne and Fermat numbers. American Mathematical Monthly, 114(8):677–687, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642303.

Szekely:2007:UPG

[2176] Gábor J. Székely and Maria L. Rizzo. The uncertainty principle of game theory. American Mathematical Monthly, 114(8):688–702, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642304.

Derksen:2007:MS

[2177] Harm Derksen, Christian Eggermont, and Arno van den Essen. Multimagic squares. American Mathematical Monthly, 114(8):703–713, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642305.

Klosinski:2007:SSW

[2178] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Seventh William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 114(8):714–724, October 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642306.

Lax:2007:CIT

[2179] Peter D. Lax. The Cauchy integral theorem. American Mathematical Monthly, 114(8):725–727, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642307.

Bornemann:2007:THS

[2180] Folkmar Bornemann and Thomas Schmelzer. Taming a hydra of singularities. American Mathematical Monthly, 114(8):727–732, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642308.

Murty:2007:PTM

[2181] M. Ram Murty and Nithum Thain. Pick’s theorem via Minkowski’s theorem. American Mathematical Monthly, 114(8):732–736, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642309.

Albiac:2007:CRS

[2182] F. Albiac and N. J. Kalton. A characterization of real C(K)-spaces. American Mathematical Monthly, 114(8):737–743, October 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642310.

Callan:2007:Pc

[2183] David Callan. Problem 11312. American Mathematical Monthly, 114(8): 744, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642311.

Lupu:2007:Pb

[2184] Cezar Lupu and Tudorel Lupu. Problem 11313. American Mathematical Monthly, 114(8):744, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642312.

Aliyev:2007:Pb

[2185] Yakub Aliyev. Problem 11314. American Mathematical Monthly, 114(8): 744, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642313.

Farrokhi:2007:Pb

[2186] M. Farrokhi. Problem 11315. American Mathematical Monthly, 114(8): 744, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642314.

Furdui:2007:Pb

[2187] Ovidiu Furdui. Problem 11316. American Mathematical Monthly, 114(8): 745, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642315.

Bataille:2007:P

[2188] Michel Bataille. Problem 11317. American Mathematical Monthly, 114(8):745, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642316.

Witula:2007:Pb

[2189] Roman Wituła and Damian Słota. Problem 11318. American Mathematical Monthly, 114(8):745, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642317.

Heath:2007:BLT

[2190] R. V. Heath and Christian Boyer. Better late than never: E 496. American Mathematical Monthly, 114(8):745–746, October 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642318.

Bornemann:2007:DRC

[2191] Folkmar Bornemann and Robert B. Israel. A disguised radius of convergence: 11194. American Mathematical Monthly, 114(8):746–747, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642319.

Faynshteyn:2007:CI

[2192] Oleg Faynshteyn and Arkady Alt. A crux inequality: 11197. American Mathematical Monthly, 114(8):747–748, October 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642320.

Zinn:2007:KLC

[2193] Joel Zinn and O. P. Lossers. Kronecker-like convergence: 11200. American Mathematical Monthly, 114(8):748–749, October 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642321.

Becker:2007:CIP

[2194] Michael S. Becker and Zuofeng Shang. Cosine infinite product: 11210. American Mathematical Monthly, 114(8):749, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642322.

Bencze:2007:MGI

[2195] Mihály Bencze and Heinz-Jürgen Seiffert. More A-G inequalities: 11211. American Mathematical Monthly, 114(8):749–750, October 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642323.

Diaz-Barrero:2007:LI

[2196] José Luis Díaz-Barrero and Eugene Herman. Limit of an integral: 11225. American Mathematical Monthly, 114(8):750, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642324.

Furdui:2007:GL

[2197] Ovidiu Furdui and Michael R. Avidon. A gamma limit: 11229. American Mathematical Monthly, 114(8):751, October 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642325.

Johnson:2007:BRB

[2198] Warren P. Johnson. Book review: Mathematics and Social Utopias in France: Olinde Rodrigues and His times, by Simon Altmann; Eduardo L. Ortiz. American Mathematical Monthly, 114(8):752–758, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642326.

Anonymous:2007:BMi

[2199] Anonymous. Back matter. American Mathematical Monthly, 114(8):??, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642327.

Anonymous:2007:FMi

[2200] Anonymous. Front matter. American Mathematical Monthly, 114(8):??, October 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642301.

Callaghan:2007:RWR

[2201] Thomas Callaghan, Peter J. Mucha, and Mason A. Porter. Random Walker ranking for NCAA division I-A football. American Mathematical Monthly, 114(9):761–777, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642330.

Starrett:2007:SDE

[2202] John Starrett. Solving differential equations by symmetry groups. American Mathematical Monthly, 114(9):778–792, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642331.

Bartnicki:2007:MCG

[2203] Tomasz Bartnicki, Jaroslaw Grytczuk, H. A. Kierstead, and Xuding Zhu. The map-coloring game. American Mathematical Monthly, 114(9):793– 803, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642332.

Montgomery:2007:BTP

[2204] Hugh L. Montgomery and Ulrike M. A. Vorhauer. Biased trigonometric polynomials. American Mathematical Monthly, 114(9):804–809, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642333.

Adelman:2007:MLP

[2205] Omer Adelman. ∑ \frac{1}{n} = ∞: a micro-lesson on probability and symmetry. American Mathematical Monthly, 114(9):809–810, November 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642334.

Lotto:2007:SG

[2206] Benjamin A. Lotto. Stacked groups. American Mathematical Monthly, 114(9):811–812, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642335.

Harrison:2007:RCP

[2207] Brett A. Harrison. On the reducibility of cyclotomic polynomials over finite fields. American Mathematical Monthly, 114(9):813–818, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642336.

Morse:2007:TE

[2208] Marston Morse. Topology and equilibria. American Mathematical Monthly, 114(9):819–834, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642337.

Beckwith:2007:P

[2209] David Beckwith. Problem 11319. American Mathematical Monthly, 114(9):835, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642338.

Knuth:2007:Pc

[2210] Donald Knuth. Problem 11320. American Mathematical Monthly, 114(9): 835, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642339.

Hillar:2007:Pb

[2211] Christopher Hillar. Problem 11321. American Mathematical Monthly, 114(9):835, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642340.

Sondow:2007:P

[2212] Jonathan Sondow. Problem 11322. American Mathematical Monthly, 114(9):835, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642341.

Lockhart:2007:P

[2213] Jody M. Lockhart and William P. Wardlaw. Problem 11323. American Mathematical Monthly, 114(9):835, November 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642342.

Kochanski:2007:P

[2214] Martin Kochanski. Problem 11324. American Mathematical Monthly, 114(9):836, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642343.

Holland:2007:Pb

[2215] Finbarr Holland and Jim Leahy. Problem 11325. American Mathematical Monthly, 114(9):836, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642344.

Diaz-Barrero:2007:P

[2216] José Luis Díaz-Barrero and Pantelimon George Popescu. Problem 11326. American Mathematical Monthly, 114(9):836, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642345.

Dalyay:2007:PC

[2217] Pál Péter Dályay and National Security Agency Problems Group. Periodic composition: 11174. American Mathematical Monthly, 114(9):836– 837, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642346.

Faynshteyn:2007:ATI

[2218] Oleg Faynshteyn and Yan loi Wong. Another triangle inequality: 11176. American Mathematical Monthly, 114(9):837–838, November 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642347.

Marghidanu:2007:X

[2219] Dorin Mǎrghidanu and Henry Ricardo. GM ≤ AMF: 11177. American Mathematical Monthly, 114(9):838, November 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642348.

Huff:2007:AUB

[2220] Barthel Wayne Huff and Gregory Keselman. Another urn and balls problem: 11191. American Mathematical Monthly, 114(9):838–839, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642349.

Stevens:2007:SSO

[2221] Brett Stevens and Richard Stong. A sudoku solution from orthogonal Latin squares: 11192. American Mathematical Monthly, 114(9):839–840, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642350.

Yu:2007:OFP

[2222] Fuxiang Yu. An old fermatian problem: 11203. American Mathematical Monthly, 114(9):840, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642351.

Hillar:2007:TSA

[2223] Christopher J. Hillar, John H. Lindsey II, and Robin Chapman. The trace of the sum of all products of two matrices: 11204. American Mathematical Monthly, 114(9):840–841, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642352.

Zhou:2007:CGS

[2224] Li Zhou, Chris Mackeprang, and Kellen Myers. Coloring graphs on sponges: 11208. American Mathematical Monthly, 114(9):842, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642353.

Nelson:2007:BRB

[2225] Edward Nelson. Book review: 18 Unconventional Essays on the Nature of Mathematics, by Reuben Hersh. American Mathematical Monthly, 114(9):843–848, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642354.

Anonymous:2007:BMj

[2226] Anonymous. Back matter. American Mathematical Monthly, 114(9):??, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642355.

Anonymous:2007:FMj

[2227] Anonymous. Front matter. American Mathematical Monthly, 114(9):??, November 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642329.

Hadlock:2007:PMP

[2228] Charles R. Hadlock. Practicing mathematics in the public arena: Challenges and outcomes in some prominent case studies. American Mathematical Monthly, 114(10):849–870, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642359.

Halmetoja:2007:DIP

[2229] Markku Halmetoja, Pentti Haukkanen, Teuvo Laurinolli, Jorma K. Merikoski, Timo Tossavainen, and Ari Virtanen. On direct and inverse proportionality. American Mathematical Monthly, 114(10):871–881, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642360.

Hales:2007:JCT

[2230] Thomas C. Hales. The Jordan curve theorem, formally and informally. American Mathematical Monthly, 114(10):882–894, December 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642361.

Levi:2007:PGB

[2231] Mark Levi and Serge Tabachnikov. The Poncelet grid and billiards in ellipses. American Mathematical Monthly, 114(10):895–908, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642362.

Connes:2007:PPI

[2232] Alain Connes and Don Zagier. A property of parallelograms inscribed in ellipses. American Mathematical Monthly, 114(10):909–914, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642363.

Wastlund:2007:EPW

[2233] Johan Wästlund. An elementary proof of the Wallis product formula for π. American Mathematical Monthly, 114(10):914–917, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642364.

Hillar:2007:AFA

[2234] Christopher J. Hillar and Darren L. Rhea. Automorphisms of finite Abelian groups. American Mathematical Monthly, 114(10):917–923, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642365.

deOliveira:2007:AM

[2235] António Guedes de Oliveira. On the adjugate of a matrix. American Mathematical Monthly, 114(10):923–924, December 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642366.

Mabry:2007:P

[2236] Rick Mabry and Debbie Shepherd. Problem 11327. American Mathematical Monthly, 114(10):925, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642367.

Vartziotis:2007:P

[2237] Dmitris Vartziotis. Problem 11328. American Mathematical Monthly, 114(10):925, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642368.

Amdeberhan:2007:P

[2238] T. Amdeberhan and Victor H. Moll. Problem 11329. American Mathematical Monthly, 114(10):925, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642369.

Tetiva:2007:Pd

[2239] Marian Tetiva. Problem 11330. American Mathematical Monthly, 114(10):925, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642370.

Bagby:2007:P

[2240] Richard Bagby. Problem 11331. American Mathematical Monthly, 114(10):926, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642371.

Tauraso:2007:P

[2241] Roberto Tauraso. Problem 11332. American Mathematical Monthly, 114(10):926, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642372.

Refolio:2007:Pb

[2242] Pablo Fernández Refolio. Problem 11333. American Mathematical Monthly, 114(10):926, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642373.

Cheng:2007:QC

[2243] Chu Cheng and Marius Stefan. A quadrilateral and its circles: 11171. American Mathematical Monthly, 114(10):926–927, December 2007. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642374.

Chao:2007:TI

[2244] Wu Wei Chao and Pál Péter Dályay. A triangle inequality: 11205. American Mathematical Monthly, 114(10):928, December 2007. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642375.

Ivan:2007:LIE

[2245] Mircea Ivan, Alexandru Lupaş, and Richard A. Stong. A limit involving Euler’s constant: 11206. American Mathematical Monthly, 114(10):928– 929, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642376.

Botsko:2007:DCS

[2246] Michael W. Botsko and Robert B. Israel. Disguised Cauchy sequence: 11207. American Mathematical Monthly, 114(10):929–930, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642377.

Dubinov:2007:CCF

[2247] Alexander Dubinov, Irina Dubinova, and Knut Dale. Can “closed form” involve lambert? 11209. American Mathematical Monthly, 114(10):930– 931, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642378.

Bataille:2007:FGL

[2248] Michel Bataille and Richard Stong. Finding the greatest lower bound: 11217. American Mathematical Monthly, 114(10):931, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642379.

Anonymous:2007:E

[2249] Anonymous. Errata: 11307. American Mathematical Monthly, 114(10): 932, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642382. See

[2155].

Holing:2007:UNL

[2250] Kent Holing. Unsolved no longer: 11154. American Mathematical Monthly, 114(10):932, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642381.

Nica:2007:RE

[2251] Bogdan Nica. Reverse engineering: 11124. American Mathematical Monthly, 114(10):932, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642380.

Henle:2007:BRB

[2252] Michael Henle. Book review: Divine Proportions: Rational Trigonometry to Universal Geometry, by N. J. Wildberger. American Mathematical Monthly, 114(10):933–937, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642383.

Anonymous:2007:BMb

[2253] Anonymous. Back matter. American Mathematical Monthly, 114(10):??, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642384.

Anonymous:2007:FMb

[2254] Anonymous. Front matter. American Mathematical Monthly, 114(10):??, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642358.

Anonymous:2007:VI

[2255] Anonymous. Volume information. American Mathematical Monthly, 114(10):??, December 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642357.

Henle:2008:SP

[2256] Frederick V. Henle and James M. Henle. Squaring the plane. American Mathematical Monthly, 115(1):3–12, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642387.

Previte:2008:NWG

[2257] Michelle Previte and Sean Yang. A novel way to generate fractals. American Mathematical Monthly, 115(1):13–32, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642388.

Komjath:2008:U

[2258] Péter Komjáth and Vilmos Totik. Ultrafilters. American Mathematical Monthly, 115(1):33–44, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642389.

Cavalcante:2008:LTD

[2259] Ray Cavalcante and Todor D. Todorov. A lost theorem: Definite integrals in an asymptotic setting. American Mathematical Monthly, 115(1):45–56, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642390.

Martin:2008:SPS

[2260] Greg Martin. A simple polynomial for a simple transposition. American Mathematical Monthly, 115(1):57–60, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642391.

Sroysang:2008:CA

[2261] Banyat Sroysang, John E. Wetzel, and Wacharin Wichiramala. Covers for angleworms. American Mathematical Monthly, 115(1):61–65, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642392.

Ucsnay:2008:NTN

[2262] Stefanie Ucsnay. A note on Tarski’s note. American Mathematical Monthly, 115(1):66–68, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642393.

Marshall:2008:ASP

[2263] Simon Marshall. Another simple proof of the high girth, high chromatic number theorem. American Mathematical Monthly, 115(1):68–70, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642394.

Bober:2008:P

[2264] Jonathan Bober and Jeffrey Lagarias. Problem 11334. American Mathematical Monthly, 115(1):71, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642395.

Gonzalez:2008:P

[2265] Juan López González. Problem 11335. American Mathematical Monthly, 115(1):71, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642396.

Knuth:2008:Pa

[2266] Donald Knuth. Problem 11336. American Mathematical Monthly, 115 (1):71, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642397.

Tetiva:2008:Pa

[2267] Marian Tetiva. Problem 11337. American Mathematical Monthly, 115(1): 71, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642398.

Furdui:2008:Pa

[2268] Ovidiu Furdui. Problem 11338. American Mathematical Monthly, 115(1): 71, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642399.

Diaz-Barrero:2008:Pa

[2269] José Luis Díaz-Barrero. Problem 11339. American Mathematical Monthly, 115(1):72, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642400.

Ciaurri:2008:P

[2270] Óscar Ciaurri and Luz Roncal. Problem 11340. American Mathematical Monthly, 115(1):72, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642401.

Mehrabi:2008:CC

[2271] Mohammad Hossein Mehrabi and Pál Péter Dályay. A condition on the commutator: 11196. American Mathematical Monthly, 115(1):72, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642402.

Russell:2008:COR

[2272] Robert Russell, Christoph Soland, William Dickinson, and William Dickenson. Can one recover a (spherical) triangle from its medial triangle? 11201. American Mathematical Monthly, 115(1):73–74, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642403.

Perfetti:2008:DDD

[2273] Paolo Perfetti and Richard Bagby. Differentiable and discontinuous densely: 11221. American Mathematical Monthly, 115(1):74–75, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642404.

Hillar:2008:SV

[2274] Christopher Hillar, Lionel Levine, and S. C. Locke. Snapshots for velocities: 11223. American Mathematical Monthly, 115(1):75–76, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642405.

Beaucoup:2008:US

[2275] Franck Beaucoup, Tamás Erdélyi, and Microsoft Research Problems Group. A unique solution: 11226. American Mathematical Monthly, 115 (1):76, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642406.

Botsko:2008:FI

[2276] Michael W. Botsko and Albert Stadler. A function inequality: 11232. American Mathematical Monthly, 115(1):77, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642408.

Tetiva:2008:TI

[2277] Marian Tetiva and Minh Can. A triangle inequality: 11228. American Mathematical Monthly, 115(1):77, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/27642407.

Downes:2008:DF

[2278] Robert Downes and Jean-Pierre Grivaux. A derivative formula: 11233. American Mathematical Monthly, 115(1):78, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642409.

Nunemacher:2008:BRB

[2279] Jeffrey Nunemacher. Book review: Fearless Symmetry: Exposing the Hidden Patterns of Numbers, by Avner Ash; Robert Gross. American Mathematical Monthly, 115(1):79–82, January 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642410.

Solomon:2008:BRB

[2280] Ron Solomon. Book review: Symmetry and the Monster, by Mark Ronan. American Mathematical Monthly, 115(1):82–85, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642411.

Anonymous:2008:BMa

[2281] Anonymous. Back matter. American Mathematical Monthly, 115(1):??, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642412.

Anonymous:2008:FMa

[2282] Anonymous. Front matter. American Mathematical Monthly, 115(1):??, January 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642386.

Hardin:2008:PCB

[2283] Christopher S. Hardin and Alan D. Taylor. A peculiar connection between the axiom of choice and predicting the future. American Mathematical Monthly, 115(2):91–96, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642415.

Balinski:2008:FMV

[2284] Michel Balinski. Fair majority voting (or how to eliminate gerrymandering). American Mathematical Monthly, 115(2):97–113, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642416.

Howards:2008:BS

[2285] Hugh Nelson Howards. Brunnian spheres. American Mathematical Monthly, 115(2):114–124, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642417.

Borwein:2008:HIW

[2286] Jonathan M. Borwein. Hilbert’s inequality and Witten’s zeta-function. American Mathematical Monthly, 115(2):125–137, February 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642418.

Ungar:2008:MG

[2287] Abraham A. Ungar. From Möbius to gyrogroups. American Mathematical Monthly, 115(2):138–144, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642419.

Melman:2008:BZD

[2288] Aaron Melman. Bounds on the zeros of the derivative of a polynomial with all real zeros. American Mathematical Monthly, 115(2):145–147, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642420.

Elias:2008:QAD

[2289] Uri Elias. Qualitative analysis of a differential equation of Abel. American Mathematical Monthly, 115(2):147–149, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642421.

Holsztynski:2008:CCS

[2290] W. Holsztyński, J. Mycielski, G. C. O’Brien, and S. Świerczkowski. Closed curves on spheres. American Mathematical Monthly, 115(2):149– 153, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642422.

Cornelius:2008:SGP

[2291] E. F. Cornelius, Jr. and Phill Schultz. Sequences generated by polynomials. American Mathematical Monthly, 115(2):154–158, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642423.

Ciesielski:2008:ONN

[2292] Krzysztof Ciesielski and Zdzisław Pogoda. On ordering the natural numbers, or, the sharkovski theorem. American Mathematical Monthly, 115(2):159–165, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642424.

Lupu:2008:Pa

[2293] Cezar Lupu. Problem 11341. American Mathematical Monthly, 115(2): 166, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642425.

Gallardo:2008:P

[2294] Luis H. Gallardo. Problem 11342. American Mathematical Monthly, 115(2):166, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642426.

Beckwith:2008:P

[2295] David Beckwith. Problem 11343. American Mathematical Monthly, 115(2):166, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642427.

Stadler:2008:P

[2296] Albert Stadler. Problem 11344. American Mathematical Monthly, 115(2): 166, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642428.

Cuculiere:2008:P

[2297] Roger Cuculière. Problem 11345. American Mathematical Monthly, 115(2):166, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642429.

Hillar:2008:Pa

[2298] Christopher Hillar. Problem 11346. American Mathematical Monthly, 115(2):166–167, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642430.

Bencze:2008:P

[2299] Mihály Bencze. Problem 11347. American Mathematical Monthly, 115(2):167, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642431.

Golomb:2008:OSS

[2300] Solomon W. Golomb and Kyle Calderhead. Orthogonal sudoku squares: 11214. American Mathematical Monthly, 115(2):167–168, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642432.

Rosset:2008:MVC

[2301] Shmuel Rosset and José Heber Nieto. Maximum velocity on a car trip: 11215. American Mathematical Monthly, 115(2):168–169, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642433.

Brennan:2008:GBB

[2302] Jim Brennan, Richard Ehrenborg, and Eugene A. Herman. Goes back to boole: 11234. American Mathematical Monthly, 115(2):169–170, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642434.

Owings:2008:CP

[2303] Jim Owings, Marian Tetiva, and Matthew Huddleston. Coloring the plane: 11236. American Mathematical Monthly, 115(2):170–172, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642435.

Dalyay:2008:ERR

[2304] Pál Péter Dályay and A. K. Shafie. Ex-to-in-radius ratio: 11240. American Mathematical Monthly, 115(2):172, February 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642436.

Harris:2008:IRC

[2305] Jolene Harris, Bogdan Suceavǎ, and Knut Dale. Iterational rate of covergence: 11244. American Mathematical Monthly, 115(2):173, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642437.

Cohen:2008:BRB

[2306] Marion Cohen. Book review: Letters to a Young Mathematician, by Ian Stewart. American Mathematical Monthly, 115(2):174–177, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642438.

Anonymous:2008:BMc

[2307] Anonymous. Back matter. American Mathematical Monthly, 115(2):??, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642439.

Anonymous:2008:FMc

[2308] Anonymous. Front matter. American Mathematical Monthly, 115(2):??, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642414.

Cohen-Corwin:2008:YGG

[2309] Amy Cohen-Corwin. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2008 to Lida K. Barrett for Distinguished Service to Mathematics. American Mathematical Monthly, 115(3):183–184, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642442.

Chung:2008:PJS

[2310] Fan Chung and Ron Graham. Primitive juggling sequences. American Mathematical Monthly, 115(3):185–194, March 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642443.

Akiyama:2008:UMB

[2311] Jin Akiyama, Hiroshi Fukuda, Chie Nara, Toshinori Sakai, and Jorge Urrutia. Universal measuring boxes with triangular bases. American Mathematical Monthly, 115(3):195–201, March 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642444.

Benardete:2008:QTS

[2312] Diego M. Benardete, V. W. Noonburg, and B. Pollina. Qualitative tools for studying periodic solutions and bifurcations as applied to the periodically harvested logistic equation. American Mathematical Monthly, 115(3):202–219, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642445.

Jacobi:2008:X

[2313] Lee W. Jacobi and Daniel J. Madden. On a^{4}+b^{4}+c^{4}+d^{4} = (a+b+c+d)^{4}. American Mathematical Monthly, 115(3):220–236, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642446.

Agoh:2008:RRB

[2314] Takashi Agoh and Karl Dilcher. Reciprocity relations for Bernoulli numbers. American Mathematical Monthly, 115(3):237–244, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642447.

Penev:2008:FP

[2315] Krassimir Penev. The Fubini principle. American Mathematical Monthly, 115(3):245–248, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642448.

Prus-Wisniowski:2008:RRT

[2316] Franciszek Prus-Wiśniowski. A refinement of Raabe’s test. American Mathematical Monthly, 115(3):249–252, March 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642449.

Alaca:2008:APB

[2317] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. Arithmetic progressions and binary quadratic forms. American Mathematical Monthly, 115(3):252–254, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642450.

Saltzman:2008:SCI

[2318] Peter W. Saltzman and Pingzhi Yuan. On sums of consecutive integral roots. American Mathematical Monthly, 115(3):254–261, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642451.

Stanley:2008:P

[2319] Richard P. Stanley. Problem 11348. American Mathematical Monthly, 115(3):262, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642452.

Lupu:2008:Pb

[2320] Cezar Lupu. Problem 11349. American Mathematical Monthly, 115(3): 262, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642453.

Deshpande:2008:P

[2321] Bhavana Deshpande and M. N. Deshpande. Problem 11350. American Mathematical Monthly, 115(3):262, March 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.jstor.org/stable/27642454.

Tetiva:2008:Pb

[2322] Marian Tetiva. Problem 11351. American Mathematical Monthly, 115 (3):262, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642455.

Reem:2008:P

[2323] Daniel Reem. Problem 11352. American Mathematical Monthly, 115(3): 263, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642456.

Schulte-Geers:2008:P

[2324] Ernst Schulte-Geers. Problem 11353. American Mathematical Monthly, 115(3):263, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642457.

Beck:2008:P

[2325] Matthias Beck and Alexander Berkovich. Problem 11354. American Mathematical Monthly, 115(3):263, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642458.

Gerber:2008:U

[2326] Hans V. Gerber and Daniel Neuenschwander. Unsolved in 1990: 6576. American Mathematical Monthly, 115(3):263–264, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642459.

Herzog:2008:DI

[2327] Gerd Herzog and Roland Lemmert. A determinant indentity: 11242. American Mathematical Monthly, 115(3):264–265, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642460.

Goldstein:2008:BFI

[2328] Lee Goldstein. A Bessel function identity: 11246. American Mathematical Monthly, 115(3):265–266, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642461.

Eckhoff:2008:SP

[2329] Jürgen Eckhoff and Li Zhou. A strip problem: 11247. American Mathematical Monthly, 115(3):266, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642462.

Cai:2008:DSI

[2330] Sun Wen Cai and Vitaly Stakhovsky. Double sum inequality: 11250. American Mathematical Monthly, 115(3):266–267, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642463.

Tetiva:2008:IPC

[2331] Marian Tetiva and Vitaly Stakhovsky. An inequality proved without computer assistance: 11251. American Mathematical Monthly, 115(3):267– 268, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642464.

Bagdasar:2008:PI

[2332] Ovidiu Bagdasar and Marian Tetiva. A productive inequality: 11252. American Mathematical Monthly, 115(3):268–269, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642465.

Struble:2008:MAI

[2333] Raimond Struble and Jean-Pierre Grivaux. A myth about infinite products: 11257. American Mathematical Monthly, 115(3):269, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642466.

Davis:2008:BRB

[2334] Martin Davis. Book review: Inexhaustibility: A Non-Exhaustive Treatment, by Torkel Franzén; Gödel’s Theorem: An Incomplete Guide to Its Use and Misuse by Torkel Franzén. American Mathematical Monthly, 115(3):270–275, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642467.

Anonymous:2008:EE

[2335] Anonymous. Editor’s endnotes. American Mathematical Monthly, 115 (3):276, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642468.

Anonymous:2008:BMd

[2336] Anonymous. Back matter. American Mathematical Monthly, 115(3):??, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642469.

Anonymous:2008:FMd

[2337] Anonymous. Front matter. American Mathematical Monthly, 115(3):??, March 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642441.

Granville:2008:PNP

[2338] Andrew Granville. Prime number patterns. American Mathematical Monthly, 115(4):279–296, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642472.

Agnew:2008:OCA

[2339] Robert A. Agnew. Optimal Congressional apportionment. American Mathematical Monthly, 115(4):297–303, April 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642473.

Krantz:2008:CKM

[2340] Steven G. Krantz. The Carathéodory and Kobayashi metrics and applications in complex analysis. American Mathematical Monthly, 115(4): 304–329, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642474.

Kalman:2008:EPM

[2341] Dan Kalman. An elementary proof of Marden’s theorem. American Mathematical Monthly, 115(4):330–338, April 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642475.

Bessenyei:2008:HHI

[2342] Mihály Bessenyei. The Hermite–Hadamard inequality on simplices. American Mathematical Monthly, 115(4):339–345, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642476.

Robertson:2008:CFA

[2343] John P. Robertson and Keith R. Matthews. A continued fractions approach to a result of feit. American Mathematical Monthly, 115(4):346– 349, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642477.

Larcher:2008:SDB

[2344] Gerhard Larcher and Friedrich Pillichshammer. The sum of distances between vertices of a convex polygon with unit perimeter. American Mathematical Monthly, 115(4):350–355, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642478.

McFarlane:2008:DSI

[2345] Cayanne McFarlane and Wm. Douglas Withers. Dynamical systems and irrational angle construction by paper-folding. American Mathematical Monthly, 115(4):355–358, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642479.

Renault:2008:LFT

[2346] Marc Renault. Lost (and found) in translation: André’s actual method and its application to the Generalized Ballot Problem. American Mathematical Monthly, 115(4):358–363, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642480.

Tripathi:2008:SPI

[2347] Amitabha Tripathi. On sums of positive integers that are not of the form ax + by. American Mathematical Monthly, 115(4):363–364, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642481.

Lagarias:2008:Pa

[2348] Jeffrey C. Lagarias. Problem 11355. American Mathematical Monthly, 115(4):365, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642482.

Poghosyan:2008:P

[2349] Michael Poghosyan. Problem 11356. American Mathematical Monthly, 115(4):365, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642483.

Sahin:2008:P

[2350] Mehmet Şahin. Problem 11357. American Mathematical Monthly, 115 (4):365, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642484.

Tetiva:2008:Pc

[2351] Marian Tetiva. Problem 11358. American Mathematical Monthly, 115(4): 365, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642485.

Monterde:2008:P

[2352] J. Monterde. Problem 11359. American Mathematical Monthly, 115(4): 365, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642486.

Lupu:2008:Pc

[2353] Cezar Lupu. Problem 11360. American Mathematical Monthly, 115(4): 365–366, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642487.

Beckwith:2008:VAS

[2354] David Beckwith, Harris Kwong, Rob Pratt, and Nicholas Singer. A vanishing alternating sum: 11212/11220. American Mathematical Monthly, 115(4):366, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642489.

Holland:2008:P

[2355] Finbarr Holland. Problem 11361. American Mathematical Monthly, 115 (4):366, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642488.

Rabinowitz:2008:SI

[2356] Stanley Rabinowitz and NSA Problems Group. Sometimes an integer: 11213. American Mathematical Monthly, 115(4):366–367, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642490.

Gordon:2008:NBL

[2357] Gary Gordon and Bruce S. Burdick. The number between 1 and n that is least prime: 11218. American Mathematical Monthly, 115(4):367–368, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642491.

Dalyay:2008:ROP

[2358] Pál Péter Dályay and Jaime Vinuesa. A Romanian Olympiad problem generalized: 11248. American Mathematical Monthly, 115(4):368–369, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642492.

Simic:2008:SI

[2359] Slavko Simic and O. P. Lossers. A symmetric inequality: 11255. American Mathematical Monthly, 115(4):369–370, April 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642493.

Holland:2008:PI

[2360] Finbarr Holland and Michel Bataille. Perimeter by integral: 11256. American Mathematical Monthly, 115(4):370–371, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642494.

Keselman:2008:CS

[2361] Gregory Keselman and E. Leonard. A choice sum: 11263. American Mathematical Monthly, 115(4):371–372, April 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642495.

Hogben:2008:BRB

[2362] Leslie Hogben and Wolfgang Kliemann. Book review: Applied Linear Algebra, by Peter J. Olver; Chehrzad Shakiban. American Mathematical Monthly, 115(4):373–378, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642496.

Anonymous:2008:BMe

[2363] Anonymous. Back matter. American Mathematical Monthly, 115(4): ??, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642497.

Anonymous:2008:FMe

[2364] Anonymous. Front matter. American Mathematical Monthly, 115(4): ??, April 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642471.

Bailey:2008:SGD

[2365] R. A. Bailey, Peter J. Cameron, and Robert Connelly. Sudoku, gerechte designs, resolutions, affine space, spreads, reguli, and Hamming codes. American Mathematical Monthly, 115(5):383–404, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642500.

Arroyo:2008:WPF

[2366] J. Arroyo, O. J. Garay, and J. J. Mencía. When is a periodic function the curvature of a closed plane curve? American Mathematical Monthly, 115(5):405–414, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642501.

Nguyen:2008:RUH

[2367] Hieu D. Nguyen. Rearing its ugly head: The cosmological constant and Newton’s greatest blunder. American Mathematical Monthly, 115(5): 415–430, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642502.

Chahal:2008:RHE

[2368] Jasbir S. Chahal and Brian Osserman. The Riemann hypothesis for elliptic curves. American Mathematical Monthly, 115(5):431–442, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642503.

Pak:2008:ICS

[2369] Igor Pak. Inflating the cube without stretching. American Mathematical Monthly, 115(5):443–445, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642504.

Kantrowitz:2008:YAP

[2370] Robert Kantrowitz and Michael M. Neumann. Yet another proof of Minkowski’s inequality. American Mathematical Monthly, 115(5):445– 447, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642505.

Ji:2008:EP

[2371] Kathy Q. Ji and Herbert S. Wilf. Extreme palindromes. American Mathematical Monthly, 115(5):447–451, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642506.

Mendivil:2008:TSF

[2372] F. Mendivil and T. D. Taylor. Thin sets with fat shadows: Projections of Cantor sets. American Mathematical Monthly, 115(5):451–456, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642507.

Eifler:2008:DPT

[2373] Larry Eifler and Noah H. Rhee. The n-dimensional Pythagorean theorem via the divergence theorem. American Mathematical Monthly, 115(5): 456–457, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642508.

Rzadkowski:2008:DEN

[2374] Grzegorz Rza̧dkowski. Derivatives and Eulerian numbers. American Mathematical Monthly, 115(5):458–460, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642509.

Callan:2008:P

[2375] David Callan. Problem 11362. American Mathematical Monthly, 115(5): 461, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642510.

Faynshteyn:2008:P

[2376] Oleh Faynshteyn. Problem 11363. American Mathematical Monthly, 115 (5):461, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642511.

Dalyay:2008:P

[2377] Pál Péter Dályay. Problem 11364. American Mathematical Monthly, 115(5):461–462, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642512.

Fraenkel:2008:P

[2378] Aviezri S. Fraenkel. Problem 11365. American Mathematical Monthly, 115(5):462, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642513.

Anghel:2008:P

[2379] Nicolae Anghel. Problem 11366. American Mathematical Monthly, 115 (5):462, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642514.

Cusumano:2008:P

[2380] Andrew Cusumano. Problem 11367. American Mathematical Monthly, 115(5):462, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642515.

Jiang:2008:P

[2381] Wei-Dong Jiang. Problem 11368. American Mathematical Monthly, 115 (5):462, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642516.

Parthasarathy:2008:GFH

[2382] P. R. Parthasarathy. Generating functions and hypergeometric series and continued fractions! 11198. American Mathematical Monthly, 115(5): 462–464, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642517.

Chinburg:2008:GRF

[2383] Ted Chinburg, Shahriar Shahriari, and Reid Huntsinger. Groups, rings, fields, and power series: 11216. American Mathematical Monthly, 115(5): 464–465, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642518.

Sondow:2008:IPB

[2384] Jonathan Sondow and GCHQ Problem Solving Group Cheltenham. An infinite product based on a base: 11222. American Mathematical Monthly, 115(5):465–466, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642519.

Trenkler:2008:SRC

[2385] Dietrich Trenkler, Götz Trenkler, and Patrick Corn. The square root of a certain matrix: 11224. American Mathematical Monthly, 115(5):466– 467, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642520.

Lupu:2008:NBR

[2386] Cezar Lupu, Tudorel Lupu, and Ronald A. Kopas. A new bound for r/R: 11245. American Mathematical Monthly, 115(5):467–468, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642521.

Perfetti:2008:ES

[2387] Paolo Perfetti and Richard Stong. Evaluate the series: 11260. American Mathematical Monthly, 115(5):468, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642522.

Roy:2008:BRB

[2388] Ranjan Roy. Book review: Euler Through Time: A New Look at Old Themes, by V. S. Varadarajan. American Mathematical Monthly, 115(5): 469–473, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642523.

Kidwell:2008:BRB

[2389] Mark E. Kidwell. Book review: King of Infinite Space: Donald Coxeter, the Man Who Saved Geometry, by Siobhan Roberts. American Mathematical Monthly, 115(5):474–476, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642524.

Anonymous:2008:BMf

[2390] Anonymous. Back matter. American Mathematical Monthly, 115(5): ??, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642525.

Anonymous:2008:FMf

[2391] Anonymous. Front matter. American Mathematical Monthly, 115(5): ??, May 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642499.

Baxter:2008:POB

[2392] Andrew M. Baxter and Ronald Umble. Periodic orbits for billiards on an equilateral triangle. American Mathematical Monthly, 115(6):479–491, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642528.

Eriksen:2008:FAC

[2393] Niklas Eriksen. The Freshman’s approach to Conway’s napkin problem. American Mathematical Monthly, 115(6):492–498, June/July 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642529.

Engelberg:2008:EDU

[2394] Shlomo Engelberg. Edge detection using Fourier coefficients. American Mathematical Monthly, 115(6):499–513, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642530.

Ali:2008:MRT

[2395] Sayel A. Ali. The mth ratio test: New convergence tests for series. American Mathematical Monthly, 115(6):514–524, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642531.

Schmelzer:2008:SCS

[2396] Thomas Schmelzer and Robert Baillie. Summing a curious, slowly convergent series. American Mathematical Monthly, 115(6):525–540, June/ July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642532.

Birmajer:2008:ARF

[2397] Daniel Birmajer and Juan B. Gil. Arithmetic in the ring of formal power series with integer coefficients. American Mathematical Monthly, 115(6):541–549, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642533.

Castryck:2008:SCP

[2398] Wouter Castryck. A shortened classical proof of the quadratic reciprocity law. American Mathematical Monthly, 115(6):550–551, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642534.

Bhatia:2008:IDG

[2399] Rajendra Bhatia and J. A. Dias da Silva. Infinite divisibility of GCD matrices. American Mathematical Monthly, 115(6):551–553, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642535.

Volovikov:2008:BUI

[2400] Alexey Yu. Volovikov. Borsuk–Ulam implies Brouwer: a direct construction revisited. American Mathematical Monthly, 115(6):553–556, June/ July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642536.

Koumoullis:2008:RNT

[2401] George Koumoullis. On the Radon–Nikodym theorem. American Mathematical Monthly, 115(6):556–558, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642537.

Kantor:2008:CGG

[2402] William M. Kantor. Cycles in graphs and groups. American Mathematical Monthly, 115(6):559–562, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642538.

Hirschhorn:2008:LFH

[2403] Michael D. Hirschhorn. A letter from fitzroy house. American Mathematical Monthly, 115(6):563–566, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642539.

Knuth:2008:Pb

[2404] Donald Knuth. Problem 11369. American Mathematical Monthly, 115(6): 567, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642540.

Goldenberg:2008:P

[2405] Michael Goldenberg and Mark Kaplan. Problem 11370. American Mathematical Monthly, 115(6):567, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642541.

Furdui:2008:Pb

[2406] Ovidiu Furdui. Problem 11371. American Mathematical Monthly, 115 (6):567–568, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642542.

Vandenbussche:2008:P

[2407] Jennifer Vandenbussche and Douglas B. West. Problem 11372. American Mathematical Monthly, 115(6):568, June/July 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642543.

Deutsch:2008:P

[2408] Emeric Deutsch. Problem 11373. American Mathematical Monthly, 115(6):568, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642544.

Flanders:2008:P

[2409] Harley Flanders and Hugh L. Montgomery. Problem 11374. American Mathematical Monthly, 115(6):568, June/July 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642545.

Lupu:2008:Pd

[2410] Cezar Lupu. Problem 11375. American Mathematical Monthly, 115(6): 568, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642546.

Marivani:2008:CSS

[2411] Syrous Marivani and Missouri State University Problem Solving Group. Consecutive squares with a square average: 11227. American Mathematical Monthly, 115(6):568–569, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642547.

Sofair:2008:ITA

[2412] Isaac Sofair and C. R. Pranesachar. Inscribed triangle area: 11261. American Mathematical Monthly, 115(6):569–571, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642548.

Moulin:2008:CI

[2413] Hervé Moulin. A combinatorial inequality: 11269. American Mathematical Monthly, 115(6):571–573, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642549.

Mihai:2008:ELA

[2414] Vasile Mihai and Sin Hitotumatu. The Euler line after all: 11272. American Mathematical Monthly, 115(6):573–574, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642550.

Vaaler:2008:BRB

[2415] Jeffrey D. Vaaler. Book review: An Invitation to Modern Number Theory, by Steven J. Miller; Bamin Takloo-Bighash. American Mathematical Monthly, 115(6):575–578, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642551.

Anonymous:2008:BMg

[2416] Anonymous. Back matter. American Mathematical Monthly, 115(6):??, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642552.

Anonymous:2008:FMg

[2417] Anonymous. Front matter. American Mathematical Monthly, 115(6):??, June/July 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642527.

Hansen:2008:SVL

[2418] Wolfhard Hansen. A strong version of Liouville’s theorem. American Mathematical Monthly, 115(7):583–595, August/September 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642555.

McSwiggen:2008:CPP

[2419] Patrick D. McSwiggen and Kenneth R. Meyer. Conjugate phase portraits of linear systems. American Mathematical Monthly, 115(7):596– 614, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642556.

Hicks:2008:PAP

[2420] Kenneth Hicks, Gary L. Mullen, Joseph L. Yucas, and Ryan Zavislak. A polynomial analogue of the 3n + 1 problem. American Mathematical Monthly, 115(7):615–622, August/September 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642557.

Khosravi:2008:WPA

[2421] Mehrdad Khosravi and Michael D. Taylor. The wedge product and analytic geometry. American Mathematical Monthly, 115(7):623– 644, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642558.

Wagon:2008:VSP

[2422] Stan Wagon and Peter Webb. Venn symmetry and prime numbers: a seductive proof revisited. American Mathematical Monthly, 115(7):645– 648, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642559.

Hanche-Olsen:2008:GPC

[2423] Harald Hanche-Olsen. On Goursat’s proof of Cauchy’s integral theorem. American Mathematical Monthly, 115(7):648–652, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642560.

Bodin:2008:NIP

[2424] Arnaud Bodin. Number of irreducible polynomials in several variables over finite fields. American Mathematical Monthly, 115(7):653– 660, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642561.

Schroer:2008:BRI

[2425] Stefan Schröer. Baer’s result: The infinite product of the integers has no basis. American Mathematical Monthly, 115(7):660–663, August/ September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642562.

Baishanski:2008:P

[2426] Bogdan M. Baishanski. Problem 11376. American Mathematical Monthly, 115(7):664, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642563.

Hillar:2008:Pb

[2427] Christopher Hillar and Lionel Levine. Problem 11377. American Mathematical Monthly, 115(7):664, August/September 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642564.

Troy:2008:P

[2428] Daniel Troy. Problem 11378. American Mathematical Monthly, 115(7): 664, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642565.

Baksalary:2008:Pa

[2429] Oskar Maria Baksalary and Götz Trenkler. Problem 11379. American Mathematical Monthly, 115(7):664, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642566.

Montgomery:2008:P

[2430] Hugh Montgomery and Harold S. Shapiro. Problem 11380. American Mathematical Monthly, 115(7):665, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642567.

Guillera:2008:P

[2431] Jesús Guillera and Jonathan Sondow. Problem 11381. American Mathematical Monthly, 115(7):665, August/September 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642568.

Tauraso:2008:P

[2432] Roberto Tauraso. Problem 11382. American Mathematical Monthly, 115 (7):665, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642569.

Hillar:2008:GWE

[2433] Christopher J. Hillar and O. P. Lossers. A group in which equations are solvable: 11231. American Mathematical Monthly, 115(7): 665–666, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642570.

Jones:2008:SP

[2434] Lenny Jones, Rachel Keller, Toni Ernvall, and Anne-Maria Ernvall-Hytönen. A special prime: 11235. American Mathematical Monthly, 115 (7):666, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642571.

Deutsch:2008:EPP

[2435] Emeric Deutsch, John H. Smith, and Albert Stadler. An equality for parts in partitions of n: 11237. American Mathematical Monthly, 115(7): 666–667, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642572.

Fraenkel:2008:SBG

[2436] Aviezri S. Fraenkel. Sequences built from the golden ratio: 11238. American Mathematical Monthly, 115(7):667–668, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642573.

Knuth:2008:PPP

[2437] Donald Knuth, Robin Chapman, and Reiner Martin. Perfect parity patterns: 11243. American Mathematical Monthly, 115(7):668–670, August/ September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642574.

Becker:2008:DIL

[2438] Michael S. Becker, David Beckwith, and Hongwei Chen. Double integral with log and sinh: 11275. American Mathematical Monthly, 115(7):670– 671, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642575.

Amghibech:2008:MI

[2439] Said Amghibech, Joanna Chachulska, and Wojciech Matysiak. A matrix inequality: 11296. American Mathematical Monthly, 115(7):671, August/ September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642576.

Saari:2008:BRB

[2440] Donald G. Saari. Book review: The Three Body Problem, by Mauri Valtonen; Hannu Karrunen. American Mathematical Monthly, 115(7): 672–676, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642577.

Anonymous:2008:BMh

[2441] Anonymous. Back matter. American Mathematical Monthly, 115(7): ??, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642578.

Anonymous:2008:FMh

[2442] Anonymous. Front matter. American Mathematical Monthly, 115(7): ??, August/September 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642554.

Minda:2008:TEC

[2443] D. Minda and S. Phelps. Triangles, ellipses, and cubic polynomials. American Mathematical Monthly, 115(8):679–689, October 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642581.

Langevin:2008:HMP

[2444] Rémi Langevin and Pewel G. Walczak. Holomorphic maps and pencils of circles. American Mathematical Monthly, 115(8):690–700, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642582.

Bashelor:2008:EAG

[2445] Andrew Bashelor, Amy Ksir, and Will Traves. Enumerative algebraic geometry of conics. American Mathematical Monthly, 115(8):701–728, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642583.

Klosinski:2008:SEW

[2446] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Eighth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 115(8):729–739, October 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642584.

Miller:2008:PPW

[2447] Steven J. Miller. A probabilistic proof of Wallis’s formula for π. American Mathematical Monthly, 115(8):740–745, October 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642585.

Passare:2008:HCS

[2448] Mikael Passare. How to compute ∑ 1/n^{2} by solving triangles. American Mathematical Monthly, 115(8):745–752, October 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642586.

vanderHofstad:2008:EPH

[2449] Remco van der Hofstad and Michael Keane. An elementary proof of the hitting time theorem. American Mathematical Monthly, 115(8):753–756, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642587.

Nyblom:2008:P

[2450] Michael Nyblom. Problem 11383. American Mathematical Monthly, 115(8):757, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642588.

Omarjee:2008:P

[2451] Moubinool Omarjee. Problem 11384. American Mathematical Monthly, 115(8):757, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642589.

Diaz-Barrero:2008:Pb

[2452] José Luis Díaz-Barrero. Problem 11385. American Mathematical Monthly, 115(8):757, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642590.

Markowsky:2008:P

[2453] Greg Markowsky. Problem 11386. American Mathematical Monthly, 115(8):757, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642591.

Baksalary:2008:Pb

[2454] Oskar Maria Baksalary. Problem 11387. American Mathematical Monthly, 115(8):758, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642592.

Farrokhi:2008:Pa

[2455] M. Farrokhi. Problem 11388. American Mathematical Monthly, 115(8): 758, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642593.

Chen:2008:P

[2456] Elizabeth R. Chen and Jeffrey C. Lagarias. Problem 11389. American Mathematical Monthly, 115(8):758, October 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642594.

De:2008:RSS

[2457] Prithwijit De and E. Omey. Recalling 11159 and sixty-sixth Putnam A5: 11277. American Mathematical Monthly, 115(8):758–759, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642595.

Simic:2008:EL

[2458] Slavko Simic and John H. Lindsey II. Entire limit: 11278. American Mathematical Monthly, 115(8):759, October 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642596.

Stakhovsky:2008:SB

[2459] Vitaly Stakhovsky. Sliding beads: 11279. American Mathematical Monthly, 115(8):759–760, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642597.

Kwong:2008:WWM

[2460] Harris Kwong and GCHQ Problem Solving Group. What we mean, not what we say: 11280. American Mathematical Monthly, 115(8):760–761, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642598.

Aliyev:2008:TR

[2461] Yakub Aliyev, Apostolis Demis, Apostolis Denis, and The GCHQ Problem Solving Group. A triangular result: 11285. American Mathematical Monthly, 115(8):761–762, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642599.

Herschkorn:2008:CB

[2462] Stephen J. Herschkorn and BSI Problems Group. Continuous blackjack: 11287. American Mathematical Monthly, 115(8):762–764, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642600.

Faynshteyn:2008:FIA

[2463] Oleh Faynshteyn and Marian Tetival. Formulas involving the angles of a triangle: 11289. American Mathematical Monthly, 115(8):764, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642601.

Kaplan:2008:BRB

[2464] Daniel T. Kaplan. Book review: Google’s PageRank and Beyond: The Science of Search Engine Rankings, by Amy N. Langville; Carl D. Meyer. American Mathematical Monthly, 115(8):765–768, October 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642602.

Lagarias:2008:EE

[2465] Jeffrey Lagarias, Colin Mallows, Allan Wilks, and Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 115(8):769–770, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642603.

Anonymous:2008:BMi

[2466] Anonymous. Back matter. American Mathematical Monthly, 115(8):??, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642604.

Anonymous:2008:FMi

[2467] Anonymous. Front matter. American Mathematical Monthly, 115(8):??, October 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642580.

Krantz:2008:CAC

[2468] Steven G. Krantz. Complex analysis as catalyst. American Mathematical Monthly, 115(9):775–794, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642607.

Apostol:2008:NDC

[2469] Tom M. Apostol and Mamikon A. Mnatsakanian. New descriptions of conics via twisted cylinders, focal disks, and directors. American Mathematical Monthly, 115(9):795–812, November 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642608.

Hackborn:2008:PMR

[2470] William W. Hackborn. Projectile motion: Resistance is fertile. American Mathematical Monthly, 115(9):813–819, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642609.

Ionascu:2008:DDK

[2471] Eugen J. Ionascu, Dan Pritikin, and Stephen E. Wright. k-dependence and domination in Kings graphs. American Mathematical Monthly, 115 (9):820–836, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642610.

Scheinerman:2008:FSC

[2472] Ed Scheinerman and Christian Blatter. Four shots for a convex quadrilateral. American Mathematical Monthly, 115(9):837–843, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642611.

Michel:2008:TPS

[2473] Reinhard Michel. The (n + 1)th proof of Stirling’s formula. American Mathematical Monthly, 115(9):844–845, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642612.

Fuglede:2008:SWC

[2474] Bent Fuglede. A sharpening of Wielandt’s characterization of the gamma function. American Mathematical Monthly, 115(9):845–850, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642613.

Guillera:2008:EPS

[2475] Jesús Guillera. Easy proofs of some Borwein algorithms for π. American Mathematical Monthly, 115(9):850–854, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642614.

Lagarias:2008:Pb

[2476] Jeffrey C. Lagarias. Problem 11390. American Mathematical Monthly, 115(9):855, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642615.

Tetiva:2008:Pd

[2477] Marian Tetiva. Problem 11391. American Mathematical Monthly, 115(9): 855, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642616.

Kouba:2008:P

[2478] Omran Kouba. Problem 11392. American Mathematical Monthly, 115(9): 855, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642617.

Pohoata:2008:P

[2479] Cosmin Pohoata. Problem 11393. American Mathematical Monthly, 115(9):856, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642618.

Bhanu:2008:P

[2480] K. S. Bhanu. Problem 11394. American Mathematical Monthly, 115(9): 856, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642619.

Farrokhi:2008:Pb

[2481] M. Farrokhi. Problem 11395. American Mathematical Monthly, 115(9): 856, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642620.

Letac:2008:P

[2482] Gérard Letac. Problem 11396. American Mathematical Monthly, 115(9): 856, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642621.

Keselman:2008:G

[2483] Gregory Keselman, Aleksandar Ilić, and O. P. Lossers. Generalizing (1 − 1)^{n} = 0: 11230. American Mathematical Monthly, 115(9):856–857, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642622.

Bataille:2008:IPT

[2484] Michel Bataille and John W. Hagood. If AB preserves A, then BA preserves B: 11239. American Mathematical Monthly, 115(9):857–858, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642623.

Tauraso:2008:FNT

[2485] Roberto Tauraso and O. P. Lossers. Fibonacci numbers and tilling a board with cuts: 11241. American Mathematical Monthly, 115(9):858– 859, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642624.

Beckwith:2008:BPD

[2486] David Beckwith, Michael R. Avidon, and Li Zhou. “the ballot problem in disguise”, in disguise: 11249. American Mathematical Monthly, 115(9):859–860, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642625.

Beckwith:2008:PLB

[2487] David Beckwith and Michel Bataille. A permanent lower bound: 11253. American Mathematical Monthly, 115(9):860–861, November 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642626.

Jones:2008:PSP

[2488] Lenny Jones and Toni Ernvall. Primes with special primitive roots: 11254. American Mathematical Monthly, 115(9):861, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642627.

Abe:2008:NS

[2489] Nobuhisa Abe and Marian Tetiva. Never a square: 11259. American Mathematical Monthly, 115(9):862, November 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642628.

Burungale:2008:GKY

[2490] Ashay Burungale, Kenneth Schilling, and Christopher Carl Heckman. Getting to know you, directly and indirectly: 11262. American Mathematical Monthly, 115(9):862, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642629.

Duren:2008:BRB

[2491] Peter Duren. Book review: Princeton Lectures in Analysis, by Elias M. Stein; Rami Shakarchi. American Mathematical Monthly, 115(9):863– 866, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642630.

Anonymous:2008:BMj

[2492] Anonymous. Back matter. American Mathematical Monthly, 115(9):??, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642631.

Anonymous:2008:FMj

[2493] Anonymous. Front matter. American Mathematical Monthly, 115(9):??, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642606.

Carlson:2008:QFT

[2494] Stephan C. Carlson. The quest for a fuzzy Tychonoff theorem. American Mathematical Monthly, 115(10):871–887, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642635.

Baillie:2008:SSS

[2495] Robert Baillie, David Borwein, and Jonathan M. Borwein. Surprising sinc sums and integrals. American Mathematical Monthly, 115(10):888– 901, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642636.

Stadje:2008:MAG

[2496] Wolfgang Stadje. The maximum average gain in a sequence of Bernoulli games. American Mathematical Monthly, 115(10):902–910, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642637.

Bookman:2008:DAM

[2497] Jack Bookman, Susan L. Ganter, and Rick Morgan. Developing assessment methodologies for quantitative literacy: a formative study. American Mathematical Monthly, 115(10):911–929, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642638.

Scheinerman:2008:ACF

[2498] Ed Scheinerman, Thomas J. Pickett, and Ann Coleman. Another continued fraction for π. American Mathematical Monthly, 115(10):930–933, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642639.

Farhi:2008:CRR

[2499] Bakir Farhi. A curious result related to Kempner’s series. American Mathematical Monthly, 115(10):933–938, December 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642640.

Vernescu:2008:SFS

[2500] Andrei Vernescu. The summation of a family of series. American Mathematical Monthly, 115(10):939–943, December 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642641.

Pfeffer:2008:DP

[2501] Washek F. Pfeffer. A Devil’s platform. American Mathematical Monthly, 115(10):943–947, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642642.

Bennett:2008:P

[2502] Grahame Bennett. Problem 11397. American Mathematical Monthly, 115(10):948, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642643.

Huang:2008:P

[2503] Stanley Huang. Problem 11398. American Mathematical Monthly, 115(10):948, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642644.

Ricceri:2008:P

[2504] Biaggi Ricceri. Problem 11399. American Mathematical Monthly, 115(10):948, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642645.

Bracken:2008:P

[2505] Paul Bracken. Problem 11400. American Mathematical Monthly, 115(10): 948, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642646.

Cavachi:2008:P

[2506] Marius Cavachi. Problem 11401. American Mathematical Monthly, 115(10):949, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642647.

Barboianu:2008:P

[2507] Catalin Barboianu. Problem 11402. American Mathematical Monthly, 115(10):949, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642648.

Yu:2008:P

[2508] Yaming Yu. Problem 11403. American Mathematical Monthly, 115(10): 949, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642649.

Kauers:2008:TFS

[2509] Manuel Kauers and Richard Stong. A telescoping Fibonacci sum: 11258. American Mathematical Monthly, 115(10):949–950, December 2008. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642650.

Herman:2008:DRT

[2510] Eugene Herman and Mark D. Meyerson. 3-D rotations and translations: 11276. American Mathematical Monthly, 115(10):950, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642651.

Glasser:2008:SS

[2511] M. L. Glasser and O. P. Lossers. A sinh a series: 11286. American Mathematical Monthly, 115(10):950–951, December 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642652.

Lupu:2008:VIV

[2512] Cezar Lupu and Kenneth F. Andersen. A variant intermediate value: 11290. American Mathematical Monthly, 115(10):951–952, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642653.

Siboni:2008:DI

[2513] Stefano Siboni and O. P. Lossers. Double integral: 11295. American Mathematical Monthly, 115(10):952–953, December 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642654.

Tetiva:2008:TM

[2514] Marian Tetiva and John H. Lindsey II. A tricky minimum: 11297. American Mathematical Monthly, 115(10):953–954, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642655.

Anonymous:2008:EEN

[2515] Anonymous. Errata and end notes for 2008. American Mathematical Monthly, 115(10):954–955, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642656.

Kontoyiannis:2008:BRB

[2516] Ioannis Kontoyiannis. Book review: Information and Complexity in Statistical Modeling, by Jorma Rissanen. American Mathematical Monthly, 115(10):956–960, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642657.

Anonymous:2008:BMb

[2517] Anonymous. Back matter. American Mathematical Monthly, 115(10):??, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642658.

Anonymous:2008:FMb

[2518] Anonymous. Front matter. American Mathematical Monthly, 115(10):??, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642634.

Anonymous:2008:VI

[2519] Anonymous. Volume information. American Mathematical Monthly, 115(10):??, December 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642633.

Grabiner:2009:WDL

[2520] Judith V. Grabiner. Why did Lagrange “prove” the parallel postulate? American Mathematical Monthly, 116(1):3–18, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642661.

Paterson:2009:O

[2521] Mike Paterson and Uri Zwick. Overhang. American Mathematical Monthly, 116(1):19–44, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642662.

Goldstein:2009:PPL

[2522] Larry Goldstein. A probabilistic proof of the Lindeberg–Feller Central Limit Theorem. American Mathematical Monthly, 116(1):45–60, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/27642663.

Cid:2009:DLR

[2523] José Ángel Cid and Rodrigo López Pouso. Does Lipschitz with respect to x imply uniqueness for the differential equation y′ = f(x, y) ? American Mathematical Monthly, 116(1):61–66, January 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642664.

Scheinerman:2009:SCA

[2524] Ed Scheinerman and Anton R. Schep. A simple complex analysis and an advanced calculus proof of the fundamental theorem of algebra. American Mathematical Monthly, 116(1):67–68, January 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/27642665.

Ferreno:2009:YAP

[2525] Natalia Casás Ferreño. Yet another proof of the irrationality of \sqrt{2}. American Mathematical Monthly, 116(1):68–69, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642666.

Panti:2009:GLT

[2526] Giovanni Panti. A general Lagrange theorem. American Mathematical Monthly, 116(1):70–74, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642667.

Mestrovic:2009:NCT

[2527] Romeo Meštrović. A note on the congruence ( \begin{matrix}n & d \\ m & d \end{matrix} ) ≡ ( \begin{matrix}n \\ m \end{matrix} r)

(mod q). American Mathematical Monthly, 116(1):75–77, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642668.

Ross:2009:SSM

[2528] Sheldon M. Ross. A simple solution to a multiple player Gambler’s ruin problem. American Mathematical Monthly, 116(1):77–81, January 2009.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642669.

Struble:2009:P

[2529] Raimond Struble. Problem 11404. American Mathematical Monthly, 116 (1):82, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642670.

Furdui:2009:Pa

[2530] Ovidiu Furdui. Problem 11405. American Mathematical Monthly, 116(1): 82, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642671.

Dzhumadildaeva:2009:P

[2531] A. A. Dzhumadiĺdaeva. Problem 11406. American Mathematical Monthly, 116(1):82, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642672.

Just:2009:Pa

[2532] Erwin Just. Problem 11407. American Mathematical Monthly, 116(1):82, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642673.

Cavachi:2009:Pa

[2533] Marius Cavachi. Problem 11408. American Mathematical Monthly, 116 (1):83, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642674.

Perfetti:2009:Pa

[2534] Paolo Perfetti. Problem 11409. American Mathematical Monthly, 116(1): 83, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642675.

Kouba:2009:Pa

[2535] Omran Kouba. Problem 11410. American Mathematical Monthly, 116 (1):83, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642676.

Jerrard:2009:SRP

[2536] Richard Jerrard, John Wetzel, and Mark D. Meyerson. A solid with the rupert property: 11291. American Mathematical Monthly, 116(1):83–84, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642677.

Refolio:2009:IP

[2537] Pablo Fernàndez Refolio and Timothy Achenbach. An infinite product: 11299. American Mathematical Monthly, 116(1):84, January 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642678.

Abel:2009:BPI

[2538] Ulrich Abel. A Bernstein polynomial integral: 11300. American Mathematical Monthly, 116(1):84–85, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642679.

Holland:2009:QI

[2539] Finbarr Holland and Byoung Tae Bae. A quartic inequality:- 11301. American Mathematical Monthly, 116(1):85–86, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642680.

Alzer:2009:SHN

[2540] Horst Alzer. A series with harmonic numbers: 11302. American Mathematical Monthly, 116(1):86–87, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/27642681.

Radulescu:2009:SST

[2541] Teodora-Liliana Rǎdulescu. Signed series terms: 11304. American Mathematical Monthly, 116(1):87–88, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642682.

Rosoiu:2009:IT

[2542] Alexandru Rosoiu and Victor Pambuccian. An inequality for triangles: 11306. American Mathematical Monthly, 116(1):88–89, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642683.

Talvila:2009:BRB

[2543] Erik Talvila. Book review: A Garden of Integrals, by Frank E. Burk. American Mathematical Monthly, 116(1):90–94, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642684.

Anonymous:2009:BMa

[2544] Anonymous. Back matter. American Mathematical Monthly, 116(1):??, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642685.

Anonymous:2009:FMa

[2545] Anonymous. Front matter. American Mathematical Monthly, 116(1):??, January 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642660.

Bartlow:2009:WWM

[2546] Thomas L. Bartlow and David E. Zitarelli. Who was miss mullikin? American Mathematical Monthly, 116(2):99–114, February 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642688.

Apostol:2009:NLC

[2547] Tom M. Apostol and Mamikon A. Mnatsakanian. A new look at the so-called trammel of archimedes. American Mathematical Monthly, 116 (2):115–133, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642689.

Morgan:2009:MDP

[2548] Frank Morgan. Manifolds with density and Perelman’s proof of the Poincaré conjecture. American Mathematical Monthly, 116(2):134–142, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642690.

Bak:2009:SDA

[2549] Joseph Bak and Pisheng Ding. Shape distortion by analytic functions. American Mathematical Monthly, 116(2):143–150, February 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642691.

Haase:2009:LPN

[2550] Christian Haase and Josef Schicho. Lattice polygons and the number 2i + 7. American Mathematical Monthly, 116(2):151–165, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642692.

Lucas:2009:ADI

[2551] Stephen K. Lucas. Approximations to π derived from integrals with nonnegative integrands. American Mathematical Monthly, 116(2):166– 172, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642693.

Pinasco:2009:NPE

[2552] Juan Pablo Pinasco. New proofs of Euclid’s and Euler’s theorems. American Mathematical Monthly, 116(2):172–174, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642694.

Januszewski:2009:NCS

[2553] Janusz Januszewski. A note on covering a square of side length 2 + ε with unit squares. American Mathematical Monthly, 116(2):174–178, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642695.

Wyn-jones:2009:P

[2554] Alun Wyn-jones. Problem 11411. American Mathematical Monthly, 116(2):179, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642696.

Kouba:2009:Pb

[2555] Omran Kouba. Problem 11412. American Mathematical Monthly, 116(2): 179, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642697.

Bencze:2009:P

[2556] Mihály Bencze. Problem 11413. American Mathematical Monthly, 116(2):179, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642698.

Tetiva:2009:Pa

[2557] Marian Tetiva. Problem 11414. American Mathematical Monthly, 116 (2):179–180, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642699.

Holland:2009:Pa

[2558] Finbar Holland. Problem 11415. American Mathematical Monthly, 116(2):180, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642700.

Yu:2009:P

[2559] Yaming Yu. Problem 11416. American Mathematical Monthly, 116(2): 180, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642701.

Lupu:2009:Pa

[2560] Cezar Lupu. Problem 11417. American Mathematical Monthly, 116(2): 180, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642702.

Butler:2009:GFE

[2561] Steve Butler and Jia Mao. Graphs with few edges, all in short cycles: 11265. American Mathematical Monthly, 116(2):180–181, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/27642703.

Tetiva:2009:PVPa

[2562] Marian Tetiva and Robin Chapman. Partitioning values of a polynomial into sets of equal sum: 11266. American Mathematical Monthly, 116(2):181–183, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642704.

Herman:2009:CT

[2563] Eugene A. Herman and Richard Stong. Commuting transformations: 11267. American Mathematical Monthly, 116(2):183, February 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642705.

Furdui:2009:IT

[2564] Ovidiu Furdui and Vitaly Stakhovsky. An inequality with trig: 11308. American Mathematical Monthly, 116(2):183–184, February 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/27642706.

Witula:2009:TTI

[2565] Roman Wituła, Damian Słta, Arman Darbinyan, and David Borwein. A trig-of-trig inequality: 11309. American Mathematical Monthly, 116 (2):184–185, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

27642707.

Herschkorn:2009:GCB

[2566] Stephen J. Herschkorn. A game with coins and balls: 11311. American Mathematical Monthly, 116(2):185–186, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/27642708.

Varadarajan:2009:BRB

[2567] V. S. Varadarajan. Book review: Introduction to Calculus and Classical Analysis, by Omar Hijab. American Mathematical Monthly, 116(2):187– 190, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/27642709.

Anonymous:2009:BMc

[2568] Anonymous. Back matter. American Mathematical Monthly, 116(2):??, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642710.

Anonymous:2009:FMc

[2569] Anonymous. Front matter. American Mathematical Monthly, 116(2):??, February 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642687.

Faires:2009:YGG

[2570] Barbara Faires. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2009 to Robert E. Megginson for Distinguished Service to Mathematics. American Mathematical Monthly, 116(3):195–196, March 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391065.

Schoenfeld:2009:WSS

[2571] Alan H. Schoenfeld. Working with schools: The story of a mathematics education collaboration. American Mathematical Monthly, 116(3):197– 217, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391066.

Gordon:2009:SII

[2572] Russell A. Gordon. Some integrals involving the Cantor function. American Mathematical Monthly, 116(3):218–227, March 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391067.

Kocik:2009:DT

[2573] Jerzy Kocik and Andrzej Solecki. Disentangling a triangle. American Mathematical Monthly, 116(3):228–237, March 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391068.

Melman:2009:OPN

[2574] Aaron Melman. Overshooting properties of Newton-like and Ostrowskilike methods. American Mathematical Monthly, 116(3):238–250, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391069.

Troubetzkoy:2009:DBF

[2575] Serge Troubetzkoy. Dual billiards, fagnano orbits, and regular polygons. American Mathematical Monthly, 116(3):251–260, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391070.

Yap:2009:PLE

[2576] Shirley Llamado Yap. The Poincaré lemma and an elementary construction of vector potentials. American Mathematical Monthly, 116(3):261– 267, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391071.

Preissmann:2009:SCA

[2577] Emmanuel Preissmann and Maurice Mischler. Seating couples around the King’s table and a new characterization of prime numbers. American Mathematical Monthly, 116(3):268–272, March 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391072.

Akbary:2009:SNR

[2578] Amir Akbary and Zachary Friggstad. Superabundant numbers and the Riemann hypothesis. American Mathematical Monthly, 116(3):273–275, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391073.

Lamb:2009:P

[2579] George Lamb. Problem 11418. American Mathematical Monthly, 116(3): 276, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391074.

Mihai:2009:P

[2580] Vasile Mihai. Problem 11419. American Mathematical Monthly, 116(3): 276, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391075.

Holing:2009:P

[2581] Kent Holing. Problem 11420. American Mathematical Monthly, 116(3): 277, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391076.

Falcon:2009:P

[2582] Sergio Falcón and Ángel Plaza. Problem 11421. American Mathematical Monthly, 116(3):277, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391077.

Hillar:2009:Pa

[2583] Christopher Hillar. Problem 11422. American Mathematical Monthly, 116(3):277, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391078.

Minton:2009:P

[2584] Gregory Minton. Problem 11423. American Mathematical Monthly, 116 (3):277, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391079.

Deutsch:2009:P

[2585] Emeric Deutsch. Problem 11424. American Mathematical Monthly, 116 (3):277, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391080.

Knuth:2009:RS

[2586] Donald Knuth and Richard Stong. Reversal by swaps: 11264 [2007, 77]. American Mathematical Monthly, 116(3):277–278, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391081.

Lupu:2009:ID

[2587] Cezar Lupu and Li Zhou. From integral to derivative: 11313 [2007, 744]. American Mathematical Monthly, 116(3):279, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391082.

Aliyev:2009:FE

[2588] Yakub Aliyev and Marian Tetiva. A floor equation: 11314 [2007, 744]. American Mathematical Monthly, 116(3):279–281, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391083.

Furdui:2009:CII

[2589] Ovidiu Furdui and National Security Agency Problems Group. Convergence of an improper integral with three parameters: 11316 [2007, 745]. American Mathematical Monthly, 116(3):281–283, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391084.

Beckwith:2009:SD

[2590] David Beckwith and GCHQ Problem Solving Group. Sum of digits: 11319 [2008, 835]. American Mathematical Monthly, 116(3):283, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391085.

Straffin:2009:BRB

[2591] Philip D. Straffin. Book review: Mr. Hopkins’ Men: Cambridge Reform and British Mathematics in the 19th Century, by Alex D. D. Craik. American Mathematical Monthly, 116(3):284–288, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391086.

Anonymous:2009:BMd

[2592] Anonymous. Back matter. American Mathematical Monthly, 116(3):??, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391087.

Anonymous:2009:FMd

[2593] Anonymous. Front matter. American Mathematical Monthly, 116(3):??, March 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391064.

Barnsley:2009:TBS

[2594] Michael F. Barnsley. Transformations between self-referential sets. American Mathematical Monthly, 116(4):291–304, April 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391089.

Barmish:2009:RRC

[2595] B. Ross Barmish and Nigel Boston. Risk and return considerations in “the weakest link”. American Mathematical Monthly, 116(4):305–315, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391090.

Fejzic:2009:DRS

[2596] H. Fejzić, C. Freiling, and D. Rinne. Descartes’ Rule of Signs, alternations of data sets, and balanced differences. American Mathematical Monthly, 116(4):316–327, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391091.

Drinen:2009:OPS

[2597] D. Drinen, K. G. Kennedy, and W. M. Priestley. An optimization problem with a surprisingly simple solution. American Mathematical Monthly, 116(4):328–341, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391092.

Bloch:2009:SPG

[2598] Ethan D. Bloch. A simple proof of a generalized no retraction theorem. American Mathematical Monthly, 116(4):342–350, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391093.

Solomon:2009:CCS

[2599] Bruce Solomon. Central cross-sections make surfaces of revolution quadric. American Mathematical Monthly, 116(4):351–355, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391094.

Mercer:2009:FPI

[2600] Idris D. Mercer. On Furstenberg’s proof of the infinitude of primes. American Mathematical Monthly, 116(4):355–356, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391095.

Koliha:2009:MMM

[2601] J. J. Koliha. Mean, meaner, and the meanest mean value theorem. American Mathematical Monthly, 116(4):356–361, April 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391096.

Alkauskas:2009:CPF

[2602] Giedrius Alkauskas. A curious proof of Fermat’s little theorem. American Mathematical Monthly, 116(4):362–364, April 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391097.

Just:2009:Pb

[2603] Erwin Just. Problem 11425. American Mathematical Monthly, 116(4): 365, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391098.

Glasser:2009:Pa

[2604] M. L. Glasser. Problem 11426. American Mathematical Monthly, 116(4): 365, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391099.

Bandila:2009:P

[2605] Viorel Bǎndilǎ. Problem 11427. American Mathematical Monthly, 116 (4):365, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391100.

Blumberg:2009:P

[2606] Walter Blumberg. Problem 11428. American Mathematical Monthly, 116 (4):365, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391101.

Lupu:2009:Pb

[2607] Cezar Lupu. Problem 11429. American Mathematical Monthly, 116(4): 365–366, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391102.

Yi:2009:P

[2608] He Yi. Problem 11430. American Mathematical Monthly, 116(4):366, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391103.

Holland:2009:Pb

[2609] Finbarr Holland and Stephen Wills. Problem 11431. American Mathematical Monthly, 116(4):366, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391104.

Tetiva:2009:PVPb

[2610] Marian Tetiva and Robin Chapman. Partitioning values of a polynomial into sets of equal sum: 11266 [2007, 77]. American Mathematical Monthly, 116(4):366–368, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391105.

Sadov:2009:DSM

[2611] Sergey Sadov and Harris Kwong. The determinant of a spiral matrix: 11270 [2007, 78]. American Mathematical Monthly, 116(4):368–369, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391106.

Bluskov:2009:CCS

[2612] Iliya Bluskov. Covering the chessboard with sets of non-attacking queens: 11271 [2007, 164]. American Mathematical Monthly, 116(4):369–370, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391107.

Sadov:2009:DAS

[2613] Sergey Sadov, Kim McInturff, and Santa Barbara. Determinant of another spiral matrix: 11293 [2007, 451]. American Mathematical Monthly, 116(4):370–371, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391108.

Jonsson:2009:OTP

[2614] Jakob Jonsson, Annie Carter, and David Callan. Oddly triangulated polygons: 11298 [2007, 547]. American Mathematical Monthly, 116(4): 371–372, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391109.

Farrokhi:2009:CPM

[2615] M. Farrokhi and T. S. Michael. A characterization of permutation matrices: 11303 [2007, 548]. American Mathematical Monthly, 116(4):372, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391110.

Vick:2009:BRB

[2616] James W. Vick. Book review: Topology Now!, by Robert Messer; Philip Straffin; Introduction to Topology: Pure and Applied by Colin Adams; Robert Franzosa. American Mathematical Monthly, 116(4):373–375, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391111.

Tapp:2009:BRB

[2617] Kristopher Tapp. Book review: Differential Geometry and Its Applications, by John Oprea. American Mathematical Monthly, 116(4):375–377, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391112.

Velleman:2009:EEa

[2618] Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 116(4):378–382, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391113.

Anonymous:2009:BMe

[2619] Anonymous. Back matter. American Mathematical Monthly, 116(4): ??, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391114.

Anonymous:2009:FMe

[2620] Anonymous. Front matter. American Mathematical Monthly, 116(4): ??, April 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391088.

Borwein:2009:EBS

[2621] Jonathan M. Borwein, Neil J. Calkin, and Dante Manna. Euler–Boole summation revisited. American Mathematical Monthly, 116(5):387–412, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391116.

Bhargava:2009:GUP

[2622] Mira Bhargava. Groups as unions of proper subgroups. American Mathematical Monthly, 116(5):413–422, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391117.

Mabry:2009:CCP

[2623] Rick Mabry and Paul Deiermann. Of cheese and crust: a proof of the Pizza Conjecture and other tasty results. American Mathematical Monthly, 116(5):423–438, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391118.

Alamo:2009:PTT

[2624] Nieves Alamo and Carlos Criado. Parallel transport and Thomas–Wigner rotation. American Mathematical Monthly, 116(5):439–446, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391119.

Heath:2009:SAG

[2625] Daniel Heath, I. M. Isaacs, John Kiltinen, and Jessica Sklar. Symmetric and alternating groups generated by a full cycle and another element. American Mathematical Monthly, 116(5):447–451, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391120.

Evans:2009:RCO

[2626] Michael J. Evans and Paul D. Humke. Revisiting a century-old characterization of Baire class one, Darboux functions. American Mathematical Monthly, 116(5):451–455, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391121.

Bernhardt:2009:PCH

[2627] Chris Bernhardt. A proof of the Cayley–Hamilton theorem. American Mathematical Monthly, 116(5):456–457, May 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391122.

Wagner:2009:NCP

[2628] Peter Wagner. A new constructive proof of the Malgrange–Ehrenpreis Theorem. American Mathematical Monthly, 116(5):457–462, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391123.

Tetiva:2009:Pb

[2629] Marian Tetiva. Problem 11432. American Mathematical Monthly, 116(5): 463, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391124.

Cavachi:2009:Pb

[2630] Marius Cavachi. Problem 11433. American Mathematical Monthly, 116 (5):463, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391125.

Simic:2009:Pa

[2631] Slavko Simic. Problem 11434. American Mathematical Monthly, 116(5): 463, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391126.

Ligouras:2009:Pa

[2632] Panagiote Ligouras. Problem 11435. American Mathematical Monthly, 116(5):463, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391127.

Aliyev:2009:Pa

[2633] Y. N. Aliyev. Problem 11436. American Mathematical Monthly, 116(5): 463–464, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391128.

Erdelyi:2009:P

[2634] Tomás Erdélyi. Problem 11437. American Mathematical Monthly, 116 (5):464, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391129.

Bailey:2009:P

[2635] David H. Bailey. Problem 11438. American Mathematical Monthly, 116 (5):464, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391130.

Tetiva:2009:SPS

[2636] Marian Tetiva and Byron Schmuland. A sequence of partitions into sets with equal sum: 11273 [2007, 164]. American Mathematical Monthly, 116(5):464–465, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391131.

Alekseyev:2009:CPT

[2637] Max Alekseyev and John Henry Steelman. Counting permutations with two displacement values: 11281 [2007, 259]. American Mathematical Monthly, 116(5):465, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391132.

Abbot:2009:IPT

[2638] John Abbot and John Wilkinson. Irreducible polynomials from a theorem of fermat: 11283 [2007, 260]. American Mathematical Monthly, 116(5): 465–466, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391133.

Oman:2009:RF

[2639] Greg Oman and Manuel Reyes. Rings that are fields: 11284 [2007, 358]. American Mathematical Monthly, 116(5):466, May 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391134.

LeBrun:2009:DD

[2640] Marc LeBrun. Descending dungeons: 11286 [2007, 358]. American Mathematical Monthly, 116(5):466–467, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391135.

Hillar:2009:PI

[2641] Christopher Hillar and Said Amghibech. A polynomial identity: 11288 [2007, 359]. American Mathematical Monthly, 116(5):468, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391136.

Callan:2009:DCB

[2642] David Callan and Robin Chapman. Divisibility of a central binomial sum: 11292 & 11307 [2007, 451 & 640]. American Mathematical Monthly, 116(5):468–470, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391137.

Farrokhi:2009:PEI

[2643] M. Farrokhi and Tuomas Hytönen. The prime exponentiation of an integer: 11315 [2007, 744]. American Mathematical Monthly, 116(5): 470, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391138.

Holder:2009:BRB

[2644] Allen Holder. Book review: Understanding and Using Linear Programming, by Jiří Matoušek; Bernd Gärtner; Introduction to Optimization by Pablo Pedregal. American Mathematical Monthly, 116(5):471–476, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391139.

Anonymous:2009:BMf

[2645] Anonymous. Back matter. American Mathematical Monthly, 116(5): ??, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391140.

Anonymous:2009:FMf

[2646] Anonymous. Front matter. American Mathematical Monthly, 116(5): ??, May 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391115.

Crans:2009:MAD

[2647] Alissa S. Crans, Thomas M. Fiore, and Ramon Satyendra. Musical actions of dihedral groups. American Mathematical Monthly, 116(6):479– 495, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391142.

Barbanel:2009:CPP

[2648] Julius B. Barbanel, Steven J. Brams, and Walter Stromquist. Cutting a pie is not a piece of cake. American Mathematical Monthly, 116(6):496– 514, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391143.

Calcut:2009:GIA

[2649] Jack S. Calcut. Gaussian integers and arctangent identities for π. American Mathematical Monthly, 116(6):515–530, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391144.

Beanland:2009:MTF

[2650] Kevin Beanland, James W. Roberts, and Craig Stevenson. Modifications of Thomae’s function and differentiability. American Mathematical Monthly, 116(6):531–535, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391145.

Abel:2009:IBO

[2651] Ulrich Abel and Mircea Ivan. Over-iterates of Bernstein’s operators: a short and elementary proof. American Mathematical Monthly, 116 (6):535–538, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391146.

Levrie:2009:EPI

[2652] Paul Levrie and Walter Daems. Evaluating the probability integral using Wallis’s product formula for π. American Mathematical Monthly, 116(6):538–541, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391147.

Kuba:2009:CFC

[2653] Gerald Kuba. Counting fields of complex numbers. American Mathematical Monthly, 116(6):541–546, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391148.

Herschkorn:2009:P

[2654] Stephen Herschkorn. Problem 11439. American Mathematical Monthly, 116(6):547, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391149.

Siboni:2009:P

[2655] Stefano Siboni. Problem 11440. American Mathematical Monthly, 116(6): 547, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391150.

Aliyev:2009:Pb

[2656] Y. N. Aliyev. Problem 11441. American Mathematical Monthly, 116(6): 547, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391151.

Diaz-Barrero:2009:P

[2657] José Luis Díaz-Barrero and José Gibergans-Báguena. Problem 11442. American Mathematical Monthly, 116(6):547–548, June/July 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391152.

Ionascu:2009:P

[2658] Eugen Ionascu. Problem 11443. American Mathematical Monthly, 116(6):548, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391153.

Tetiva:2009:Pc

[2659] Marian Tetiva. Problem 11444. American Mathematical Monthly, 116(6): 548, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391154.

ShahAli:2009:P

[2660] H. A. ShahAli. Problem 11445. American Mathematical Monthly, 116(6): 548, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391155.

Knuth:2009:SPS

[2661] Donald Knuth and Julian Hook. Sums and powers, set counting, and coefficient tracking: 11274 [2007, 165]. American Mathematical Monthly, 116(6):548–549, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391156.

Dokovic:2009:UM

[2662] Dragomir Ž. – Doković and Richard Stong. A unipotent matrix: 11282 [2007, 260]. American Mathematical Monthly, 116(6):549–550, June/ July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391157.

Mkrtchyan:2009:DMS

[2663] Vahan Mkrtchyan and Codrut Grosu. Disjoint maximum-sized matchings: 11305 [2007, 640]. American Mathematical Monthly, 116(6):550– 551, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391158.

Just:2009:ECC

[2664] Erwin Just and Tom Kezlan. An exponential condition for commutativity: 11310 [2007, 641]. American Mathematical Monthly, 116(6):551–552, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391159.

Callan:2009:UDV

[2665] David Callan and Marc Renault. Uniformly distributed valleys: 11312 [2007, 744]. American Mathematical Monthly, 116(6):552–553, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391160.

Bataille:2009:QT

[2666] Michel Bataille and John Cade. From quadrilateral to trapezoid: 11317 [2007, 745]. American Mathematical Monthly, 116(6):553–554, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391161.

Grunbaum:2009:BRB

[2667] Branko Grünbaum. Book review: The Symmetries of Things, by John H. Conway; Heidi Burgiel; Chaim Goodman-Strauss. American Mathematical Monthly, 116(6):555–562, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391162.

Anonymous:2009:BMg

[2668] Anonymous. Back matter. American Mathematical Monthly, 116(6):??, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391163.

Anonymous:2009:FMg

[2669] Anonymous. Front matter. American Mathematical Monthly, 116(6):??, June/July 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391141.

Baruah:2009:RSS

[2670] Nayandeep Deka Baruah, Bruce C. Berndt, and Heng Huat Chan. Ramanujan’s series for 1/π: a survey. American Mathematical Monthly, 116(7):567–587, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391165.

Martin:2009:AAI

[2671] Greg Martin and Erick B. Wong. Almost all integer matrices have no integer eigenvalues. American Mathematical Monthly, 116(7):588– 597, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391166.

Apostol:2009:NIC

[2672] Tom M. Apostol and Mamikon A. Mnatsakanian. New insight into cycloidal areas. American Mathematical Monthly, 116(7):598–611, August/ September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391167.

Kenyon:2009:BP

[2673] Richard Kenyon and Peter Winkler. Branched polymers. American Mathematical Monthly, 116(7):612–628, August/September 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391168.

Hou:2009:TA

[2674] Shui-Hung Hou. On a theorem of Abel. American Mathematical Monthly, 116(7):629–630, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391169.

Sondow:2009:RPB

[2675] Jonathan Sondow. Ramanujan primes and Bertrand’s postulate. American Mathematical Monthly, 116(7):630–635, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391170.

Hedman:2009:LSD

[2676] Shawn Hedman and David Rose. Light subsets of N with dense quotient sets. American Mathematical Monthly, 116(7):635–641, August/ September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391171.

Sabia:2009:LPC

[2677] Juan Sabia and Susana Tesauri. The least prime in certain arithmetic progressions. American Mathematical Monthly, 116(7):641–643, August/ September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391172.

Deiser:2009:SCB

[2678] Oliver Deiser. A simple continuous bijection from natural sequences to dyadic sequences. American Mathematical Monthly, 116(7):643– 646, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391173.

Hillar:2009:Pb

[2679] Christopher Hillar. Problem 11446. American Mathematical Monthly, 116(7):647, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391174.

Furdui:2009:Pb

[2680] Ovidiu Furdui and Campia Turzii. Problem 11447. American Mathematical Monthly, 116(7):647, August/September 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391175.

Jiang:2009:P

[2681] Wei-Dong Jiang. Problem 11448. American Mathematical Monthly, 116 (7):647, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391176.

Bataille:2009:Pa

[2682] Michel Bataille. Problem 11449. American Mathematical Monthly, 116 (7):647, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391177.

Pohoata:2009:Pa

[2683] Cosmin Pohoata. Problem 11450. American Mathematical Monthly, 116 (7):647, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391178.

Oman:2009:P

[2684] Greg Oman. Problem 11451. American Mathematical Monthly, 116(7): 648, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391179.

Knuth:2009:P

[2685] Donald Knuth. Problem 11452. American Mathematical Monthly, 116 (7):648, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391180.

Witula:2009:IC

[2686] Roman Witula, Damian Stota, and Szeged Problem Group ‘Fejéntaláltuka’. An iteration convergence: 11318 [2007, 745]. American Mathematical Monthly, 116(7):648–649, August/September 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391181.

Knuth:2009:SRB

[2687] Donald Knuth and O. P. Lossers. Solving a recurrence by binary expansion: 11320 [2007, 835]. American Mathematical Monthly, 116(7): 649, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391182.

Hillar:2009:PAS

[2688] Christopher Hillar, Bruce S. Burdick, and Roger Williams. Polynomials not associated with symmetric matrices: 11321 [2007, 835]. American Mathematical Monthly, 116(7):649–650, August/September 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391183.

Sondow:2009:DI

[2689] Jonathan Sondow and George Lamb Tucson. A double integral: 11322 [2007, 835]. American Mathematical Monthly, 116(7):650, August/ September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391184.

Lockhart:2009:SIP

[2690] Jody M. Lockhart, William P. Wardlaw, and O. P. Lossers. On the sum of integer powers of a matrix: 11323 [2007, 835]. American Mathematical Monthly, 116(7):650–651, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391185.

Kochanski:2009:MGR

[2691] Martin Kochanski and Richard Stong. Measuring the growth rate of a Bernoulli polynomial: 11324 [2007, 836]. American Mathematical Monthly, 116(7):651–653, August/September 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391186.

Holland:2009:TI

[2692] Finbarr Holland and Michael Vowe. A triangle inequality from 1953: 11325 [2007, 836]. American Mathematical Monthly, 116(7):653– 654, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391187.

Diaz-Barrero:2009:PRI

[2693] José Luis Díaz-Barrero and Pál Péter Dályay. A problem rescued from incorrectness: 11326 [2007, 836]. American Mathematical Monthly, 116(7): 654–655, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391188.

Mabry:2009:CF

[2694] Rick Mabry, Debbie Shepherd, and Kenneth Schilling. A chancy function: 11327 [2007, 925]. American Mathematical Monthly, 116(7):655– 656, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391189.

Vartziotis:2009:VAT

[2695] Dmitris Vartziotis and Michel Bataille. Van Aubel’s theorem: 11328 [2007, 925]. American Mathematical Monthly, 116(7):656, August/ September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391190.

Falkner:2009:BRB

[2696] Neil Falkner. Book review: Basic Real Analysis, by Anthony W. Knapp; Advanced Real Analysis by Anthony W. Knapp. American Mathematical Monthly, 116(7):657–664, August/September 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391191.

Anonymous:2009:BMh

[2697] Anonymous. Back matter. American Mathematical Monthly, 116(7): ??, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391192.

Anonymous:2009:FMh

[2698] Anonymous. Front matter. American Mathematical Monthly, 116(7): ??, August/September 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391164.

Daugherty:2009:VSG

[2699] Zajj Daugherty, Alexander K. Eustis, Gregory Minton, and Michael E. Orrison. Voting, the symmetric group, and representation theory. American Mathematical Monthly, 116(8):667–687, October 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391194.

Sam:2009:BPT

[2700] Steven V. Sam. A bijective proof for a theorem of ehrhart. American Mathematical Monthly, 116(8):688–701, October 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391195.

Provan:2009:SSV

[2701] J. Scott Provan. Sudoku: Strategy versus structure. American Mathematical Monthly, 116(8):702–707, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/40391196.

Bukatin:2009:PMS

[2702] Michael Bukatin, Ralph Kopperman, Steve Matthews, and Homeira Pajoohesh. Partial metric spaces. American Mathematical Monthly, 116(8):708–718, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391197.

Klosinski:2009:SNW

[2703] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Sixty-Ninth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 116(8):719–726, October 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391198.

Brown:2009:W

[2704] Ezra Brown and Nicholas Loehr. Why is PSL(2, 7) ≅ GL(3, 2)? American Mathematical Monthly, 116(8):727–732, October 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391199.

Feldman:2009:AP

[2705] David V. Feldman and Daniel A. Klain. Angles as probabilities. American Mathematical Monthly, 116(8):732–735, October 2009. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/40391200.

Csiszar:2009:JHS

[2706] Villö Csiszár. Jump home and shift: An acyclic operation on permutations. American Mathematical Monthly, 116(8):735–738, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391201.

Lee:2009:PME

[2707] Hojoo Lee. Ptolemy meets Erdős and Mordell again. American Mathematical Monthly, 116(8):738–742, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/40391202.

Bonciocat:2009:IPL

[2708] Anca I. Bonciocat and Nicolae C. Bonciocat. On the irreducibility of polynomials with leading coefficient divisible by a large prime power. American Mathematical Monthly, 116(8):743–745, October 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391203.

Bataille:2009:Pb

[2709] Michel Bataille. Problem 11449. American Mathematical Monthly, 116(8):746, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391204.

Stanley:2009:P

[2710] Richard Stanley. Problem 11453. American Mathematical Monthly, 116(8):746, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391205.

Kerimov:2009:P

[2711] Azer Kerimov. Problem 11454. American Mathematical Monthly, 116(8): 746, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391206.

Blatter:2009:P

[2712] Christian Blatter. Problem 11455. American Mathematical Monthly, 116(8):746, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391207.

Mortini:2009:P

[2713] Raymond Mortini. Problem 11456. American Mathematical Monthly, 116(8):747, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391208.

Glasser:2009:Pb

[2714] M. L. Glasser. Problem 11457. American Mathematical Monthly, 116(8): 747, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391209.

Lupu:2009:Pc

[2715] Cezar Lupu. Problem 11458. American Mathematical Monthly, 116(8): 747, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391210.

Dalyay:2009:P

[2716] Pál Péter Dályay. Problem 11459. American Mathematical Monthly, 116(8):747, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391211.

Amdeberhan:2009:TLG

[2717] T. Amdeberhan, Victor H. Moll, and Hongwei Chen. Two log gamma integrals: 11329 [2007, 925]. American Mathematical Monthly, 116(8):747– 748, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391212.

Tetiva:2009:NLB

[2718] Marian Tetiva and João Guerreiro. A new lower bound for the sum of the altitudes: 11330 [2007, 925]. American Mathematical Monthly, 116(8):748–749, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391213.

Tetiva:2009:BI

[2719] Marian Tetiva and Peter Nüesch. A bisector inequality: 11337 [2008, 71]. American Mathematical Monthly, 116(8):749–750, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391214.

Furdui:2009:GL

[2720] Ovidiu Furdui and O. P. Lossers. A gamma limit: 11338 [2008, 71]. American Mathematical Monthly, 116(8):750, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391215.

Ciaurri:2009:RFS

[2721] Óscar Ciaurri, Luz Roncal, and Nora Thornber. Raindrops flying from a spinning umbrella: 11340 [2008, 72]. American Mathematical Monthly, 116(8):750–752, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391216.

Lupu:2009:API

[2722] Cezar Lupu and Marian Tetiva. An application of Popovidu’s inequality: 11341 [2008, 166]. American Mathematical Monthly, 116(8):752–753, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391217.

Cuculiere:2009:FE

[2723] Roger Cuculière and Richard Stong. A functional equation: 11345 [2008, 166]. American Mathematical Monthly, 116(8):753, October 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391218.

Holdener:2009:BRB

[2724] Judy Holdener. Book review: Roots to Research: A Vertical Development of Mathematical Problems, by Judith D. Sally; Paul J. Sally, Jr. American Mathematical Monthly, 116(8):754–758, October 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391219.

Anonymous:2009:BMi

[2725] Anonymous. Back matter. American Mathematical Monthly, 116(8):??, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391220.

Anonymous:2009:FMi

[2726] Anonymous. Front matter. American Mathematical Monthly, 116(8):??, October 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391193.

Paterson:2009:MO

[2727] Mike Paterson, Yuval Peres, Mikkel Thorup, Peter Winkler, and Uri Zwick. Maximum overhang. American Mathematical Monthly, 116 (9):763–787, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391297.

Diaconis:2009:CSA

[2728] Persi Diaconis and Jason Fulman. Carries, shuffling, and an amazing matrix. American Mathematical Monthly, 116(9):788–803, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391298.

Verde-Star:2009:RF

[2729] Luis Verde-Star. Rational functions. American Mathematical Monthly, 116(9):804–827, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391299.

deJong:2009:LMF

[2730] Theo de Jong. Lagrange multipliers and the fundamental theorem of algebra. American Mathematical Monthly, 116(9):828–830, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391300.

Einstein:2009:SDB

[2731] D. M. Einstein, C. C. Heckman, and T. S. Norfolk. On Sara’s Dove bar habit. American Mathematical Monthly, 116(9):831–835, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391301.

Farhi:2009:IIL

[2732] Bakir Farhi. An identity involving the least common multiple of binomial coefficients and its application. American Mathematical Monthly, 116(9):836–839, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391302.

DeAngelis:2009:SSR

[2733] Valerio De Angelis. Stirling’s series revisited. American Mathematical Monthly, 116(9):839–843, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391303.

Pohoata:2009:Pb

[2734] Cosmin Pohoaţǎ. Problem 11460. American Mathematical Monthly, 116(9):844, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391304.

Ligouras:2009:Pb

[2735] Panagiote Ligouras. Problem 11461. American Mathematical Monthly, 116(9):844, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391305.

Alexandrova:2009:P

[2736] Nadezhda Alexandrova. Problem 11462. American Mathematical Monthly, 116(9):844, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391306.

Chang:2009:Pa

[2737] Xiang Qian Chang. Problem 11463. American Mathematical Monthly, 116(9):844, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391307.

Beckwith:2009:P

[2738] David Beckwith. Problem 11464. American Mathematical Monthly, 116(9):845, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391308.

Popescu:2009:P

[2739] Pantelimon George Popescu. Problem 11465. American Mathematical Monthly, 116(9):845, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391309.

Yongge:2009:P

[2740] Tian Yongge. Problem 11466. American Mathematical Monthly, 116(9): 845, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391310.

Bagby:2009:WSR

[2741] Richard Bagby, David Beckwith, Sag Harbor, and Douglas B. Tyler. A weighted sample of the Riemann zeta function: 11331 [2007, 926]. American Mathematical Monthly, 116(9):845–846, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391311.

Tauraso:2009:IS

[2742] Roberto Tauraso and Kenneth Schilling. Irrational symmetry: 11332 [2007, 926]. American Mathematical Monthly, 116(9):846–847, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391312.

Refolio:2009:BPF

[2743] Pablo Fernández Refolio and John B. Zacharias. Baking a pi with factorials: 11333 [2007, 926]. American Mathematical Monthly, 116(9):847, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391313.

Bober:2009:VCN

[2744] Jonathan Bober, Jeffrey Lagarias, and Byron Schmuland. Very composite numbers: 11334 [2008, 71]. American Mathematical Monthly, 116(9):847–848, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391314.

Knuth:2009:NBC

[2745] Donald Knuth and GCHQ Problem Solving Group. Near-de Bruijn cycles: 11336 [2008, 71]. American Mathematical Monthly, 116(9):848–849, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391315.

Bencze:2009:MAI

[2746] Mihály Bencze and Julien Grivaux. More AGM inequalities: 11347 [2008,167]. American Mathematical Monthly, 116(9):849–851, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391316.

Katz:2009:BRB

[2747] Victor J. Katz. Book review: Mathematical Masterpieces: Further Chronicles by the Explorers, by Arthur Knoebel; Reinhard Laubenbacher; Jerry Lodder; David Pengelley. American Mathematical Monthly, 116(9):852–856, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391317.

Velleman:2009:EEb

[2748] Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 116(9):857–858, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391318.

Anonymous:2009:BMj

[2749] Anonymous. Back matter. American Mathematical Monthly, 116(9):??, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391319.

Anonymous:2009:FMj

[2750] Anonymous. Front matter. American Mathematical Monthly, 116(9):??, November 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391296.

Lax:2009:RLI

[2751] Peter D. Lax. Rethinking the Lebesgue integral. American Mathematical Monthly, 116(10):863–881, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391247.

Laubenbacher:2009:CAS

[2752] Reinhard Laubenbacher and Bernd Sturmfels. Computer algebra in systems biology. American Mathematical Monthly, 116(10):882–891, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391248.

Palais:2009:DLE

[2753] Bob Palais, Richard Palais, and Stephen Rodi. A disorienting look at Euler’s theorem on the axis of a rotation. American Mathematical Monthly, 116(10):892–909, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391249.

Richmond:2009:HRP

[2754] Bettina Richmond and Tom Richmond. How to recognize a parabola. American Mathematical Monthly, 116(10):910–922, December 2009. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/40391250.

Beals:2009:OSA

[2755] Robert Beals. On orders of subgroups in Abelian groups: An elementary solution of an exercise of Herstein. American Mathematical Monthly, 116(10):923–926, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391251.

Holroyd:2009:NSC

[2756] Alexander E. Holroyd and Terry Soo. A nonmeasurable set from coin flips. American Mathematical Monthly, 116(10):926–928, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/40391252.

Brillhart:2009:NEF

[2757] John Brillhart. A note on Euler’s factoring problem. American Mathematical Monthly, 116(10):928–931, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/40391253.

Keller:2009:V

[2758] Joseph B. Keller and Ravi Vakil. π_{p}, the value of π in ℓ_{p}. American Mathematical Monthly, 116(10):931–935, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391254.

Berrone:2009:APL

[2759] Lucio R. Berrone. The associativity of the Pythagorean Law. American Mathematical Monthly, 116(10):936–939, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/40391255.

Chang:2009:Pb

[2760] Xiang Qian Chang. Problem 11467. American Mathematical Monthly, 116(10):940, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391256.

Pohoata:2009:Pc

[2761] Cosmin Pohoata. Problem 11468. American Mathematical Monthly, 116(10):940, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391257.

Simic:2009:Pb

[2762] Slavko Simic. Problem 11469. American Mathematical Monthly, 116(10): 940, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391258.

Tetiva:2009:Pd

[2763] Marian Tetiva. Problem 11470. American Mathematical Monthly, 116(10):941, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391259.

Holland:2009:Pc

[2764] Finbarr Holland. Problem 11471. American Mathematical Monthly, 116(10):941, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391260.

Makhul:2009:P

[2765] Mahdi Makhul. Problem 11472. American Mathematical Monthly, 116(10):941, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391261.

Perfetti:2009:Pb

[2766] Paolo Perfetti. Problem 11473. American Mathematical Monthly, 116(10):941, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391262.

Gonzalez:2009:MWS

[2767] Juan López González and O. P. Lossers. Multiples of 6 whose successors have greater divisor sum: 11335 [2008, 71]. American Mathematical Monthly, 116(10):941–942, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391263.

Diaz-Barrero:2009:FI

[2768] José Luis Díaz-Barrero and Mihai Cipu. A Fibonacci inequality: 11339 [2008, 72]. American Mathematical Monthly, 116(10):942–943, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391264.

Gallardo:2009:SFF

[2769] Luis H. Gallardo and Sam Vandervelde. A square in a finite field: 11342 [2008, 166]. American Mathematical Monthly, 116(10):943–944, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391265.

Beckwith:2009:NBS

[2770] David Beckwith and CMC 328. The number of balanced senates from n states: 11343 [2008, 166]. American Mathematical Monthly, 116(10):944– 945, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://www.jstor.org/stable/40391266.

Stadler:2009:MFI

[2771] Albert Stadler and Allen Stenger. A Möbius function identity: 11344 [2008, 166]. American Mathematical Monthly, 116(10):945–946, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391267.

Anonymous:2009:EEN

[2772] Anonymous. Errata and end notes for 2008. American Mathematical Monthly, 116(10):947, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/40391268.

Henle:2009:BRB

[2773] Michael Henle. Book review: Music: A Mathematical Offering, by David J. Benson; Musimathics: The Mathematical Foundations of Music by Gareth Loy. American Mathematical Monthly, 116(10):948–955, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/40391269.

Anonymous:2009:BMb

[2774] Anonymous. Back matter. American Mathematical Monthly, 116(10):??, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391270.

Anonymous:2009:FMb

[2775] Anonymous. Front matter. American Mathematical Monthly, 116(10):??, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/40391246.

Anonymous:2009:VI

[2776] Anonymous. Volume information. American Mathematical Monthly, 116(10):??, December 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/stable/

40391245.